### 29 Jan 2024 v2

# An Elliott intertwining approach to classifying actions of $\mathrm{C}^{*}$-tensor categories

Abstract.
We introduce a categorical approach to classifying actions of $*{}^{*}$-tensor categories $\mathcal{C}$ on $\mathrm{C}^{*}$-algebras up to cocycle conjugacy. We show that, in this category, inductive limits exist and there is a natural notion of approximate unitary equivalence. Then, we generalise classical Elliott intertwining results to the $\mathcal{C}$-equivariant case, in the same fashion as done by Szabó for the group equivariant case in [40].
Introduction
The study of existence and classification of symmetries on operator algebras has been a ubiquitous theme in the field. In the case of von Neumann algebras, this can be traced back to the work of Connes ([7, 9]), where he classified automorphisms of the hyperfinite $\mathrm{II}_{1}$ factor $\mathcal{R}$ up to outer conjugacy. Subsequently, Jones classified actions of finite groups on $\mathcal{R}$ ([24]), while Ocneanu generalised these results to the case of amenable groups acting on $\mathcal{R}$ ([32]).
Later, Popa classified amenable subfactors $N\subset\mathcal{R}$ by their standard invariant ([36]). Following subsequent reformulations ([29, 30]), the standard invariant of a finite index subfactor $N\subset\mathcal{R}$ can be understood as a pair $(F,Q)$ with $F$ an action of a unitary tensor category $\mathcal{C}$ on $N$ and $Q$ a special object in $\mathcal{C}$ that allows one to recover the inclusion $N\subset\mathcal{R}$ as a generalised crossed product. [^1] Precisely, $Q$ is a $Q$-system as defined by Longo in [29]. The general setting of $\mathrm{C}^{*}$-tensor categories represents a unifying framework for studying group actions and more general quantum symmetries arising from subfactor theory. Following the recent spectacular classification results for group actions on Kirchberg algebras by group equivariant KK-theory ([17]), and the development of a $\mathcal{C}$-equivariant KK-theory ([1]), it is natural to consider to what extent one can classify actions of more general $*{}^{*}$-tensor categories on simple, amenable $*{}^{*}$-algebras.
Prior and closely linked to the classification of symmetries of operator algebras was the classification of the operator algebras themselves. First, Connes classified amenable factors in [8] with the exception of one type which was later completed by Haagerup ([22]). In the case of $*{}^{*}$-algebras, Elliott classified $*{}^{*}$-inductive limits of finite dimensional $*{}^{*}$-algebras by their ordered K-theory in [10]. Elliott’s methods laid the foundations for a roadmap to a classification of more general simple, amenable $*{}^{*}$-algebras. Following Elliott’s strategy and building on decades of work by many mathematicians, the classification programme of simple, amenable $\mathrm{C}^{*}$-algebras successfully culminated in [27, 35, 20, 21, 41, 5].
The general strategy for a classification of operator algebras is to first achieve existence and uniqueness results for morphisms with respect to the proposed classification invariant. Loosely speaking, to construct an isomorphism between two objects $A$ and $B$ in the category of $\mathrm{C}^{*}$-algebras, by virtue of an existence type statement, one obtains morphisms $\phi:A\to B$ and $\psi:B\to A$ which induce mutually inverse elements at the level of the chosen abstract invariant. Then, a suitable uniqueness result will give that $\psi\circ\phi$ is equivalent to $\mathrm{id}_{A}$ and $\phi\circ\psi$ is equivalent to $\mathrm{id}_{B}$. In the category of separable $\mathrm{C}^{*}$-algebras, the required notion of equivalence is approximate unitary equivalence. The final step is to perform a so-called approximate intertwining argument. Essentially, one can tweak the morphisms $\phi$ and $\psi$ by unitaries until they become mutually inverse isomorphisms (see [38, Corollary 2.3.4] and the introduction of [40] for a more detailed breakdown of this roadmap to classification). Furthermore, in [12], Elliott proposes a similar classification strategy for arbitrary objects in more general categories that have a suitable notion of inner automorphisms.
Instances of the implementation of Elliott’s strategy for the classification of compact (quantum) group actions on $*{}^{*}$-algebras appear in [18, 2], where the actions are assumed to have the Rokhlin property. In [40], Szabó articulates Elliott’s proposed strategy in the generality of actions of locally compact groups as an alternative to the Evans–Kishimoto intertwining type arguments (see [13]). In his work, Szabó defines an appropriate category of $\Gamma$$*{}^{*}$-algebras, that is $*{}^{*}$-algebras carrying an action of $\Gamma$, and produces a vast collection of intertwining results for this category. Szabó’s construction is a key ingredient facilitating the recent groundbreaking classification of amenable actions on Kirchberg algebras in [17].
In this paper, we will develop the necessary techniques to perform approximate intertwining arguments for $*{}^{*}$-algebras carrying an action of a tensor category $\mathcal{C}$ ($\mathcal{C}$-$\mathrm{C}^{*}$-algebras). Glimpses of these techniques can be found in [6], where exact intertwining arguments are performed to classify inductive limit actions of fusion categories on AF-algebras. However, for classification, it is often necessary to develop a more general framework which allows to perform approximate intertwining arguments. In the setting of $\mathrm{C}^{*}$-algebras, Elliott used exact intertwining arguments to classify AF-algebras ([10]), but more refined approximate versions were needed to classify simple $A\mathbb{T}$-algebras in [11]. Approximate intertwining arguments in the setting of actions of tensor category have already appeared in the literature; see [42] for an adaptation of the Evans-Kishimoto intertwining argument introduced in [13].
Primarily, the categorical framework developed by Szabó in [40] provides the conceptual skeleton for our construction. However, unlike in the group action case, a tensor category might not act by automorphisms, so our techniques differ. In [40], Szabó defines a cocycle morphism as a pair consisting of an extendible $*{}^{*}$-homomorphism and a unitary cocycle. In the setting of tensor categories, the action on a $\mathrm{C}^{*}$-algebra is given by a family of bimodules acting compatibly with respect to the tensor product. Therefore, the cocycles will be given by certain bimodule maps satisfying a family of commuting diagrams. Hence, it is not apparent how to adapt Szabó’s arguments.
[…]
In $\mathrm{C}^{*}_{\mathcal{C}}$, we are able to construct inductive limits and define a suitable topology on the space of morphisms. An important challenge in constructing a category of cocycle morphisms is the composition. Composing two cocycle morphisms might not give a cocycle morphism. Therefore, we need to introduce a slightly different composition which will allow us to obtain a category. If the morphisms are non-degenerate, this new composition agrees with the canonical one, and with the one introduced by Szabó in the group action case ([40, Proposition 1.15]). Moreover, to avoid the need of restricting to non-degenerate morphisms, the maps $\mathbbm{v}_{X}$ are assumed to be isometries. This assumption allows us, unlike in [40], to work in the possibly non-extendible setting, generalising the notion of a cocycle morphism from [40] when restricting to twisted group actions.
An essential observation for our construction is that we can encode the information of a cocycle morphism into a family of linear maps satisfying certain conditions (see [6, Lemma 3.8] where this is done for unital injective cocycle morphisms). Therefore, a cocycle morphism can be equivalently defined as follows.
Definition A (Lemma 3.11).
Let $\mathcal{C}$ be a $\mathrm{C}^{*}$-tensor category acting on $\mathrm{C}^{*}$-algebras $A$ and $B$ via $(A,F,J)$ and $(B,G,I)$ respectively. Then a cocycle morphism $(\phi,h):(A,F,J)\rightarrow(B,G,I)$ is given by a $*{}^{*}$-homomorphism $\phi:A\to B$ and a family of linear maps:
$h=\{h^{X}:F(X)\rightarrow G(X)\}_{X\in\mathcal{C}}$
such that for any $X,Y\in\mathcal{C}$
  1. (i)
    $h^{X}(a\rhd x\lhd a^{\prime})=\phi(a)\rhd h^{X}(x)\lhd\phi(a^{\prime})$ for any $a,a^{\prime}\in A$ and any $x\in F(X)$;
  2. (ii)
    for any morphism $f\in\mathrm{Hom}(X,Y)$, $G(f)\circ h^{X}=h^{Y}\circ F(f)$;
  3. (iii)
    $\phi(\langle x,y\rangle_{A})=\langle h^{X}(x),h^{X}(y)\rangle_{B}$ for any $x,y\in F(X)$;
  4. (iv)
    the following diagram commutes:
[…]
  5. (v)
[…]
In the case of a group action, we give an explicit formula for the family of linear maps $\{h^{X}\}_{X\in\mathcal{C}}$ in Example 3.14. Compared with the usual definition of a cocycle morphism given by pairs $(\phi,\mathbbm{v})$ consisting of a $*{}^{*}$-homomorphism and a family of bimodule maps $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ (see Definition 3.1), Definition A provides a more direct framework for setting up the intertwining arguments. In particular, to compose two cocycle morphisms, one composes the $*{}^{*}$-homomorphisms, as well as the corresponding linear maps. Furthermore, the natural topology on cocycle morphisms is obtained by considering pointwise differences of the linear maps. When the acting category is semisimple and has countably many isomorphism classes of simple objects (which is assumed for the rest of the introduction), we can phrase convergence in this topology purely in terms of the linear maps.
Definition B (Lemma 4.9).
Let $(\phi_{\lambda},h_{\lambda}),(\phi,h):(A,F,J)\to(B,G,I)$ be cocycle morphisms. Then $(\phi_{\lambda},h_{\lambda})$ converges to $(\phi,h)$ if and only if for any $X\in\mathrm{Irr}(\mathcal{C})$, $h_{\lambda}^{X}$ converges pointwise to $h^{X}$ in the norm induced by the right inner product. [^4] The collection $\mathrm{Irr}(\mathcal{C})$ is a choice of representatives for isomorphism classes of simple objects in $\mathcal{C}$.
In general, this topology is coarser than the one used by Szabó in the group action case ([40, Definition 2.5]) (see Example 4.10). However, it is the same when restricted to non-degenerate cocycle morphisms. With this topology in hand, we can then define approximate unitary conjugation for cocycle morphisms. Moreover, unlike Szabó’s definition in [40, Definition 2.8], this is a symmetric relation (see Lemma 4.16), which simplifies the intertwining arguments. This can be formulated as follows.
Definition C (Lemma 4.14).
If $(\phi,h),(\psi,l):(A,F,J)\to(B,G,I)$ are cocycle morphisms, then $(\psi,l)$ is approximately unitarily equivalent to $(\phi,h)$ if and only if there exists a net of unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
[…]
for any $X\in\mathrm{Irr}(\mathcal{C})$ and any $x\in F(X)$.
With the introduced topology at the level of morphism spaces and notion of unitary equivalence, we can place the subcategory of $\mathrm{C}^{*}_{\mathcal{C}}$ consisting of separable $\mathcal{C}$-$\mathrm{C}^{*}$-algebras and extendible cocycle morphisms in Elliott’s abstract classification framework from [12]. Precisely, the quotient category of $\mathrm{C}^{*}_{\mathcal{C}}$ by approximate unitary equivalence is a classification category in the sense of [12]. Hence, we get the following theorem.
Theorem D (Corollary 5.6).
Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras. Let
$(\phi,h):(A,F,J)\to(B,G,I)\quad and\quad(\psi,l):(B,G,I)\to(A,F,J)$
be two extendible cocycle morphisms such that the compositions $(\psi,l)\circ(\phi,h)$ and $(\phi,h)\circ(\psi,l)$ are approximately inner. Then $(\phi,h)$ and $(\psi,l)$ are approximately unitarily equivalent to mutually inverse cocycle conjugacies.
Along with Theorem D, in Section 5 we use our construction to obtain various $\mathcal{C}$-equivariant versions of classical intertwining arguments. For instance, we establish a general Elliott two-sided intertwining argument (see [38, Proposition 2.3.2] for a non-equivariant version) as well as an intertwining through reparametrisation in Theorem 5.9.
In Section 6, the techniques we develop allow us to perfom a $\mathcal{C}$-equivariant one-sided intertwining argument (see Theorem 6.1), generalising the classical one-sided intertwining arguments of [38, Proposition 2.3.5] and [40, Proposition 4.3]. This result is a key ingredient in the work of Evington, C. Jones, and the first named author in obtaining a McDuff-type characterisation of equivariant $\mathcal{D}$-stability of an action of a unitary tensor category, for strongly self-absorbing $\mathcal{D}$ ([15]). Precisely, $\mathcal{D}$-stability of an action is equivalent to the existence of a unital embedding of $\mathcal{D}$ into an appropriate subalgebra of the Kirchberg central sequence algebra of $A$ (in the case of a group action, this subalgebra coincides with the fixed point subalgebra of Kirchberg’s central sequence algebra). This generalises the classical result of [43, Theorem 2.2] (see also [27]) and its group equivariant counterpart (see [39, Corollary 3.8]). The remaining part of Section 6 is concerned with asymptotic versions of the results in Section 5.
1. Preliminaries
1.1. Hilbert bimodules
In this subsection, we collect a few basics on the theory of Hilbert bimodules. We refer the reader to [28] for a more detailed exposition. First, we start by recalling the definition of a Hilbert module as introduced by Paschke in [34].
Definition 1.1.
Let $X$ be a vector space over $\mathbb{C}$ and $B$ be a $\mathrm{C}^{*}$-algebra. We say that $X$ is a (right-)Hilbert $B$-module if $X$ is a right $B$-module equipped with a function $\langle\cdot,\cdot\rangle_{B}:X\times X\to B$ satisfying the following properties:
  1. (i)
    $\langle\cdot,\cdot\rangle_{B}$ is left conjugate linear and right linear.
  2. (ii)
    For any $x,y\in X$ and $b\in B$, one has that $\langle x,yb\rangle_{B}=\langle x,y\rangle_{B}b$.
  3. (iii)
    For any $x\in X$, $\langle x,x\rangle_{B}\geq 0$ and $\langle x,x\rangle_{B}=0$ if and only if $x=0$.
  4. (iv)
    For any $x,y\in X$, $\langle x,y\rangle_{B}=\langle y,x\rangle_{B}^{*}$.
  5. (v)
    $X$ is complete with respect to the norm induced by $\|\langle x,x\rangle_{B}\|^{1/2}$.
If $X$ only satisfies the properties (i)-(iv) above, then we say $X$ is a pre-Hilbert $B$-module. For $x\in X$, we denote by $|x|^{2}$ the positive element of $B$ given by $\langle x,x\rangle_{B}$. To obtain a right-Hilbert $A$-$B$-bimodule, we need a compatible left action of a $\mathrm{C}^{*}$-algebra $A$.
Definition 1.2.
([25, Definition 2.1]) Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras and $X={}_{A}X_{B}$ a bimodule over the complex algebras $A$ and $B$. We say that $X$ is a (right-)Hilbert $A$-$B$-bimodule if $X$ is a right-Hilbert $B$-module, and for all $a\in A$, the map $\phi(a):x\in X\mapsto ax\in X$ is adjointable, with adjoint $\phi(a)^{*}=\phi(a^{*})$.
In particular, $\phi:a\in A\to\phi(a)\in\mathcal{L}(X)$ is a $*{}^{*}$-homomorphism from $A$ to the $*{}^{*}$-algebra $\mathcal{L}(X)$ of adjointable maps on $X$. The map $\phi$ will be referred to as the left action of $A$ on $X$. For a Hilbert $A$-$B$-bimodule $X$, we often write the left action of an element $a\in A$ on a vector $x\in X$ by $a\rhd x$ and the right action of an element $b\in B$ on $x$ by $x\lhd b$.
Example 1.3.
For a $*{}^{*}$-algebra $B$, the vector space $X=B$ is a Hilbert $B$-module with the right $B$-module structure given by multiplication and the inner product $\langle a,b\rangle_{B}=a^{*}b$ for $a,b\in B$. In this case, any $*{}^{*}$-homomorphism $\phi:A\rightarrow M(B)$ induces a right-Hilbert $A$-$B$-bimodule which we denote by ${}_{\phi}B$.
Let $A,B,C$ be $\mathrm{C}^{*}$-algebras. If $X$ is a Hilbert $A$-$B$-bimodule and $E$ is a Hilbert $B$-$C$-bimodule we may form their internal tensor product $X\boxtimes E$ that is an Hilbert $A$-$C$-bimodule. We sketch this construction and refer to [28, Section 4] for details. To perform the internal tensor product one starts by considering the algebraic tensor product of vector spaces $X\odot E$. We identify the elements of the form $x\lhd b\odot y$ with $x\odot b\rhd y$ to form the quotient
$V=X\odot E/\mathrm{span}\{x\lhd b\odot y-x\odot b\rhd y:\ x\in X,\ y\in E,\ b\in B\}.$
We denote the image of the elementary tensor $x\odot y$ of $X\odot E$ in $V$ under the canonical quotient map by $x\boxtimes y$. One may define a right $C$-action and a right $C$-inner product on $V$ by
$(x\boxtimes y)\lhd c=x\boxtimes(y\lhd c),$
$\langle x\boxtimes y,z\boxtimes w\rangle_{C}=\langle x,\langle y,z\rangle_{B}\rhd w\rangle_{C},$
for any $x,z\in X$ and $z,w\in E$. It follows that $V$ equipped with this $C$-action and $C$-valued inner product satisfies (i)-(iv) of Definition 1.1. We produce a Hilbert $C$-module $X\boxtimes E$ by completing $V$ under the norm defined by the inner product. Moreover, one can induce a left $A$ action on $X\boxtimes E$ through
$a\rhd(x\boxtimes y)=(a\rhd x)\boxtimes y$
for all $a\in A$, $x\in X$ and $y\in E$. This equips $X\boxtimes E$ with the structure of a Hilbert $A$-$C$-bimodule.
If the bimodules are given by $*{}^{*}$-homomorphisms into the multiplier algebra as in Example 1.3, one may get greater insight into the structure of their tensor product. First we recall the definition of non-degenerate and extendible $*{}^{*}$-homomorphisms.
Definition 1.4.
Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras and $\phi:A\to\mathcal{M}(B)$ a $*{}^{*}$-homomorphism. Then $\phi$ is said to be non-degenerate if $\phi(A)B$ is dense in $B$.
If $\phi:A\to\mathcal{M}(B)$ is non-degenerate, then $\phi$ extends to a unital $*{}^{*}$-homomorphism $\phi:\mathcal{M}(A)\to\mathcal{M}(B)$ ([28, Proposition 2.5]).
Definition 1.5.
Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras. A $*{}^{*}$-homomorphism $\phi:A\to\mathcal{M}(B)$ is called extendible if for any increasing approximate unit $e_{\lambda}\in A$, the net $\phi(e_{\lambda})\in\mathcal{M}(B)$ converges strictly to a projection $p\in\mathcal{M}(B)$. [^5] Any non-degenerate $*{}^{*}$-homomorphism $\phi:A\rightarrow M(B)$ is extendible as for an approximate unit $e_{\lambda}\in A$ the net $\phi(e_{\lambda})$ converges strictly to $1_{\mathcal{M}(B)}$ (see e.g. [3, II.7.3.8.]). As discussed in [40, Definition 1.7], $\phi$ then factorises through $\mathcal{M}(pBp)\cong p\mathcal{M}(B)p\subseteq\mathcal{M}(B)$, with the $*{}^{*}$-homomorphism $\phi_{p}:A\to\mathcal{M}(pBp)$ now non-degenerate. In this case, following [40], we let $\phi^{\dagger}:\mathcal{U}(\mathcal{M}(A))\to\mathcal{U}(\mathcal{M}(B))$ be the unital group homomorphism defined by $\phi^{\dagger}(u)=\phi_{p}(u)+(1_{M(B)}-p)$ for all $u\in\mathcal{U}(\mathcal{M}(A))$. Note that for any $a\in A$ and $u\in\mathcal{U}(\mathcal{M}(A))$ one has that $\phi(ua)=\phi^{\dagger}(u)\phi(a)$.
Remark 1.6.
Let $\phi:A\rightarrow\mathcal{M}(B)$, $\psi:B\rightarrow\mathcal{M}(C)$ be $*{}^{*}$-homomorphisms with $\psi$ extendible, and $q\in\mathcal{M}(C)$ be the corresponding projection to $\psi$. One may compose them by the $*{}^{*}$-homomorphism $\psi_{q}\circ\phi:A\rightarrow\mathcal{M}(C)$ that we simply denote by $\psi\circ\phi$ (see [40, Remark 1.9]). If $\phi$ is also extendible, then the composition $\psi\circ\phi$ remains extendible and $(\psi\circ\phi)^{\dagger}=\psi^{\dagger}\circ\phi^{\dagger}$.
The proposition below yields a more explicit form for the tensor products of bimodules arising from $*{}^{*}$-homomorphisms into the multiplier algebra as in Example 1.3. If the homomorphisms are non-degenerate, then the result is folklore (see [4]).
Proposition 1.7.
Suppose that $A,B$, and $C$ are $\mathrm{C}^{*}$-algebras and $\phi:A\to\mathcal{M}(B)$, $\psi:B\to\mathcal{M}(C)$ are $*{}^{*}$-homomorphisms with $\psi$ extendible, but possibly degenerate. Then ${}_{\phi}B\boxtimes{}_{\psi}C\cong{}_{\psi\circ\phi}\overline{\psi(B)C}={}_{\psi\circ\phi}\psi(B)C$.
Proof.
First, note that $\overline{\psi(B)C}=\psi(B)C$ by Cohen’s factorisation (see [37, Proposition 2.33]). Let $T:{}_{\phi}B\boxtimes{}_{\psi}C\to{}_{\psi\circ\phi}\psi(B)C$ be the continuous linear map given by $T(b\boxtimes c)=\psi(b)c$ for any $b\in B$ and $c\in C$. We claim that $T$ is a bimodule isomorphism. Using the definition of the inner product, a standard check shows that $T$ is a well-defined bimodule map. Taking $(\eta_{\lambda})_{\lambda\in\Lambda}$ to be an approximate unit for $B$ let $S:{}_{\psi\circ\phi}\psi(B)C\to{}_{\phi}B\boxtimes{}_{\psi}C$ be given by $S(c)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes c$ for any $c\in C$. The map $S$ is well-defined precisely because $\psi$ is extendible. Indeed, for $\lambda,\mu\in\Lambda$
$|(\eta_{\lambda}-\eta_{\mu})\boxtimes c|^{2}$ $=\langle c,\langle\eta_{\lambda}-\eta_{\mu},\eta_{\lambda}-\eta_{\mu}\rangle_{B}\rhd c\rangle_{C}$
$=\langle c,(\eta_{\lambda}-\eta_{\mu})(\eta_{\lambda}-\eta_{\mu})\rhd c\rangle$
$=\langle\psi(\eta_{\lambda}-\eta_{\mu})c,\psi(\eta_{\lambda}-\eta_{\mu})c\rangle.$
Therefore, $S$ is well defined if and only if $\psi(\eta_{\lambda}-\eta_{\mu})c$ converges to 00 for any $c\in C$, which is equivalent to saying that $\psi(\eta_{\lambda})$ converges strictly in $\mathcal{M}(C)$. Hence $S$ is well-defined. Moreover, it can be seen that $S$ is the adjoint of $T$ and hence continuous. Now for any $b\in B$ and $c\in C$,
$S(T(b\boxtimes c))=\lim\limits_{n\to\infty}\eta_{n}\boxtimes\psi(b)c=\lim\limits_{n\to\infty}\eta_{n}\boxtimes b\rhd c=\lim\limits_{n\to\infty}\eta_{n}b\boxtimes c=b\boxtimes c.$
Then, by continuity, $S\circ T$ is the identity on the domain of $T$. Therefore, $T$ is injective. As $T$ is surjective by construction, it follows that $T$ is a bimodule isomorphism. ∎
Remark 1.8.
By uniqueness of the adjoint, the map $S:{}_{\psi\circ\phi}\psi(B)C\to{}_{\phi}B\boxtimes{}_{\psi}C$ of the proof of Proposition 1.7 given by $S(c)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes c$ for any $c\in C$ and a choice of approximate unit $(\eta_{\lambda})_{\lambda\in\Lambda}$ for $B$, does not depend on the choice of the approximate unit $\eta_{\lambda}$.
Definition 1.9.
Let $A$ be a $*{}^{*}$-algebra. A Hilbert $A$-$B$-bimodule $X$ is called non-degenerate if $XB$ is dense in $X$ and $AX$ is dense in $X$.
Remark 1.10.
Note that if $X$ is a right-Hilbert $B$-module then $XB$ is always dense in $X$ by the argument in [25, Proposition 2.6]. Therefore, a Hilbert $A$-$B$-bimodule $E$ may only fail to be non-degenerate if $AX$ is not dense in $X$.
We end this subsection by showing that if $A$ is a $\mathrm{C}^{*}$-algebra and $E$ is a non-degenerate Hilbert $A$-bimodule, then we can obtain a well-defined action of $\mathcal{M}(A)$ on $E$.
Let $(\eta_{\lambda})_{\lambda\in\Lambda}$ be an approximate unit of $A$ and let $L_{E}:E\to A\boxtimes E$ and $R_{E}:E\to E\boxtimes A$ be the $A$-bimodule maps given by $L_{E}(x)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes x$ and $R_{E}(x)=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}$ for all $x\in E$. Note that, since $E$ is non-degenerate, $L_{E}$ and $R_{E}$ are unitary bimodule isomorphisms. This follows similarly to the proof of Proposition 1.7. The inverses of $L_{E}$ and $R_{E}$ are given by their adjoints $L_{E}^{-1}(a\boxtimes x)=a\rhd x$ and $R_{E}^{-1}(x\boxtimes a)=x\lhd a$, for all $a\in A$ and $x\in E$ respectively.
Lemma 1.11.
Let $A$ be a $\mathrm{C}^{*}$-algebra and $E$ be a non-degenerate Hilbert $A$-bimodule. Then one may extend the left and right actions of $A$ on $E$ to left and right actions of $\mathcal{M}(A)$ on $E$. These extended actions equip $E$ with the structure of a Hilbert $\mathcal{M}(A)$-bimodule.
Proof.
For any $v\in\mathcal{M}(A)$ and any $x\in E$, let us define
(1.1) $v\rhd x=L_{E}^{-1}(v\rhd L_{E}(x))\quad\mathrm{and}\quad x\lhd v=R_{E}^{-1}(R_{E}(x)\lhd v).$
We claim that these formulae define left and right actions of $\mathcal{M}(A)$ on $E$. Moreover, it is clear that (1.1) restricted to $A$ coincides with the $A$-bimodule structure of $E$.
First, note that the left action of $\mathcal{M}(A)$ on $A\boxtimes E$ is given by left multiplication on $A$. Similarly, the right action of $\mathcal{M}(A)$ on $E\boxtimes A$ is given by right multiplication on $A$. Since $L_{E},R_{E}$, and their inverses are bimodule maps, it is straightforward to see that the formulae in (1.1) define left and right actions of $\mathcal{M}(A)$ on $E$. To see that $E$ with its right $A$-valued inner product is a Hilbert $\mathcal{M}(A)$-bimodule it suffices to check that the right $\mathcal{M}(A)$-action commutes with the inner product and the left $\mathcal{M}(A)$-action consists of adjointable operators. First, for $x,y\in E$ and $v\in\mathcal{M}(A)$ we have that
$\langle y,x\lhd v\rangle_{A}$ $=\langle y,R_{E}^{-1}(R_{E}(x)\lhd v)\rangle_{A}$
$=\lim\limits_{\lambda}\lim\limits_{\mu}\langle y\boxtimes\eta_{\lambda},x\boxtimes\eta_{\mu}v\rangle_{A}$
$=\lim\limits_{\lambda}\langle\eta_{\lambda},\langle y,x\rangle_{A}v\rangle_{A}$
$=\langle y,x\rangle_{A}v,$
so the right $\mathcal{M}(A)$-action commutes with the right inner product. Moreover, the operator of left multiplication by $v\in\mathcal{M}(A)$ has as an adjoint the operator of left multiplication by $v^{*}$. Indeed left multiplication by $v^{*}$ is the adjoint of the operator of left multiplication by $v$ on the Hilbert $A$-module $A$ so for $x,y\in E$
$\langle v\rhd x,y\rangle_{A}=\langle L_{E}^{-1}(v\rhd L_{E}(x)),y\rangle_{A}=\langle L_{E}(x),v^{*}\rhd L_{E}(y)\rangle_{A}=\langle x,v^{*}\rhd y\rangle_{A}$
as required. ∎
1.2. Correspondences
A reformulation of the theory of Hilbert bimodules is the language of $*{}^{*}$-correspondences.
Definition 1.12.
Let $A,B$ be $\mathrm{C}^{*}$-algebras. An $A$-$B$ correspondence is a $*{}^{*}$-homomorphism $\phi:A\to\mathcal{L}(X_{B})$ into the adjointable operators of a Hilbert $B$-module $X_{B}$. We denote the collection of $A$-$B$ correspondences by $\mathrm{Corr}(A,B)$.
Remark 1.13.
Note that any $A$-$B$ correspondence induces a right-Hilbert $A$-$B$-bimodule and vice versa. If $\phi:A\to\mathcal{L}(X_{B})$ is an $A$-$B$ correspondence, then $X_{B}$ becomes a right-Hilbert $A$-$B$-bimodule, with the left action given by $\phi$. We will often denote this bimodule by ${}_{\phi}X$, forgetting the right $B$-action. Conversely, given a right-Hilbert $A$-$B$-bimodule $X$, the left action by $A$ induces an $A$-$B$ correspondence. Therefore, we will freely flip between the two pictures.
One may compose correspondences through the tensor product of bimodules. Precisely, let $X$ be a Hilbert $A$-$B$ bimodule inducing $\phi\in\mathrm{Corr}(A,B)$, and let $E$ be a Hilbert $B$-$C$ bimodule inducing $\psi\in\mathrm{Corr}(B,C)$. Then $X\boxtimes E$ gives a Hilbert $A$-$C$ bimodule which induces an element in $\mathrm{Corr}(A,C)$ denoted by $\psi\circ\phi$. Although the theories of bimodules and correspondences are equivalent, we sometimes choose to work with correspondences as the composition resembles composition of $*{}^{*}$-homomorphisms between $\mathrm{C}^{*}$-algebras in a covariant manner.
1.3. $\mathrm{C}^{*}$-tensor categories
Throughout this section we will assume that the reader is familiar with the standard language of category theory. For a category $\mathcal{C}$ we will use capital letters e.g. $X,Y,$ and $Z$ to denote objects of the category. The space of morphisms between two objects $X,Y\in\mathcal{C}$ will be denoted by $\mathrm{Hom}(X,Y)$. All categories in this section will be $\mathbb{C}$-linear, that is and additive category such that its space of morphisms $\mathrm{Hom}(X,Y)$ between any two objects $X,Y$ is a $\mathbb{C}$-vector space and the composition of morphisms yields a bilinear map $\mathrm{Hom}(X,Y)\times\mathrm{Hom}(Y,Z)\rightarrow\mathrm{Hom}(X,Z)$ for any $X,Y,Z\in\mathcal{C}$. We start by defining $\mathrm{C}^{*}$-categories.
Definition 1.14.
(see [19]) A $\mathbb{C}$-linear category $\mathcal{C}$ is said to be a $\rm{C}^{*}$-category if it is equipped with conjugate linear mappings :*${}^{*}\!:\mathrm{Hom}(X,Y)\rightarrow\mathrm{Hom}(Y,X)$ for every $X,Y\in\mathcal{C}$ such that
  1. (i)
    $\phi^{**}=\phi$ for all $\phi\in\mathrm{Hom}(X,Y);$
  2. (ii)
    $(\phi\circ\psi)^{*}=\psi^{*}\circ\phi^{*}$ for all $\psi\in\mathrm{Hom}(X,Y),\phi\in\mathrm{Hom}(Y,Z)$;
  3. (iii)
    The function $\|\cdot\|:\mathrm{Hom}(X,Y)\rightarrow[0,\infty]$ given by
    $\|\phi\|^{2}=\sup\{\lambda>0:\phi^{*}\circ\phi-\lambda\mathrm{id}_{X}\mbox{ isnot invertible}\}$
    is a complete norm on $\mathrm{Hom}(X,Y)$;
  4. (iv)
    $\|\phi\circ\psi\|\leq\|\phi\|\|\psi\|$ for all $\psi\in\mathrm{Hom}(X,Y),\phi\in\mathrm{Hom}(Y,Z)$;
  5. (v)
    $\|\phi^{*}\circ\phi\|=\|\phi\|^{2}$ for all $\phi\in\mathrm{Hom}(X,Y)$;
  6. (vi)
    For all $\phi\in\mathrm{Hom}(X,Y)$, $\phi^{*}\circ\phi$ is a positive element of the $\rm{C}^{*}$-algebra $\mathrm{Hom}(X,X)$.
Let $\mathcal{C}$ and $\mathcal{D}$ be $\mathrm{C}^{*}$-categories. A functor $F:\mathcal{C}\rightarrow D$ is called a $\mathrm{C}^{*}$-functor if the induced mappings $\mathrm{Hom}(X,Y)\mapsto\mathrm{Hom}(F(X),F(Y))$ are $\mathbb{C}$-linear and $*{}^{*}$-preserving. A natural transformation $\nu:F\rightarrow G$ between $\mathrm{C}^{*}$-functors is called an isometry if $\nu_{X}^{*}\nu_{X}=\mathrm{id}_{F(X)}$ for all $X\in\mathcal{C}$. Moreover, $\nu$ is called a unitary if it is a surjective isometry in which case $\nu_{X}\nu_{X}^{*}=\mathrm{id}_{G(x)}$ for all $X\in\mathcal{C}$. We are interested in $*{}^{*}$-categories that admit tensor product structures.
Definition 1.15.
(see for example [23]) A $\rm{C}^{*}$-tensor category is a $\rm{C}^{*}$-category $\mathcal{C}$ together with a $\mathbb{C}$-linear bifunctor $-\otimes-:\mathcal{C}\times\mathcal{C}\rightarrow\mathcal{C}$, a distinguished object $1_{\mathcal{C}}\in\mathcal{C}$ and unitary natural isomorphisms
(1.2) $\alpha_{X,Y,Z}$ $:(X\otimes Y)\otimes Z\rightarrow X\otimes(Y\otimes Z),$
$\lambda_{X}$ $:(1_{\mathcal{C}}\otimes X)\rightarrow X,$
$\rho_{X}$ $:(X\otimes 1_{\mathcal{C}})\rightarrow X,$
such that $(\phi\otimes\psi)^{*}=(\phi^{*}\otimes\psi^{*})$ and the following diagrams commute for any $X,Y,Z,W\in\mathcal{C}$
[…]
and $\mathrm{Hom}(1_{\mathcal{C}},1_{\mathcal{C}})\cong\mathbb{C}$. Moreover, $\mathcal{C}$ is said to be semisimple if there exists a collection of objects $X_{i}\in\mathcal{C}$ with $\mathrm{Hom}(X_{i},X_{j})\cong\delta_{ij}\mathbb{C}$ and any object $X\in\mathcal{C}$ can be decomposed uniquely as a finite direct sum $X\cong\bigoplus_{i}X_{i}$.
We call the structure morphism $\alpha$ as in Definition 1.15 associated to a $*{}^{*}$-tensor category its associator and the maps $\lambda$ and $\rho$ the unitors. We call an object $X\in\mathcal{C}$ such that $\mathrm{Hom}(X,X)\cong\mathbb{C}$ irreducible or simple. For a $*{}^{*}$-tensor category $\mathcal{C}$, we will denote by $\mathrm{Irr}(\mathcal{C})$ a collection of isomorphism class representatives for simple objects in $\mathcal{C}$.
Example 1.16.
Let $\Gamma$ be a countable, discrete group (or more generally a countable, discrete monoid). We denote by $\mathrm{Hilb}(\Gamma)$ the semisimple $*{}^{*}$-tensor category whose objects are finite-dimensional $\Gamma$-graded Hilbert spaces, i.e. finite-dimensional Hilbert spaces $\mathcal{H}$ endowed with a decomposition $\mathcal{H}=\bigoplus_{g\in\Gamma}\mathcal{H}_{g}$. The morphisms are linear maps that preserve the $\Gamma$-grading. The tensor product is the usual Hilbert space tensor product with the grading defined by
$(\mathcal{H}\otimes\mathcal{K})_{g}=\bigoplus_{hk=g}\mathcal{H}_{h}\otimes\mathcal{K}_{k}.$
The isomorphism classes of simple objects in this category are indexed by group elements in $\Gamma$. We denote these graded Hilbert spaces by $\mathbb{C}_{g}$ where
$(\mathbb{C}_{g})_{h}=\begin{cases}\mathbb{C}\ \text{if}\ g=h,\\ 0\ \text{otherwise}.\end{cases}$
If $\Gamma$ is a group and $\omega\in Z^{3}(\Gamma,\mathbb{T})$ is a normalised $3$-cocycle, the category $\mathrm{Hilb}(\Gamma,\omega)$ is defined exactly as is $\mathrm{Hilb}(\Gamma)$ but with associators now given by
$\alpha_{\mathcal{H},\mathcal{K},\mathcal{L}}:(\xi\boxtimes\eta)\boxtimes\mu\mapsto\omega(g,h,k)\xi\boxtimes(\eta\boxtimes\mu)$
for $\xi\in\mathcal{H}_{g},\eta\in\mathcal{K}_{h}$ and $\mu\in\mathcal{H}_{k}$.
For the remainder of this paper we will only make use of this example when $\Gamma$ is a group. However, our results also apply to the case of monoids.
We now introduce the most important example for our purposes.
Example 1.17.
Let $A$ be a $*{}^{*}$-algebra. We denote by $\mathrm{Corr}_{0}(A)$ the $*{}^{*}$-tensor category whose objects are non-degenerate $A$-$A$ correspondences and whose morphisms are adjointable intertwiners of the underlying Hilbert bimodules. The tensor product of two $A$-$A$ correspondences $\varphi$ with $\psi$ is given by their composition $\varphi\circ\psi$. The tensor identity of $\mathrm{Corr}_{0}(A)$ is given by the identity homomorphism $\mathrm{id}_{A}$, the associator is given by the rebracketing morphism associated to the underlying tensor product of Hilbert bimodules.
Remark 1.18.
In general, $\mathrm{Corr}(A)$ is not a $*{}^{*}$-tensor category as there is no tensor unit on degenerate correspondences; $A$ does not act as a unit on a degenerate correspondence. It is a non-unital $*{}^{*}$-tensor category, this is a weakening of Definition 1.15 which omits the necessity of a tensor unit.
1.4. Szabó’s cocycle category
Before we begin our discussion on actions of $\mathrm{C}^{*}$-tensor categories, we recall the case of a twisted action by a second-countable locally compact group $\Gamma$. We shall later identify this, in the case when $\Gamma$ is countable discrete, with actions of the $*{}^{*}$- tensor category $\mathrm{Hilb}(\Gamma)$.
Definition 1.19 (cf. [40, Definition 1.1]).
Let $\Gamma$ be a locally compact group and $A$ be a $\mathrm{C}^{*}$-algebra. A twisted action of $\Gamma$ on $A$ is a pair $(\alpha,\mathfrak{u})$, where $\alpha:\Gamma\to\mathrm{Aut}(A)$ is a point-norm continuous map, and $\mathfrak{u}:\Gamma\times\Gamma\to\mathcal{U}(\mathcal{M}(A))$ is a strictly continuous map satisfying
(1.5) $\alpha_{1}=\mathrm{id}_{A},\quad\mathrm{Ad}(\mathfrak{u}_{s,t})\circ\alpha_{s}\circ\alpha_{t}=\alpha_{st}$
and
(1.6) $\mathfrak{u}_{s,1}=\mathfrak{u}_{1,s}=1,\quad\mathfrak{u}_{r,st}\alpha_{r}(\mathfrak{u}_{s,t})=\mathfrak{u}_{rs,t}\mathfrak{u}_{r,s}$
for all $r,s,t\in\Gamma$.
Remark 1.20.
Note that the formulae above differ slightly from the definition of a twisted action in [40]. In fact, the sole difference is that our unitary cocycles $\mathfrak{u}_{s,t}$ for $s,t\in\Gamma$ are the adjoints of the cocycles in [40]. The reason for this change of conventions will be discussed in Section 2.
A triple $(A,\alpha,\mathfrak{u})$ as above is called a twisted $\Gamma$-$\mathrm{C}^{*}$-algebra. If the cocycle is trivial i.e. $\mathfrak{u}_{s,t}=\mathfrak{1}$ for all $s,t\in\Gamma$, then $(A,\alpha)$ is said to be a $\Gamma$-$\mathrm{C}^{*}$-algebra.
Definition 1.21 (cf. [40, Definition 1.10]).
Let $(\alpha,\mathfrak{u}):\Gamma\curvearrowright A$ and $(\beta,\mathfrak{v}):\Gamma\curvearrowright B$ be two twisted actions on $\mathrm{C}^{*}$-algebras $A$ and $B$ respectively.
  1. (i)
    A cocycle representation from $(A,\alpha,\mathfrak{u})$ to $(B,\beta,\mathfrak{v})$ is a pair $(\psi,\mathbbm{v})$, where $\psi:A\to\mathcal{M}(B)$ is an extendible $*{}^{*}$-homomorphism and $\mathbbm{v}:\Gamma\to\mathcal{U}(\mathcal{M}(B))$ is a strictly continuous map such that
    (1.7) $\beta_{g}\circ\psi=\mathrm{Ad}(\mathbbm{v}_{g})\circ\psi\circ\alpha_{g}$
    and
    (1.8) $\psi^{\dagger}(\mathfrak{u}_{g,h})=\mathbbm{v}_{gh}^{*}\mathfrak{v}_{g,h}\beta_{g}(\mathbbm{v}_{h})\mathbbm{v}_{g}$
    for all $g,h\in\Gamma$.
  2. (ii)
    A cocycle morphism from $(A,\alpha,\mathfrak{u})$ to $(B,\beta,\mathfrak{v})$ is a cocycle representation $(\psi,\mathbbm{v})$ as above, with the additional requirement that $\psi(A)\subseteq B$.
Remark 1.22.
Due to our change of conventions when defining twisted actions we need to change the definition of a cocycle representation. The formula in (1.8) differs from [40, Definition 1.10] precisely by taking adjoints.
As shown in [40], there exists a category with objects being twisted $\Gamma$-$\mathrm{C}^{*}$-algebras and morphisms being cocycle morphisms with composition defined in [40, Proposition 1.15]. This category is denoted by $\mathrm{C}^{*}_{\Gamma,t}$ ([40, Definition 1.16]).
Our attention will now focus on generalising this construction to actions of semisimple $\mathrm{C}^{*}$-tensor categories.
2. Actions of $\mathrm{C}^{*}$-tensor categories
For $\mathrm{C}^{*}$-tensor categories $(\mathcal{C},\otimes)$ and $(\mathcal{D},\boxtimes)$, $F:\mathcal{C}\to\mathcal{D}$ is said to be a $\mathrm{C}^{*}$-tensor functor if it is a $\mathrm{C}^{*}$-functor such that $F(1_{\mathcal{C}})=1_{\mathcal{D}}$ there exists a unitary natural isomorphism $J_{X,Y}:F(X)\boxtimes F(Y)\to F(X\otimes Y)$ such that
[…]
commutes for all $X,Y,Z\in\mathcal{C}$.
In the following definition we denote by $\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$ the full subcategory of $\mathrm{Corr}_{0}(A)$ consisting of bimodules with countable dense subsets.
Definition 2.1.
[…]
Remark 2.2.
In the literature, the main interest is actions of unitary tensor categories on $*{}^{*}$-algebras (see e.g. [1, 6, 23, 25, 42, 33]). This is because unitary tensor categories axiomatise the standard invariant in subfactor theory and can be thought of as the mathematical objects encoding the symmetry in finite index inclusions of $*{}^{*}$-algebras. If $\mathcal{C}$ is a unitary tensor category, $A$ is separable, and $(F,J):\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$ is a $*{}^{*}$-tensor functor, then the bimodule associated to the correspondence $F(X)$ for any $X\in\mathcal{C}$ has a countable dense subset by [25, Corollary 2.24] and hence $(F,J)$ automatically falls into $\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$.
Example 2.3.
Let $\Gamma$ be a a countable discrete group and let $(\alpha,\mathfrak{u}):\Gamma\rightarrow\mathrm{Aut}(A)$ be a twisted action. The pair $(\alpha,\mathfrak{u})$ will induce a $*{}^{*}$-tensor functor $(\alpha,\mathfrak{u}):\mathrm{Hilb}(\Gamma)\rightarrow\mathrm{Corr}_{0}(A)$ by setting
(2.2) $\alpha(\mathbb{C}_{g})={}_{\alpha_{g}}A,$
(2.3) $\mathfrak{u}_{\mathbb{C}_{g},\mathbb{C}_{h}}(a\boxtimes b)=\mathfrak{u}_{g,h}^{*}\alpha_{g}(a)b.$
The functor may then be extended by linearity to all of $\mathrm{Hilb}(\Gamma)$ in a similar manner to [14, Proposition 5.6]. In general, actions of $\mathrm{Hilb}(\Gamma)$ on $A$ correspond to twisted actions of $\Gamma$ on $A\otimes\mathbb{K}$.
If $A$ is a $\mathrm{C}^{*}$-algebra, then its sequence algebra $A_{\infty}$ is defined by
$A_{\infty}=\ell^{\infty}(\mathbb{N},A)/\Big{\{}(a_{n})_{n\geq 1}:\lim\limits_{n\to\infty}\|a_{n}\|=0\Big{\}}.$
[…]
For any $X\in\mathcal{C}$, we can view $F(X)$ as a non-degenerate Hilbert $A$-bimodule. Define
(2.4) $F_{\infty}(X)=\ell^{\infty}(\mathbb{N},F(X))/\Big{\{}(\xi_{n})_{n\geq 1}:\lim_{n\to\infty}\|\xi_{n}\|=0\Big{\}},$
with the left and right actions of $A_{\infty}$ given pointwise. Precisely, for any $a\in A_{\infty}$ and any $\xi\in F_{\infty}(X)$ represented by $(a_{n})_{n\geq 1}$ and $(\xi_{n})_{n\geq 1}$, we let
(2.5) $a\rhd\xi=(a_{n}\rhd\xi_{n})_{n\geq 1}\quad\mathrm{and}\quad\xi\lhd a=(\xi_{n}\lhd a_{n})_{n\geq 1}.$
Similarly, for any $\xi,\eta\in F_{\infty}(X)$ we define
(2.6) $\langle\xi,\eta\rangle_{A_{\infty}}=\big{(}\langle\xi_{n},\eta_{n}\rangle_{A}\big{)}_{n\geq 1}.$
Lemma 2.4.
$F_{\infty}(X)$ is a right-Hilbert $A_{\infty}$-module.
Proof.
First, we need to check that the formulae in (2.5) are well-defined. For this, suppose the sequences $(a_{n})_{n\geq 1}$ and $(a_{n}^{\prime})_{n\geq 1}$ induce the same element in $A_{\infty}$, and let $\xi\in F_{\infty}(X)$ be represented by the sequence $(\xi_{n})_{n\geq 1}$. Then, a direct calculation shows that
$\|\langle\xi_{n}\lhd(a_{n}-a_{n}^{\prime}),\xi_{n}\lhd(a_{n}-a_{n}^{\prime})\rangle_{A}\|\leq\|\langle\xi_{n},\xi_{n}\rangle_{A}\|\|a_{n}-a_{n}^{\prime}\|^{2},$
which converges to 00 as $n\to\infty$. Exactly the same calculation shows that if $(\xi_{n})_{n\geq 1}$ and $(\xi_{n}^{\prime})_{n\geq 1}$ induce the same element in $F_{\infty}(X)$ and $(a_{n})_{n\geq 1}\in A_{\infty}$, then $(\xi_{n}\lhd a_{n})_{n\geq 1}=(\xi_{n}^{\prime}\lhd a_{n})_{n\geq 1}$ as elements in $F_{\infty}(X)$. Thus, (2.5) gives a well-defined right action of $A_{\infty}$.
We now check that (2.6) gives a well-defined right inner product. First, if $(\xi_{n})_{n\geq 1}$ and $(\xi_{n}^{\prime})_{n\geq 1}$ induce $\xi$, and $(\eta_{n})_{n\geq 1}$ and $(\eta_{n})_{n\geq 1}^{\prime}$ induce $\eta$, then
$\langle\xi_{n},\eta_{n}\rangle_{A}-\langle\xi_{n}^{\prime},\eta_{n}^{\prime}\rangle_{A}=\langle\xi_{n}-\xi_{n}^{\prime},\eta_{n}\rangle_{A}+\langle\xi_{n}^{\prime},\eta_{n}-\eta_{n}^{\prime}\rangle_{A},$
which converges to 00 by Cauchy-Schwarz. Moreover, the sequences $\big{(}\langle\xi_{n},\eta_{n}\rangle_{A}\big{)}_{n\geq 1}$ and $\big{(}\langle\xi_{n}^{\prime},\eta_{n}^{\prime}\rangle_{A}\big{)}_{n\geq 1}$ are bounded, so induce elements in $A_{\infty}$.
Therefore, the function $\langle\cdot,\cdot\rangle_{A_{\infty}}:F_{\infty}(X)\times F_{\infty}(X)\to A_{\infty}$ is well-defined. It is now straightforward to check that this function is right linear, left conjugate linear, and antisymmetric since all this properties are satisfied pointwise for each $n\in\mathbb{N}$. Finally, it is clear that $\langle\xi,\xi\rangle_{A_{\infty}}\geq 0$ and that $\langle\xi,\xi\rangle_{A_{\infty}}=0$ if and only if $\langle\xi_{n},\xi_{n}\rangle_{A}$ converges to 00 i.e. $\xi=0$ in $F_{\infty}(X)$. Since, $F_{\infty}(X)$ is complete with respect to the inner product defined in (2.6), it is a right-Hilbert $A_{\infty}$-bimodule. ∎
Lemma 2.5.
$F_{\infty}(X)$ is a non-degenerate Hilbert $A_{\infty}$-bimodule.
Proof.
As in the proof of Lemma 2.4 we have that (2.5) gives a well-defined left action by $A_{\infty}$. Since the left action of $A$ on $F(X)$ is adjointable, the left action of $(a_{n})_{n\geq 1}\in A_{\infty}$ is an adjointable operator with adjoint given by the action of $(a_{n}^{*})_{n\geq 1}$. Moreover, $F_{\infty}(X)$ is non-degenerate for any $X\in\mathcal{C}$ as $F(X)$ is. ∎
Consider the functor $F_{\infty}:\mathcal{C}\to\mathrm{Corr}_{0}(A_{\infty})$ defined by sending any $X\in\mathcal{C}$ to the correspondence induced by the Hilbert $A_{\infty}$-bimodule $F_{\infty}(X)$, and any $f\in\mathrm{Hom}(X,Y)$ for $X,Y\in\mathcal{C}$ to the intertwiner defined on $F_{\infty}(X)$ by $F_{\infty}(f)((\xi_{n})_{n\geq 1})=(F(f)(\xi_{n}))_{n\geq 1}$. Also, one can define $J_{X,Y}^{\infty}:F_{\infty}(Y)\boxtimes F_{\infty}(X)\to F_{\infty}(X\otimes Y)$ as the unique continuous extension of the map
$J_{X,Y}^{\infty}(\xi\boxtimes\eta)=J_{X,Y}(\xi_{n}\boxtimes\eta_{n}),$
for any $\xi\in F_{\infty}(Y)$, any $\eta\in F_{\infty}(X)$ (that this is well defined follows in a similar fashion as the arguments in the proof of Lemma 2.4). Note that $J^{\infty}_{X,Y}$ is isometric for all $X,Y\in\mathcal{C}$.
Lemma 2.6.
Let $F:\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$ be an action of a $*{}^{*}$-tensor category $\mathcal{C}$ on a unital $*{}^{*}$-algebra $A$ such that $F(X)$ is algebraically finitely generated as a right $A$-module for all $X\in\mathcal{C}$. Then the map $F_{\infty}:\mathcal{C}\to\mathrm{Corr}_{0}(A_{\infty})$ defined above is a $C^{*}$-tensor functor and so $\mathcal{C}$ acts on $A_{\infty}$ via the triple $(A_{\infty},F_{\infty},J^{\infty})$.
Proof.
It follows from construction that $F_{\infty}$ is a $*{}^{*}$-functor. Moreover, the naturality of $J^{\infty}$ and that $J^{\infty}$ satisfies commuting diagrams as in (2.1) follows from direct computations. It remains to show that $J_{X,Y}^{\infty}$ is surjective. Let $\iota_{X,Y}:F_{\infty}(Y)\boxtimes F_{\infty}(X)\rightarrow(F(Y)\boxtimes F(X))_{\infty}$ be the canonical inclusion maps and $T_{X,Y}:(F(Y)\boxtimes F(X))_{\infty}\rightarrow F_{\infty}(X\otimes Y)$ be the bounded bimodule map defined by $T(\xi)_{n}=J_{X,Y}(\xi_{n})$ for $\xi=(\xi_{n})_{n\geq 0}\in(F(Y)\boxtimes F(X))_{\infty}$. The composition $T_{X,Y}\circ\iota_{X,Y}$ coincides with $J_{X,Y}^{\infty}$ so it suffices to show the surjectivity of both $T_{X,Y}$ and $\iota_{X,Y}$.
Firstly $T_{X,Y}$ is an adjointable unitary with adjoint defined by the mapping $(\xi_{n})_{n\geq 0}\mapsto(J_{X,Y}^{*}(\xi_{n}))_{n\geq 0}$ for $(\xi_{n})_{n\geq 0}\in F_{\infty}(X\otimes Y)$. It remains to show the surjectivity of $\iota_{X,Y}$. As each $F(X)$ is algebraically finitely generated, then $F(X)$ is also projective ([44, Corollary 15.4.8]). Equivalently for each $X\in\mathcal{C}$ the bimodule $F(X)$ has a right $A$-basis, i.e. there exist a finite set $I_{X}$ and $u_{i}^{X}\in F(X)$ for $i\in I_{X}$ such that $\sum_{i\in I_{X}}u_{i}^{X}\lhd\langle u_{i}^{X},\xi\rangle_{A}=\xi$ for all $\xi\in F(X)$ (see for example [26, Lemma 1.11]). Similarly $(F(Y)\boxtimes F(X))_{\infty}$ is also finitely generated projective as a right $A_{\infty}$-module with right $A_{\infty}$-basis $u_{i}^{Y}\boxtimes u_{j}^{X}\in(F(Y)\boxtimes F(X))_{\infty}$ for $i\in I_{Y}$ and $j\in I_{X}$. As $\iota_{X,Y}(u_{i}^{Y}\boxtimes u_{j}^{X})=u_{i}^{Y}\boxtimes u_{j}^{X}$ and $u_{i}^{Y}\boxtimes u_{j}^{X}$ algebraically generate $(F(Y)\boxtimes F(X))_{\infty}$ as a right $A_{\infty}$-module it follows that $\iota_{X,Y}$ is surjective. ∎
Remark 2.7.
If $F:\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$ is an action of a unitary tensor category (in the sense of [6] for example) on a unital $*{}^{*}$-algebra $A$, then it follows from [25, 26] that $F(X)$ is algebraically finitely generated as a right $A$-module for each $X\in\mathcal{C}$.
3. The generalised cocycle category
In this section we introduce the category of $\mathcal{C}$-$\mathrm{C}^{*}$-algebras for which we will later perform intertwining arguments. Throughout the rest of this paper $\mathcal{C}$ is always assumed to be a $\mathrm{C}^{*}$-tensor category. Recall the definition of an action of $\mathcal{C}$ on a $\mathrm{C}^{*}$-algebra $A$ from Definition 2.1.
Definition 3.1.
[…]
  1. (i)
    A correspondence morphism from $(A,F,J)$ to $(B,G,I)$ is a pair $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}})$, where $\phi:A\to\mathcal{L}(E)$ is an $A$-$B$ correspondence and $\{\mathbbm{v}_{X}:\phi\circ F(X)\to G(X)\circ\phi\}_{X\in\mathcal{C}}$ is a natural family of $A$-$B$-bimodule maps such that $\mathbbm{v}_{X}$ is an isometry (not necessarily adjointable) for any $X\in\mathcal{C}$ [^6] By an isometry, we mean a map which preserves the norm. By the polarisation identity, it is equivalent to assume that it preserves the inner product. Moreover, for all $X,Y\in\mathcal{C}$ the following pentagon diagram commutes
    (3.1) {tikzcd}
    and $\mathbbm{v}_{1_{\mathcal{C}}}:A\boxtimes{}_{\phi}E\to{}_{\phi}E\boxtimes B\cong{}_{\phi}E$ is given by $\mathbbm{v}_{1_{\mathcal{C}}}(a\boxtimes x)=\phi(a)x$ for any $a\in A$ and $x\in{}_{\phi}E$. For convenience, we write $(\phi,\mathbbm{v})$, where $\mathbbm{v}$ denotes the collection of maps $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$. [^7] The notation of the maps in (3.1) denotes the tensor product of bimodules, which is equivalent to composition of correspondences.
  2. (ii)
    A cocycle representation $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ is a correspondence morphism for which we further require that $\phi:A\to\mathcal{M}(B)$ is a $*{}^{*}$-homomorphism.
  3. (iii)
    A cocycle morphism $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ is a cocycle representation for which we further require that $\phi:A\to B$ is a $*{}^{*}$-homomorphism.
Remark 3.2.
By naturality, if $\mathcal{C}$ is semisimple, any $\mathcal{C}$-equivariant structure is uniquely determined by its values on $\mathrm{Irr}(\mathcal{C})$. In particular, for any cocycle morphism, the family of maps $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ is uniquely determined by the family of maps $\{\mathbbm{v}_{X}\}_{X\in\mathrm{Irr}(\mathcal{C})}$.
In the case of group actions, Definition 3.1 recovers Szabó’s notion of a cocycle morphism (see Definition 1.21).
Example 3.3.
Suppose $(A,\alpha)$ and $(B,\beta)$ are actions of a group $\Gamma$ on $\mathrm{C}^{*}$-algebras $A$ and $B$. Consider them as actions of $\mathrm{Hilb}(\Gamma)$ as in Example 2.3. Let $(\phi,\mathbbm{v}):(A,\alpha)\to(B,\beta)$ be an extendible cocycle morphism as in Definition 3.1 with $\mathbbm{v}_{g}$ being adjointable for all $g\in\Gamma$. Fix $g\in\Gamma$ and let $f_{g}:=T_{g}^{\prime}\circ\mathbbm{v}_{\mathbb{C}_{g}}\circ S_{g}:{}_{\phi\circ\alpha_{g}}B\to{}_{\beta_{g}\circ\phi}B$, where $S_{g}:{}_{\phi\circ\alpha_{g}}B\to{}_{\alpha_{g}}A\boxtimes{}_{\phi}B$ is given by $S_{g}(b)=\lim\limits_{\lambda}\xi_{\lambda}\boxtimes b$ for any $b\in B$, where $\xi_{\lambda}$ is an approximate unit for $A$ and $T_{g}^{\prime}:{}_{\phi}B\boxtimes{}_{\beta_{g}}B\rightarrow{}_{\beta_{g}\circ\phi}B$ is given by $T_{g}^{\prime}(b\boxtimes c)=\beta_{g}(b)c$ for $b,c\in B$ [^8] Note that $S_{g}$ is well-defined by Proposition 1.7. Moreover, let $T_{g}:{}_{\alpha_{g}}A\boxtimes{}_{\phi}B\rightarrow{}_{\phi\circ\alpha_{g}}B$ be given by $T_{g}(a\boxtimes b)=\phi(a)b$ for any $a\in A$ and any $b\in B$.
Since $f_{g}$ is an adjointable bimodule map, it follows that $f_{g}(b)=\mathbbm{u}_{g}b$ for any $b\in B$, for some $\mathbbm{u}_{g}\in\mathcal{M}(B)$. [^9] Note that $T_{g}$ is adjointable as $\phi$ is extendible (see Proposition 1.7). Thus, $f_{g}$ is a composition of adjointable maps. In particular, the equality $f_{g}(a\rhd b)=a\rhd f_{g}(b)$ implies that
$\beta_{g}(\phi(a))\mathbbm{u}_{g}=\mathbbm{u}_{g}\phi(\alpha_{g}(a))$
for all $a\in A$. This gives (1.7), although $\mathbbm{u}_{g}$ might not be a unitary.
As $S_{g}\circ T_{g}$ is the identity map, it follows that the diagram
[…]
commutes. Hence,
$\mathbbm{v}_{\mathbb{C}_{g}}(a\boxtimes b)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes\mathbbm{u}_{g}\phi(a)b,$
for any $a\in A,b\in B$ and $\eta_{\lambda}$ an approximate unit for $B$. Following the pentagon diagram for $\mathbbm{v}$, one gets (1.8).
Conversely, if $(\phi,\mathbbm{u}):(A,\alpha)\to(B,\beta)$ is a cocycle morphism as in Definition 1.21, then define $f_{g}:{}_{\phi\circ\alpha_{g}}B\rightarrow{}_{\beta_{g}\circ\phi}B$ by $f_{g}(b)=\mathbbm{u}_{g}b$ for all $b\in B$ and $\mathbbm{v}_{\mathbb{C}_{g}}$ be given by (3.2). Note that (1.7) implies that $\mathbbm{v}_{\mathbb{C}_{g}}$ is a bimodule map, while (1.8) gives the pentagon diagram for $\mathbbm{v}_{\mathbb{C}_{g}}$. Hence $(\phi,\mathbbm{v})$ yields a cocycle morphism in the sense of Definition 3.1.
Remark 3.4.
Note that, unlike in [40, Definition 1.10], we do not require $\phi\in\mathrm{Corr}(A,B)$ to be extendible. In fact, this is precisely the reason why we consider the maps $\mathbbm{v}_{X}$ to be isometries (possibly non-adjointable) instead of unitaries. For example, the map $\mathbbm{v}_{1_{\mathcal{C}}}:A\boxtimes{}_{\phi}B\to{}_{\phi}B\boxtimes B\cong{}_{\phi}B$ is given by $\mathbbm{v}_{1_{\mathcal{C}}}(a\boxtimes b)=\phi(a)b$ for all $a\in A$ and $b\in B$. Therefore, it is not surjective unless $\phi$ is non-degenerate. Moreover, it might not be adjointable if $\phi$ is not extendible as seen in the proof of Proposition 1.7.
Furthermore, the morphisms of Definition 3.1 fit into the $\mathcal{C}$-equivariant $\mathrm{KK}$-theory developed in [1] (see [1, Example 3.3]). In [1] a correspondence morphism is instead called a $\mathcal{C}$-Hilbert $A$-$B$-bimodule and a cocycle morphism is called a cocycle-$\mathcal{C}$-***-homomorphism.
[…]
(3.3) $(\mathbbm{w}\circ\mathbbm{v})_{X}=(\mathrm{id}_{\phi}\boxtimes\mathbbm{w}_{X})\circ(\mathbbm{v}_{X}\boxtimes\mathrm{id}_{\psi}).$
By combining the pentagon diagrams for $\mathbbm{v}$ and $\mathbbm{w}$, one obtains the following lemma.
Lemma 3.5.
With the notation above, $(\psi\circ\phi,\mathbbm{w}\circ\mathbbm{v}):(A,F,J)\to(C,H,K)$ is a correspondence morphism which denotes the composition of $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$.
However, if $\phi$ and $\psi$ are possibly degenerate cocycle morphisms, the composition formula above will not give a cocycle morphism. This problem arises as the bimodule ${}_{\phi}B\boxtimes{}_{\psi}C$ may no longer be isomorphic to $C$ as a right $C$ module. Therefore, to form a category, we introduce a slightly different composition on cocycle morphisms.
Definition 3.6.
Let $(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ and $(\psi,\mathbbm{w}):(B,G,I)\to(C,H,K)$ be cocycle morphisms. Let $\mathbbm{w}*\mathbbm{v}$ be the collection of isometries $\{(\mathbbm{w}*\mathbbm{v})_{X}\}_{X\in\mathcal{C}}$ given by
[…]
Here $S_{X}(x\boxtimes c)=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}\boxtimes c$ and $T(b\boxtimes c)=\psi(b)c$ for any $X\in\mathcal{C}$, $x\in F(X)$, $b\in B$, and $c\in C$, with $\eta_{\lambda}$ being an approximate unit for $B$ [^10] We have not shown that $S_{X}$ is well-defined at this point.
Lemma 3.7.
The continuous linear map $S_{X}:F(X)\boxtimes{}_{\psi\circ\phi}C\to F(X)\boxtimes{}_{\phi}B\boxtimes{}_{\psi}C$ given by $S_{X}(x\boxtimes c)=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}\boxtimes c$ for any $x\in F(X)$, any $c\in C$, and some approximate unit $\eta_{\lambda}$ of $B$ is a well-defined isometric bimodule isomorphism for any $X\in\mathcal{C}$.
Proof.
We will show that for any $x\in F(X)$ and any $c\in C$, the net $x\boxtimes\eta_{\lambda}\boxtimes c$ is Cauchy with respect to the norm induced by the right inner product. By definition, we have that $\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c,x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c\rangle_{C}=\langle c,\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu}),x\boxtimes(\eta_{\lambda}-\eta_{\mu})\rangle_{B}\rhd c\rangle_{C}.$ Then, a direct computation shows that
$\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu}),x\boxtimes(\eta_{\lambda}-\eta_{\mu})\rangle_{B}$ $=\langle\eta_{\lambda}-\eta_{\mu},\langle x,x\rangle_{A}\rhd(\eta_{\lambda}-\eta_{\mu})\rangle_{B}$
$=(\eta_{\lambda}-\eta_{\mu})\phi(\langle x,x\rangle_{A})(\eta_{\lambda}-\eta_{\mu}).$
Therefore,
$\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c,x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c\rangle_{C}=\langle c,\psi((\eta_{\lambda}-\eta_{\mu})\phi(\langle x,x\rangle_{A})(\eta_{\lambda}-\eta_{\mu}))c\rangle_{C}.$
By the Cauchy-Schwarz inequality and since $(\eta_{\lambda}-\eta_{\mu})\phi(\langle x,x\rangle_{A})^{1/2}$ converges to 00, it is readily seen that $\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c,x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c\rangle_{C}$ converges to 00, so $S_{X}$ is well-defined. Moreover, commutation with the left and right actions are immediate.
Let $T_{X}:F(X)\boxtimes{}_{\phi}B\boxtimes{}_{\psi}C\to F(X)\boxtimes{}_{\psi\circ\phi}C$ be the continuous linear map given by $T_{X}(x\boxtimes b\boxtimes c)=x\boxtimes\psi(b)c$. Then,
$S_{X}(T_{X}(x\boxtimes b\boxtimes c))$ $=S_{X}(x\boxtimes\psi(b)c)$
$=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}\boxtimes\psi(b)c$
$=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}\boxtimes b\rhd c$
$=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}\lhd b\boxtimes c$
$=x\boxtimes b\boxtimes c.$
Therefore, by linearity and continuity of both $S_{X}$ and $T_{X}$, it follows that $S_{X}$ is surjective and $T_{X}$ is injective. It now suffices to show that $T_{X}$ is surjective. This will imply that $S_{X}$ is invertible with $T_{X}$ being the inverse. Let $x\in F(X)$ and $c\in C$. Since $F(X)$ is a non-degenerate $A$-bimodule, recall from Lemma 1.11 that the map $R_{X}^{-1}:F(X)\boxtimes A\to F(X)$ given by $R_{X}^{-1}(x\boxtimes a)=x\lhd a$ is a bimodule isomorphism. Then, let $y\in F(X)$ and $a\in A$ such that $y\lhd a=x$. A straightforward calculation shows that $T_{X}(y\boxtimes\phi(a)\boxtimes c)=x\boxtimes c$, so $T_{X}$ is surjective. Hence, $S_{X}=T_{X}^{-1}$ is an isomorphism for any $X\in\mathcal{C}$.
Finally, for any $X\in\mathcal{C}$, $T_{X}=\mathrm{id}_{F(X)}\boxtimes T$, where $T:{}_{\phi}B\boxtimes{}_{\psi}C\to{}_{\psi\circ\phi}C$ is given by $T(b\boxtimes c)=\psi(b)c$ for any $b\in B$ and $c\in C$. Since $T$ is an isometry, we conclude that $T_{X}$, and hence $S_{X}$ are isometric maps. ∎
Note that the proof of Lemma 3.7 also shows that the map $S_{X}$ does not depend on the choice of approximate unit.
Lemma 3.8.
Let $(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ and $(\psi,\mathbbm{w}):(B,G,I)\to(C,H,K)$ be cocycle morphisms. Then $(\psi\circ\phi,\mathbbm{w}*\mathbbm{v})$ is a cocycle morphism.
Proof.
Note that $\psi\circ\phi:A\to C$ is a $*{}^{*}$-homomorphism. Then, $(\mathbbm{w}*\mathbbm{v})_{X}$ is an isometry as it is a composition of isometries, and it is also an $A$-$C$-bimodule map as it is a composition of bimodule maps. Moreover, naturality of $\mathbbm{w}*\mathbbm{v}$ follows from naturality of $\mathbbm{w}\circ\mathbbm{v}$. Let us now check that the family of maps $\{(\mathbbm{w}*\mathbbm{v})_{X}\}$ satisfies the diagram
(3.5) {tikzcd}
for all $X,Y\in\mathcal{C}$.
Fix $X,Y\in\mathcal{C}$. By Definition 3.6, composing the two rightmost arrows in (3.5) yields that
(3.6) $(\mathbbm{w}*\mathbbm{v})_{X\otimes Y}\circ(J_{X,Y}\boxtimes\mathrm{id}_{C})$
$=(T\boxtimes\mathrm{id}_{H(X\otimes Y)})\circ(\mathbbm{w}\circ\mathbbm{v})_{X\otimes Y}\circ S_{X\otimes Y}\circ(J_{X,Y}\boxtimes\mathrm{id}_{C})$
(3.7) $=(T\boxtimes\mathrm{id}_{H(X\otimes Y)})\circ(\mathbbm{w}\circ\mathbbm{v})_{X\otimes Y}\circ(J_{X,Y}\boxtimes\mathrm{id}_{B}\boxtimes\mathrm{id}_{C})\circ(\mathrm{id}_{F(Y)}\boxtimes S_{X}).$
Then, by Lemma 3.5,
(3.8) $(\mathbbm{w}\circ\mathbbm{v})_{X\otimes Y}\circ(J_{X,Y}\boxtimes\mathrm{id}_{B}\boxtimes\mathrm{id}_{C})$
(3.9) $=(\mathrm{id}_{B}\boxtimes\mathrm{id}_{C}\boxtimes K_{X,Y})\circ\big{(}(\mathbbm{w}\circ\mathbbm{v})_{Y}\boxtimes\mathrm{id}_{H(X)}\big{)}\circ\big{(}\mathrm{id}_{F(Y)}\boxtimes(\mathbbm{w}\circ\mathbbm{v})_{X}\big{)}.$
Substituting (3.8) into (3.6), and using Definition 3.6, it follows that the diagram in (3.5) commutes.
It remains to check that $(\mathbbm{w}*\mathbbm{v})_{1_{\mathcal{C}}}:A\boxtimes{}_{\psi\circ\phi}C\to{}_{\psi\circ\phi}C\boxtimes C$ is given by $(\mathbbm{w}*\mathbbm{v})_{1_{\mathcal{C}}}(a\boxtimes c)=\psi(\phi(a))\boxtimes c.$ By definition,
$(\mathbbm{w}*\mathbbm{v})_{1_{\mathcal{C}}}=(T\boxtimes\mathrm{id}_{C})\circ(\mathrm{id}_{B}\boxtimes\mathbbm{w}_{1_{\mathcal{C}}})\circ(\mathbbm{v}_{1_{\mathcal{C}}}\boxtimes\mathrm{id}_{C})\circ S_{1_{\mathcal{C}}}.$
Let $\eta_{\lambda}$ be an approximate unit for $B$. Then, a direct calculation shows that
$(\mathbbm{w}*\mathbbm{v})_{1_{\mathcal{C}}}(a\boxtimes c)$ $=\lim\limits_{\lambda}(T\boxtimes\mathrm{id}_{C})\circ(\mathrm{id}_{B}\boxtimes\mathbbm{w}_{1_{\mathcal{C}}})\circ(\mathbbm{v}_{1_{\mathcal{C}}}\boxtimes\mathrm{id}_{C})(a\boxtimes\eta_{\lambda}\boxtimes c)$
$=\lim\limits_{\lambda}(T\boxtimes\mathrm{id}_{C})\circ(\mathrm{id}_{B}\boxtimes\mathbbm{w}_{1_{\mathcal{C}}})(\phi(a)\boxtimes\eta_{\lambda}\boxtimes c)$
$=\lim\limits_{\lambda}(T\boxtimes\mathrm{id}_{C})(\phi(a)\boxtimes\psi(\eta_{\lambda})\boxtimes c)$
$=\lim\limits_{\lambda}\psi(\phi(a))\psi(\eta_{\lambda})\boxtimes c$
$=\psi(\phi(a))\boxtimes c,$
for any $a\in A$ and any $c\in C$. Hence, $(\psi\circ\phi,\mathbbm{w}*\mathbbm{v})$ is a cocycle morphism. ∎
Remark 3.9.
With the same notation as in Definition 3.6, if $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are extendible cocycle representations, then it also follows that $(\psi\circ\phi,\mathbbm{w}*\mathbbm{v})$ is a well-defined cocycle representation.
Remark 3.10.
If the cocycle morphisms in Definition 3.6 are assumed to be non-degenerate, then the maps $S_{X}$ and $T$ are bimodule isomorphisms for any $X\in\mathcal{C}$. Therefore, the composition in Definition 3.6 (denoted ***) corresponds canonically to the composition of correspondence morphisms (denoted $\circ$).
The composition considered in Definition 3.6 defines a category. To show this we first reformulate the notion of cocycle morphism. Roughly speaking, all the information carried by the collection of isometries $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ can be encoded into a collection of linear maps $\{h^{X}:F(X)\to G(X)\}_{X\in\mathcal{C}}$ satisfying some conditions. This viewpoint will facilitate our constructions and proofs in later sections. Our approach is motivated by [6, Lemma 3.8] which introduces this alternative viewpoint in the unital setting.
First, we recall a way of extending an action $(F,J)$ of a $*{}^{*}$-tensor category $\mathcal{C}$ on $A$ to its matrix amplification $M_{n}(A)$. Consider the functor $F^{\mathrm{amp}(n)}:\mathcal{C}\rightarrow\mathrm{Corr}(M_{n}(A))$ that maps objects $X\in\mathcal{C}$ to the correspondence with underlying bimodule $F(X)\otimes M_{n}(\mathbb{C})$ with the right inner product defined by $\langle(x_{ij}),(y_{ij})\rangle_{M_{n}(A)}=(\sum_{l}\langle x_{li},y_{lj}\rangle_{A})$, right $M_{n}(A)$ action given by $(x_{ij})\lhd(a_{ij})=(\sum_{l}x_{il}\lhd a_{lj})$ and with the left action given by $(a_{ij})\rhd(x_{ij})=(\sum_{l}a_{il}\rhd x_{lj})$ for $a_{ij}\in A$ and $x_{ij},y_{ij}\in F(X)$ for $1\leq i,j\leq n$ [^11] This bimodule can be identified with the external tensor product of $F(X)$ and $M_{n}(\mathbb{C})$. Hence the inner product defines a Hilbert $M_{n}(A)$-bimodule (see [28]). For a morphism $T\in\mathrm{Hom}(X,Y)$ we let $F^{\mathrm{amp}(n)}(T)=T\otimes\mathrm{id}_{M_{n}(\mathbb{C})}$. Moreover, letting $J_{X,Y}^{\mathrm{amp}(n)}((y_{ij})\boxtimes(x_{ij})))=(\sum_{l}J_{X,Y}(y_{il}\boxtimes x_{lj}))$ for $x_{ij}\in F(X)$ and $y_{ij}\in F(Y)$ for $1\leq i,j\leq n$, it is a straightforward calculation that $(F^{\mathrm{amp}(n)},J^{\mathrm{amp}(n)})$ is an action of $\mathcal{C}$ on $M_{n}(A)$.
Lemma 3.11.
Let $\mathcal{C}$ be a $\mathrm{C}^{*}$-tensor category acting on $\mathrm{C}^{*}$-algebras $A$ and $B$ via $(A,F,J)$ and $(B,G,I)$ respectively, and let $\phi:A\rightarrow B$ be a $*{}^{*}$-homomorphism. Then there is a bijection between the families $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ corresponding to a cocycle morphism $(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ and families of linear maps:
$\{h^{X}:F(X)\rightarrow G(X)\}_{X\in\mathcal{C}}$
such that for any $X,Y\in\mathcal{C}$
  1. (i)
    $h^{X}(a\rhd x\lhd a^{\prime})=\phi(a)\rhd h^{X}(x)\lhd\phi(a^{\prime})$ for any $a,a^{\prime}\in A$;
  2. (ii)
    for any morphism $f\in\mathrm{Hom}(X,Y)$, $G(f)\circ h^{X}=h^{Y}\circ F(f)$;
  3. (iii)
    $\phi(\langle x,y\rangle_{A})=\langle h^{X}(x),h^{X}(y)\rangle_{B}$ for any $x,y\in F(X)$;
  4. (iv)
    the diagram:
[…]
    commutes;
  5. (v)
    $h^{1_{\mathcal{C}}}:A\to B$ is given by $h^{1_{\mathcal{C}}}(a)=\phi(a)$ for any $a\in A$.
Proof.
Suppose we are given a collection of linear maps $\{h^{X}\}$ satisfying the conditions listed above. Fix $\zeta_{\lambda}$ an approximate unit for $A$. For all $X\in\mathcal{C}$ let $\mathbbm{v}_{X}:F(X)\boxtimes{}_{\phi}B\to{}_{\phi}B\boxtimes G(X)$ be given by
(3.10) $\mathbbm{v}_{X}(x\boxtimes b)=\lim\limits_{\lambda}\phi(\zeta_{\lambda})\boxtimes h^{X}(x)\lhd b,\quad x\in F(X),\ b\in B.$
First, we need to show that $\mathbbm{v}_{X}$ is well-defined. It suffices to show that the net $\phi(\zeta_{\lambda})\boxtimes h^{X}(x)\lhd b$ is Cauchy. A straightforward computation using the definition of the inner product and (i) gives that
$\|\phi(\zeta_{\lambda}-\zeta_{\mu})\boxtimes h^{X}(x)\lhd b\|^{2}$ $=\|\phi(\zeta_{\lambda}-\zeta_{\mu})\rhd h^{X}(x)\lhd b\|^{2}$
(3.11) $=\|h^{X}(\zeta_{\lambda}\rhd x-\zeta_{\mu}\rhd x)\lhd b\|^{2}.$
Since the bimodule $F(X)$ is non-degenerate, denoting the left action by $\sigma:A\to\mathcal{L}(F(X))$, we have that $\sigma(\zeta_{\lambda})$ converges strictly to $1_{\mathcal{L}(F(X))}$. Then, for any $x\in F(X)$, $\zeta_{\lambda}\rhd x=\sigma(\zeta_{\lambda})x$ converges to $x$. Moreover, condition (iii) implies that $h^{X}$ is continuous, so the right hand side of (3.11) converges to zero and hence $\|\phi(\zeta_{\lambda}-\zeta_{\mu})\boxtimes h^{X}(x)\lhd b\|^{2}$ converges to 00. Therefore the formula in (3.10) gives a well-defined map $\mathbbm{v}_{X}$ for all $X\in\mathcal{C}$.
The maps $\mathbbm{v}_{X}$ are linear by linearity of $h^{X}$ and naturality follows by using naturality of the family $\{h^{X}\}$ given by condition (ii). It is straightforward to see that $\mathbbm{v}_{X}$ commutes with the right $B$-action, and for any $a\in A$, $b\in B$, $X\in\mathcal{C}$ and $x\in F(X)$
$\mathbbm{v}_{X}(a\rhd(x\boxtimes b))$ $=\lim\limits_{\lambda}\phi(\zeta_{\lambda})\boxtimes h^{X}(a\rhd x)\lhd b$
$=\lim\limits_{\lambda}\phi(\zeta_{\lambda})\boxtimes\phi(a)\rhd h^{X}(x)\lhd b$
$=\lim\limits_{\lambda}\phi(\zeta_{\lambda})\phi(a)\boxtimes h^{X}(x)\lhd b$
$=\phi(a)\boxtimes h^{X}(x)\lhd b$
$=a\rhd\mathbbm{v}_{X}(x\boxtimes b).$
Therefore, $\mathbbm{v}_{X}$ is an $A$-$B$-bimodule map. Moreover, using (v),
$\mathbbm{v}_{1_{\mathcal{C}}}(a\boxtimes b)$ $=\lim\limits_{\lambda}\phi(\zeta_{\lambda})\boxtimes\phi(a)b$
$=\lim\limits_{\lambda}\phi(\zeta_{\lambda})\phi(a)\boxtimes b$
$=\phi(a)\boxtimes b,\ \forall a\in A,\ b\in B.$
To prove that $(\phi,\mathbbm{v})$ defines a cocycle morphism, it remains to show that each map $\mathbbm{v}_{X}$ is an isometry and the family $\{\mathbbm{v}_{X}\}$ is such that diagram (3.1) commutes. Let us first show that each map $\mathbbm{v}_{X}$ is an isometry. For any $x\in F(X)$ and any $b\in B$, using (i), we have
$|\mathbbm{v}_{X}(x\boxtimes b)|^{2}$ $=\lim\limits_{\lambda}\langle\phi(\zeta_{\lambda})\rhd h^{X}(x)\lhd b,\phi(\zeta_{\lambda})\rhd h^{X}(x)\lhd b\rangle$
$=\lim\limits_{\lambda}\langle h^{X}(\zeta_{\lambda}\rhd x)\lhd b,h^{X}(\zeta_{\lambda}\rhd x)\lhd b\rangle$
$=\lim\limits_{\lambda}b^{*}\langle h^{X}(\zeta_{\lambda}\rhd x),h^{X}(\zeta_{\lambda}\rhd x)\rangle b.$
On the other hand, (iii) yields that
$|x\boxtimes b|^{2}$ $=\langle b,\langle x,x\rangle_{A}\rhd b\rangle$
$=\langle\phi(|x|)b,\phi(|x|)b\rangle$
$=b^{*}\phi(\langle x,x\rangle_{A})b$
$=b^{*}\langle h^{X}(x),h^{X}(x)\rangle_{B}b.$
Similarly, as $h^{X}$ is continuous by (iii) and $F(X)$ is non-degenerate, it follows that $h^{X}(\zeta_{\lambda}\rhd x)$ converges to $h^{X}(x)$, which shows that $\mathbbm{v}_{X}$ is an isometry when restricted to elementary tensors. To show that $\mathbbm{v}_{X}$ acts as an isometry on sums of the form $\sum_{i=1}^{n}x_{i}\boxtimes b_{i}$ for $x_{i}\in F(X)$, $b_{i}\in B$ and $X\in\mathcal{C}$ we consider the amplified actions $(F^{\mathrm{amp}(n)},J^{\mathrm{amp}(n)})$ and $(G^{\mathrm{amp}(n)},I^{\mathrm{amp}(n)})$ on $M_{n}(A)$ and $M_{n}(B)$ respectively. It follows from a direct computation that the family of linear maps $h^{X,\mathrm{amp}(n)}:F^{\mathrm{amp}(n)}(X)\rightarrow G^{\mathrm{amp}(n)}(X)$ defined by $(x_{ij})\mapsto(h^{X}(x_{ij}))$ for $X\in\mathcal{C}$ and $x_{ij}\in F(X)$ satisfies conditions (i)-(v) with the amplified homomorphism $\phi:M_{n}(A)\rightarrow M_{n}(B)$. Therefore $\mathbbm{v}_{X}^{\mathrm{amp}(n)}$ defined as in (3.10) but instead with the pair $(\phi,h^{X,\mathrm{amp}(n)})$ is an isometry when restricted to elementary tensors. Choose $\mathbf{X}$ in $F^{\mathrm{amp}(n)}(X)$ with first row given by the vector $(x_{1},x_{2},\ldots,x_{n})$ and zero elsewhere and $\mathbf{B}$ in $M_{n}(B)$ have first column $(b_{1},b_{2},\ldots,b_{n})$ and zero elsewhere. Now, by definition
$\|\mathbf{X}\boxtimes\mathbf{B}\|^{2}$ $=\|\langle\mathbf{B},\langle\mathbf{X},\mathbf{X}\rangle_{M_{n}(A)}\rhd\mathbf{B}\rangle\|$
$=\|\sum_{i=1}^{n}x_{i}\boxtimes b_{i}\|^{2}$
and similarly through a direct computation
$\|\mathbbm{v}_{X}^{\mathrm{amp}(n)}(\mathbf{X}\boxtimes\mathbf{B})\|^{2}=\|\mathbbm{v}_{X}(\sum_{i=1}^{n}x_{i}\boxtimes b_{i})\|^{2}.$
As $\mathbbm{v}_{X}^{\mathrm{amp}(n)}$ is an isometry when restricted to elementary tensors, it follows that $\mathbbm{v}_{X}$ is an isometry.
It remains to check that the diagram
(3.12) {tikzcd}
commutes for all $X,Y\in\mathcal{C}$.
Starting with an elementary tensor $y\boxtimes x\boxtimes b$ with $X,Y\in\mathcal{C},y\in F(Y),x\in F(X)$ and $b\in B$ and following the two rightmost maps of the diagram, we get that
$y\boxtimes x\boxtimes b\mapsto J_{X,Y}(y\boxtimes x)\boxtimes b\mapsto\lim\limits_{\lambda}\phi(\zeta_{\lambda})\boxtimes h^{X\otimes Y}(J_{X,Y}(y\boxtimes x))\lhd b.$
Moreover, using that the family of linear maps satisfies condition (iv), this composition coincides with the mapping
$y\boxtimes x\boxtimes b\mapsto\lim\limits_{\lambda}\phi(\zeta_{\lambda})\boxtimes I_{X,Y}(h^{Y}(y)\boxtimes h^{X}(x))\lhd b.$
Again, starting with $y\boxtimes x\boxtimes b$ but now following the three leftmost arrows in diagram (3.12) we get
$y\boxtimes x\boxtimes b$ $\mapsto\lim\limits_{\lambda}y\boxtimes\phi(\zeta_{\lambda})\boxtimes h^{X}(x)\lhd b$
$\mapsto\lim\limits_{\lambda}\lim\limits_{\mu}\phi(\zeta_{\mu})\boxtimes h^{Y}(y)\lhd\phi(\zeta_{\lambda})\boxtimes h^{X}(x)\lhd b$
$=\lim\limits_{\mu}\phi(\zeta_{\mu})\boxtimes h^{Y}(y)\boxtimes h^{X}(x)\lhd b$
$\mapsto\lim\limits_{\mu}\phi(\zeta_{\mu})\boxtimes I_{X,Y}(h^{Y}(y)\boxtimes h^{X}(x)\lhd b)$
$=\lim\limits_{\mu}\phi(\zeta_{\mu})\boxtimes I_{X,Y}(h^{Y}(y)\boxtimes h^{X}(x))\lhd b,$
where the first equality holds since $h^{Y}(y)\lhd\phi(\zeta_{\lambda})=h^{Y}(y\lhd\zeta_{\lambda})$ converges to $h^{Y}(y)$. So (3.12) commutes and $(\phi,\mathbbm{v})$ is a cocycle morphism.
Now, consider the map $\Psi:\{h^{X}\}\to\{\mathbbm{v}_{X}\}$ given by the formula in (3.10). We claim that $\Psi$ is independent of the choice of approximate unit. Indeed let $\zeta_{\lambda}$ and $\xi_{\lambda}$ be two approximate units for $A$. Similarly as in (3.11) we have that, $\|\phi(\zeta_{\lambda}-\xi_{\lambda})\boxtimes h^{X}(x)\lhd b\|=\|h^{X}(\zeta_{\lambda}\rhd x-\xi_{\lambda}\rhd x)\lhd b\|$, which converges to 00. Hence, the map $\mathbbm{v}_{X}$ is independent of the choice of approximate unit, and so $\Psi$ is well-defined.
Conversely, suppose we have a cocycle morphism $(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ and for each $X\in\mathcal{C}$ let $h^{X}:F(X)\to G(X)$ be given by
[…]
where $\iota(x)=\lim_{\lambda}x\boxtimes\eta_{\lambda}$ for some approximate unit $\eta_{\lambda}$ of $B$ and all $x\in F(X)$, and $f$ is the map given by $f(b\boxtimes y)=b\rhd y$ for all $b\in B$ and $y\in G(X)$. Note that $f$ is an $A$-$B$-bimodule isomorphism if we see $G(X)$ as a left $A$-module through $\phi$ (i.e. $\phi(a)\rhd_{B}f(y)=a\rhd_{A}f(y)$ for all $y\in{}_{\phi}B\boxtimes G(X)$ and $a\in A$).
To check that $\iota$ is well-defined, we show that the net $x\boxtimes\eta_{\lambda}$ is Cauchy for all $X\in\mathcal{C}$ and $x\in F(X)$. Precisely, one has that $\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu}),x\boxtimes(\eta_{\lambda}-\eta_{\mu})\rangle_{B}=\langle\eta_{\lambda}-\eta_{\mu},\langle x,x\rangle_{A}\rhd(\eta_{\lambda}-\eta_{\mu})\rangle_{B}=|\phi(\langle x,x\rangle_{A})^{1/2}(\eta_{\lambda}-\eta_{\mu})|^{2}$. This converges to 00 since the image of $\phi$ is contained in $B$ and $\eta_{\lambda}$ is an approximate unit for $B$.
We now check that the family $\{h^{X}\}$ defined above satisfies the required compatibility conditions. Since each of the maps in (3.13) are linear, we get that $h^{X}$ is linear. To see (i), note that $f$, $\mathbbm{v}_{X}$ and $\iota$ are left module maps so $\phi(a)\rhd h^{X}(x)=h^{X}(a\rhd x)$ for all $a\in A$ and $x\in F(X)$. Moreover, as $f$ and $\mathbbm{v}_{X}$ are right $B$-module maps and $\iota$ is a right $A$-module map
$h^{X}(x)\lhd\phi(a)$ $=\lim\limits_{\lambda}f(\mathbbm{v}_{X}(x\boxtimes\eta_{\lambda}\phi(a)))=h^{X}(x\lhd a).$
Hence, $h^{X}$ satisfies (i). It is straightforward to see that $h^{X}$ satisfies (ii) by naturality of $\mathbbm{v}$.
Note that $\mathbbm{v}_{X}$ and $f$ are isometries. So one has that for any $x\in F(X)$, $\langle h^{X}(x),h^{X}(x)\rangle_{B}=\langle\iota(x),\iota(x)\rangle_{B}=\lim_{\lambda}\langle\eta_{\lambda},\phi(\langle x,x\rangle_{A})\eta_{\lambda}\rangle_{B}=\phi(\langle x,x\rangle_{A})$ and (iii) follows from the polarisation identity. Condition (iv) follows from the fact that the maps $\mathbbm{v}_{X}$ satisfy the diagram in (3.1). Finally,
$h^{1_{\mathcal{C}}}(a)=\lim\limits_{\lambda}f(\mathbbm{v}_{1_{\mathcal{C}}}(a\boxtimes\eta_{\lambda}))=\lim\limits_{\lambda}f(\phi(a)\boxtimes\eta_{\lambda})=\phi(a).$
Now, let $\Phi:\{\mathbbm{v}_{X}\}\to\{h^{X}\}$ be the map induced by the formula in (3.13). Note that $\iota$ and hence $\Phi$ is independent of the choice of approximate unit. Indeed, let $\eta_{\lambda}$ and $\xi_{\lambda}$ be two approximate units for $B$. We show that the net $x\boxtimes(\eta_{\lambda}-\xi_{\lambda})$ converges to 00 for any $x\in F(X)$. Note that
$|x\boxtimes(\eta_{\lambda}-\xi_{\lambda})|^{2}$ $=\langle(\eta_{\lambda}-\xi_{\lambda}),\langle x,x\rangle_{A}\rhd(\eta_{\lambda}-\xi_{\lambda})\rangle$
$=|\phi(\langle x,x\rangle_{A})^{1/2}(\eta_{\lambda}-\xi_{\lambda})|^{2},$
which converges to 00.
We claim that $\Phi$ and $\Psi$ are inverses to each other. First we show that $\Phi\circ\Psi$ is the identity map. For any $X\in\mathcal{C}$ and $x\in F(X)$, it follows that
$\Phi(\Psi(h^{X}))(x)=f\Big{(}\Psi(h^{X})\Big{(}\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}\Big{)}\Big{)}=f\Big{(}\lim\limits_{\mu}\lim\limits_{\lambda}\phi(\zeta_{\mu})\boxtimes h^{X}(x)\lhd\eta_{\lambda}\Big{)}.$
Since $\eta_{\lambda}$ is an approximate unit for $B$ and $G(X)$ is non-degenerate, it follows that
$\Phi(\Psi(h^{X}))(x)=f\Big{(}\lim\limits_{\mu}\phi(\zeta_{\mu})\boxtimes h^{X}(x)\Big{)}=\lim\limits_{\mu}\phi(\zeta_{\mu})\rhd h^{X}(x).$
As $h^{X}$ satisfies condition (i), $\Phi(\Psi(h^{X}))(x)=\lim\limits_{\mu}h^{X}(\zeta_{\mu}\rhd x)$. Thus, it suffices to show that $\|h^{X}(\zeta_{\mu}\rhd x-x)\|\to 0$. This follows by continuity of $h^{X}$ and that $F(X)$ is non-degenerate.
To prove that $\Psi\circ\Phi$ is the identity map, note that
$\Psi(\Phi(\mathbbm{v}_{X}))(x\boxtimes b)$ $=\lim\limits_{\mu}\phi(\zeta_{\mu})\boxtimes\Phi(\mathbbm{v}_{X})(x)\lhd b$
$=\lim\limits_{\mu}\phi(\zeta_{\mu})\boxtimes f\Big{(}\mathbbm{v}_{X}\Big{(}\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}\Big{)}\Big{)}\lhd b$
$=\lim\limits_{\mu}\phi(\zeta_{\mu})\boxtimes f\Big{(}\mathbbm{v}_{X}\Big{(}\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}b\Big{)}\Big{)}$
$=\lim\limits_{\mu}\phi(\zeta_{\mu})\boxtimes f(\mathbbm{v}_{X}(x\boxtimes b)).$
Applying $f$ to both sides
$f(\Psi(\Phi(\mathbbm{v}_{X}))(x\boxtimes b))$ $=\lim\limits_{\mu}\phi(\zeta_{\mu})\rhd_{B}f(\mathbbm{v}_{X}(x\boxtimes b))$
$=\lim\limits_{\mu}\zeta_{\mu}\rhd_{A}f(\mathbbm{v}_{X}(x\boxtimes b))$
$=\lim\limits_{\mu}f(\mathbbm{v}_{X}((\zeta_{\mu}\rhd x)\boxtimes b))$
$=f(\mathbbm{v}_{X}(x\boxtimes b))$
as $\zeta_{\mu}\rhd x$ converges to $x$. Hence, we reach the conclusion by composing with $f^{-1}:G(X)\to{}_{\phi}B\boxtimes G(X)$ given by $f^{-1}(x)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes x$. ∎
Note that this alternative picture only holds for cocycle morphisms. In the generality of cocycle representations, the maps $h^{X}$ may not be well-defined. That is because if $(\phi,\mathbbm{v})$ is a cocycle representation, $\phi$ can land in $M(B)\setminus B$. As $\eta_{\lambda}$ is an approximate unit for $B$, $\|\phi(\langle x,x\rangle_{A})^{1/2}(\eta_{\lambda}-\eta_{\mu})\|^{2}$ need not converge to 00.
Remark 3.12.
As in Remark 3.2, if the acting category $\mathcal{C}$ is semisimple, then the family of linear maps $\{h^{X}\}_{X\in\mathcal{C}}$ is uniquely determined by the family $\{h^{X}\}_{X\in\mathrm{Irr}(\mathcal{C})}$. Precisely, if $X\cong\bigoplus_{i}X_{i}\in\mathcal{C}$ is the decomposition as a direct sum of elements in $\mathrm{Irr}(\mathcal{C})$, then $F(X)$ is naturally isomorphic to $\bigoplus_{i}F(X_{i})$ and $G(X)$ is naturally isomorphic to $\bigoplus_{i}G(X_{i})$. Then, the map $h^{X}$ is $\bigoplus_{i}h^{X_{i}}$. In particular, it suffices to check that a family of linear maps $\{h^{X}:F(X)\rightarrow G(X)\}_{X\in\mathrm{Irr}(\mathcal{C})}$ satisfy the conditions of Lemma 3.11 to yield a cocycle morphism, with understanding condition (iv) as $I_{X,Y}\circ(h^{Y}\boxtimes h^{X})=\bigoplus_{i}h^{X_{i}}\circ J_{X,Y}$ for all $X,Y\in\mathrm{Irr}(\mathcal{C})$ and $X\otimes Y\cong\bigoplus_{i}X_{i}$ is the irreducible decomposition.
Remark 3.13.
When $\phi:A\rightarrow B$ is extendible and $(\phi,\{h^{X}\}_{X\in\mathcal{C}})$ is a cocycle morphism, condition (i) also follows for $a,a^{\prime}\in\mathcal{M}(A)$. Precisely, if $a,a^{\prime}\in\mathcal{M}(A)$ then $h^{X}(a\rhd x\lhd a^{\prime})=\phi_{p}(a)\rhd h^{X}(x)\lhd\phi_{p}(a^{\prime})$ for any $X\in\mathcal{C}$ and $x\in F(X)$ where $p\in\mathcal{M}(B)$ is the projection associated to $\phi$. Similarly, for any $u,u^{\prime}\in\mathcal{U}(\mathcal{M}(A))$ it follows that $h^{X}(u\rhd x\lhd u^{\prime})=\phi^{\dagger}(u)\rhd h^{X}(x)\lhd\phi^{\dagger}(u^{\prime})$. Indeed, for an approximate unit $e_{\lambda}$ of $A$, $\phi(e_{\lambda})$ converges to $p$ and
$p\rhd h^{X}(x)=\lim\limits_{\lambda}\phi(e_{\lambda})\rhd h^{X}(x)=\lim\limits_{\lambda}h^{X}(e_{\lambda}\rhd x)=h^{X}(x).$
So $\phi^{\dagger}(u)\rhd h^{X}(x)\lhd\phi^{\dagger}(u^{\prime})=\phi_{p}(u)\rhd h^{X}(x)\lhd\phi_{p}(u^{\prime})=h^{X}(u\rhd x\lhd u^{\prime})$.
Example 3.14.
Suppose $\mathcal{C}=\mathrm{Hilb}(\Gamma)$ and $(\phi,\{h^{g}\}_{g\in\Gamma}):(A,\alpha,J)\to(B,\beta,I)$ is an extendible cocycle morphism. Recall from Example 3.3 that $\mathbbm{v}_{g}(a\boxtimes b)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes\mathbbm{u}_{g}\phi(a)b,$ where $\eta_{\lambda}$ is an approximate unit for $B$ and $(\mathbbm{u}_{g})_{g\in\Gamma}\subseteq\mathcal{M}(B)$. Then for any $g\in\Gamma$ and $a\in A$,
$h^{g}(a)=f\left(\mathbbm{v}_{g}\left(\lim\limits_{\lambda}a\boxtimes\eta_{\lambda}\right)\right)=f\left(\lim\limits_{\lambda}\eta_{\lambda}\boxtimes\mathbbm{u}_{g}\phi(a)\right)=\mathbbm{u}_{g}\phi(a).$
For convenience, we will denote a cocycle morphism by $(\phi,h)$, where $h$ denotes the collection of linear maps $\{h^{X}\}_{X\in\mathcal{C}}$. We now discuss their composition.
Lemma 3.15.
If $(\phi,h):(A,F,J)\to(B,G,I)$ and $(\psi,l):(B,G,I)\to(C,H,K)$ are cocycle morphisms, then $(\psi\circ\phi,l\circ h):(A,F,J)\to(C,H,K)$ is a cocycle morphism and coincides with the composition of $(\phi,h)$ and $(\psi,l)$.
Proof.
Clearly $\psi\circ\phi:A\to C$ is a $*{}^{*}$-homomorphism and $\{l^{X}\circ h^{X}:F(X)\to H(X)\}_{X\in\mathcal{C}}$ is a family of linear maps. Conditions (i), (ii), (iii), and (v) are immediate. The compatibility with the tensor product follows by stacking the diagrams in (iv) of Lemma 3.11 for $h$ and $l$. Thus, $(\psi\circ\phi,l\circ h):(A,F,J)\to(C,H,K)$ induces a cocycle morphism by Lemma 3.11.
Suppose that $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are cocycle morphisms associated to the families of linear maps $\{h^{X}\}_{X\in\mathcal{C}}$ and $\{l^{X}\}_{X\in\mathcal{C}}$ respectively. We claim that the cocycle morphism $(\psi\circ\phi,\mathbbm{w}*\mathbbm{v})$ has the associated family of linear maps $\{l^{X}\circ h^{X}\}_{X\in\mathcal{C}}$.
Recall from Definition 3.6 that
$(\mathbbm{w}*\mathbbm{v})_{X}=(T\boxtimes\mathrm{id}_{H(X)})\circ(\mathbbm{w}\circ\mathbbm{v})_{X}\circ S_{X},$
where $(\mathbbm{w}\circ\mathbbm{v})_{X}=(\mathrm{id}_{\phi}\boxtimes\mathbbm{w}_{X})\circ(\mathbbm{v}_{X}\boxtimes\mathrm{id}_{\psi})$ (see (3.3)). Start with an elementary tensor $x\boxtimes c$ and let $\zeta_{\mu}$ be an approximate unit of $A$ and $\eta_{\lambda}$ be an approximate unit of $B$. Following the composition of maps defining $(\mathbbm{w}*\mathbbm{v})_{X}$, we get that
$x\boxtimes c$ $\mapsto\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}\boxtimes c$
$\mapsto\lim\limits_{\lambda}\lim\limits_{\mu}\phi(\zeta_{\mu})\boxtimes h^{X}(x)\lhd\eta_{\lambda}\boxtimes c$
$=\lim\limits_{\mu}\phi(\zeta_{\mu})\boxtimes h^{X}(x)\boxtimes c$
$\mapsto\lim\limits_{\mu}\lim\limits_{\lambda}\phi(\zeta_{\mu})\boxtimes\psi(\eta_{\lambda})\boxtimes l^{X}(h^{X}(x))\lhd c$
$\mapsto\lim\limits_{\mu}\lim\limits_{\lambda}\psi(\phi(\zeta_{\mu}))\psi(\eta_{\lambda})\boxtimes l^{X}(h^{X}(x))\lhd c$
$=\lim\limits_{\mu}\psi(\phi(\zeta_{\mu}))\boxtimes l^{X}(h^{X}(x))\lhd c,$
where the last equality follows since $\eta_{\lambda}$ is an approximate unit, and in particular fixes $\phi(\zeta_{\mu})$ in the limit. But this is precisely the formula in (3.10) corresponding to the family of linear maps $l\circ h$, so the two compositions agree. ∎
Lemma 3.11 shows that any cocycle morphism can be equivalently represented by a pair $(\phi,h)$. Moreover Lemma 3.15 shows that in the latter picture the composition of cocycle morphisms translates to the composition of the underlying $*{}^{*}$-homomorphisms and linear maps. From now on we will freely identify these two pictures.
Lemma 3.16.
The class of $\mathcal{C}$-$\mathrm{C}^{*}$-algebras $(A,F,J)$, together with cocycle morphisms $(\phi,h):(A,F,J)\to(B,G,I)$ defines a category with respect to the composition in Lemma 3.15.
Proof.
The composition in Lemma 3.15 is easily seen to be associative. Moreover, for any cocycle morphism $(\phi,h)$, the cocycle morphisms $(\mathrm{id}_{A},\{\mathrm{id}_{F(X)}\})$ and $(\mathrm{id}_{B},\{\mathrm{id}_{G(X)}\})$ are left and right identities respectively. ∎
In the spirit of [40, Definition 1.16], we have constructed a category of $\mathcal{C}$-$\mathrm{C}^{*}$-algebras.
Definition 3.17.
The generalised cocycle category $\mathrm{C}^{*}_{\mathcal{C}}$ is defined as the category whose objects are $\mathcal{C}$-$\mathrm{C}^{*}$-algebras and whose morphisms are cocycle morphisms. Composition of morphisms in $\mathrm{C}^{*}_{\mathcal{C}}$ is defined in Lemma 3.15. On any object $(A,F,J)$, the identity morphism in this category is given by $(\mathrm{id}_{A},\{\mathrm{id}_{F(X)}\}_{X\in\mathcal{C}})$. A cocycle morphism $(\phi,h):(A,F,J)\to(B,G,I)$ is invertible in this category if and only if $\phi:A\to B$ is an isomorphism and $h^{X}$ is bijective for any $X\in\mathcal{C}$, in which case the inverse is given by $(\phi^{-1},h^{-1})$. Following the terminology in [40], we say that an invertible morphism in this category is a cocycle conjugacy.
4. Inductive limits
In this section, we construct inductive limits in $\mathrm{C}^{*}_{\mathcal{C}}$ for semisimple $\mathcal{C}$. This is done in [6, Proposition 4.4], when restricted to unital $\mathrm{C}^{*}$-algebras and unital, injective connecting maps. Our approach is slightly different and does not need these assumptions. Before starting our construction, let us set up some notation.
Let $\mathcal{C}$ be a semisimple $\mathrm{C}^{*}$-tensor category, and $A_{n}$ be a sequence of separable $\mathrm{C}^{*}$-algebras on which $\mathcal{C}$ acts via the pair $(F_{n},J^{(n)})$. Then, let
(4.1) $(\phi_{n},h_{n}):(A_{n},F_{n},J^{(n)})\to(A_{n+1},F_{n+1},J^{(n+1)})$
be a sequence of cocycle morphisms. Recall that the $\mathrm{C}^{*}$-inductive limit $A=\lim\limits_{\longrightarrow}\{A_{n},\phi_{n}\}$ is defined as the completion of
$A^{(0)}=\left\{(a_{n})_{n\geq 1}\in\frac{\bigoplus_{\ell^{\infty}}A_{n}}{\bigoplus_{c_{0}}A_{n}}:\exists N\in\mathbb{N}\ \text{s.t.}\ \forall n\geq N,\phi_{n}(a_{n})=a_{n+1}\right\}$
with respect to the topology induced by the norm $\|(a_{n})_{n\geq 1}\|=\lim\limits_{n\to\infty}\|a_{n}\|_{A_{n}}$. Recall that the connecting maps $\phi_{n,\infty}:A_{n}\to A$ are given by
(4.2) $(\phi_{n,\infty}(a_{n}))_{k}=\begin{cases}\phi_{n,k}(a_{n}),\ k\geq n\\ 0,\quad\quad\quad k<n\end{cases}$
for all $n\geq 1$, where we adopt the standard notation $\phi_{n,m}:=\phi_{m-1}\circ\ldots\circ\phi_{n}$ for any $m>n\geq 1$.
Similarly, for any $X\in\mathcal{C}$ and any $m>n\geq 1$, we consider the natural family $h_{n,m}^{X}:F_{n}(X)\to F_{m}(X)$ obtained by composition, with the convention that $h_{n,n+1}^{X}=h_{n}^{X}$.
To build an action on $A$, we start by constructing bimodules that will form the image of the functor. Essentially, for any $X\in\mathcal{C}$, we can build a Hilbert $A$-$A$-bimodule as an inductive limit of $F_{n}(X)$. The construction is very similar to the one in (2.4).
Define
$F^{(0)}(X)=\Big{\{}(x_{n})_{n\geq 1}\in\frac{\bigoplus_{\ell^{\infty}}F_{n}(X)}{\bigoplus_{c_{0}}F_{n}(X)}:\exists N\in\mathbb{N}\ \text{s.t.}\ \forall n\geq N,h_{n}^{X}(x_{n})=x_{n+1}\Big{\}},$
where the norm on $F_{n}(X)$ is induced by the right inner product.
For any $x=(x_{n})_{n\geq 1},y=(y_{n})_{n\geq 1}\in F^{(0)}(X)$ and any $a=(a_{n})_{n\geq 1}\in A^{(0)}$, we can define
(4.3) $x\lhd a=(x_{n}\lhd a_{n})_{n\geq 1}$
and
(4.4) $\langle x,y\rangle_{A^{(0)}}=(\langle x_{n},y_{n}\rangle_{A_{n}})_{n\geq 1}.$
Lemma 4.1.
For any $X\in\mathcal{C}$, $F^{(0)}(X)$ equipped with the structure in (4.3) and (4.4) is a right pre-Hilbert-$A^{(0)}$-module.
Proof.
We start by checking that the right action is well-defined. Firstly, if $x=(x_{n})_{n\geq 1}\in F^{(0)}(X)$ and $a=(a_{n})_{n\geq 1}\in A^{(0)}$, then $(x_{n}\lhd a_{n})_{n\geq 1}$ induces an element in $F^{(0)}(X)$. By (i) of Lemma 3.11, $h_{n}^{X}(x_{n}\lhd a_{n})=h_{n}^{X}(x_{n})\lhd\phi_{n}(a_{n})$ for any $n\in\mathbb{N},a_{n}\in A_{n}$, and $x_{n}\in F_{n}(X)$. Moreover, since $x\in F^{(0)}(X)$ and $a\in A^{(0)}$, there exists $N\in\mathbb{N}$ such that $h_{n}^{X}(x_{n})=x_{n+1}$ and $\phi_{n}(a_{n})=a_{n+1}$ for all $n\geq N$. Hence,
$h_{n}^{X}(x_{n}\lhd a_{n})=x_{n+1}\lhd a_{n+1}$
for all $n\geq N$, as required.
That (4.3) is independent of the choice of representative sequences, follows exactly as in the proof of Lemma 2.4. Thus, (4.3) gives a well-defined right action of $A^{(0)}$.
We now check that (4.4) gives a well-defined right pre-inner product. Firstly, if $x=(x_{n})_{n\geq 1},y=(y_{n})_{n\geq 1}\in F^{(0)}(X)$, then $\langle x,y\rangle$ is an element of $A^{(0)}$. By (iii) of Lemma 3.11 applied to each map $h_{n}^{X}$, we have that
$\phi_{n}(\langle x_{n},y_{n}\rangle_{A_{n}})=\langle h_{n}^{X}(x_{n}),h_{n}^{X}(y_{n})\rangle_{A_{n+1}}.$
Moreover, since $x,y\in F^{(0)}(X)$, there exists $N\in\mathbb{N}$ such that for any $n\geq N$, $\langle h_{n}^{X}(x_{n}),h_{n}^{X}(y_{n})\rangle_{A_{n+1}}=\langle x_{n+1},y_{n+1}\rangle_{A_{n+1}}$. Hence,
$\phi_{n}(\langle x_{n},y_{n}\rangle_{A_{n}})=\langle x_{n+1},y_{n+1}\rangle_{A_{n+1}}$
for all $n\geq N$, as required.
That (4.4) is independent of the choice of representative sequences, follows as in the proof of Lemma 2.4. Thus, (4.4) gives a well-defined $A^{(0)}$-valued map. It is now straightforward to check that this function is right linear, left conjugate linear, and antisymmetric. Finally, it is clear that $\langle x,x\rangle_{A^{(0)}}\geq 0$ and $\langle x,x\rangle_{A^{(0)}}=0$ if and only if $\langle x_{n},x_{n}\rangle_{A_{n}}$ converges to 00 i.e. $x=0$ in $F^{(0)}(X)$. Thus, the conclusion follows. ∎
For any $X\in\mathcal{C}$, since $A^{(0)}$ is a dense $*{}^{*}$-subalgebra of $A$, combining Lemma 4.1 and [37, Lemma 2.16], we form the completion of $F^{(0)}(X)$, denoted by $F(X)$. This is a right-Hilbert $A$-module.
Lemma 4.2.
For any $X\in\mathcal{C}$, with the notation above, $F(X)$ is a non-degenerate right Hilbert $A$-$A$-bimodule.
Proof.
We start by defining a left $A$-action on $F(X)$. For any $x=(x_{n})_{n\geq 1}\in F^{(0)}(X)$ and any $a=(a_{n})_{n\geq 1}\in A^{(0)}$, we can define
(4.5) $a\rhd x=(a_{n}\rhd x_{n})_{n\geq 1}.$
The fact that (4.5) is a well-defined left $A^{(0)}$-action on $F^{(0)}(X)$ follows as in the proof of Lemma 4.1. Moreover, the left $A^{(0)}$-action is adjointable, since the left action of $A_{n}$ on $F_{n}(X)$ is adjointable for any $n\geq 1$. Thus, using [37, Lemma 2.16], it extends to a left action of $A$ on $F(X)$ by density. ∎
We now show that the assignment $X\to F(X)$ extends to an action of $\mathcal{C}$ on $A$.
Lemma 4.3.
With the notation above, $F:\mathcal{C}\to\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$ is a $\mathrm{C}^{*}$-functor. Moreover, there exists a unitary natural isomorphism
$J:=\{J_{X,Y}:F(Y)\boxtimes F(X)\to F(X\otimes Y):X,Y\in\mathcal{C}\}$
such that the pair $(F,J)$ is an action of $\mathcal{C}$ on $A$.
Proof.
Since $F_{n}(X)\in\mathrm{Corr}_{0}^{\mathrm{sep}}(A_{n})$ for any $X\in\mathcal{C}$ and any $n\in\mathbb{N}$, it follows that $F(X)\in\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$ for any $X\in\mathcal{C}$.
Suppose $X,Y\in\mathcal{C}$ and $f:X\to Y$ is a morphism in $\mathcal{C}$. We define $F(f)(x)=(F_{n}(f)(x_{n}))_{n\geq 1}$ for any $x\in F^{(0)}(X)$. As $\mathcal{C}$ is semisimple, by decomposing $X$ and $Y$ into irreducibles one has that $f=\bigoplus\lambda_{i}\mathrm{id}_{X_{i}}$ for finitely many non-zero scalars $\lambda_{i}$ and $X_{i}\in\mathrm{Irr}(\mathcal{C})$. Therefore, as each $F_{n}$ is a $*{}^{*}$-functor, $F_{n}(f)$ is uniformly bounded by the supremum of $\lambda_{i}$ and so $F(f)$ is bounded on $F^{(0)}(X)$. Therefore, one may extend $F(f)$ to a bounded $A$-$A$-bimodule map on $F(X)$. This operator is further adjointable as each $F_{n}(f)$ is adjointable. It is straightforward to check that $F$ defines a $\mathrm{C}^{*}$-functor.
Furthermore, for any $X,Y\in\mathcal{C}$, $x=(x_{n})_{n\geq 1}\in F^{(0)}(X)$, and $y=(y_{n})_{n\geq 1}\in F^{(0)}(Y)$, let
$J_{X,Y}(y\boxtimes x)=(J_{X,Y}^{(n)}(y_{n}\boxtimes x_{n}))_{n\geq 1}.$
We claim that $J_{X,Y}$ is well-defined. Firstly, note that $J_{X,Y}$ is independent of the choice of representative sequences by continuity of $J_{X,Y}^{(n)}$ for any $n\in\mathbb{N}$.
Secondly, there exists $N\in\mathbb{N}$ such that for any $n\geq N$,
$h_{n}^{X}(x_{n})=x_{n+1}\ \text{and}\ h_{n}^{Y}(y_{n})=y_{n+1}.$
Then, Lemma 3.11 gives that
$h_{n}^{X\otimes Y}(J_{X,Y}^{(n)}(y_{n}\boxtimes x_{n}))=J_{X,Y}^{(n+1)}(h_{n}^{Y}(y_{n})\boxtimes h_{n}^{X}(x_{n}))=J_{X,Y}^{(n+1)}(y_{n+1}\boxtimes x_{n+1}),$
so the image of $J_{X,Y}$ is indeed contained in $F^{(0)}(X\otimes Y)$. By density, we can now extend $J_{X,Y}$ to $F(Y)\boxtimes F(X)$.
Moreover, since each $J^{(n)}$ is a unitary natural isomorphism, so is $J$. Finally, since each family $\{J_{X,Y}^{(n)}:X,Y\in\mathcal{C}\}$ satisfies a commuting diagram as in (2.1), so does the collection of maps $\{J_{X,Y}:X,Y\in\mathcal{C}\}$. Hence, the pair $(F,J)$ gives an action of $\mathcal{C}$ on $A$. ∎
We now show the existence of sequential inductive limits in $\mathrm{C}^{*}_{\mathcal{C}}$.
Proposition 4.4.
The triple $(A,F,J)$ constructed above defines the inductive limit of the inductive system in (4.1).
Proof.
For any $n\geq 1$, we need to define cocycle morphisms
$(\phi,h)_{n,\infty}:(A_{n},F_{n},J^{(n)})\to(A,F,J)$
such that for any $n\geq 1$, the diagram
[…]
commutes.
For any $n\geq 1$, we have a $*{}^{*}$-homomorphism $\phi_{n,\infty}:A_{n}\to A$. To define a cocycle morphism from $(A_{n},F_{n},J^{(n)})$ to $(A,F,J)$, it suffices to find a collection of linear maps $\{h_{n,\infty}^{X}:F_{n}(X)\to F(X)\}_{X\in\mathcal{C}}$ satisfying the conditions in Lemma 3.11. If $X\in\mathcal{C}$ and $x_{n}\in F_{n}(X)$ for $n\in\mathbb{N}$, we define
(4.7) $(h_{n,\infty}^{X}(x_{n}))_{k}=\begin{cases}h_{n,k}^{X}(x_{n})&\ k\geq n\\ 0&k<n.\end{cases}$
By convention, $h_{n,n}^{X}(x_{n})=x_{n}$ for any $X\in\mathcal{C}$ and any $x_{n}\in F_{n}(X)$. Then, $h_{n,\infty}^{X}$ is a linear map into $F(X)$. Moreover, the conditions of Lemma 3.11 are checked pointwise since $(\phi_{n,k},h_{n,k})$ is a cocycle morphism for any $k>n$. Thus, for any $n\geq 1$, the pair $(\phi_{n,\infty},h_{n,\infty})$ defines a cocycle morphism from $(A_{n},F_{n},J^{(n)})$ to $(A,F,J)$.
We now check that the triple $(A,F,J)$ is the inductive limit of the sequence of triples $(A_{n},F_{n},J^{(n)})$, together with connecting morphisms $(\phi_{n},h_{n})$. Firstly, we claim that (4.6) commutes. Since $A$ is an inductive limit for the system $(A_{n},\phi_{n})_{n\geq 1}$, it follows that for any $n\geq 1$, $\phi_{n,\infty}=\phi_{n+1,\infty}\circ\phi_{n}$. Therefore, it suffices to check that for any $X\in\mathcal{C}$ and any $n\geq 1$, we have that $h_{n,\infty}^{X}=h_{n+1,\infty}^{X}\circ h_{n}^{X}$. This follows by (4.7) since for any $k>n+1$, $h_{n,k}^{X}=h_{n+1,k}^{X}\circ h_{n}^{X}$.
It remains to check that $(A,F,J)$ satisfies the universal property. Precisely, let $B$ be a $\mathrm{C}^{*}$-algebra and $(B,G,I)$ defining an action of $\mathcal{C}$ on $B$. Suppose there exists a sequence of cocycle morphisms $(\psi,l)_{n,\infty}:(A_{n},F_{n},J^{(n)})\to(B,G,I)$ such that the diagram
[…]
commutes for any $m>n\geq 1$. We claim that there exists a unique cocycle morphism $(\Phi,r):(A,F,J)\to(B,G,I)$ such that for any $n\geq 1$,
$(\Phi,r)\circ(\phi,h)_{n,\infty}=(\psi,l)_{n,\infty}.$
By the universal property of the inductive limit $A$, there exists a unique $*{}^{*}$-homomorphism $\Phi:A\to B$ such that $\Phi\circ\phi_{n,\infty}=\psi_{n,\infty}$ for all $n\geq 1$. Then, for any $X\in\mathcal{C}$, the union $\bigcup_{k\geq 1}h_{k,\infty}^{X}(F_{k}(X))$ is dense in $F(X)$. Since $l_{k,\infty}^{X}$ is uniformly bounded in $k$, (4.8) yields that $\mathrm{Ker}\ h_{k,\infty}^{X}\subseteq\mathrm{Ker}\ l_{k,\infty}^{X}$ for any $X\in\mathcal{C}$ and any $k\in\mathbb{N}$. Therefore, for any $x_{k}\in F_{k}(X)$, define $r^{X}(h_{k,\infty}^{X}(x_{k}))=l_{k,\infty}^{X}(x_{k})$. As $l_{k,\infty}^{X}$ is uniformly bounded in $k$, we may extend $r^{X}$ to a well-defined linear map $r^{X}:F(X)\to G(X)$. The fact that $r^{X}$ satisfies the conditions appearing in Lemma 3.11 is routine and essentially follows from the fact that for each $k\geq 1$, $(\psi_{k,\infty},l_{k,\infty})$ and $(\phi_{k,\infty},h_{k,\infty})$ are cocycle morphisms. For the convenience of the reader, we check condition (iii) of Lemma 3.11. Let $X\in\mathcal{C}$, $x_{k},y_{k}\in F_{k}(X)$. Then,
$\langle r^{X}(h_{k,\infty}^{X}(x_{k})),r^{X}(h_{k,\infty}^{X}(y_{k}))\rangle$ $=\langle l_{k,\infty}^{X}(x_{k}),l_{k,\infty}^{X}(y_{k})\rangle$
$=\psi_{k,\infty}(\langle x_{k},y_{k}\rangle)$
$=\Phi\circ\phi_{k,\infty}(\langle x_{k},y_{k}\rangle)$
$=\Phi(\langle h_{k,\infty}^{X}(x_{k}),h_{k,\infty}^{X}(y_{k})\rangle).$
The remaining conditions follow similarly. Therefore, $(\Phi,r)$ is the unique cocycle morphism such that $(\Phi,r)\circ(\phi,h)_{n,\infty}=(\psi,l)_{n,\infty}$. Hence, the triple $(A,F,J)$ is the inductive limit of the system in (4.1). ∎
Remark 4.5.
Note that if the connecting maps $\phi_{n}:A_{n}\to A_{n+1}$ of the inductive system are extendible, then the inductive limit connecting map $\phi_{k,\infty}:A_{k}\to A$ is extendible for any $k\geq 1$.
4.1. Topology induced on cocycle representations
[…]
Remark 4.6.
Recall that for any fixed $X\in\mathcal{C}$, $\mathbbm{v}_{X}:F(X)\boxtimes{}_{\phi}B\to{}_{\phi}B\boxtimes G(X)$ and $\mathbbm{w}_{X}:F(X)\boxtimes{}_{\psi}B\to{}_{\psi}B\boxtimes G(X)$. By the construction of the tensor product, both $F(X)\boxtimes{}_{\phi}B$ and $F(X)\boxtimes{}_{\psi}B$ are completions of quotients of the algebraic tensor product $F(X)\odot B$. To compare the difference between the maps, it suffices to do so on their respective images of elementary tensors $x\odot b$ in the right Hilbert $B$-module $B\boxtimes G(X)$.
Definition 4.7.
For any finite set $K\subset\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$ and any compact sets $\mathcal{F}^{B}\subset B$, and $\mathcal{F}^{X}\subset F(X)$ for any $X\in K$, denote $\mathcal{F}=\mathcal{F}^{B}\times\left(\coprod_{X\in K}\mathcal{F}^{X}\right)$. Then, we define the pseudometric
$d_{\mathcal{F}}\big{(}(\phi,\mathbbm{v}),(\psi,\mathbbm{w})\big{)}$ $=\max\limits_{(b,\xi_{X})\in\mathcal{F}}\|\mathbbm{v}_{X}(\xi_{X}\boxtimes_{\phi}b)-\mathbbm{w}_{X}(\xi_{X}\boxtimes_{\psi}b)\|,$
where the norm is induced by the right inner product on the right Hilbert $B$-module $B\boxtimes G(X)$.
Remark 4.8.
Note that $F(1_{\mathcal{C}})=A$ and for any $a\in A$ and $b\in B$ $\mathbbm{v}_{1_{\mathcal{C}}}(a\boxtimes_{\phi}b)=\phi(a)b$ and $\mathbbm{w}_{1_{\mathcal{C}}}(a\boxtimes_{\psi}b)=\psi(a)b$. Therefore, the family of pseudometrics $d_{\mathcal{F}}$ bounds the pointwise distance between $\phi$ and $\psi$ in the strict topology.
We will only use the topology generated by the family of pseudometrics $d_{\mathcal{F}}$ in the setting of cocycle morphisms, so let us derive the relevant subset topology.
Lemma 4.9.
Let $(\phi_{\lambda},\mathbbm{v}_{\lambda}),(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ be cocycle morphisms with associated families of linear maps $\{h_{\lambda}^{X}\}$ and $\{h^{X}\}$. Then $(\phi_{\lambda},\mathbbm{v}_{\lambda})$ converges to $(\phi,\mathbbm{v})$ if and only if for any $X\in\mathrm{Irr}(\mathcal{C})$, $h_{\lambda}^{X}$ converges pointwise to $h^{X}$in the norm induced by the right inner product.
Proof.
Suppose that for any $X\in\mathrm{Irr}(\mathcal{C})$, $h_{\lambda}^{X}$ converges pointwise to $h^{X}$ in the norm induced by the right inner product. Let $\zeta_{\mu}$ be an approximate unit of $A$ and recall from (3.10) that for any $X\in\mathcal{C}$, $x\in F(X)$, and $b\in B$
$\mathbbm{v}_{X}(x\boxtimes_{\phi}b)=\lim\limits_{\mu}\phi(\zeta_{\mu})\boxtimes h^{X}(x)\lhd b$
and
$(\mathbbm{v}_{\lambda})_{X}(x\boxtimes_{\phi_{\lambda}}b)=\lim\limits_{\mu}\phi_{\lambda}(\zeta_{\mu})\boxtimes h_{\lambda}^{X}(x)\lhd b.$
Fix $X\in\mathcal{C}$, $x\in F(X)$, and $b\in B$ and recall that since $F(X)$ is non-degenerate, there exists $a\in A$ and $y\in F(X)$ such that $a\rhd y=x$. In particular, (i) of Lemma 3.11 yields that $h^{X}(x)=\phi(a)\rhd h^{X}(y)$ and $h^{X}_{\lambda}(x)=\phi_{\lambda}(a)\rhd h^{X}_{\lambda}(y)$. Therefore, by the definition of the tensor product, one has that
(4.9) $\mathbbm{v}_{X}(x\boxtimes_{\phi}b)=\lim\limits_{\mu}\phi(\zeta_{\mu})\lhd\phi(a)\boxtimes h^{X}(y)\lhd b=\phi(a)\boxtimes h^{X}(y)\lhd b.$
Similarly,
(4.10) $(\mathbbm{v}_{\lambda})_{X}(x\boxtimes_{\phi_{\lambda}}b)=\phi_{\lambda}(a)\boxtimes h^{X}_{\lambda}(y)\lhd b.$
Moreover, by (v) of Lemma 3.11 $h^{1_{\mathcal{C}}}(a)=\phi(a)$ and $h^{1_{\mathcal{C}}}_{\lambda}(a)=\phi_{\lambda}(a)$ for any $a\in A$. By (4.9) and (4.10), it follows that for any $X\in\mathrm{Irr}(\mathcal{C})$, $(\mathbbm{v}_{\lambda})_{X}$ converges pointwise to $\mathbbm{v}_{X}$. Therefore, $(\phi_{\lambda},\mathbbm{v}_{\lambda})$ converges to $(\phi,\mathbbm{v})$ by definition of the pseudometrics in Definition 4.7.
Conversely, suppose that $(\phi_{\lambda},\mathbbm{v}_{\lambda})$ converges to $(\phi,\mathbbm{v})$. Then $(\phi_{\lambda},\mathbbm{v}_{\lambda})$ satisfies the Cauchy criterion with respect to every pseudometric $d_{\mathcal{F}}$ above. Recall from (3.13) that $h_{\lambda}^{X}(x)=f((\mathbbm{v}_{\lambda})_{X}(\iota(x))$ for any $X\in\mathcal{C}$ and $x\in F(X)$, where $f$ and $\iota$ are defined in (3.13). Since $(\mathbbm{v}_{\lambda})_{X}$ is Cauchy in the point-norm topology and $f$ is continuous, it follows that $h_{\lambda}^{X}$ is Cauchy for any $X\in\mathrm{Irr}(\mathcal{C})$. But the norm induced by the right inner product is complete, so $h_{\lambda}^{X}$ converges pointwise to $h^{X}$ for any $X\in\mathrm{Irr}(\mathcal{C})$. ∎
Example 4.10.
Let $\Gamma$ be a countable discrete group. Let $(\alpha,\mathfrak{u}):\Gamma\curvearrowright A$ and $(\beta,\mathfrak{v}):\Gamma\curvearrowright B$ be two twisted actions on $\mathrm{C}^{*}$-algebras and $(\phi,\mathbbm{v}^{\prime})$, $(\psi,\mathbbm{w}^{\prime}):(A,\alpha,\mathfrak{u})\to(B,\beta,\mathfrak{v})$ be two extendible cocycle morphisms as in Definition 1.21. Recall from Example 3.3 that $(\phi,\mathbbm{v}^{\prime})$ and $(\psi,\mathbbm{w}^{\prime})$ induce cocycle morphisms $(\phi,\mathbbm{v}),(\psi,\mathbbm{w}):(A,\alpha,\mathfrak{u})\to(B,\beta,\mathfrak{v})$ in the sense of Definition 3.1, where we view $(\alpha,\mathfrak{u})$ and $(\beta,\mathfrak{v})$ as $\mathrm{Hilb}(\Gamma)$-actions. So,
(4.11) $\mathbbm{v}_{g}([a\boxtimes b]_{\phi})=\lim\limits_{\mu}\eta_{\mu}\boxtimes\mathbbm{v}_{g}^{\prime}\phi(a)b,\quad\mathbbm{w}_{g}([a\boxtimes b]_{\psi})=\lim\limits_{\mu}\eta_{\mu}\boxtimes\mathbbm{w}_{g}^{\prime}\psi(a)b,$
for all $a\in A$, $b\in B$, $g\in\Gamma$, and $\eta_{\mu}$ an approximate unit for $B$. In this case, the topology is induced by the pseudometrics
(4.12) $d_{\mathcal{F}}\big{(}(\phi,\mathbbm{v}),(\psi,\mathbbm{w})\big{)}=\max\limits_{(b,a_{g})\in\mathcal{F}}\|\mathbbm{v}_{g}^{\prime}\phi(a_{g})b-\mathbbm{w}_{g}^{\prime}\psi(a_{g})b\|,$
where $\mathcal{F}=\mathcal{F}^{B}\times\left(\coprod_{g\in K}\mathcal{F}^{A}\right)$, for compact sets $\mathcal{F}^{A}\subset A,\mathcal{F}^{B}\subset B$, and finite $K\subset\Gamma$ containing $1_{\Gamma}$. Equivalently, by taking an approximate unit for $B$, the topology is generated by the pseudometrics
(4.13) $d_{\mathcal{F}^{A}\times K}\big{(}(\phi,\mathbbm{v}),(\psi,\mathbbm{w})\big{)}=\max\limits_{a_{g}\in\mathcal{F}}\|\mathbbm{v}_{g}^{\prime}\phi(a_{g})-\mathbbm{w}_{g}^{\prime}\psi(a_{g})\|.$
In [40, Definition 2.5], Szabó defines a topology on cocycle morphisms, which using the conventions of this paper, is generated by the family of pseudometrics
(4.14) $d_{\mathcal{F}}\big{(}(\phi,\mathbbm{v}^{\prime}),(\psi,\mathbbm{w}^{\prime})\big{)}=\max\limits_{a\in\mathcal{F}^{A}}\|\phi(a)-\psi(a)\|+\max\limits_{g\in K}\max\limits_{b\in\mathcal{F}^{B}}\|b((\mathbbm{v}_{g}^{\prime})^{*}-(\mathbbm{w}_{g}^{\prime})^{*})\|.$
The convergence with respect to the family of pseudometrics in (4.14) implies convergence with respect to the family of pseudometrics in (4.12). If $\phi$ and $\psi$ are non-degenerate, then $\phi(A)B$ and $\psi(A)B$ are dense in $B$. Moreover, the case $g=1$ in (4.12) recovers the pointwise difference of the morphisms. Thus, convergence with respect to the family of pseudometrics in (4.12) implies convergence with respect to the family of pseudometrics in (4.14). However, these topologies are different outside the non-degenerate setting, with the topology induced by the family of pseudometrics in (4.12) being coarser than the topology induced by the family of pseudometrics in (4.14). Finally, we would like to point out that the topology induced by (4.13) coincides with that used in [31].
We finish our discussion by noticing that the composition of cocycle morphisms is jointly continuous with respect to the topology defined above. This fact will be used in Section 6 in the context of asymptotic unitary equivalence.
Lemma 4.11.
Let $(\phi,h):(A,F,J)\to(B,G,I)$ and $(\psi,l):(B,G,I)\to(C,H,K)$ be two cocycle morphisms. Then the composition map given by
$[(\phi,h),(\psi,l)]\mapsto(\psi,l)\circ(\phi,h)$
is jointly continuous.
Proof.
Let $(\phi_{\lambda},h_{\lambda}):(A,F,J)\to(B,G,I)$ and $(\psi_{\lambda},l_{\lambda}):(B,G,I)\to(C,H,K)$ be two convergent nets with limits $(\phi,h)$ and $(\psi,l)$ respectively. First, note that for any $X\in\mathrm{Irr}(\mathcal{C})$, $l^{X}\circ h^{X}=\lim\limits_{\lambda\to\infty}l_{\lambda}^{X}\circ h_{\lambda}^{X}$ in the point-norm topology. Then, the case $X=1_{\mathcal{C}}$ yields that $\psi\circ\phi=\lim\limits_{\lambda\to\infty}\psi_{\lambda}\circ\phi_{\lambda}$ holds in the point-norm topology. Using the composition formula in Lemma 3.15 and Lemma 4.9, the conclusion follows. ∎
4.2. Approximate unitary equivalence
We can now introduce a notion of approximate unitary equivalence that will be crucial to perform equivariant Elliott intertwining arguments. We start by defining unitary equivalence for cocycle morphisms and then use the topology on the space of morphisms to obtain an approximate notion of unitary equivalence.
Suppose $\mathcal{C}$ is a semisimple $\mathrm{C}^{*}$-tensor category with countably many isomorphism classes of simple objects acting on a $\mathrm{C}^{*}$-algebra $B$. We denote this action by the triple $(B,G,I)$. For any $u$ unitary in $\mathcal{M}(B)$, we consider $\mathrm{Ad}(u):B\to B$ to be the $*{}^{*}$-homomorphism given by $b\mapsto ubu^{*}$. Then $\mathrm{Ad}(u)$ induces a Hilbert $B$-bimodule ${}_{\mathrm{Ad}(u)}B$. The map $T_{u}:{}_{\mathrm{Ad}(u)}B\to B$ given by $b\mapsto u^{*}b$ is a bimodule isomorphism. But for any $X\in\mathcal{C}$, $G(X)\boxtimes B\cong B\boxtimes G(X)$, so there exists a unitary isomorphism $(\mathbbm{v}_{u})_{X}:G(X)\boxtimes{}_{\mathrm{Ad}(u)}B\to{}_{\mathrm{Ad}(u)}B\boxtimes G(X)$. It follows that $(\mathrm{Ad}(u),\mathbbm{v}_{u}):(B,G,I)\to(B,G,I)$ is a cocycle morphism. We denote by $h_{u}=\{h_{u}^{X}:G(X)\to G(X)\}_{X\in\mathcal{C}}$ the collection of linear maps corresponding to the cocycle morphism induced by $\mathrm{Ad}(u)$.
Lemma 4.12.
[…]
Proof.
For any $X\in\mathcal{C}$, consider the bimodule maps $L_{X}:G(X)\to B\boxtimes G(X)$ and $R_{X}:G(X)\to G(X)\boxtimes B$ given by $L_{X}(x)=\lim\limits_{\mu}\eta_{\mu}\boxtimes x$ and $R_{X}(x)=\lim\limits_{\mu}x\boxtimes\eta_{\mu}$ for some $\eta_{\mu}$ quasicentral approximate unit of $B$ with respect to $\mathcal{M}(B)$.
As the bimodule $G(X)$ is non-degenerate for any $X\in\mathcal{C}$, it follows from Lemma 1.11 that
$u\rhd x=L_{X}^{-1}(u\rhd L_{X}(x))\quad\mathrm{and}\quad x\lhd u^{*}=R_{X}^{-1}(R_{X}(x)\lhd u^{*}).$
Then, consider the map
$\iota:G(X)\to G(X)\boxtimes{}_{\mathrm{Ad}(u)}B\to G(X)\boxtimes B$
given by $x\mapsto\lim\limits_{\mu}x\boxtimes\eta_{\mu}\mapsto\lim\limits_{\mu}x\boxtimes u^{*}\eta_{\mu}$. Since $\eta_{\mu}$ is quasicentral,
$\iota(x)=\lim\limits_{\mu}x\boxtimes\eta_{\mu}u^{*}=R_{X}(x)\lhd u^{*}.$
Then, $L_{X}\circ R_{X}^{-1}\circ\iota:G(X)\to B\boxtimes G(X)$ is the map given by
$L_{X}(R_{X}^{-1}(\iota(x)))=L_{X}(x\lhd u^{*}).$
Finally, we consider the map
$f:B\boxtimes G(X)\to{}_{\mathrm{Ad}(u)}B\boxtimes G(X)\to G(X)$
given by $f(b\boxtimes x)=L_{X}^{-1}(u\rhd(b\boxtimes x))$.
Following the construction in Lemma 3.11 and since the family of maps $\{h_{u}^{X}\}$ is independent of the chosen approximate unit, it follows that
$h_{u}^{X}=f\circ L_{X}\circ R_{X}^{-1}\circ\iota.$
Therefore, for any $X\in\mathcal{C}$ and any $x\in G(X)$,
$h_{u}^{X}(x)=f(L_{X}(x\lhd u^{*}))=L_{X}^{-1}(u\rhd L_{X}(x\lhd u^{*}))=u\rhd x\lhd u^{*},$
which finishes the proof. ∎
Definition 4.13.
  1. (i)
[…]
    $(\mathrm{Ad}(u),\mathbbm{v}_{u})\circ(\phi,\mathbbm{v})=(\psi,\mathbbm{w}).$
  2. (ii)
    Let $(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ be a cocycle representation and $u\in\mathcal{M}(B)$ a unitary. We say that $(\psi,\mathbbm{w})$ is an approximate unitary conjugate of $(\phi,\mathbbm{v})$ if there exists a net of unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
    $\psi(a)=\lim\limits_{\lambda}u_{\lambda}\phi(a)u_{\lambda}^{*},$
    and
[…]
    for all $a\in A$, $b\in B$, $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$. We denote this by $(\psi,\mathbbm{w})\precapprox_{u}(\phi,\mathbbm{v})$. [^14] Note that this means precisely that $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})\circ(\phi,\mathbbm{v})$ converges to $(\psi,\mathbbm{w})$ with respect to the topology in Definition 4.7.
  3. (iii)
    We say that $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are approximately unitarily equivalent if $(\psi,\mathbbm{w})\precapprox_{u}(\phi,\mathbbm{v})$ and $(\phi,\mathbbm{v})\precapprox_{u}(\psi,\mathbbm{w})$. This will be denoted by $(\phi,\mathbbm{v})\approx_{u}(\psi,\mathbbm{w})$.
In general, we will be interested in these notions when $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are cocycle morphisms, so let us record the following lemma.
Lemma 4.14.
If $(\phi,\mathbbm{v}),(\psi,\mathbbm{w}):(A,F,J)\to(B,G,I)$ are cocycle morphisms with associated families of linear maps $\{h^{X}\}$ and $\{l^{X}\}$ respectively, then $(\psi,\mathbbm{w})$ is an approximate unitary conjugate of $(\phi,\mathbbm{v})$ if and only if there exists a net of unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
[…]
for any $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$.
Proof.
By definition, $(\psi,\mathbbm{w})$ is an approximate unitary conjugate of $(\phi,\mathbbm{v})$ if and only if there exists a net of unitaries $u_{\lambda}$ in $\mathcal{M}(B)$ inducing a net of cocycle morphisms $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})$ such that $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})\circ(\phi,\mathbbm{v})$ converges to $(\psi,\mathbbm{w})$. Suppose the cocycle morphisms $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})$ have associated families of linear maps $\{h_{u_{\lambda}}^{X}\}$. By Lemma 4.9, $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})\circ(\phi,\mathbbm{v})$ converges to $(\psi,\mathbbm{w})$ if and only if $h_{u_{\lambda}}^{X}\circ h^{X}$ converges pointwise to $l^{X}$ for any $X\in\mathrm{Irr}(\mathcal{C})$. Moreover, by Lemma 4.12,
$h_{u_{\lambda}}^{X}(h^{X}(x))=u_{\lambda}\rhd h^{X}(x)\lhd u_{\lambda}^{*}$
for all $X\in\mathcal{C}$ and all $x\in F(X)$, so the conclusion follows. ∎
Remark 4.15.
There is a key difference to Szabó’s notion of approximate unitary conjugacy in [40, Definition 2.8]. The notion of approximate unitary conjugacy in Definition 4.13 is symmetric, even when restricted to possibly degenerate morphisms. In Lemma 4.16, we show that approximate unitary conjugacy implies approximate unitary equivalence. This is a consequence of the difference in topologies (see Example 4.10).
Lemma 4.16.
Let $(\phi,h),(\psi,l):(A,F,J)\to(B,G,I)$ be cocycle morphisms. If $(\psi,l)$ is an approximate unitary conjugate of $(\phi,h)$, then $(\phi,h)$ and $(\psi,l)$ are approximately unitarily equivalent.
Proof.
Suppose there exist unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
[…]
for any $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$. We show that
[…]
for any $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$.
Note that for any $X\in\mathcal{C}$ and any $x\in F(X)$, we have that
$h_{u_{\lambda}^{*}}^{X}(h_{u_{\lambda}}^{X}(h^{X}(x)))=u_{\lambda}^{*}\rhd h_{u_{\lambda}}^{X}(h^{X}(x))\lhd u_{\lambda}=1_{\mathcal{M}(B)}\rhd h^{X}(x)\lhd 1_{\mathcal{M}(B)}.$
But $G(X)$ is non-degenerate, so $1_{\mathcal{M}(B)}\rhd h^{X}(x)\lhd 1_{\mathcal{M}(B)}=h^{X}(x)$ by Lemma 1.11. Therefore, for any $X\in\mathcal{C}$ and any $x\in F(X)$,
$\|h_{u_{\lambda}^{*}}^{X}\circ l^{X}(x)-h^{X}(x)\|=\|h_{u_{\lambda}^{*}}^{X}(l^{X}(x)-h_{u_{\lambda}}^{X}\circ h^{X}(x))\|.$
Take $y_{\lambda}=l^{X}(x)-h_{u_{\lambda}}^{X}\circ h^{X}(x)$ and note that
$\langle h_{u_{\lambda}^{*}}^{X}(y_{\lambda}),h_{u_{\lambda}^{*}}^{X}(y_{\lambda})\rangle_{B}=\mathrm{Ad}(u_{\lambda}^{*})(\langle y_{\lambda},y_{\lambda}\rangle_{B}).$
[…]
We finish this section by defining asymptotic unitary equivalence for cocycle representations.
Definition 4.17.
[…]
$\psi(a)=\lim\limits_{t\to\infty}u_{t}\phi(a)u_{t}^{*},$
and
$\|\mathbbm{w}_{X}(x\boxtimes_{\psi}b)-(\mathbbm{v}_{u_{t}}*\mathbbm{v})_{X}(x\boxtimes_{\phi}b)\|\overset{t\to\infty}{\longrightarrow}0$
for all $a\in A$, $b\in B$, $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$. This will be denoted by $(\phi,\mathbbm{v})\cong_{u}(\psi,\mathbbm{w})$. [^15] Following Definition 4.13, we could have defined asymptotic unitary conjugacy. However, this implies asymptotic unitary equivalence as in Lemma 4.16.
The same argument as in Lemma 4.14 gives the following equivalent characterisation for asymptotic unitary equivalence of cocycle morphisms.
Lemma 4.18.
If $(\phi,\mathbbm{v}),(\psi,\mathbbm{w}):(A,F,J)\to(B,G,I)$ are cocycle morphisms with associated families of linear maps $\{h^{X}\}_{X\in\mathcal{C}}$ and $\{l^{X}\}_{X\in\mathcal{C}}$ respectively, then $(\psi,\mathbbm{w})$ is asymptotically unitarily equivalent to $(\phi,\mathbbm{v})$ if and only if there exists a strictly continuous map $u:[0,\infty)\to\mathcal{U}(\mathcal{M}(B))$ such that
$\|l^{X}(x)-u_{t}\rhd h^{X}(x)\lhd u_{t}^{*}\|\overset{t\to\infty}{\longrightarrow}0$
for any $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$.
Remark 4.19.
Let $(\phi_{1},h_{1}):(A,F,J)\to(B,G,I),(\psi_{1},l_{1}):(B,G,I)\to(C,H,K)$ be two cocycle morphisms. Suppose that $(\phi_{1},h_{1})$ is unitarily equivalent to $(\phi_{2},h_{2})$ and $(\psi_{1},l_{1})$ is unitarily equivalent to $(\psi_{2},l_{2})$. If $\psi_{1}$ is extendible, then $(\psi_{1}\circ\phi_{1},l_{1}\circ h_{1})$ is unitarily equaivalent to $(\psi_{2}\circ\phi_{2},l_{2}\circ h_{2})$. Indeed, let $u\in\mathcal{U}(\mathcal{M}(B))$ and $v\in\mathcal{U}(\mathcal{M}(C))$ be such that $u\rhd h_{1}^{X}(x)\lhd u^{*}=h_{2}^{X}(x)$ and $v\rhd l_{1}^{X}(y)\lhd v^{*}=l_{2}^{X}(y)$ for any $X\in\mathcal{C}$, $x\in F(X)$, and $y\in G(X)$. Then, $v\psi_{1}^{\dagger}(u)\rhd l_{1}^{X}(h_{1}^{X}(x))\lhd(v\psi_{1}^{\dagger}(u))^{*}=l_{2}^{X}(h_{2}^{X}(x))$. In general, the vertical composition in the $2$-category of correspondences is not well defined outside of the extendible setting.
We will finish by recording the following lemma which will be used in Section 6.
Lemma 4.20.
Let $(\phi,h):(A,F,J)\to(B,G,I)$ and $(\psi,l):(B,G,I)\to(C,H,K)$ be two cocycle morphisms which are asymptotically unitarily equivalent to cocycle conjugacies. Then their composition $(\psi\circ\phi,l\circ h)$ is asymptotically unitarily equivalent to a cocycle conjugacy.
Proof.
Suppose that $(\Phi,H)$ and $(\Psi,L)$ are cocycle conjugacies such that
$(\Phi,H)\cong_{u}(\phi,h)\quad\mathrm{and}\quad(\Psi,L)\cong_{u}(\psi,l).$
By Lemma 4.11, composition is jointly continuous. Moreover, since $\Psi$ is extendible, $(\Psi\circ\Phi,L\circ H)\cong_{u}(\psi\circ\phi,l\circ h)$ by Remark 4.19. ∎
5. Two-sided Elliott intertwining
With our setup we may now use Elliott’s abstract framework from [12] to show Theorem D.
Proof.
We check the conditions of the second theorem in [12]. Our underlying category is the subcategory of $\mathrm{C}^{*}_{\mathcal{C}}$ of separable $*{}^{*}$-algebras, together with extendible cocycle morphisms (see Definition 3.17). In this category, Lemma 4.12 gives us a notion of inner automorphisms in the sense of [12]. Moreover, the topology of the morphism spaces in this category, as in Lemma 4.9, is induced by a complete metric. Indeed, let $K_{n}$ be an increasing sequence of finite sets containing $1_{\mathcal{C}}$ such that $\bigcup_{n\in\mathbb{N}}K_{n}=\mathrm{Irr}(\mathcal{C})$. For any $X\in\mathrm{Irr}(\mathcal{C})$, choose a sequence of contractions $\mu_{n}^{X}$ which is dense in the unit ball of $F(X)$. Then, the assignment
$\left((\phi,h),(\psi,l)\right)\mapsto\sum_{n=1}^{\infty}2^{-n}\max_{X\in K_{n}}\max_{m\leq n}\|h^{X}(\mu_{m}^{X})-l^{X}(\mu_{m}^{X})\|.$
yields a complete metric recovering the topology on the space of morphisms. Furthermore, composition with an inner automorphism is isometric. Since the composition of cocycle morphisms is jointly continuous by Lemma 4.11, all the conditions of the second theorem in [12] are satisfied. Hence, the conclusion follows. ∎
However, we also decide to give the technical argument in full, picking up some more general results in the process. We start by introducing the setup for performing approximate intertwining arguments in the spirit of [40, Definition 3.3].
Definition 5.1.
Let $\mathcal{C}$ be a semisimple $\mathrm{C}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F_{n},J^{(n)}):\mathcal{C}\curvearrowright A_{n}$ and $(G_{n},I^{(n)}):\mathcal{C}\curvearrowright B_{n}$ be sequences of actions on separable $\mathrm{C}^{*}$-algebras. Let
$(\phi_{n},\{h_{n}^{X}\}_{X\in\mathcal{C}}):(A_{n},F_{n},J^{(n)})\to(A_{n+1},F_{n+1},J^{(n+1)})$
and
$(\psi_{n},\{l_{n}^{X}\}_{X\in\mathcal{C}}):(B_{n},G_{n},I^{(n)})\to(B_{n+1},G_{n+1},I^{(n+1)})$
be sequences of cocycle morphisms which we view as two inductive systems in the category $\mathrm{C}^{*}_{\mathcal{C}}$. [^16] As in previous sections, we will denote these morphisms by $(\phi_{n},h_{n})$ and $(\psi_{n},l_{n})$ for ease of notation.
Consider two sequences of cocycle morphisms
$(\kappa_{n},\{r_{n}^{X}\}_{X\in\mathcal{C}}):(B_{n},G_{n},I^{(n)})\to(A_{n},F_{n},J^{(n)})$
and
$(\theta_{n},\{s_{n}^{X}\}_{X\in\mathcal{C}}):(A_{n},F_{n},J^{(n)})\to(B_{n+1},G_{n+1},I^{(n+1)})$
fitting into the family of not necessarily commutative diagrams
[…]
We will call the collection of diagrams (5.1) an approximate cocycle intertwining, if the following hold: There exist an increasing sequence of finite sets $K_{n}\subset\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$, finite sets $\mathcal{F}_{n}^{X}\subset F_{n}(X)$ and $\mathcal{G}_{n}^{X}\subset G_{n}(X)$ for any $X\in K_{n}$, and numbers $\delta_{n}>0$ satisfying
  1. (i)
    $\|l_{n}^{X}(x)-s_{n}^{X}(r_{n}^{X}(x))\|\leq\delta_{n}$ for all $X\in K_{n}$ and all $x\in\mathcal{G}_{n}^{X}$;
  2. (ii)
    $\|h_{n}^{X}(x)-r_{n+1}^{X}(s_{n}^{X}(x))\|\leq\delta_{n}$ for all $X\in K_{n}$ and all $x\in\mathcal{F}_{n}^{X}$;
  3. (iii)
    $h_{n}^{X}(\mathcal{F}_{n}^{X})\subseteq\mathcal{F}_{n+1}^{X}$, $l_{n}^{X}(\mathcal{G}_{n}^{X})\subseteq\mathcal{G}_{n+1}^{X}$, $r_{n}^{X}(\mathcal{G}_{n}^{X})\subseteq\mathcal{F}_{n}^{X}$, and $s_{n}^{X}(\mathcal{F}_{n}^{X})\subseteq\mathcal{G}_{n+1}^{X}$;
  4. (iv)
    $\bigcup\limits_{m>n}(h_{n,m}^{X})^{-1}(\mathcal{F}_{m}^{X})\subset F_{n}(X)$ and $\bigcup\limits_{m>n}(l_{n,m}^{X})^{-1}(\mathcal{G}_{m}^{X})\subset G_{n}(X)$ are dense for all $X\in K_{n}$ and all $n$;
  5. (v)
    $\bigcup\limits_{n\in\mathbb{N}}K_{n}=\mathrm{Irr}(\mathcal{C})$;
  6. (vi)
    $\sum\limits_{n\in\mathbb{N}}\delta_{n}<\infty$.
Remark 5.2.
The conditions listed above are in the spirit of Elliott approximate intertwining arguments, as seen for example in [38, Definition $2.3.1$] or [40, Definition $3.3$]. Note that applying $X=1_{\mathcal{C}}$ in conditions (i)-(iv) above, together with the (vi) recover the usual assumptions in the two-sided Elliott intertwining argument. However, to boost the Elliott intertwining in the equivariant setting, one needs to preserve the equivariant structure. This is why conditions (i)-(iv) above are phrased for irreducible elements in the category. These conditions resemble the assumptions needed in [40, Definition 3.3] in the case of a twisted group action, although recall that the topology on cocycle morphisms is different.
Moreover, (i)-(ii) encode that the diagrams in 5.1 approximately commute in the topology described in Lemma 4.9. Furthermore, in (iv) we are making crucial use of our definition of the functors $F_{n}$ and $G_{n}$ to assume that $F_{n}(X)$ and $G_{n}(X)$ have a countable dense subset for any $X\in\mathcal{C}$ and any $n\in\mathbb{N}$.
Note that, in the light of Remark 3.12, it is enough to state the conditions above for $X\in\mathrm{Irr}(\mathcal{C})$. This will uniquely determine the family of linear maps $\{h^{X}\}_{X\in\mathcal{C}}$.
Theorem 5.3.
Let $(A,F,J)$ and $(B,G,I)$ be inductive limits in $\mathrm{C}^{*}_{\mathcal{C}}$ given by
$(A,F,J)=\lim\limits_{\longrightarrow}\Big{\{}(A_{n},F_{n},J^{(n)}),(\phi_{n},\{h_{n}^{X}\}_{X\in\mathcal{C}})\Big{\}}$
and
$(B,G,I)=\lim\limits_{\longrightarrow}\Big{\{}(B_{n},G_{n},I^{(n)}),(\psi_{n},\{l_{n}^{X}\}_{X\in\mathcal{C}})\Big{\}}.$
Let
$(\kappa_{n},\{r_{n}^{X}\}_{X\in\mathcal{C}}):(B_{n},G_{n},I^{(n)})\to(A_{n},F_{n},J^{(n)})$
and
$(\theta_{n},\{s_{n}^{X}\}_{X\in\mathcal{C}}):(A_{n},F_{n},J^{(n)})\to(B_{n+1},G_{n+1},I^{(n+1)})$
be sequences of cocycle morphisms making (5.1) an approximate cocycle intertwining.
Then there exist mutually inverse cocycle conjugacies $(\theta,\{s^{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ and $(\kappa,\{r^{X}\}_{X\in\mathcal{C}}):(B,G,I)\to(A,F,J)$ given by the formulae
(5.2) $\theta(\phi_{n,\infty}(a))=\lim_{k\to\infty}(\psi_{k+1,\infty}\circ\theta_{k}\circ\phi_{n,k})(a),\quad a\in A_{n}$
(5.3) $s^{X}(h_{n,\infty}^{X}(x))=\lim_{k\to\infty}(l_{k+1,\infty}^{X}\circ s_{k}^{X}\circ h_{n,k}^{X})(x),\quad X\in\mathcal{C},\quad x\in F_{n}(X)$
and
(5.4) $\kappa(\psi_{n,\infty}(b))=\lim_{k\to\infty}(\phi_{k,\infty}\circ\kappa_{k}\circ\psi_{n,k})(b),\quad b\in B_{n},$
(5.5) $r^{X}(l_{n,\infty}^{X}(x))=\lim_{k\to\infty}(h_{k,\infty}^{X}\circ r_{k}^{X}\circ l_{n,k}^{X})(x),\quad X\in\mathcal{C},\quad x\in G_{n}(X).$
The limits in the formulae (5.3) and (5.5) are taken in the topologies induced by the respective right inner products.
Remark 5.4.
To deduce the formulae above, one can work backwards. If the cocycle conjugacy $\theta$ was to exist, the fact that $A$ and $B$ are inductive limits would force the diagrams
[…]
to commute for all $k\in\mathbb{N}$. This immediately gives (5.2). Considering the associated families of linear maps for the cocycle morphisms in (5.2), we would need a similar diagram involving the linear maps to commute, which gives (5.3). The formulae in (5.4) and (5.5) can be obtained in exactly the same fashion.
Proof.
We will show that the limits in (5.2) and (5.3) exist and that the pair $(\theta,\{s^{X}\}_{X\in\mathcal{C}})$ is a cocycle morphism. Firstly, note that the limit in (5.2) exists and $\theta:A\to B$ is a $*{}^{*}$-homomorphism by the general Elliott intertwining, as can be found for example in [38, Proposition 2.3.2].
We first show that the limit in (5.3) exists for all $X\in\mathrm{Irr}(\mathcal{C}).$ We can employ a similar argument to the one in [38, Proposition 2.3.2]. Let $X\in K_{n}$ and given $x_{n}\in F_{n}(X)$, by condition (iv) in Definition 5.1, we may assume $x_{n}$ is contained in $\bigcup_{m>n}(h_{n,m}^{X})^{-1}(\mathcal{F}_{m}^{X})$. Moreover, we may assume that $h_{n,m}^{X}(x_{n})\in\mathcal{F}_{m}^{X}$ for all $m$ greater than or equal to some $m_{0}$.
Consider the possibly non-commutative diagram
[…]
Using triangle inequality, it follows that
$\|(s_{m+1}^{X}\circ h_{m}^{X})(x_{m})-(l_{m+1}^{X}\circ s_{m}^{X})(x_{m})\|$ $\leq\|s_{m+1}^{X}(h_{m}^{X}(x_{m})-(r_{m+1}^{X}\circ s_{m}^{X})(x_{m}))\|$
$+\|(s_{m+1}^{X}\circ r_{m+1}^{X}-l_{m+1}^{X})(s_{m}^{X}(x_{m}))\|.$
Also, by (iii) $s_{m}^{X}(\mathcal{F}_{m}^{X})\subset\mathcal{G}_{m+1}^{X}$, so we may combine (i), (ii), and that $s_{m+1}^{X}$ is contractive to get that
$\|(s_{m+1}^{X}\circ h_{m}^{X})(x_{m})-(l_{m+1}^{X}\circ s_{m}^{X})(x_{m})\|\leq\delta_{m}+\delta_{m+1},\quad\forall\quad x_{m}\in\mathcal{F}_{m}^{X}\subseteq F_{m}(X).$
Moreover, since $l_{m+2,\infty}^{X}$ is contractive, it follows that
$\|(l_{m+2,\infty}^{X}\circ s_{m+1}^{X}\circ h_{n,m+1}^{X})(x_{n})-(l_{m+1,\infty}^{X}\circ s_{m}^{X}\circ h_{n,m}^{X})(x_{n})\|=$
$\|l_{m+2,\infty}^{X}((s_{m+1}^{X}\circ h_{m}^{X})(h_{n,m}^{X}(x_{n}))-(l_{m+1}^{X}\circ s_{m}^{X})(h_{n,m}^{X}(x_{n})))\|$
$\leq\delta_{m}+\delta_{m+1},$
for all $m\geq m_{0}$ and all $x_{n}\in F_{n}(X)$. Using (vi), the sequence in (5.3) is Cauchy and therefore convergent. Moreover (v) yields that $s^{X}:F(X)\to G(X)$ is a well-defined linear map for all $X\in\mathrm{Irr}(\mathcal{C})$. Extending by linearity we may define $s^{X}$ for any $X\in\mathcal{C}$.
We claim that $(\theta,\{s^{X}\}_{X\in\mathrm{Irr}(\mathcal{C})})$ induces a cocycle morphism. For this, we check the conditions in Lemma 3.11. For any $n\in\mathbb{N}$, $X\in\mathrm{Irr}(\mathcal{C})$, $x_{n}\in F_{n}(X)$, and any $a_{n},a_{n}^{\prime}\in A_{n}$, as $(\phi_{n,\infty},h_{n,\infty})$ is a cocycle morphism, it follows that
$s^{X}(\phi_{n,\infty}(a_{n})\rhd h_{n,\infty}^{X}(x_{n})\lhd\phi_{n,\infty}(a_{n}^{\prime}))=s^{X}(h_{n,\infty}^{X}(a_{n}\rhd x_{n}\lhd a_{n}^{\prime})).$
Since the limit in (5.3) exists and $l_{k+1,\infty}^{X},s_{k}^{X},h_{n,k}^{X}$ are cocycle morphisms, a direct calculation shows that
$s^{X}(h_{n,\infty}^{X}(a_{n}\rhd x_{n}\lhd a_{n}^{\prime}))=\theta(\phi_{n,\infty}(a_{n}))\rhd s^{X}(h_{n,\infty}^{X}(x_{n}))\lhd\theta(\phi_{n,\infty}(a_{n}^{\prime})).$
Given $X,Y\in\mathrm{Irr}(\mathcal{C})$, consider a morphism $f\in\mathrm{Hom}(X,Y)$. Since $\mathrm{Hom}(X,Y)$ is 00 if $X\neq Y$ and $\mathrm{Hom}(X,X)\cong\mathbb{C}$, it follows immediately that $G(f)\circ s^{X}=s^{Y}\circ F(f)$.
Next, we want to show that $\theta(\langle x,y\rangle_{A})=\langle s^{X}(x),s^{X}(y)\rangle_{B}$ for any $x,y\in F(X)$. It suffices to check this condition on the dense subset $\bigcup_{n\geq 1}h_{n,\infty}^{X}(F_{n}(X))$. If $x_{n},y_{n}\in F_{n}(X)$ and $k>n$, as $l_{k+1,\infty}^{X}\circ s_{k}^{X}\circ h_{n,k}^{X}$ satisfies condition (iii) of Lemma 3.11 one has
$\langle(l_{k+1,\infty}^{X}\circ s_{k}^{X}\circ h_{n,k}^{X})(x_{n}),(l_{k+1,\infty}^{X}\circ s_{k}^{X}\circ h_{n,k}^{X})(y_{n})\rangle=(\psi_{k+1,\infty}\circ\theta_{k}\circ\phi_{n,k})(\langle x_{n},y_{n}\rangle).$
Taking the limit as $k$ goes to infinity and using the formulae in (5.2) and (5.3),
$\theta(\phi_{n,\infty}(\langle x_{n},y_{n}\rangle))=\langle s^{X}(h_{n,\infty}^{X}(x_{n})),s^{X}(h_{n,\infty}^{X}(y_{n}))\rangle.$
Moreover, $\phi_{n,\infty}(\langle x_{n},y_{n}\rangle)=\langle h_{n,\infty}^{X}(x_{n}),h_{n,\infty}^{X}(y_{n})\rangle$, so $\theta(\langle x,y\rangle)=\langle s^{X}(x),s^{X}(y)\rangle$ for any $x,y$ in the dense subset $\bigcup_{n\geq 1}h_{n,\infty}^{X}(F_{n}(X))$.
Then, we check (iv) of Lemma 3.11. Let $X,Y\in\mathcal{C}$. Since $(\psi_{k+1,\infty}\circ\theta_{k}\circ\phi_{n,k},l_{k+1,\infty}\circ s_{k}\circ h_{n,k})$ is a cocycle morphism, the diagram
[…]
commutes. Then, by taking the limit as $k$ goes to infinity, it follows that $I_{X,Y}\circ((s^{Y}\circ h_{n,\infty}^{Y})\boxtimes(s^{X}\circ h_{n,\infty}^{X}))=s^{X\otimes Y}\circ h_{n,\infty}^{X\otimes Y}\circ J_{X,Y}^{(n)}.$ Note that by the construction of the map $J_{X,Y}$ in Lemma 4.3,
$J_{X,Y}\circ(h_{n,\infty}^{Y}\boxtimes h_{n,\infty}^{X})=h_{n,\infty}^{X\otimes Y}\circ J_{X,Y}^{(n)}.$
Thus,
$I_{X,Y}\circ(s^{Y}\boxtimes s^{X})\circ(h_{n,\infty}^{Y}\boxtimes h_{n,\infty}^{X})=s^{X\otimes Y}\circ J_{X,Y}\circ(h_{n,\infty}^{Y}\boxtimes h_{n,\infty}^{X}).$
By density it now follows that $I_{X,Y}\circ(s^{Y}\boxtimes s^{X})=s^{X\otimes Y}\circ J_{X,Y}$. Finally, it follows from (5.2) that for any $a\in A$, $s^{1_{\mathcal{C}}}(a)=\theta(a)$. Hence, as $\mathcal{C}$ is semisimple, $(\theta,\{s^{X}\}_{X\in\mathrm{Irr}(\mathcal{C})})$ induces a cocycle morphism (see Remark 3.12). This follows as the map $s^{X}$ will be given by a direct sum of linear maps corresponding to irreducible objects. Since the same holds for $h_{n,\infty}^{X},l_{k+1,\infty}^{X},s_{k}^{X},$ and $h_{n,k}^{X}$, and the limit preserves this decomposition, the formula in (5.3) holds for any $X\in\mathcal{C}$.
It follows in the same way that $(\kappa,\{r^{X}\}_{X\in\mathcal{C}})$ given by the formulae in (5.4) and (5.5) yields a well-defined cocycle morphism. Moreover, the fact that $\theta$ and $\kappa$ are mutually inverse isomorphisms follows from [38, Proposition 2.3.2].
Finally, it remains to check that for any $X\in\mathcal{C}$, $r^{X}\circ s^{X}=\mathrm{id}_{F(X)}$ and $s^{X}\circ r^{X}=\mathrm{id}_{G(X)}$. It suffices to show that for any $n\geq 1$, $r^{X}\circ s^{X}\circ h_{n,\infty}^{X}=h_{n,\infty}^{X}$. For any $k>n$, by (5.5), we have that
(5.7) $r^{X}\circ l_{k+1,\infty}^{X}\circ s_{k}^{X}\circ h_{n,k}^{X}=\lim\limits_{m\to\infty}(h_{m,\infty}^{X}\circ r_{m}^{X}\circ l_{k+1,m}^{X}\circ s_{k}^{X}\circ h_{n,k}^{X}).$
Using conditions (i) and (ii) of Definition 5.1, it follows that the right hand side evaluated at $x\in(h_{n,k}^{X})^{-1}(\mathcal{F}_{k}^{X})$ converges to $h_{n,\infty}(x)$ uniformly as $k$ tends to infinity. Now, taking the limit as $k$ goes to infinity in (5.7) and using condition (iv) of Definition 5.1, we get that $r^{X}\circ s^{X}\circ h_{n,\infty}^{X}=h_{n,\infty}^{X}$. Similarly, $s^{X}\circ r^{X}=\mathrm{id}_{G(X)}$, so $(\theta,\{s^{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ and $(\kappa,\{r^{X}\}_{X\in\mathcal{C}}):(B,G,I)\to(A,F,J)$ are mutually inverse cocycle conjugacies. ∎
We now use Theorem 5.3 to show that if we assume that the diagrams in (5.1) commute up to approximate unitary equivalence, then there exist mutually inverse cocycle conjugacies as in Theorem 5.3. The proof follows in a similar fashion to [40, Theorem 3.6] and [38, Corollary 2.3.3].
Theorem 5.5.
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F_{n},J^{(n)}):\mathcal{C}\curvearrowright A_{n}$ and $(G_{n},I^{(n)}):\mathcal{C}\curvearrowright B_{n}$ be sequences of actions on separable $\mathrm{C}^{*}$-algebras. Let
$(\phi_{n},\{h_{n}^{X}\}_{X\in\mathcal{C}}):(A_{n},F_{n},J^{(n)})\to(A_{n+1},F_{n+1},J^{(n+1)})$
and
$(\psi_{n},\{l_{n}^{X}\}_{X\in\mathcal{C}}):(B_{n},G_{n},I^{(n)})\to(B_{n+1},G_{n+1},I^{(n+1)})$
be sequences of cocycle morphisms, in the sense of Lemma 3.11, which we view as two inductive systems in the category $\mathrm{C}^{*}_{\mathcal{C}}$.
Consider two sequences of extendible cocycle morphisms
$(\kappa_{n},\{r_{n}^{X}\}_{X\in\mathcal{C}}):(B_{n},G_{n},I^{(n)})\to(A_{n},F_{n},J^{(n)})$
and
$(\theta_{n},\{s_{n}^{X}\}_{X\in\mathcal{C}}):(A_{n},F_{n},J^{(n)})\to(B_{n+1},G_{n+1},I^{(n+1)})$
fitting into the not necessarily commutative collection of diagrams
[…]
Suppose that
$(\psi_{n},l_{n})\approx_{u}(\theta_{n},s_{n})\circ(\kappa_{n},r_{n})\quad and\quad(\phi_{n},h_{n})\approx_{u}(\kappa_{n+1},r_{n+1})\circ(\theta_{n},s_{n})$
for all $n\in\mathbb{N}$. Then there exist mutually inverse cocycle conjugacies $(\theta,\{s^{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ and $(\kappa,\{r^{X}\}_{X\in\mathcal{C}}):(B,G,I)\to(A,F,J)$. Moreover, if $\phi_{n}$ and $\psi_{n}$ are extendible for any $n\in\mathbb{N}$, then
$(\theta,s)\circ(\phi_{n,\infty},h_{n,\infty})\approx_{u}(\psi_{n+1,\infty},l_{n+1,\infty})\circ(\theta_{n},s_{n})$
and
$(\kappa,r)\circ(\psi_{n,\infty},l_{n,\infty})\approx_{u}(\phi_{n,\infty},h_{n,\infty})\circ(\kappa_{n},r_{n}).$
Proof.
This will follow as an application of Theorem 5.3. For this, it suffices to show that we can obtain a collection of diagrams as in Definition 5.1. The strategy is to replace the families of cocycle morphisms $(\kappa_{n},r_{n})$ and $(\theta_{n},s_{n})$ by unitary perturbations $(\eta_{n},R_{n})$ and $(\zeta_{n},S_{n})$ respectively such that the diagrams in (5.8) become an approximate cocycle intertwining.
Let $\delta_{n}=2^{-n}$ for any $n\in\mathbb{N}$ and $K_{n}$ an increasing sequence of finite sets containing $1_{\mathcal{C}}$ such that $\bigcup_{n\in\mathbb{N}}K_{n}=\mathrm{Irr}(\mathcal{C})$. For any $n\in\mathbb{N}$, and any $X\in K_{n}$, we can choose $t_{n}^{X}\in\mathbb{N}$ and finite sets $\{f_{m,n}^{X}\}_{1\leq m\leq t_{n}^{X}}\subset F_{n}(X)$ and $\{g_{m,n}^{X}\}_{1\leq m\leq t_{n}^{X}}\subset G_{n}(X)$ such that the inclusions
(5.9) $\bigcup\limits_{k>n}(h_{n,k}^{X})^{-1}(\{f_{m,k}^{X}\}_{1\leq m\leq t_{k}^{X}})\subset F_{n}(X)\quad\mathrm{and}\quad\bigcup\limits_{k>n}(l_{n,k}^{X})^{-1}(\{g_{m,k}^{X}\}_{1\leq m\leq t_{k}^{X}})\subset G_{n}(X)$
are dense for all $n\in\mathbb{N}$. To simplify notation, we will write $\{f_{m,n}^{X}\}$ to denote the set $\{f_{m,n}^{X}\}_{1\leq m\leq t_{n}^{X}}$.
Set $(\eta_{1},R_{1})=(\kappa_{1},r_{1})$, $\mathcal{G}_{1}^{X}=\{g_{m,1}^{X}\}\subset G_{1}(X)$, and $\mathcal{F}_{1}^{X}=\{f_{m,1}^{X}\}\cup r_{1}^{X}(\mathcal{G}_{1}^{X})$ for any $X\in K_{1}$. Since
$(\psi_{1},l_{1})\approx_{u}(\theta_{1},s_{1})\circ(\kappa_{1},r_{1})=(\theta_{1},s_{1})\circ(\eta_{1},R_{1}),$
we can find a unitary $u_{1}\in\mathcal{U}(\mathcal{M}(B_{2}))$ such that if we set $(\zeta_{1},S_{1})=\mathrm{Ad}(u_{1})\circ(\theta_{1},s_{1})$, one has that for any $X\in K_{1}$
$\max\limits_{x\in\mathcal{G}_{1}^{X}}\|l_{1}^{X}(x)-S_{1}^{X}(R_{1}^{X}(x))\|\leq\delta_{1}.$
At the next stage, let $\mathcal{G}_{2}^{X}=\{g_{m,2}^{X}\}\cup S_{1}^{X}(\mathcal{F}_{1}^{X})\cup l_{1}^{X}(\mathcal{G}_{1}^{X})$ for any $X\in K_{1}$ and $\mathcal{G}_{2}^{X}=\{g_{m,2}^{X}\}$ for any $X\in K_{2}\setminus K_{1}$. Using the assumption $(\phi_{1},h_{1})\approx_{u}(\kappa_{2},r_{2})\circ(\theta_{1},s_{1})$, that $(\theta_{1},s_{1})$ is unitarily equivalent to $(\zeta_{1},S_{1})$, and that $\kappa_{2}$ is extendible, it follows that $(\phi_{1},h_{1})\approx_{u}(\kappa_{2},r_{2})\circ(\zeta_{1},S_{1})$ (see Remark 4.19). Therefore, there exists a unitary $v_{2}\in\mathcal{U}(\mathcal{M}(A_{2}))$ such that if we set $(\eta_{2},R_{2})=\mathrm{Ad}(v_{2})\circ(\kappa_{2},r_{2})$, we have that for any $X\in K_{1}$
$\max\limits_{x\in\mathcal{F}_{1}^{X}}\|h_{1}^{X}(x)-R_{2}^{X}(S_{1}^{X}(x))\|\leq\delta_{1}.$
Then, we let $\mathcal{F}_{2}^{X}=\{f_{m,2}^{X}\}\cup R_{2}^{X}(\mathcal{G}_{2}^{X})\cup h_{1}^{X}(\mathcal{F}_{1}^{X})$ for any $X\in K_{1}$ and $\mathcal{F}_{2}^{X}=\{f_{m,2}^{X}\}\cup R_{2}^{X}(\mathcal{G}_{2}^{X})$ for any $X\in K_{2}\setminus K_{1}$.
By continuing inductively we construct finite sets
(5.10) $\mathcal{G}_{n}^{X}=\{g_{m,n}^{X}\}\cup S_{n-1}^{X}(\mathcal{F}_{n-1}^{X})\cup l_{n-1}^{X}(\mathcal{G}_{n-1}^{X})\subset G_{n}(X),$
(5.11) $\mathcal{F}_{n}^{X}=\{f_{m,n}^{X}\}\cup R_{n}^{X}(\mathcal{G}_{n}^{X})\cup h_{n-1}^{X}(\mathcal{F}_{n-1}^{X})\subset F_{n}(X)$
for any $X\in K_{n-1}$ and
(5.12) $\mathcal{G}_{n}^{X}=\{g_{m,n}^{X}\}\subset G_{n}(X),$
(5.13) $\mathcal{F}_{n}^{X}=\{f_{m,n}^{X}\}\cup R_{n}^{X}(\mathcal{G}_{n}^{X})\subset F_{n}(X)$
for any $X\in K_{n}\setminus K_{n-1}$ and unitaries $v_{n}\in\mathcal{U}(\mathcal{M}(A_{n}))$ and $u_{n}\in\mathcal{U}(\mathcal{M}(B_{n+1}))$ such that setting $(\eta_{n},R_{n})=\mathrm{Ad}(v_{n})\circ(\kappa_{n},r_{n})$ and $(\zeta_{n},S_{n})=\mathrm{Ad}(u_{n})\circ(\theta_{n},s_{n})$ one has for any $X\in K_{n}$
(5.14) $\max\limits_{x\in\mathcal{G}_{n}^{X}}\|l_{n}^{X}(x)-S_{n}^{X}(R_{n}^{X}(x))\|\leq\delta_{n}$
and
(5.15) $\max\limits_{x\in\mathcal{F}_{n}^{X}}\|h_{n}^{X}(x)-R_{n+1}^{X}(S_{n}^{X}(x))\|\leq\delta_{n}.$
We claim that the diagram
[…]
is an approximate cocycle intertwining in the sense of Definition 5.1. Conditions (i) and (ii) is ensured by (5.14) and (5.15) respectively. Then, condition (iii) follows by (5.10), (5.11),(5.12), and (5.13). Moreover, condition (iv) follows by (5.10), (5.11),(5.12), (5.13), and the choice of finite sets $\{f_{m,n}^{X}\},\{g_{m,n}^{X}\}$ from (5.9). Then, (v) is ensured by the choice of finite sets $K_{n}$, while (vi) is satisfied because the sum $\sum_{n}2^{-n}$ converges.
Hence, by Theorem 5.3 applied to the family of diagrams in (5.16), there exists mutually inverse cocycle conjugacies $(\theta,\{s^{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ and $(\kappa,\{r^{X}\}_{X\in\mathcal{C}}):(B,G,I)\to(A,F,J)$. Moreover, Theorem 5.3 also gives that
(5.17) $\theta(\phi_{n,\infty}(a))=\lim_{k\to\infty}(\psi_{k+1,\infty}\circ\zeta_{k}\circ\phi_{n,k})(a),\quad a\in A_{n}$
and
(5.18) $s^{X}(h_{n,\infty}^{X}(x))=\lim_{k\to\infty}(l_{k+1,\infty}^{X}\circ S_{k}^{X}\circ h_{n,k}^{X})(x),\quad X\in\mathcal{C},\quad x\in F_{n}(X).$
Now, assume that $\phi_{n}$ and $\psi_{n}$ are extendible. Recall that
$(\psi_{m},l_{m})\approx_{u}(\theta_{m},s_{m})\circ(\kappa_{m},r_{m})\ \text{and}\ (\phi_{m},h_{m})\approx_{u}(\kappa_{m+1},r_{m+1})\circ(\theta_{m},s_{m})$
and $(\zeta_{m},S_{m})=\mathrm{Ad}(u_{m})\circ(\theta_{m},s_{m})$ for any $m\in\mathbb{N}$. Then, for all $n\geq 1$ and all $k>n$ we have that
$\begin{array}[]{cl}(\psi_{n+1,k+1},l_{n+1,k+1})\circ(\theta_{n},s_{n})\hfil\\ \approx_{u}&(\theta_{k},s_{k})\circ(\kappa_{k},r_{k})\circ(\theta_{k-1},s_{k-1})\circ\ldots\circ(\kappa_{n+1},r_{n+1})\circ(\theta_{n},s_{n})\\ \approx_{u}&(\theta_{k},s_{k})\circ(\phi_{n,k},h_{n,k})\\ \approx_{u}&(\zeta_{k},S_{k})\circ(\phi_{n,k},h_{n,k}).\end{array}$
Composing with $(\psi_{k+1,\infty},l_{k+1,\infty})$ yields that
$(\psi_{n+1,\infty},l_{n+1,\infty})\circ(\theta_{n},s_{n})\approx_{u}(\psi_{k+1,\infty},l_{k+1,\infty})\circ(\zeta_{k},S_{k})\circ(\phi_{n,k},h_{n,k}).$
Then, by (5.17) and (5.18), one gets that
$(\theta,s)\circ(\phi_{n,\infty},h_{n,\infty})\approx_{u}(\psi_{n+1,\infty},l_{n+1,\infty})\circ(\theta_{n},s_{n}).$
The fact that
$(\kappa,r)\circ(\psi_{n,\infty},l_{n,\infty})\approx_{u}(\phi_{n,\infty},h_{n,\infty})\circ(\kappa_{n},r_{n})$
follows analogously. ∎
As a corollary of Theorem 5.5, we obtain the following result.
Corollary 5.6.
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras. Let
$(\phi,h):(A,F,J)\to(B,G,I)\quad and\quad(\psi,l):(B,G,I)\to(A,F,J)$
be two extendible cocycle morphisms such that
$\mathrm{id}_{A}\approx_{u}(\psi,l)\circ(\phi,h)\quad and\quad\mathrm{id}_{B}\approx_{u}(\phi,h)\circ(\psi,l).$
Then there exist mutually inverse cocycle conjugacies
$(\Phi,H):(A,F,J)\to(B,G,I)\quad and\quad(\Psi,L):(B,G,I)\to(A,F,J)$
such that
$(\Phi,H)\approx_{u}(\phi,h)\quad and\quad(\Psi,L)\approx_{u}(\psi,l).$
Proof.
This is a direct application of Theorem 5.5 with $A_{n}=A$, $B_{n}=B$, $\phi_{n}=\mathrm{id}_{A}$, $\psi_{n}=\mathrm{id}_{B}$, $k_{n}=\psi$, and $\theta_{n}=\phi$ for all $n\in\mathbb{N}$. ∎
Remark 5.7.
Note that in Corollary 5.6, if the approximate unitary equivalences are realised by unitaries in the minimal unitisations, then we may drop the assumption of extendibility on the connecting morphisms. This follows as extendibility in Theorem 5.5 is only required to evaluate a morphism on a given unitary in the multiplier algebra. However, there is no issue in doing so when the unitary is in the minimal unitisation.
5.1. Intertwining through reparametrisation
Intertwining through reparametrisation is a type of intertwining argument commonly employed in the classification programme of $\mathrm{C}^{*}$-algebras and $\mathrm{C}^{*}$-dynamics. In broad terms, if we want to prove that a morphism $\theta:A\to B_{\infty}$ is unitarily equivalent to a morphism $\psi:A\to B_{\infty}$ which factors through $B$, then it suffices to check that $\theta$ is invariant under reparametrisations. This type of result appears for example in [16, 40] and it is used in successful classification results in [16, 17, 5].
If $A$ is a separable $C^{*}$-algebra and $\eta:\mathbb{N}\to\mathbb{N}$ is any map with $\lim\limits_{n\to\infty}\eta(n)=\infty$, then it induces an endomorphism $\eta^{*}$ on $A_{\infty}$ via $\eta^{*}((a_{n})_{n})=(a_{\eta(n)})_{n}$. Moreover, if a $*{}^{*}$-tensor category $\mathcal{C}$ acts on a unital $*{}^{*}$-algebra $A$ by the triple $(A,F,J)$ with $F(X)$ algebraically finitely generated for all $X\in\mathcal{C}$, it follows from Lemma 2.6 that there exists an induced action on $A_{\infty}$ by the triple $(A_{\infty},F_{\infty},J^{\infty})$ (see also Remark 2.7).
Then, a straightforward checking of the conditions in Lemma 3.11 shows that $\eta$ induces a cocycle morphism $(\eta^{*},r):(A_{\infty},F_{\infty},J^{\infty})\to(A_{\infty},F_{\infty},J^{\infty})$, where $r^{X}:F_{\infty}(X)\to F_{\infty}(X)$ is given by $r^{X}((\xi_{n})_{n})=((\xi_{\eta(n)})_{n})$ for any $X\in\mathcal{C}$ and any $(\xi_{n})_{n}\in F_{\infty}(X)$. Using this construction, we will prove an intertwining argument concerning maps into sequence algebras. First we need a preparatory lemma.
Lemma 5.8.
Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions of a $*{}^{*}$-tensor category $\mathcal{C}$ on separable $\mathrm{C}^{*}$-algebras with $B$ unital and $G(X)$ be algebraically finitely generated as a right $B$-module for all $X\in\mathcal{C}$. Suppose that $(\phi,h):(A,F,J)\to(B_{\infty},G_{\infty},I^{\infty})$ is a cocycle morphism such that for any map $\eta:\mathbb{N}\to\mathbb{N}$ with $\lim\limits_{n\to\infty}\eta(n)=\infty$, the cocycle morphisms $(\phi,h)$ and $(\eta^{*},r)\circ(\phi,h)$ are approximately unitarily equivalent. For each $X\in\mathcal{C}$, let $(h_{n}^{X})_{n}$ be any lift of $h^{X}$. Then, for every finite set $K\subseteq\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$, finite sets $\mathcal{F}^{X}\subseteq F(X)$ for $X\in K$, $\epsilon>0$, and $m\in\mathbb{N}$, there is an integer $k\geq m$ such that for every integer $n\geq k$ there is a unitary $u\in B$ for which
$\|u\rhd h_{n}^{X}(x)\lhd u^{*}-h_{k}^{X}(x)\|<\epsilon,\quad X\in K,\quad x\in\mathcal{F}^{X}.$
Proof.
We prove this by contradiction. Suppose that there exists a finite set $K\subseteq\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$, finite sets $\mathcal{F}^{X}\subseteq F(X)$ for $X\in K$, $\epsilon>0$, and $m\in\mathbb{N}$ such that for every $k\geq m$, there exists $n_{k}\geq k$ for which
(5.19) $\max\limits_{x\in\mathcal{F}^{X}}\|u_{k}\rhd h_{n_{k}}^{X}(x)\lhd u_{k}-h_{k}^{X}(x)\|\geq\epsilon,$
for any $X\in K$, $x\in\mathcal{F}^{X}$, and every unitary $u_{k}\in B$. Let $\eta:\mathbb{N}\to\mathbb{N}$ be the map $\eta(k)=n_{k}$ whenever $k\geq m$ and $\eta(k)=1$ for $k<m$. As $\eta_{k}\geq k$ for $k\geq m$, it follows that $\lim\limits_{k\to\infty}\eta(k)=\infty$. Moreover, $(\phi,h)$ and $(\eta^{*},r)\circ(\phi,h)$ are approximately unitarily equivalent, so there exists a unitary $u\in B_{\infty}$ for which
$\|u\rhd r^{X}(h^{X}(x))\lhd u^{*}-h^{X}(x)\|<\epsilon,\quad X\in K,\quad x\in\mathcal{F}^{X}.$
If we let $(u_{k})_{k\geq 1}$ to be a sequence of unitaries lifting $u$, then
$\limsup\limits_{k\to\infty}\|u_{k}\rhd h_{n_{k}}^{X}(x)\lhd u_{k}^{*}-h_{k}^{X}(x)\|<\epsilon$
for all $X\in K$ and $x\in\mathcal{F}^{X}$. But this contradicts (5.19), so we reach the conclusion. ∎
Theorem 5.9.
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras with $B$ unital and $G(X)$ be algebraically finitely generated as a right $B$-module for all $X\in\mathcal{C}$. Let $(\phi,h):(A,F,J)\to(B_{\infty},G_{\infty},I^{\infty})$ be a cocycle morphism. Then the following are equivalent:
  1. (i)
    $(\phi,h)$ is unitarily equivalent to a cocycle morphism $(\psi,l):(A,F,J)\to(B,G,I)$; [^17] We view $\psi$ as a cocycle morphism into $B_{\infty}$ after composing with the canonical inclusion $\iota:B\rightarrow B_{\infty}$, which itself is canonically a cocycle morphism.
  2. (ii)
    for any map $\eta:\mathbb{N}\to\mathbb{N}$ with $\lim\limits_{n\to\infty}\eta(n)=\infty$, the cocycle morphisms $(\phi,h)$ and $(\eta^{*},r)\circ(\phi,h)$ are approximately unitarily equivalent.
Proof.
Let us first show that (i) implies (ii). Suppose that $u\in B_{\infty}$ is a unitary such that $(\mathrm{Ad}(u),h_{u})\circ(\psi,l)=(\phi,h)$. [^18] Recall the definition of the family of linear maps $h_{u}$ from Lemma 4.12. Let $\eta:\mathbb{N}\to\mathbb{N}$ be any map with $\lim\limits_{n\to\infty}\eta(n)=\infty$. Since $r^{X}\circ l^{X}=l^{X}$ for any $X\in\mathcal{C}$, it follows that
$r^{X}\circ h^{X}=r^{X}\circ h_{u}^{X}\circ l^{X}=h_{\eta^{*}(u)}^{X}\circ r^{X}\circ l^{X}=h_{\eta^{*}(u)u^{*}}^{X}\circ h^{X}.$
Since this holds for any $X\in\mathcal{C}$, we get that
$(\mathrm{Ad}(\eta^{*}(u)u^{*}),h_{\eta^{*}(u)u^{*}})\circ(\phi,h)=(\eta^{*},r)\circ(\phi,h).$
Hence, $(\phi,h)$ and $(\eta^{*},r)\circ(\phi,h)$ are unitarily equivalent.
Suppose now that $(\phi,h)$ and $(\eta^{*},r)\circ(\phi,h)$ are approximately unitarily equivalent for any $\eta:\mathbb{N}\rightarrow\mathbb{N}$ such that $\eta(n)$ converges to infinity. Let $K_{n}\subseteq\mathrm{Irr}(\mathcal{C})$ be an increasing sequence of finite sets containing $1_{\mathcal{C}}$ such that $\bigcup_{n=1}^{\infty}K_{n}=\mathrm{Irr}(\mathcal{C})$, and $\mathcal{F}_{n}^{X}\subseteq F(X)$ be an increasing sequence of finite sets such that $\bigcup_{n=1}^{\infty}\mathcal{F}_{n}^{X}$ is dense in $F(X)$ for any $X\in\mathrm{Irr}(\mathcal{C})$. Recursively applying Lemma 5.8 to $K=K_{n}$, $\mathcal{F}^{X}=\mathcal{F}_{n}^{X}$, $\epsilon=\frac{1}{2^{n}}$, one may pick $k_{0}=1<k_{1}<k_{2}<\ldots$ and unitaries $u_{n}\in B$ for $n\in\mathbb{N}$ such that
(5.20) $\|u_{n}\rhd h_{k_{n}}^{X}(x)\lhd u_{n}^{*}-h_{k_{n-1}}^{X}(x)\|<\frac{1}{2^{n}},\quad X\in K_{n},\quad x\in\mathcal{F}_{n}^{X}.$
Let $v_{n}=u_{1}u_{2}\ldots u_{n}$, we claim that $l^{X}(x):=\lim\limits_{n\to\infty}v_{n}\rhd h_{k_{n}}^{X}(x)\lhd v_{n}^{*}$ is well-defined for any $X\in\mathrm{Irr}(\mathcal{C})$ and any $x\in F(X)$. First note that $h^{X}$ is contractive by (iii) of Lemma 3.11, so $\limsup_{n\to\infty}\|h^{X}_{n}(x)\|\leq\|x\|$ for any $x\in F(X)$. In particular, the sequence $h^{X}_{k_{n}}$ is norm bounded.
We will now show that the sequence $(v_{n}\rhd h_{k_{n}}^{X}(x)\lhd v_{n}^{*})_{n\geq 1}$ is Cauchy for any $X\in\mathrm{Irr}(\mathcal{C})$ and any $x\in F(X)$. Given $\epsilon>0$, $n\in\mathbb{N}$, $X\in K_{n}$, and $x\in F(X)$, there exists $m\in\mathbb{N}$ and $y\in\mathcal{F}_{m}^{X}$ such that $\|x-y\|<\epsilon/3$. Then, for $j>l\geq m$, we may use (5.20) repeatedly, together with triangle inequality, to achieve that
$\|v_{j}\rhd h_{k_{j}}^{X}(y)\lhd v_{j}^{*}-v_{l}\rhd h_{k_{l}}^{X}(y)\lhd v_{l}^{*}\|$ $\leq\sum\limits_{i=l+1}^{j}2^{-i}<\epsilon/3$
for any $l$ greater than some large enough $N$. Moreover, since $h_{k_{l}}^{X}$ is a contractive linear map for any $l\in\mathbb{N}$ and any $X\in\mathrm{Irr}(\mathcal{C})$, it follows that $\|h_{k_{l}}^{X}(x)-h_{k_{l}}^{X}(y)\|<\epsilon/3$ Therefore,
$\|v_{j}\rhd h_{k_{j}}^{X}(x)\lhd v_{j}^{*}-v_{l}\rhd h_{k_{l}}^{X}(x)\lhd v_{l}^{*}\|<\epsilon,$
for all $l>N$, so the sequence is Cauchy. Thus, $l^{X}:F(X)\to G(X)$ is a well-defined map for every $X\in\mathrm{Irr}(\mathcal{C})$ and it is linear.
We let $\psi:A\to B$ be the linear map given by $\psi=l^{1_{\mathcal{C}}}$. If $v\in B_{\infty}$ is the unitary given by $(v_{n})_{n}$ and $\eta:\mathbb{N}\to\mathbb{N}$ is the map $\eta(n)=k_{n}$, we have that
$(\psi,l^{X})=(\mathrm{Ad}(v),h_{v}^{X})\circ(\eta^{*},r^{X})\circ(\phi,h^{X})$
for any $X\in\mathrm{Irr}(\mathcal{C})$. Hence, $\psi$ is a $*{}^{*}$-homomorphism and the pair $(\psi,l)$ induces a cocycle morphism. Extending by linearity, we can define $l^{X}$ in the same way for any $X\in\mathcal{C}$ and it is straightforward to see that these maps are linear. Moreover, $(\psi,l)$ is approximately unitarily equivalent to $(\phi,h)$, which finishes the proof. ∎
Remark 5.10.
As noted in Remark 2.7 the assumption that $G(X)$ are algebraically finitely generated in Theorem 5.9 is immediate whenever the acting category $\mathcal{C}$ is a unitary tensor category.
6. One sided intertwining arguments
We start this section by showing a tensor category equivariant adaptation of the classical one-sided intertwining argument (see [38, Proposition 2.3.5]). A group equivariant version of this result can be found in [40, Proposition 4.3]. We end this section by proving an asymptotic Elliott two-sided intertwining (see Theorem 6.4).
Theorem 6.1.
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $*{}^{*}$-algebras and $(\varphi,h):(A,F,J)\rightarrow(B,G,I)$ an injective cocycle morphism. Then $(\varphi,h)$ is asymptotically unitarily equivalent to a cocycle conjugacy if and only if:
For all $\varepsilon>0$ and finite sets $K\subset\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$, $\mathcal{F}^{X}\subset F(X)$ and $\mathcal{G}^{X}\subset G(X)$ there exists a strictly continuous path $z:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ with $z_{0}=1$ such that
  1. A.1
    $\sup\limits_{0\leq t\leq 1}\|z_{t}\rhd h^{X}(x)\lhd z_{t}^{*}-h^{X}(x)\|\leq\varepsilon$ for all $X\in K,x\in\mathcal{F}^{X}$,
  2. A.2
    $\mathrm{dist}(z_{1}^{*}\rhd y\lhd z_{1},h^{X}(F(X)))\leq\varepsilon$ for all $X\in K$ and $y\in\mathcal{G}^{X}$.
Proof.
The proof of this Theorem will follow the proof of [40, Proposition 4.3] closely. We start by showing the “only if” statement. Let $(\Phi,H)$ be a cocycle conjugacy such that $(\Phi,H)\cong_{u}(\varphi,h)$. By Lemma 4.18 there exists a strictly continuous map $w:[0,\infty)\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that
(6.1) $H^{X}(x)=\lim\limits_{t\rightarrow\infty}w_{t}\rhd h^{X}(x)\lhd w_{t}^{*}$
for all $X\in\mathrm{Irr}(\mathcal{C})$ and $x\in F(X)$. Therefore, for any finite sets $K\subset\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$, $\mathcal{F}^{X}\subset F(X)$, and $\mathcal{G}^{X}\subset G(X)$ we may choose $n_{1}\geq 1$ sufficiently large such that
(6.2) $\sup\limits_{t\geq n_{1}}\|H^{X}(x)-w_{t}\rhd h^{X}(x)\lhd w_{t}^{*}\|\leq\varepsilon/2,\ X\in K,\ x\in\mathcal{F}^{X}.$
Similarly, one may pick $n_{2}>n_{1}$ such that
(6.3) $\sup\limits_{t\geq n_{2}}\|H^{X}(x)-w_{t}\rhd h^{X}(x)\lhd w_{t}^{*}\|\leq\varepsilon,\ X\in K,\ x\in(H^{X})^{-1}(w_{n_{1}}\rhd\mathcal{G}^{X}\lhd w_{n_{1}}^{*}).$
We claim that the unitary path $z_{t}=w_{n_{1}}^{*}w_{(1-t)n_{1}+tn_{2}}$ satisfies A.1 and A.2. Indeed, it is a strictly continuous path $z:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ with $z_{0}=1$. Moreover, for $X\in K$, $t\in[0,1]$ and $x\in\mathcal{F}^{X}$, using (6.2) we have that
$\|h^{X}(x)-z_{t}$ $*\displaystyle\rhd h^{X}(x)\lhd z_{t}^{*}\|$
$=\|h^{X}(x)-w_{n_{1}}^{*}w_{(1-t)n_{1}+tn_{2}}\rhd h^{X}(x)\lhd w_{(1-t)n_{1}+tn_{2}}^{*}w_{n_{1}}\|$
$=\|w_{n_{1}}\rhd h^{X}(x)\lhd w_{n_{1}}^{*}-w_{(1-t)n_{1}+tn_{2}}\rhd h^{X}(x)\lhd w_{(1-t)n_{1}+tn_{2}}^{*}\|$
$\leq\|w_{n_{1}}\rhd h^{X}(x)\lhd w_{n_{1}}^{*}-H^{X}(x)\|$
$\quad+\|H^{X}(x)-w_{(1-t)n_{1}+tn_{2}}\rhd h^{X}(x)\lhd w_{(1-t)n_{1}+tn_{2}}^{*}\|$
$\leq\varepsilon.$
This shows condition A.1. We now turn to condition A.2. Let $X\in K$, $y\in\mathcal{G}^{X}$, and $x=(H^{X})^{-1}(w_{n_{1}}\rhd y\lhd w_{n_{1}}^{*})$. Then, we get that
$\|z_{1}^{*}\rhd y\lhd z_{1}-h^{X}(x)\|$
$=\|w_{n_{2}}^{*}\rhd(w_{n_{1}}\rhd y\lhd w_{n_{1}}^{*})\lhd w_{n_{2}}-h^{X}(x)\|$
$=\|H^{X}(H^{X})^{-1}(w_{n_{1}}\rhd y\lhd w_{n_{1}}^{*})-w_{n_{2}}\rhd h^{X}(x)\lhd w_{n_{2}}^{*}\|$
$=\|H^{X}(x)-w_{n_{2}}\rhd h^{X}(x)\lhd w_{n_{2}}^{*}\|$
(6.3)ϵ.superscriptitalic-(6.3italic-)absentitalic-${{\leq}}\epsilon.$
As $x\in F(X)$, condition A.2 holds.
We now turn to the if direction. To prove this statement we will use A.1-A.2 to construct a strictly continuous path of unitaries $v_{t}\in\mathcal{U}(\mathcal{M}(B))$ for $t\in[0,\infty)$ such that $(\mathrm{Ad}(v_{t}),h_{v_{t}})\circ(\varphi,h)$ converges to a cocycle conjugacy.
Let $\{x_{n}^{X}\}_{n\in\mathbb{N}}\subset F(X)$, $\{y_{n}^{X}\}_{n\in\mathbb{N}}\subset G(X)$ be countable dense subsets for any $X\in\mathrm{Irr}(\mathcal{C})$ and $K_{n}$ increasing finite subsets of $\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$ such that $\mathrm{Irr}(\mathcal{C})=\bigcup_{n\in\mathbb{N}}K_{n}$. Firstly, use A.1-A.2 to find $x_{1,1}^{X}\in F(X)$ for $X\in K_{1}$ and $z^{(1)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that $z_{0}^{(1)}=1$ and for $0\leq t\leq 1$
  1. (i)
    $\|z_{t}^{(1)}\rhd h^{X}(x_{1}^{X})\lhd(z_{t}^{(1)})^{*}-h^{X}(x_{1}^{X})\|\leq 1/2$ for $X\in K_{1}$,
  2. (ii)
    $\|(z_{1}^{(1)})^{*}\rhd y_{1}^{X}\lhd z_{1}^{(1)}-h^{X}(x_{1,1}^{X})\|\leq 1/2$ for $X\in K_{1}$.
Again use A.1-A.2 to find $x_{2,1}^{X}$, $x_{2,2}^{X}$ in $F(X)$ for $X\in K_{2}$, $z^{(2)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that $z^{(2)}_{0}=1$ and for every $0\leq t\leq 1$
  1. (i)
    $\|z_{t}^{(2)}\rhd h^{X}(x_{j}^{X})\lhd(z_{t}^{(2)})^{*}-h^{X}(x_{j}^{X})\|\leq 1/4$ for $X\in K_{2}$ and $1\leq j\leq 2$,
  2. (ii)
    $\|z_{t}^{(2)}\rhd h^{X}(x_{1,1}^{X})\lhd(z_{t}^{(2)})^{*}-h^{X}(x_{1,1}^{X})\|\leq 1/4$ for $X\in K_{2}$,
  3. (iii)
    $\|(z_{1}^{(2)})^{*}\rhd((z_{1}^{(1)})^{*}\rhd y_{j}^{X}\lhd z_{1}^{(1)})\lhd z_{1}^{(2)}-h^{X}(x_{2,j}^{X})\|\leq 1/4$ for $X\in K_{2}$ and $1\leq j\leq 2$.
Now suppose you have $z^{(k)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ for $1\leq k\leq n$ with $z_{0}^{(k)}=1$ and $x_{m,j}\in F(X)$ for $X\in K_{m}$ with $1\leq j\leq m\leq n$ such that for any $t\in[0,1]$
  1. S.1
    $\|z_{t}^{(n)}\rhd h^{X}(x_{j}^{X})\lhd(z_{t}^{(n)})^{*}-h^{X}(x_{j}^{X})\|\leq 2^{-n}$ for $X\in K_{n}$ and $1\leq j\leq n$,
  2. S.2
    $\|z_{t}^{(n)}\rhd h^{X}(x_{m,j}^{X})\lhd(z_{t}^{(n)})^{*}-h^{X}(x_{m,j}^{X})\|\leq 2^{-n}$ for $X\in K_{n}$ and $1\leq j\leq m<n$,
  3. S.3
    $\|(z_{1}^{(n)})^{*}\ldots(z_{1}^{(1)})^{*}\rhd y_{j}^{X}\lhd z_{1}^{(1)}\ldots z_{1}^{(n)}-h^{X}(x_{n,j}^{X})\|\leq 2^{-n}$ for $X\in K_{n}$ and $1\leq j\leq n$.
Then use A.1-A.2 to get $\{x_{n+1,j}^{X}\}_{j\leq n+1}\in F(X)$ for $X\in K_{n+1}$ and $z^{(n+1)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that for all $t\in[0,1]$
  1. (i)
    $\|z_{t}^{(n+1)}\rhd h^{X}(x_{j}^{X})\lhd(z_{t}^{(n+1)})^{*}-h^{X}(x_{j}^{X})\|\leq 2^{-(n+1)}$ for $X\in K_{n+1}$ and $11\leq j\leq n+$ 1,
  2. (ii)
    $\|z_{t}^{(n+1)}\rhd h^{X}(x_{m,j}^{X})\lhd(z_{t}^{(n+1)})^{*}-h^{X}(x_{m,j}^{X})\|\leq 2^{-(n+1)}$ for $X\in K_{n+1}$ and $11\leq j\leq m<n+$ 1,
  3. (iii)
    $\|(z_{1}^{(n+1)})^{*}\ldots(z_{1}^{(1)})^{*}\rhd y_{j}^{X}\lhd z_{1}^{(1)}\ldots z_{1}^{(n+1)}-h^{X}(x_{n+1,j}^{X})\|\leq 2^{-(n+1)}$ for $X\in K_{n+1}$ and $11\leq j\leq n+$ 1.
We carry on inductively to construct $z^{(n)}$ and $x_{m,j}$ for every $n,m\in\mathbb{N}$ and $j\leq m$ satisfying S.1-S.3. We may now define the path $v_{t}:[0,\infty)\rightarrow\mathcal{U}(\mathcal{M}(B))$ by $v_{t}=z_{1}^{(1)}\ldots z_{1}^{(n)}z_{t-n}^{(n+1)}$ for every $t\in[n,n+1]$. This path is strictly continuous on every open interval $(n,n+1)$ for $n\in\mathbb{N}$ as the paths $z^{(k)}$ are strictly continuous for each $k\in\mathbb{N}$. Moreover, the path $v_{t}$ is strictly continuous at each $n\in\mathbb{N}$ as $z_{0}^{(k)}=1$ for every $k\in\mathbb{N}$. Adjoining by $v_{t}$ we obtain a continuous path of cocycle morphisms $(\mathrm{Ad}(v_{t}),h_{v_{t}})\circ(\varphi,h)=(\psi_{t},h_{t})$ where $\psi_{t}=\mathrm{Ad}(v_{t})\circ\varphi$ and $h_{t}^{X}=h_{v_{t}}^{X}\circ h^{X}$ for any $t\in[0,\infty)$ and $X\in\mathrm{Irr}(\mathcal{C})$.
We claim that as $t\rightarrow\infty$ the path $\psi_{t}$ converges to an isomorphism $\Psi$ and that the path $h_{t}^{X}$ converges to a bijective linear map $H^{X}$ for all $X\in\mathrm{Irr}(\mathcal{C})$ such that the pair $(\Psi,H)$ induces a cocycle morphism (recall Remark 3.12). For any $X\in\mathrm{Irr}(\mathcal{C})$ and $j\in\mathbb{N}$ the net $(h_{t}^{X}(x_{j}^{X}))_{t\geq 0}$ is Cauchy by S.1. Since the set $\{x_{j}^{X}\}_{j\in\mathbb{N}}$ is dense in $F(X)$ for any $X\in\mathrm{Irr}(\mathcal{C})$ a standard triangle inequality argument shows that the net $(h_{t}^{X}(x))_{t\geq 0}$ converges for any $x\in F(X)$ and $X\in\mathrm{Irr}(\mathcal{C})$. We may hence define $H^{X}(x)=\lim_{t\rightarrow\infty}h_{t}^{X}(x)$ for all $x\in F(X)$ and $X\in\mathrm{Irr}(\mathcal{C})$. The maps $H^{X}:F(X)\rightarrow G(X)$ are linear as inherited by the linearity of $h_{t}^{X}$. Letting $X=1_{\mathcal{C}}$ the map $h^{1_{\mathcal{C}}}$ coincides with $\varphi$ and we get that $\Psi(a):=\lim_{t\rightarrow\infty}\mathrm{Ad}(v_{t})\circ\varphi(a)=\lim_{t\rightarrow\infty}\psi_{t}(a)$ exists for all $a\in A$. As each $\psi_{t}$ is a $*{}^{*}$-homomorphism, then so is $\Psi$. In light of Remark 3.12, it suffices to check that the family of maps $\{H^{X}\}_{X\in\mathrm{Irr}(\mathcal{C})}$ satisfies the conditions (i)-(v) of Lemma 3.11 to conclude that $(\Psi,H)$ is a cocycle morphism. Condition (v) follows by construction, whereas conditions (i)-(iv) are easily verified as $(\psi_{t},h_{t})$ is a cocycle morphism for every $t\in[0,\infty)$ and one may approximate $\Psi$ pointwise by $\psi_{t}$ and $H$ pointwise by $h_{t}$ for some large enough $t$.
It remains to check that $\Phi$ is bijective and $H^{X}$ is bijective for every $X\in\mathcal{C}$. By Remark 3.12 and by taking $X=1_{\mathcal{C}}$, it suffices to check that $H^{X}$ is bijective for every $X\in\mathrm{Irr}(\mathcal{C})$. Firstly, each $H^{X}$ is isometric. To see this, note that $\Psi$ is an injective $*{}^{*}$-homomorphism and hence isometric, so as $(\Psi,H)$ is a cocycle morphism
$\|H^{X}(x)\|^{2}$ $=\|\langle H^{X}(x),H^{X}(x)\rangle_{B}\|$
$=\|\Psi(\langle x,x\rangle_{A})\|$
$=\|\langle x,x\rangle_{A}\|$
$=\|x\|^{2}$
for any $X\in\mathcal{C}$ and $x\in F(X)$. Therefore $H^{X}$ is injective for every $X\in\mathcal{C}$. We now turn to surjectivity. As $H^{X}$ are isometric, it suffices to show that $H^{X}$ have dense image. Fix $X\in\mathrm{Irr}(\mathcal{C})$ and $j\in\mathbb{N}$. There is a large enough $n_{0}\in\mathbb{N}$ satisfying $n_{0}>j$ and $X\in K_{n}$ for all $n\geq n_{0}$. So by S.3 $\|h_{n}^{X}(x_{n,j}^{X})-y_{j}^{X}\|\leq 2^{-n}$ for any $n\geq n_{0}$. Moreover, by S.2
$\|H^{X}(x_{n,j}^{X})-h_{n}^{X}(x_{n,j})\|$ $\leq\sum_{k=n}^{\infty}\|h_{k+1}^{X}(x_{n,j}^{X})-h_{k}^{X}(x_{n,j}^{X})\|$
$\leq\sum_{k=n}^{\infty}\|z_{1}^{(k+1)}\rhd h^{X}(x_{n,j}^{X})\lhd(z_{1}^{(k+1)})^{*}-h^{X}(x_{n,j}^{X})\|$
$\leq\sum_{k=n}^{\infty}2^{-(k+1)}$
$\leq 2^{1-n},$
for any $n\geq n_{0}$. So $\|y_{j}^{X}-H^{X}(x_{n,j}^{X})\|\leq 2^{1-n}+2^{-n}$. Therefore, as $n$ may be chosen arbitrarily and $\{y_{j}^{X}\}_{j\in\mathbb{N}}$ is dense in $G(X)$, it follows that $H^{X}$ is surjective. ∎
Theorem 6.1 also holds in the setting of approximate unitary equivalence by replacing the path of unitaries with a single unitary and dropping the assumption that $z_{0}=\mathbf{1}$.
Theorem 6.2.
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $*{}^{*}$-algebras and $(\varphi,h):(A,F,J)\rightarrow(B,G,I)$ an injective cocycle morphism. Then $(\varphi,h)$ is approximately unitarily equivalent to a cocycle conjugacy if and only if:
For all $\varepsilon>0$ and finite sets $K\subset\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$, $\mathcal{F}^{X}\subset F(X)$ and $\mathcal{G}^{X}\subset G(X)$ there exists a unitary $z\in\mathcal{M}(B)$ such that
  1. A.1
    $\|z\rhd h^{X}(x)\lhd z^{*}-h^{X}(x)\|\leq\varepsilon$ for all $X\in K$ and $x\in\mathcal{F}^{X}$,
  2. A.2
    $\mathrm{dist}(z^{*}\rhd y\lhd z,h^{X}(F(X)))\leq\varepsilon$ for all $X\in K$ and $y\in\mathcal{G}^{X}$.
Lemma 6.3.
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F_{n},J^{(n)}):\mathcal{C}\curvearrowright A_{n}$ be a sequence of actions on separable $\mathrm{C}^{*}$-algebras. Let
$(\phi_{n},h_{n}):(A_{n},F_{n},J^{(n)})\to(A_{n+1},F_{n+1},J^{(n+1)})$
be a sequence of injective and extendible cocycle morphisms with inductive limit $(A,F,J)=\lim\limits_{\longrightarrow}(A_{n},F_{n},J^{(n)})$. Suppose that for every $n\geq 1$, $(\phi_{n},h_{n})$ is asymptotically unitarily equivalent to a cocycle conjugacy. Then it follows that
$(\phi_{1,\infty},h_{1,\infty}):(A_{1},F_{1},J^{(1)})\to(A,F,J)$
is asymptotically unitarily equivalent to a cocycle conjugacy.
Proof.
Let $\varepsilon>0$ and finite sets $K\subset\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$, $\mathcal{F}_{1}^{X}\subset F_{1}(X),\ \mathcal{F}^{X}\subset F(X)$. We will check the conditions in Theorem 6.1 for $(\phi_{1,\infty},h_{1,\infty})$. Perturbing $\mathcal{F}^{X}$ by an arbitarily small tolerance, we may assume that there exists $n\geq 1$ large enough and a finite set $\mathcal{F}_{n}^{X}\subset F_{n}(X)$ such that $\mathcal{F}^{X}=h_{n,\infty}^{X}(\mathcal{F}_{n}^{X})$ for any $X\in K$.
By Lemma 4.20, $(\phi_{1,n},h_{1,n})$ is asymptotically unitarily equivalent to a cocycle conjugacy. Then, by Theorem 6.1, there exists a unitary path $y:[0,1]\to\mathcal{U}(\mathcal{M}(A_{n}))$ such that $y_{0}=1$ and
  1. (i)
    $\sup\limits_{0\leq t\leq 1}\|y_{t}\rhd h_{1,n}^{X}(\mu)\lhd y_{t}^{*}-h_{1,n}^{X}(\mu)\|\leq\varepsilon$ for any $X\in K,\mu\in\mathcal{F}_{1}^{X}$,
  2. (ii)
    $\mathrm{dist}(y_{1}^{*}\rhd\eta\lhd y_{1},h_{1,n}^{X}(F_{1}(X)))\leq\varepsilon$ for any $X\in K$, $\eta\in\mathcal{F}_{n}^{X}$
Let $z_{t}=\phi_{n,\infty}^{\dagger}(y_{t})$ be a unitary in $\mathcal{M}(A)$ for any $t\in[0,1]$, where $\phi_{n,\infty}^{\dagger}:\mathcal{U}(\mathcal{M}(A_{n}))\to\mathcal{U}(\mathcal{M}(A))$ is as in Definition 1.5. Now we claim that the path of unitaries $z_{t}$ satisfies the conditions of Theorem 6.1 for $(\phi_{1,\infty},h_{1,\infty})$.
By (i) and (iii) of Lemma 3.11 (see also Remark 3.13) we get that
$\|z_{t}\rhd h_{1,\infty}^{X}(\mu)\lhd z_{t}^{*}-h_{1,\infty}^{X}(\mu)\|\leq\|y_{t}\rhd h_{1,n}^{X}(\mu)\lhd y_{t}^{*}-h_{1,n}^{X}(\mu)\|\leq\varepsilon$
for any $X\in K$, $\mu\in\mathcal{F}_{1}^{X}$, and any $t\in[0,1]$. Moreover, recall that $\mathcal{F}^{X}=h_{n,\infty}^{X}(\mathcal{F}_{n}^{X})$ and
$\mathrm{dist}(y_{1}^{*}\rhd\eta\lhd y_{1},h_{1,n}^{X}(F_{1}(X)))\leq\varepsilon$
for any $X\in K$ and any $\eta\in\mathcal{F}_{n}^{X}$. Hence, by applying $h_{n,\infty}^{X}$ to the formula above, it follows that for any $X\in K$, $\xi\in\mathcal{F}^{X}$,
$\mathrm{dist}(z_{1}^{*}\rhd\xi\lhd z_{1},h_{1,\infty}^{X}(F_{1}(X)))\leq\varepsilon.$
Therefore, $(\phi_{1,\infty},h_{1,\infty})$ is asymptotically unitarily equivalent to a cocycle conjugacy by Theorem 6.1. ∎
The following result can be seen as the asymptotic version of Corollary 5.6.
Theorem 6.4.
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras. Let
$(\phi,h):(A,F,J)\to(B,G,I)\quad and\quad(\psi,l):(B,G,I)\to(A,F,J)$
be two extendible cocycle morphisms such that
$\mathrm{id}_{A}\cong_{u}(\psi,l)\circ(\phi,h)\quad and\quad\mathrm{id}_{B}\cong_{u}(\phi,h)\circ(\psi,l).$
Then there exist mutually inverse cocycle conjugacies
$(\Phi,H):(A,F,J)\to(B,G,I)\quad and\quad(\Psi,L):(B,G,I)\to(A,F,J)$
such that
$(\Phi,H)\cong_{u}(\phi,h)\quad and\quad(\Psi,L)\cong_{u}(\psi,l).$
Proof.
Consider the cocycle morphisms
$(\kappa,r)=(\psi,l)\circ(\phi,h):(A,F,J)\to(A,F,J)$
and
$(\theta,s)=(\phi,h)\circ(\psi,l):(B,G,I)\to(B,G,I)$
fitting into the family of commuting diagrams
[…]
We can then form the inductive limits
$(A^{(\infty)},F^{(\infty)},J^{(\infty)})=\lim\limits_{\longrightarrow}\{(A,F,J),(\kappa,r)\}$
and
$(B^{(\infty)},G^{(\infty)},I^{(\infty)})=\lim\limits_{\longrightarrow}\{(B,G,I),(\theta,s)\}$
in the category $\mathrm{C}^{*}_{\mathcal{C}}$. Consider the universal embeddings
$(\kappa_{\infty},r_{\infty}):(A,F,J)\to(A^{(\infty)},F^{(\infty)},J^{(\infty)})$
and
$(\theta_{\infty},s_{\infty}):(B,G,I)\to(B^{(\infty)},G^{(\infty)},I^{(\infty)}).$
Since the collection of diagrams in (6.4) commutes, the universal properties of both inductive limits yield that there exist mutually inverse cocycle conjugacies
$(\phi_{\infty},h_{\infty}):(A^{(\infty)},F^{(\infty)},J^{(\infty)})\to(B^{(\infty)},G^{(\infty)},I^{(\infty)})$
and
$(\psi_{\infty},l_{\infty}):(B^{(\infty)},G^{(\infty)},I^{(\infty)})\to(A^{(\infty)},F^{(\infty)},J^{(\infty)})$
such that $(\phi_{\infty},h_{\infty})\circ(\kappa_{\infty},r_{\infty})=(\theta_{\infty},s_{\infty})\circ(\phi,h)$ and $(\psi_{\infty},l_{\infty})\circ(\theta_{\infty},s_{\infty})=(\kappa_{\infty},r_{\infty})\circ(\psi,l)$.
Furthermore, the cocycle morphism $(\kappa,r)$ is asymptotically unitarily equivalent to the cocycle morphism induced by the identity map on $A$, so Lemma 6.3 gives that $(\kappa_{\infty},r_{\infty})$ is asymptotically unitarily equivalent to a cocycle conjugacy $(K,R):(A,F,J)\to(A^{(\infty)},F^{(\infty)},J^{(\infty)})$. Likewise, $(\theta_{\infty},s_{\infty})$ is asymptotically unitarily equivalent to a cocycle conjugacy $(\Theta,S):(B,G,I)\to(B^{(\infty)},G^{(\infty)},I^{(\infty)})$.
Then, taking
$(\Phi,H)=(\Theta,S)^{-1}\circ(\phi_{\infty},h_{\infty})\circ(K,R)$
yields that
$(\Phi,H)$ $\cong_{u}(\Theta,S)^{-1}\circ(\phi_{\infty},h_{\infty})\circ(\kappa_{\infty},r_{\infty})$
$=(\Theta,S)^{-1}\circ(\theta_{\infty},s_{\infty})\circ(\phi,h)$
$\cong_{u}(\phi,h).$
Similarly, if we take $(\Psi,L)=(\Phi,H)^{-1}=(K,R)^{-1}\circ(\psi_{\infty},l_{\infty})\circ(\Theta,S)$, we get that $(\Psi,L)\cong_{u}(\psi,l)$, which finishes the proof. ∎
Acknowledgements
We would like to thank Stuart White and Samuel Evington for their supervision on this project. We would also like to thank George Elliott and Gabór Szabó for useful discussions on the topic of this paper. Part of this work was completed during the authors’ stay at the Fields Institute for Research in Mathematical Sciences for the ’Thematic Program on Operator Algebras and Applications’ in Autumn 2023. We thank the Fields Institute and the organisers for the hospitality.
The first named author was supported by the Ioan and Rosemary James Scholarship awarded by St John’s College and the Mathematical Institute, University of Oxford, as well as by project G085020N funded by the Research Foundation Flanders (FWO). The second named author was supported by the EPSRC grant EP/R513295/1. The authors’ stay at the Fields Institute was partially funded by a Special Grant awarded by St John’s College, Oxford. The first named author’s stay was also partially supported by the Fields Institute while the travel costs of the second named author were supported by the EPSRC grant EP/X026647/1.

### 08 Apr 2026 v3

@@
 Abstract.
-We introduce a categorical approach to classifying actions of $*{}^{*}$-tensor categories $\mathcal{C}$ on $\mathrm{C}^{*}$-algebras up to cocycle conjugacy.
+We introduce a categorical approach to classifying actions of C-tensor categories on $\mathrm{C}^{*}$-algebras up to cocycle conjugacy.
 We show that, in this category, inductive limits exist and there is a natural notion of approximate unitary equivalence.
-Then, we generalise classical Elliott intertwining results to the $\mathcal{C}$-equivariant case, in the same fashion as done by Szabó for the group equivariant case in [40].
+Then, we generalise classical Elliott intertwining results to the tensor category equivariant case, in the same fashion as done by Szabó for the group equivariant case.
+RN funded by the European Union.
+Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council.
+Neither the EU nor the ERC can be held responsible for them.
 Introduction
@@
 Later, Popa classified amenable subfactors $N\subset\mathcal{R}$ by their standard invariant ([38]).
-Following subsequent reformulations ([29, 30]), the standard invariant of a finite index subfactor $N\subset\mathcal{R}$ can be understood as a pair $(F,Q)$ with $F$ an action of a unitary tensor category $\mathcal{C}$ on $N$ and $Q$ a special object in $\mathcal{C}$ that allows one to recover the inclusion $N\subset\mathcal{R}$ as a generalised crossed product.
+Following subsequent reformulations ([31, 32]), the standard invariant of a finite index subfactor $N\subset\mathcal{R}$ can be understood as a pair $(F,Q)$ with $F$ an action of a unitary tensor category $\mathcal{C}$ on NN and $Q$ a special object in $\mathcal{C}$ that allows one to recover the inclusion $N\subset\mathcal{R}$ as a generalised crossed product.
 [^1] Precisely, $Q$ is a $Q$-system as defined by Longo in [31].
 The general setting of $\mathrm{C}^{*}$-tensor categories represents a unifying framework for studying group actions and more general quantum symmetries arising from subfactor theory.
-Following the recent spectacular classification results for group actions on Kirchberg algebras by group equivariant KK-theory ([17]), and the development of a $\mathcal{C}$-equivariant KK-theory ([1]), it is natural to consider to what extent one can classify actions of more general $*{}^{*}$-tensor categories on simple, amenable $*{}^{*}$-algebras.
+Following the recent spectacular classification results for group actions on Kirchberg algebras by group equivariant $K$-theory ([17]), and the development of a $\mathcal{C}$-equivariant $K$-theory ([1]), it is natural to consider to what extent one can classify actions of more general C-tensor categories on simple, amenable C-algebras.
 Prior and closely linked to the classification of symmetries of operator algebras was the classification of the operator algebras themselves.
 First, Connes classified amenable factors in [10] with the exception of one type which was later completed by Haagerup ([23]).
-In the case of $*{}^{*}$-algebras, Elliott classified $*{}^{*}$-inductive limits of finite dimensional $*{}^{*}$-algebras by their ordered K-theory in [10].
-Elliott’s methods laid the foundations for a roadmap to a classification of more general simple, amenable $*{}^{*}$-algebras.
+In the case of C-algebras, Elliott classified C-inductive limits of finite dimensional C-algebras by their ordered K-theory in [11].
+Elliott’s methods laid the foundations for a roadmap to a classification of more general simple, amenable C-algebras.
 Following Elliott’s strategy and building on decades of work by many mathematicians, the classification programme of simple, amenable $\mathrm{C}^{*}$-algebras successfully culminated in [28, 37, 21, 22, 42, 6].
 The general strategy for a classification of operator algebras is to first achieve existence and uniqueness results for morphisms with respect to the proposed classification invariant.
-Loosely speaking, to construct an isomorphism between two objects $A$ and $B$ in the category of $\mathrm{C}^{*}$-algebras, by virtue of an existence type statement, one obtains morphisms $\phi:A\to B$ and $\psi:B\to A$ which induce mutually inverse elements at the level of the chosen abstract invariant.
+Loosely speaking, to construct an isomorphism between two objects AA and $B$ in the category of $\mathrm{C}^{*}$-algebras, by virtue of an existence type statement, one obtains morphisms $\phi:A\to B$ and $\psi:B\to A$ which induce mutually inverse elements at the level of the chosen abstract invariant.
 Then, a suitable uniqueness result will give that $\psi\circ\phi$ is equivalent to $\mathrm{id}_{A}$ and $\phi\circ\psi$ is equivalent to $\mathrm{id}_{B}$.
@@
 The final step is to perform a so-called approximate intertwining argument.
-Essentially, one can tweak the morphisms $\phi$ and $\psi$ by unitaries until they become mutually inverse isomorphisms (see [38, Corollary 2.3.4] and the introduction of [40] for a more detailed breakdown of this roadmap to classification).
+Essentially, one can tweak the morphisms $\phi$ and $\psi$ by unitaries until they become mutually inverse isomorphisms (see [40, Corollary 2.3.4] and the introduction of [41] for a more detailed breakdown of this roadmap to classification).
 Furthermore, in [13], Elliott proposes a similar classification strategy for arbitrary objects in more general categories that have a suitable notion of inner automorphisms.
-Instances of the implementation of Elliott’s strategy for the classification of compact (quantum) group actions on $*{}^{*}$-algebras appear in [18, 2], where the actions are assumed to have the Rokhlin property.
+Instances of the implementation of Elliott’s strategy for the classification of compact (quantum) group actions on C-algebras appear in [19, 2], where the actions are assumed to have the Rokhlin property.
 In [41], Szabó articulates Elliott’s proposed strategy in the generality of actions of locally compact groups as an alternative to the Evans–Kishimoto intertwining type arguments (see [14]).
-In his work, Szabó defines an appropriate category of $\Gamma$$*{}^{*}$-algebras, that is $*{}^{*}$-algebras carrying an action of $\Gamma$, and produces a vast collection of intertwining results for this category.
+In his work, Szabó defines an appropriate category of $\Gamma$-C-algebras, that is C-algebras carrying an action of $\Gamma$, and produces a vast collection of intertwining results for this category.
 Szabó’s construction is a key ingredient facilitating the recent groundbreaking classification of amenable actions on Kirchberg algebras in [17].
-In this paper, we will develop the necessary techniques to perform approximate intertwining arguments for $*{}^{*}$-algebras carrying an action of a tensor category $\mathcal{C}$ ($\mathcal{C}$-$\mathrm{C}^{*}$-algebras).
+In this paper, we will develop the necessary techniques to perform approximate intertwining arguments for C-algebras carrying an action of a tensor category $\mathcal{C}$ ($\mathcal{C}$-$\mathrm{C}^{*}$-algebras).
 Glimpses of these techniques can be found in [7], where exact intertwining arguments are performed to classify inductive limit actions of fusion categories on AF-algebras.
@@
 However, unlike in the group action case, a tensor category might not act by automorphisms, so our techniques differ.
-In [40], Szabó defines a cocycle morphism as a pair consisting of an extendible $*{}^{*}$-homomorphism and a unitary cocycle.
+In [41], Szabó defines a cocycle morphism as a pair consisting of an extendible -homomorphism and a unitary cocycle.
 In the setting of tensor categories, the action on a $\mathrm{C}^{*}$-algebra is given by a family of bimodules acting compatibly with respect to the tensor product.
@@
 Hence, it is not apparent how to adapt Szabó’s arguments.
-[…]
+To perform intertwining techniques, we first need to introduce a category whose objects are $\mathcal{C}$-C-algebras.
+Multiple notions of morphisms between $\mathcal{C}$-C-algebras have appeared in the literature (see [1, Definition 2.11] or [7, Definition 3.2] for example).
+Any such notion has a common flavour: Given a C-tensor category $\mathcal{C}$ acting on $\mathrm{C}^{*}$-algebras AA and $B$ via tensor functors $(F,J):\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}(A)$ and $(G,I):\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}(B)$ respectively, a morphism between $(A,F,J)$ and $(B,G,I)$ is given by an AA-$B$-correspondence $E$ and a family of bimodule maps $\mathbbm{v}_{X}:F(X)\boxtimes E\rightarrow E\boxtimes G(X)$ satisfying coherence diagrams.
+[^2] See Section 2 for our conventions on actions of tensor categories.
+The intricacies between the different definitions lie in how general correspondences we allow and what structure we expect the maps $\mathbbm{v}_{X}$ to have.
+We choose to work with correspondences which arise from (possibly degenerate) -homomorphisms $\varphi:A\rightarrow B$ and with (possibly non-adjointable) isometries $\mathbbm{v}_{X}$.
+We will say this data yields a cocycle morphism and we denote the category consisting of $\mathcal{C}$-$\mathrm{C}^{*}$-algebras and cocycle morphisms by $\mathrm{C}^{*}_{\mathcal{C}}$.
+When the acting category is $\mathrm{Hilb}(\Gamma)$, and the action factors through automorphisms, we give an explicit formula for the family of bimodule maps $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ in Example $3.$
 In $\mathrm{C}^{*}_{\mathcal{C}}$, we are able to construct inductive limits and define a suitable topology on the space of morphisms.
@@
 Therefore, we need to introduce a slightly different composition which will allow us to obtain a category.
-If the morphisms are non-degenerate, this new composition agrees with the canonical one, and with the one introduced by Szabó in the group action case ([40, Proposition 1.15]).
+If the morphisms are non-degenerate, this new composition agrees with the canonical one, and with the one introduced by Szabó in the group action case ([41, Proposition 1.15]).
 Moreover, to avoid the need of restricting to non-degenerate morphisms, the maps $\mathbbm{v}_{X}$ are assumed to be isometries.
 This assumption allows us, unlike in [41], to work in the possibly non-extendible setting, generalising the notion of a cocycle morphism from [41] when restricting to twisted group actions.
-An essential observation for our construction is that we can encode the information of a cocycle morphism into a family of linear maps satisfying certain conditions (see [6, Lemma 3.8] where this is done for unital injective cocycle morphisms).
+An essential observation for our construction is that we can encode the information of a cocycle morphism into a family of linear maps satisfying certain conditions (see [7, Lemma 3.8] where this is done for unital injective cocycle morphisms).
 Therefore, a cocycle morphism can be equivalently defined as follows.
 Definition A (Lemma 3.10).
-Let $\mathcal{C}$ be a $\mathrm{C}^{*}$-tensor category acting on $\mathrm{C}^{*}$-algebras $A$ and $B$ via $(A,F,J)$ and $(B,G,I)$ respectively.
-Then a cocycle morphism $(\phi,h):(A,F,J)\rightarrow(B,G,I)$ is given by a $*{}^{*}$-homomorphism $\phi:A\to B$ and a family of linear maps:
+Let $\mathcal{C}$ be a $\mathrm{C}^{*}$-tensor category acting on $\mathrm{C}^{*}$-algebras AA and $B$ via $(A,F,J)$ and $(B,G,I)$ respectively.
+Then a cocycle morphism $(\phi,h):(A,F,J)\rightarrow(B,G,I)$ is given by a -homomorphism $\phi:A\to B$ and a family of linear maps:
 $h=\{h^{X}:F(X)\rightarrow G(X)\}_{X\in\mathcal{C}}$
@@
   1. (i)
-    $h^{X}(a\rhd x\lhd a^{\prime})=\phi(a)\rhd h^{X}(x)\lhd\phi(a^{\prime})$ for any $a,a^{\prime}\in A$ and any $x\in F(X)$;
+    for any morphism $f\in\mathrm{Hom}(X,Y)$, $G(f)\circ h^{X}=h^{Y}\circ F(f)$;
   2. (ii)
-    for any morphism $f\in\mathrm{Hom}(X,Y)$, $G(f)\circ h^{X}=h^{Y}\circ F(f)$;
+    $\phi(\langle x,y\rangle_{A})=\langle h^{X}(x),h^{X}(y)\rangle_{B}$ for any $x,y\in F(X)$;
   3. (iii)
-    $\phi(\langle x,y\rangle_{A})=\langle h^{X}(x),h^{X}(y)\rangle_{B}$ for any $x,y\in F(X)$;
+    the following diagram commutes:
+    $F(Y)\boxtimes F(X)$$F(X\otimes Y)$$G(Y)\boxtimes G(X)$$G(X\otimes Y);$$\scriptstyle{h^{Y}\boxtimes h^{X}}$$\scriptstyle{J_{X,Y}}$$\scriptstyle{h^{X\otimes Y}}$$\scriptstyle{I_{X,Y}}$
   4. (iv)
-    the following diagram commutes:
-[…]
-  5. (v)
-[…]
+    $h^{1_{\mathcal{C}}}:A\to B$ is given by $h^{1_{\mathcal{C}}}(a)=\phi(a)$ for any $a\in A$.
 In the case of a group action, we give an explicit formula for the family of linear maps $\{h^{X}\}_{X\in\mathcal{C}}$ in Example 3.14.
-Compared with the usual definition of a cocycle morphism given by pairs $(\phi,\mathbbm{v})$ consisting of a $*{}^{*}$-homomorphism and a family of bimodule maps $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ (see Definition 3.1), Definition A provides a more direct framework for setting up the intertwining arguments.
-In particular, to compose two cocycle morphisms, one composes the $*{}^{*}$-homomorphisms, as well as the corresponding linear maps.
+Compared with the usual definition of a cocycle morphism given by pairs $(\phi,\mathbbm{v})$ consisting of a -homomorphism and a family of bimodule maps $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ (see Definition 3.1), Definition A provides a more direct framework for setting up the intertwining arguments.
+In particular, to compose two cocycle morphisms, one composes the -homomorphisms, as well as the corresponding linear maps.
 Furthermore, the natural topology on cocycle morphisms is obtained by considering pointwise differences of the linear maps.
@@
 [^3] The collection $\mathrm{Irr}(\mathcal{C})$ is a choice of representatives for isomorphism classes of simple objects in $\mathcal{C}$.
-In general, this topology is coarser than the one used by Szabó in the group action case ([40, Definition 2.5]) (see Example 4.10).
+In general, this topology is coarser than the one used by Szabó in the group action case ([41, Definition 2.5]) (see Example 5.5).
 However, it is the same when restricted to non-degenerate cocycle morphisms.
 With this topology in hand, we can then define approximate unitary conjugation for cocycle morphisms.
-Moreover, unlike Szabó’s definition in [40, Definition 2.8], this is a symmetric relation (see Lemma 4.16), which simplifies the intertwining arguments.
+Moreover, unlike Szabó’s definition in [41, Definition 2.8], this is a symmetric relation (see Lemma 5.9), which simplifies the intertwining arguments.
 This can be formulated as follows.
@@
 If $(\phi,h),(\psi,l):(A,F,J)\to(B,G,I)$ are cocycle morphisms, then $(\psi,l)$ is approximately unitarily equivalent to $(\phi,h)$ if and only if there exists a net of unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
+$\|l^{X}(x)-u_{\lambda}\rhd h^{X}(x)\lhd u_{\lambda}^{*}\|\overset{\lambda}{\longrightarrow}0$
 […]
-for any $X\in\mathrm{Irr}(\mathcal{C})$ and any $x\in F(X)$.
 With the introduced topology at the level of morphism spaces and notion of unitary equivalence, we can place the subcategory of $\mathrm{C}^{*}_{\mathcal{C}}$ consisting of separable $\mathcal{C}$-$\mathrm{C}^{*}$-algebras and extendible cocycle morphisms in Elliott’s abstract classification framework from [13].
@@
 Along with Theorem D, in Section 6 we use our construction to obtain various $\mathcal{C}$-equivariant versions of classical intertwining arguments.
-For instance, we establish a general Elliott two-sided intertwining argument (see [38, Proposition 2.3.2] for a non-equivariant version) as well as an intertwining through reparametrisation in Theorem 5.9.
-In Section 6, the techniques we develop allow us to perfom a $\mathcal{C}$-equivariant one-sided intertwining argument (see Theorem 6.1), generalising the classical one-sided intertwining arguments of [38, Proposition 2.3.5] and [40, Proposition 4.3].
+For instance, we establish a general Elliott two-sided intertwining argument (see [40, Proposition 2.3.2] for a non-equivariant version) as well as an intertwining through reparametrisation in Theorem 6.8.
+These results are essential ingredients in our upcoming classification results for actions of unitary tensor category on Kirchberg algebras in joint work with Kitamura.
+In Section 7, the techniques we develop allow us to perfom a $\mathcal{C}$-equivariant one-sided intertwining argument (see Theorem 7.1), generalising the classical one-sided intertwining arguments of [40, Proposition 2.3.5] and [41, Proposition 4.3].
 This result is a key ingredient in the work of Evington, C.
 Jones, and the first named author in obtaining a McDuff-type characterisation of equivariant $\mathcal{D}$-stability of an action of a unitary tensor category, for strongly self-absorbing $\mathcal{D}$ ([15]).
-Precisely, $\mathcal{D}$-stability of an action is equivalent to the existence of a unital embedding of $\mathcal{D}$ into an appropriate subalgebra of the Kirchberg central sequence algebra of $A$ (in the case of a group action, this subalgebra coincides with the fixed point subalgebra of Kirchberg’s central sequence algebra).
-This generalises the classical result of [43, Theorem 2.2] (see also [27]) and its group equivariant counterpart (see [39, Corollary 3.8]).
 The remaining part of Section 7 is concerned with asymptotic versions of the results in Section 6.
 1. Preliminaries
-1.1. Hilbert bimodules
+$1.$ Hilbert bimodules
 In this subsection, we collect a few basics on the theory of Hilbert bimodules.
@@
 First, we start by recalling the definition of a Hilbert module as introduced by Paschke in [36].
-Definition 1.1.
+Definition $1.$
 Let $X$ be a vector space over $\mathbb{C}$ and $B$ be a $\mathrm{C}^{*}$-algebra.
@@
   3. (iii)
-    For any $x\in X$, $\langle x,x\rangle_{B}\geq 0$ and $\langle x,x\rangle_{B}=0$ if and only if $x=0$.
+[…]
   4. (iv)
@@
 For $x\in X$, we denote by $|x|^{2}$ the positive element of $B$ given by $\langle x,x\rangle_{B}$.
-To obtain a right-Hilbert $A$-$B$-bimodule, we need a compatible left action of a $\mathrm{C}^{*}$-algebra $A$.
-Definition 1.2.
-([25, Definition 2.1]) Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras and $X={}_{A}X_{B}$ a bimodule over the complex algebras $A$ and $B$.
-We say that $X$ is a (right-)Hilbert $A$-$B$-bimodule if $X$ is a right-Hilbert $B$-module, and for all $a\in A$, the map $\phi(a):x\in X\mapsto ax\in X$ is adjointable, with adjoint $\phi(a)^{*}=\phi(a^{*})$.
-In particular, $\phi:a\in A\to\phi(a)\in\mathcal{L}(X)$ is a $*{}^{*}$-homomorphism from $A$ to the $*{}^{*}$-algebra $\mathcal{L}(X)$ of adjointable maps on $X$.
-The map $\phi$ will be referred to as the left action of $A$ on $X$.
-For a Hilbert $A$-$B$-bimodule $X$, we often write the left action of an element $a\in A$ on a vector $x\in X$ by $a\rhd x$ and the right action of an element $b\in B$ on $x$ by $x\lhd b$.
+We denote the adjointable operators of $X$ by $\mathcal{L}(X)$.
+A (right-)Hilbert AA-$B$-bimodule, is a right Hilbert $B$-module with a left AA action by adjointable operators.
+For a Hilbert AA-$B$-bimodule $X$, we often write the left action of an element $a\in A$ on a vector $x\in X$ by $a\rhd x$ and the right action of an element $b\in B$ on $x$ by $x\lhd b$.
+The morphisms we consider between Hilbert AA-$B$-bimodules are adjointable linear maps which commute with the left AA-action.
+In particular, Hilbert AA-$B$-bimodules are called isomorphic if there is a unitary bimodule map between them.
 Example 1.2.
-For a $*{}^{*}$-algebra $B$, the vector space $X=B$ is a Hilbert $B$-module with the right $B$-module structure given by multiplication and the inner product $\langle a,b\rangle_{B}=a^{*}b$ for $a,b\in B$.
-In this case, any $*{}^{*}$-homomorphism $\phi:A\rightarrow M(B)$ induces a right-Hilbert $A$-$B$-bimodule which we denote by ${}_{\phi}B$.
+For a C-algebra $B$, the vector space $X=B$ is a Hilbert $B$-module with the right $B$-module structure given by multiplication and the inner product $\langle a,b\rangle_{B}=a^{*}b$ for $a,b\in B$.
+In this case, any -homomorphism $\phi:A\rightarrow M(B)$ induces a right-Hilbert AA-$B$-bimodule which we denote by $_{\phi}B$.
 Let $A,B,C$ be $\mathrm{C}^{*}$-algebras.
-If $X$ is a Hilbert $A$-$B$-bimodule and $E$ is a Hilbert $B$-$C$-bimodule we may form their internal tensor product $X\boxtimes E$ that is an Hilbert $A$-$C$-bimodule.
-We sketch this construction and refer to [28, Section 4] for details.
+If $X$ is a Hilbert AA-$B$-bimodule and $E$ is a Hilbert $B$-$C$-bimodule we may form their internal tensor product $X\boxtimes E$ that is a Hilbert AA-$C$-bimodule.
+We sketch this construction and refer to [29, Section 4] for details.
 To perform the internal tensor product one starts by considering the algebraic tensor product of vector spaces $X\odot E$.
@@
 $(x\boxtimes y)\lhd c=x\boxtimes(y\lhd c),$
-$\langle x\boxtimes y,z\boxtimes w\rangle_{C}=\langle x,\langle y,z\rangle_{B}\rhd w\rangle_{C},$
-for any $x,z\in X$ and $z,w\in E$.
-It follows that $V$ equipped with this $C$-action and $C$-valued inner product satisfies (i)-(iv) of Definition 1.1.
-We produce a Hilbert $C$-module $X\boxtimes E$ by completing $V$ under the norm defined by the inner product.
-Moreover, one can induce a left $A$ action on $X\boxtimes E$ through
+$\langle x\boxtimes y,z\boxtimes w\rangle_{C}=\langle y,\langle x,z\rangle_{B}\rhd w\rangle_{C},$
+for any $x,z\in X$ and $y,w\in E$.
+It follows that $V$ equipped with this $C$-action and $C$-valued inner product satisfies (i)-(iv) of Definition $1.$ We produce a Hilbert $C$-module $X\boxtimes E$ by completing $V$ under the norm defined by the inner product.
+Moreover, one can induce a left AA action on $X\boxtimes E$ through
 $a\rhd(x\boxtimes y)=(a\rhd x)\boxtimes y$
-for all $a\in A$, $x\in X$ and $y\in E$. This equips $X\boxtimes E$ with the structure of a Hilbert $A$-$C$-bimodule.
-If the bimodules are given by $*{}^{*}$-homomorphisms into the multiplier algebra as in Example 1.3, one may get greater insight into the structure of their tensor product.
-First we recall the definition of non-degenerate and extendible $*{}^{*}$-homomorphisms.
+for all $a\in A$, $x\in X$ and $y\in E$. This equips $X\boxtimes E$ with the structure of a Hilbert AA-$C$-bimodule.
+If the bimodules are given by -homomorphisms into the multiplier algebra as in Example 1.2, one may get greater insight into the structure of their tensor product.
+First we recall the definition of non-degenerate and extendible -homomorphisms.
 Definition 1.3.
-Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras and $\phi:A\to\mathcal{M}(B)$ a $*{}^{*}$-homomorphism. Then $\phi$ is said to be non-degenerate if $\phi(A)B$ is dense in $B$.
-If $\phi:A\to\mathcal{M}(B)$ is non-degenerate, then $\phi$ extends to a unital $*{}^{*}$-homomorphism $\phi:\mathcal{M}(A)\to\mathcal{M}(B)$ ([28, Proposition 2.5]).
+Let AA and $B$ be $\mathrm{C}^{*}$-algebras and $\phi:A\to\mathcal{M}(B)$ a -homomorphism. Then $\phi$ is said to be non-degenerate if $\phi(A)B$ is dense in $B$.
+If $\phi:A\to\mathcal{M}(B)$ is non-degenerate, then $\phi$ extends to a unital -homomorphism $\phi:\mathcal{M}(A)\to\mathcal{M}(B)$ ([29, Proposition 2.5]).
 Definition 1.4.
-Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras.
-A $*{}^{*}$-homomorphism $\phi:A\to\mathcal{M}(B)$ is called extendible if for any increasing approximate unit $e_{\lambda}\in A$, the net $\phi(e_{\lambda})\in\mathcal{M}(B)$ converges strictly to a projection $p\in\mathcal{M}(B)$.
-[^5] Any non-degenerate $*{}^{*}$-homomorphism $\phi:A\rightarrow M(B)$ is extendible as for an approximate unit $e_{\lambda}\in A$ the net $\phi(e_{\lambda})$ converges strictly to $1_{\mathcal{M}(B)}$ (see e.g.
+Let AA and $B$ be $\mathrm{C}^{*}$-algebras.
+A -homomorphism $\phi:A\to\mathcal{M}(B)$ is called extendible if for any increasing approximate unit $e_{\lambda}\in A$ (throughout this paper all approximate units are assumed to be increasing), the net $\phi(e_{\lambda})\in\mathcal{M}(B)$ converges strictly to a projection $p\in\mathcal{M}(B)$.
+[^5] Any non-degenerate -homomorphism $\phi:A\rightarrow\mathcal{M}(B)$ is extendible as for an approximate unit $e_{\lambda}\in A$ the net $\phi(e_{\lambda})$ converges strictly to $1_{\mathcal{M}(B)}$ (see e.g.
 [3, II.7.3.8.]).
-As discussed in [40, Definition 1.7], $\phi$ then factorises through $\mathcal{M}(pBp)\cong p\mathcal{M}(B)p\subseteq\mathcal{M}(B)$, with the $*{}^{*}$-homomorphism $\phi_{p}:A\to\mathcal{M}(pBp)$ now non-degenerate.
-In this case, following [40], we let $\phi^{\dagger}:\mathcal{U}(\mathcal{M}(A))\to\mathcal{U}(\mathcal{M}(B))$ be the unital group homomorphism defined by $\phi^{\dagger}(u)=\phi_{p}(u)+(1_{M(B)}-p)$ for all $u\in\mathcal{U}(\mathcal{M}(A))$.
+As discussed in [41, Definition 1.7], $\phi$ then factorises through $\mathcal{M}(pBp)\cong p\mathcal{M}(B)p\subseteq\mathcal{M}(B)$, with the -homomorphism $\phi_{p}:A\to\mathcal{M}(pBp)$ now non-degenerate.
+In this case, following [41], we let $\phi^{\dagger}:\mathcal{U}(\mathcal{M}(A))\to\mathcal{U}(\mathcal{M}(B))$ be the unital group homomorphism defined by $\phi^{\dagger}(u)=\phi_{p}(u)+(1_{\mathcal{M}(B)}-p)$ for all $u\in\mathcal{U}(\mathcal{M}(A))$.
 Note that for any $a\in A$ and $u\in\mathcal{U}(\mathcal{M}(A))$ one has that $\phi(ua)=\phi^{\dagger}(u)\phi(a)$.
 Remark 1.5.
-Let $\phi:A\rightarrow\mathcal{M}(B)$, $\psi:B\rightarrow\mathcal{M}(C)$ be $*{}^{*}$-homomorphisms with $\psi$ extendible, and $q\in\mathcal{M}(C)$ be the corresponding projection to $\psi$.
-One may compose them by the $*{}^{*}$-homomorphism $\psi_{q}\circ\phi:A\rightarrow\mathcal{M}(C)$ that we simply denote by $\psi\circ\phi$ (see [40, Remark 1.9]).
+Let $\phi:A\rightarrow\mathcal{M}(B)$, $\psi:B\rightarrow\mathcal{M}(C)$ be -homomorphisms with $\psi$ extendible, and $q\in\mathcal{M}(C)$ be the projection corresponding to $\psi$.
+One may compose them to get the -homomorphism $\psi_{q}\circ\phi:A\rightarrow\mathcal{M}(C)$ that we simply denote by $\psi\circ\phi$ (see [41, Remark 1.9]).
 If $\phi$ is also extendible, then the composition $\psi\circ\phi$ remains extendible and $(\psi\circ\phi)^{\dagger}=\psi^{\dagger}\circ\phi^{\dagger}$.
-The proposition below yields a more explicit form for the tensor products of bimodules arising from $*{}^{*}$-homomorphisms into the multiplier algebra as in Example 1.3.
+The proposition below yields a more explicit form for tensor products of bimodules arising from -homomorphisms into the multiplier algebra as in Example 1.2.
 If the homomorphisms are non-degenerate, then the result is folklore (see [5]).
 Proposition 1.6.
-Suppose that $A,B$, and $C$ are $\mathrm{C}^{*}$-algebras and $\phi:A\to\mathcal{M}(B)$, $\psi:B\to\mathcal{M}(C)$ are $*{}^{*}$-homomorphisms with $\psi$ extendible, but possibly degenerate.
+Suppose that $A,B$, and $C$ are $\mathrm{C}^{*}$-algebras and $\phi:A\to\mathcal{M}(B)$, $\psi:B\to\mathcal{M}(C)$ are -homomorphisms with $\psi$ extendible, but possibly degenerate.
 Then $_{\phi}B\boxtimes{}_{\psi}C\cong{}_{\psi\circ\phi}\overline{\psi(B)C}={}_{\psi\circ\phi}\psi(B)C$.
 Proof.
-First, note that $\overline{\psi(B)C}=\psi(B)C$ by Cohen’s factorisation (see [37, Proposition 2.33]).
+First, note that $\overline{\psi(B)C}=\psi(B)C$ by Cohen’s factorisation (see [39, Proposition 2.33]).
 Let $T:{}_{\phi}B\boxtimes{}_{\psi}C\to{}_{\psi\circ\phi}\psi(B)C$ be the continuous linear map given by $T(b\boxtimes c)=\psi(b)c$ for any $b\in B$ and $c\in C$.
@@
 $=\langle\psi(\eta_{\lambda}-\eta_{\mu})c,\psi(\eta_{\lambda}-\eta_{\mu})c\rangle.$
-Therefore, $S$ is well defined if and only if $\psi(\eta_{\lambda}-\eta_{\mu})c$ converges to 00 for any $c\in C$, which is equivalent to saying that $\psi(\eta_{\lambda})$ converges strictly in $\mathcal{M}(C)$.
+Therefore, $S$ is well defined if and only if $\psi(\eta_{\lambda}-\eta_{\mu})c$ converges to 0 for any $c\in C$, which is equivalent to saying that $\psi(\eta_{\lambda})$ converges strictly in $\mathcal{M}(C)$.
 Hence $S$ is well-defined.
@@
 Definition 1.8.
-Let $A$ be a $*{}^{*}$-algebra. A Hilbert $A$-$B$-bimodule $X$ is called non-degenerate if $XB$ is dense in $X$ and $AX$ is dense in $X$.
+Let AA be a C-algebra. A Hilbert AA-$B$-bimodule $X$ is called non-degenerate if XBXB is dense in $X$ and AXAX is dense in $X$.
 Remark 1.9.
-Note that if $X$ is a right-Hilbert $B$-module then $XB$ is always dense in $X$ by the argument in [25, Proposition 2.6].
-Therefore, a Hilbert $A$-$B$-bimodule $E$ may only fail to be non-degenerate if $AX$ is not dense in $X$.
-We end this subsection by showing that if $A$ is a $\mathrm{C}^{*}$-algebra and $E$ is a non-degenerate Hilbert $A$-bimodule, then we can obtain a well-defined action of $\mathcal{M}(A)$ on $E$.
-Let $(\eta_{\lambda})_{\lambda\in\Lambda}$ be an approximate unit of $A$ and let $L_{E}:E\to A\boxtimes E$ and $R_{E}:E\to E\boxtimes A$ be the $A$-bimodule maps given by $L_{E}(x)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes x$ and $R_{E}(x)=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}$ for all $x\in E$.
+Note that if $X$ is a right-Hilbert $B$-module then XBXB is always dense in $X$ by the argument in [26, Proposition 2.16].
+Therefore, a Hilbert AA-$B$-bimodule $E$ may only fail to be non-degenerate if AXAX is not dense in $X$.
+We end this subsection by showing that if AA is a $\mathrm{C}^{*}$-algebra and $E$ is a non-degenerate Hilbert AA-bimodule, then we can obtain a well-defined action of $\mathcal{M}(A)$ on $E$.
+Let $(\eta_{\lambda})_{\lambda\in\Lambda}$ be an approximate unit of AA and let $L_{E}:E\to A\boxtimes E$ and $R_{E}:E\to E\boxtimes A$ be the AA-bimodule maps given by $L_{E}(x)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes x$ and $R_{E}(x)=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}$ for all $x\in E$.
 Note that, since $E$ is non-degenerate, $L_{E}$ and $R_{E}$ are unitary bimodule isomorphisms.
@@
 Lemma 1.10.
-Let $A$ be a $\mathrm{C}^{*}$-algebra and $E$ be a non-degenerate Hilbert $A$-bimodule.
-Then one may extend the left and right actions of $A$ on $E$ to left and right actions of $\mathcal{M}(A)$ on $E$.
+Let AA be a $\mathrm{C}^{*}$-algebra and $E$ be a non-degenerate Hilbert AA-bimodule.
+Then one may extend the left and right actions of AA on $E$ to left and right actions of $\mathcal{M}(A)$ on $E$.
 These extended actions equip $E$ with the structure of a Hilbert $\mathcal{M}(A)$-bimodule.
@@
 For any $v\in\mathcal{M}(A)$ and any $x\in E$, let us define
-(1.1) $v\rhd x=L_{E}^{-1}(v\rhd L_{E}(x))\quad\mathrm{and}\quad x\lhd v=R_{E}^{-1}(R_{E}(x)\lhd v).$
-We claim that these formulae define left and right actions of $\mathcal{M}(A)$ on $E$. Moreover, it is clear that (1.1) restricted to $A$ coincides with the $A$-bimodule structure of $E$.
-First, note that the left action of $\mathcal{M}(A)$ on $A\boxtimes E$ is given by left multiplication on $A$.
-Similarly, the right action of $\mathcal{M}(A)$ on $E\boxtimes A$ is given by right multiplication on $A$.
+[…]
+We claim that these formulae define left and right actions of $\mathcal{M}(A)$ on $E$. Moreover, it is clear that (1.1) restricted to AA coincides with the AA-bimodule structure of $E$.
+First, note that the left action of $\mathcal{M}(A)$ on $A\boxtimes E$ is given by left multiplication on AA.
+Similarly, the right action of $\mathcal{M}(A)$ on $E\boxtimes A$ is given by right multiplication on AA.
 Since $L_{E},R_{E}$, and their inverses are bimodule maps, it is straightforward to see that the formulae in (1.1) define left and right actions of $\mathcal{M}(A)$ on $E$.
-To see that $E$ with its right $A$-valued inner product is a Hilbert $\mathcal{M}(A)$-bimodule it suffices to check that the right $\mathcal{M}(A)$-action commutes with the inner product and the left $\mathcal{M}(A)$-action consists of adjointable operators.
+To see that $E$ with its right AA-valued inner product is a Hilbert $\mathcal{M}(A)$-bimodule it suffices to check that the right $\mathcal{M}(A)$-action commutes with the inner product and the left $\mathcal{M}(A)$-action consists of adjointable operators.
 First, for $x,y\in E$ and $v\in\mathcal{M}(A)$ we have that
@@
 Moreover, the operator of left multiplication by $v\in\mathcal{M}(A)$ has as an adjoint the operator of left multiplication by $v^{*}$.
-Indeed left multiplication by $v^{*}$ is the adjoint of the operator of left multiplication by $v$ on the Hilbert $A$-module $A$ so for $x,y\in E$
+Indeed left multiplication by $v^{*}$ is the adjoint of the operator of left multiplication by $v$ on the Hilbert AA-module AA so for $x,y\in E$
 $\langle v\rhd x,y\rangle_{A}=\langle L_{E}^{-1}(v\rhd L_{E}(x)),y\rangle_{A}=\langle L_{E}(x),v^{*}\rhd L_{E}(y)\rangle_{A}=\langle x,v^{*}\rhd y\rangle_{A}$
@@
 1.2. Correspondences
-A reformulation of the theory of Hilbert bimodules is the language of $*{}^{*}$-correspondences.
+A reformulation of the theory of Hilbert bimodules is the language of C-correspondences.
 Definition 1.11.
 Let $A,B$ be $\mathrm{C}^{*}$-algebras.
-An $A$-$B$ correspondence is a $*{}^{*}$-homomorphism $\phi:A\to\mathcal{L}(X_{B})$ into the adjointable operators of a Hilbert $B$-module $X_{B}$.
-We denote the collection of $A$-$B$ correspondences by $\mathrm{Corr}(A,B)$.
+An AA-$B$-correspondence is a Hilbert $B$-module $X_{B}$ together with a -homomorphism $\phi:A\to\mathcal{L}(X_{B})$.
+We often simply denote a correspondence by its underlying -homomorphism.
+Moreover, we denote the collection of AA-$B$-correspondences by $\mathrm{Corr}(A,B)$.
 Remark 1.12.
-Note that any $A$-$B$ correspondence induces a right-Hilbert $A$-$B$-bimodule and vice versa.
-If $\phi:A\to\mathcal{L}(X_{B})$ is an $A$-$B$ correspondence, then $X_{B}$ becomes a right-Hilbert $A$-$B$-bimodule, with the left action given by $\phi$.
+Note that any AA-$B$-correspondence induces a right-Hilbert AA-$B$-bimodule and vice versa.
+If $\phi:A\to\mathcal{L}(X_{B})$ is an AA-$B$-correspondence, then $X_{B}$ becomes a right-Hilbert AA-$B$-bimodule, with the left action given by $\phi$.
 We will often denote this bimodule by $_{\phi}X$, forgetting the right $B$-action.
-Conversely, given a right-Hilbert $A$-$B$-bimodule $X$, the left action by $A$ induces an $A$-$B$ correspondence.
+Conversely, given a right-Hilbert AA-$B$-bimodule $X$, the left action by AA induces an AA-$B$-correspondence.
 Therefore, we will freely flip between the two pictures.
 One may compose correspondences through the tensor product of bimodules.
-Precisely, let $X$ be a Hilbert $A$-$B$ bimodule inducing $\phi\in\mathrm{Corr}(A,B)$, and let $E$ be a Hilbert $B$-$C$ bimodule inducing $\psi\in\mathrm{Corr}(B,C)$.
-Then $X\boxtimes E$ gives a Hilbert $A$-$C$ bimodule which induces an element in $\mathrm{Corr}(A,C)$ denoted by $\psi\circ\phi$.
-Although the theories of bimodules and correspondences are equivalent, we sometimes choose to work with correspondences as the composition resembles composition of $*{}^{*}$-homomorphisms between $\mathrm{C}^{*}$-algebras in a covariant manner.
+Precisely, let $X$ be a Hilbert AA-$B$-bimodule inducing $\phi\in\mathrm{Corr}(A,B)$, and let $E$ be a Hilbert $B$-$C$-bimodule inducing $\psi\in\mathrm{Corr}(B,C)$.
+Then $X\boxtimes E$ gives a Hilbert AA-$C$-bimodule which induces an element in $\mathrm{Corr}(A,C)$ denoted by $\psi\circ\phi$.
+Although the theories of bimodules and correspondences are equivalent, we sometimes choose to work with correspondences as the composition resembles composition of -homomorphisms between $\mathrm{C}^{*}$-algebras in a covariant manner.
 1.3. $\mathrm{C}^{*}$-tensor categories
@@
 The space of morphisms between two objects $X,Y\in\mathcal{C}$ will be denoted by $\mathrm{Hom}(X,Y)$.
-All categories in this section will be $\mathbb{C}$-linear, that is and additive category such that its space of morphisms $\mathrm{Hom}(X,Y)$ between any two objects $X,Y$ is a $\mathbb{C}$-vector space and the composition of morphisms yields a bilinear map $\mathrm{Hom}(X,Y)\times\mathrm{Hom}(Y,Z)\rightarrow\mathrm{Hom}(X,Z)$ for any $X,Y,Z\in\mathcal{C}$.
-We start by defining $\mathrm{C}^{*}$-categories.
-Definition 1.14.
-(see [19]) A $\mathbb{C}$-linear category $\mathcal{C}$ is said to be a $\rm{C}^{*}$-category if it is equipped with conjugate linear mappings :*${}^{*}\!:\mathrm{Hom}(X,Y)\rightarrow\mathrm{Hom}(Y,X)$ for every $X,Y\in\mathcal{C}$ such that
-  1. (i)
-    $\phi^{**}=\phi$ for all $\phi\in\mathrm{Hom}(X,Y);$
-  2. (ii)
-    $(\phi\circ\psi)^{*}=\psi^{*}\circ\phi^{*}$ for all $\psi\in\mathrm{Hom}(X,Y),\phi\in\mathrm{Hom}(Y,Z)$;
-  3. (iii)
-    The function $\|\cdot\|:\mathrm{Hom}(X,Y)\rightarrow[0,\infty]$ given by
-    $\|\phi\|^{2}=\sup\{\lambda>0:\phi^{*}\circ\phi-\lambda\mathrm{id}_{X}\mbox{ isnot invertible}\}$
-    is a complete norm on $\mathrm{Hom}(X,Y)$;
-  4. (iv)
-    $\|\phi\circ\psi\|\leq\|\phi\|\|\psi\|$ for all $\psi\in\mathrm{Hom}(X,Y),\phi\in\mathrm{Hom}(Y,Z)$;
-  5. (v)
-    $\|\phi^{*}\circ\phi\|=\|\phi\|^{2}$ for all $\phi\in\mathrm{Hom}(X,Y)$;
-  6. (vi)
-    For all $\phi\in\mathrm{Hom}(X,Y)$, $\phi^{*}\circ\phi$ is a positive element of the $\rm{C}^{*}$-algebra $\mathrm{Hom}(X,X)$.
+All categories in this section will be $\mathbb{C}$-linear, that is an additive category such that its space of morphisms $\mathrm{Hom}(X,Y)$ between any two objects $X,Y$ is a $\mathbb{C}$-vector space and the composition of morphisms yields a bilinear map $\mathrm{Hom}(X,Y)\times\mathrm{Hom}(Y,Z)\rightarrow\mathrm{Hom}(X,Z)$ for any $X,Y,Z\in\mathcal{C}$.
+We invite the reader to recall the definition of a $\mathrm{C}^{*}$-category from [20].
 Let $\mathcal{C}$ and $\mathcal{D}$ be $\mathrm{C}^{*}$-categories.
-A functor $F:\mathcal{C}\rightarrow D$ is called a $\mathrm{C}^{*}$-functor if the induced mappings $\mathrm{Hom}(X,Y)\mapsto\mathrm{Hom}(F(X),F(Y))$ are $\mathbb{C}$-linear and $*{}^{*}$-preserving.
+A functor $F:\mathcal{C}\rightarrow\mathcal{D}$ is called a $\mathrm{C}^{*}$-functor if the induced mappings $\mathrm{Hom}(X,Y)\to\mathrm{Hom}(F(X),F(Y))$ are $\mathbb{C}$-linear and -preserving.
 A natural transformation $\nu:F\rightarrow G$ between $\mathrm{C}^{*}$-functors is called an isometry if $\nu_{X}^{*}\nu_{X}=\mathrm{id}_{F(X)}$ for all $X\in\mathcal{C}$.
 Moreover, $\nu$ is called a unitary if it is a surjective isometry in which case $\nu_{X}\nu_{X}^{*}=\mathrm{id}_{G(X)}$ for all $X\in\mathcal{C}$.
-We are interested in $*{}^{*}$-categories that admit tensor product structures.
+We are interested in C-categories that admit tensor product structures.
 Definition 1.13.
-(see for example [23]) A $\rm{C}^{*}$-tensor category is a $\rm{C}^{*}$-category $\mathcal{C}$ together with a $\mathbb{C}$-linear bifunctor $-\otimes-:\mathcal{C}\times\mathcal{C}\rightarrow\mathcal{C}$, a distinguished object $1_{\mathcal{C}}\in\mathcal{C}$ and unitary natural isomorphisms
+(see for example [24]) A $\rm{C}^{*}$-tensor category is a $\rm{C}^{*}$-category $\mathcal{C}$ together with a $\mathbb{C}$-linear bifunctor $\otimes-:\mathcal{C}\times\mathcal{C}\rightarrow\mathcal{C}$, a distinguished object $1_{\mathcal{C}}\in\mathcal{C}$ and unitary natural isomorphisms
 (1.2) $\alpha_{X,Y,Z}$ $:(X\otimes Y)\otimes Z\rightarrow X\otimes(Y\otimes Z),$
@@
 such that $(\phi\otimes\psi)^{*}=(\phi^{*}\otimes\psi^{*})$ and the following diagrams commute for any $X,Y,Z,W\in\mathcal{C}$
-[…]
-[…]
+(1.3) $((W\otimes X)\otimes Y)\otimes Z$$(W\otimes(X\otimes Y))\otimes Z$$(W\otimes X)\otimes(Y\otimes Z)$$W\otimes((X\otimes Y)\otimes Z)$$W\otimes(X\otimes(Y\otimes Z)),$$\scriptstyle{\alpha_{W,X,Y}\otimes\operatorname{id_{Z}}}$$\scriptstyle{\alpha_{W\otimes X,Y,Z}}$$\scriptstyle{\alpha_{W,X\otimes Y,Z}}$$\scriptstyle{\alpha_{W,X,Y\otimes Z}}$$\scriptstyle{\operatorname{id_{W}}\otimes\alpha_{X,Y,Z}}$
+(1.4) $(X\otimes 1_{\mathcal{C}})\otimes Y$$X\otimes(1_{\mathcal{C}}\otimes Y)$$X\otimes Y.$$\scriptstyle{\alpha_{X,1,Y}}$$\scriptstyle{\rho_{X}\otimes\mathrm{id}_{Y}}$$\scriptstyle{\mathrm{id}_{X}\otimes\lambda_{Y}}$
 and $\mathrm{Hom}(1_{\mathcal{C}},1_{\mathcal{C}})\cong\mathbb{C}$.
-Moreover, $\mathcal{C}$ is said to be semisimple if there exists a collection of objects $X_{i}\in\mathcal{C}$ with $\mathrm{Hom}(X_{i},X_{j})\cong\delta_{ij}\mathbb{C}$ and any object $X\in\mathcal{C}$ can be decomposed uniquely as a finite direct sum $X\cong\bigoplus_{i}X_{i}$.
-We call the structure morphism $\alpha$ as in Definition 1.15 associated to a $*{}^{*}$-tensor category its associator and the maps $\lambda$ and $\rho$ the unitors.
+Moreover, $\mathcal{C}$ is said to be semisimple if there exists a collection of objects $X_{i}\in\mathcal{C}$ with $\mathrm{Hom}(X_{i},X_{j})\cong\delta_{ij}\mathbb{C}$ and any object $X\in\mathcal{C}$ can be decomposed uniquely as a finite direct sum $X\cong\bigoplus_{i}X_{i}^{\oplus m_{i}}$, where $m_{i}$ is the multiplicity of each irreducible $X_{i}$.
+We call the structure morphism $\alpha$ as in Definition 1.13 associated to a C-tensor category its associator and the maps $\lambda$ and $\rho$ the unitors.
 We call an object $X\in\mathcal{C}$ such that $\mathrm{Hom}(X,X)\cong\mathbb{C}$ irreducible or simple.
-For a $*{}^{*}$-tensor category $\mathcal{C}$, we will denote by $\mathrm{Irr}(\mathcal{C})$ a collection of isomorphism class representatives for simple objects in $\mathcal{C}$.
+For a C-tensor category $\mathcal{C}$, we will denote by $\mathrm{Irr}(\mathcal{C})$ a collection of isomorphism class representatives for simple objects in $\mathcal{C}$.
 Example 1.14.
-Let $\Gamma$ be a countable, discrete group (or more generally a countable, discrete monoid).
-We denote by $\mathrm{Hilb}(\Gamma)$ the semisimple $*{}^{*}$-tensor category whose objects are finite-dimensional $\Gamma$-graded Hilbert spaces, i.e.
+Let $\Gamma$ be a countable discrete group (or more generally a countable, discrete monoid).
+We denote by $\mathrm{Hilb}(\Gamma)$ the semisimple C-tensor category whose objects are finite-dimensional $\Gamma$-graded Hilbert spaces, i.e.
 finite-dimensional Hilbert spaces $\mathcal{H}$ endowed with a decomposition $\mathcal{H}=\bigoplus_{g\in\Gamma}\mathcal{H}_{g}$.
@@
 $(\mathbb{C}_{g})_{h}=\begin{cases}\mathbb{C}\ \text{if}\ g=h,\\ 0\ \text{otherwise}.\end{cases}$
-If $\Gamma$ is a group and $\omega\in Z^{3}(\Gamma,\mathbb{T})$ is a normalised $3$-cocycle, the category $\mathrm{Hilb}(\Gamma,\omega)$ is defined exactly as is $\mathrm{Hilb}(\Gamma)$ but with associators now given by
+If $\Gamma$ is a countable discrete group and $\omega\in Z^{3}(\Gamma,\mathbb{T})$ is a normalised 33-cocycle (see [4, Section III.1] for definitions), the category $\mathrm{Hilb}(\Gamma,\omega)$ is defined exactly as is $\mathrm{Hilb}(\Gamma)$ but with associators now given by
 $\alpha_{\mathcal{H},\mathcal{K},\mathcal{L}}:(\xi\boxtimes\eta)\boxtimes\mu\mapsto\omega(g,h,k)\xi\boxtimes(\eta\boxtimes\mu)$
 for $\xi\in\mathcal{H}_{g},\eta\in\mathcal{K}_{h}$ and $\mu\in\mathcal{H}_{k}$.
-For the remainder of this paper we will only make use of this example when $\Gamma$ is a group. However, our results also apply to the case of monoids.
+For the remainder of this paper we will only make use of this example when $\Gamma$ is a countable discrete group. However, our results also apply to the case of monoids.
 We now introduce the most important example for our purposes.
 Example 1.15.
-Let $A$ be a $*{}^{*}$-algebra.
-We denote by $\mathrm{Corr}_{0}(A)$ the $*{}^{*}$-tensor category whose objects are non-degenerate $A$-$A$ correspondences and whose morphisms are adjointable intertwiners of the underlying Hilbert bimodules.
-The tensor product of two $A$-$A$ correspondences $\varphi$ with $\psi$ is given by their composition $\varphi\circ\psi$.
+Let AA be a C-algebra.
+We denote by $\mathrm{Corr}_{0}(A)$ the C-tensor category whose objects are non-degenerate AA-AA-correspondences and whose morphisms are adjointable bimodule maps between the underlying Hilbert bimodules.
+The tensor product of two AA-AA-correspondences $\varphi$ with $\psi$ is given by their composition $\varphi\circ\psi$.
 The tensor identity of $\mathrm{Corr}_{0}(A)$ is given by the identity homomorphism $\mathrm{id}_{A}$, the associator is given by the rebracketing morphism associated to the underlying tensor product of Hilbert bimodules.
 Remark 1.16.
-In general, $\mathrm{Corr}(A)$ is not a $*{}^{*}$-tensor category as there is no tensor unit on degenerate correspondences; $A$ does not act as a unit on a degenerate correspondence.
-It is a non-unital $*{}^{*}$-tensor category, this is a weakening of Definition 1.15 which omits the necessity of a tensor unit.
+In general, $\mathrm{Corr}(A)$ is not a C-tensor category as there is no tensor unit on degenerate correspondences; AA does not act as a unit on a degenerate correspondence.
+It is a non-unital C-tensor category; this is a weakening of Definition 1.13 which omits the necessity of a tensor unit.
 1.4. Szabó’s cocycle category
 Before we begin our discussion on actions of $\mathrm{C}^{*}$-tensor categories, we recall the case of a twisted action by a second-countable locally compact group $\Gamma$.
-We shall later identify this, in the case when $\Gamma$ is countable discrete, with actions of the $*{}^{*}$- tensor category $\mathrm{Hilb}(\Gamma)$.
-Definition 1.19 (cf. [40, Definition 1.1]).
-Let $\Gamma$ be a locally compact group and $A$ be a $\mathrm{C}^{*}$-algebra.
-A twisted action of $\Gamma$ on $A$ is a pair $(\alpha,\mathfrak{u})$, where $\alpha:\Gamma\to\mathrm{Aut}(A)$ is a point-norm continuous map, and $\mathfrak{u}:\Gamma\times\Gamma\to\mathcal{U}(\mathcal{M}(A))$ is a strictly continuous map satisfying
-(1.5) $\alpha_{1}=\mathrm{id}_{A},\quad\mathrm{Ad}(\mathfrak{u}_{s,t})\circ\alpha_{s}\circ\alpha_{t}=\alpha_{st}$
+We shall later identify this, in the case when $\Gamma$ is countable discrete, with actions of the C- tensor category $\mathrm{Hilb}(\Gamma)$.
+Definition 1.17 (cf. [41, Definition 1.1]).
+Let $\Gamma$ be a locally compact group and AA be a $\mathrm{C}^{*}$-algebra.
+A twisted action of $\Gamma$ on AA is a pair $(\alpha,\mathfrak{u})$, where $\alpha:\Gamma\to\mathrm{Aut}(A)$ is a point-norm continuous map, and $\mathfrak{u}:\Gamma\times\Gamma\to\mathcal{U}(\mathcal{M}(A))$ is a strictly continuous map satisfying
+(1.5) $\alpha_{1}=\mathrm{id}_{A},\quad\mathrm{Ad}(\mathfrak{u}_{g,h})\circ\alpha_{g}\circ\alpha_{h}=\alpha_{gh}$
 and
-(1.6) $\mathfrak{u}_{s,1}=\mathfrak{u}_{1,s}=1,\quad\mathfrak{u}_{r,st}\alpha_{r}(\mathfrak{u}_{s,t})=\mathfrak{u}_{rs,t}\mathfrak{u}_{r,s}$
-for all $r,s,t\in\Gamma$.
+(1.6) $\mathfrak{u}_{g,1}=\mathfrak{u}_{1,g}=1,\quad\mathfrak{u}_{k,gh}\alpha_{k}(\mathfrak{u}_{g,h})=\mathfrak{u}_{kg,h}\mathfrak{u}_{k,g}$
+for all $k,g,h\in\Gamma$.
 Remark 1.18.
 Note that the formulae above differ slightly from the definition of a twisted action in [41].
-In fact, the sole difference is that our unitary cocycles $\mathfrak{u}_{s,t}$ for $s,t\in\Gamma$ are the adjoints of the cocycles in [40].
+In fact, the sole difference is that our unitary cocycles $\mathfrak{u}_{g,h}$ for $g,h\in\Gamma$ are the adjoints of the cocycles in [41].
 The reason for this change of conventions will be discussed in Section 2.
@@
 If the cocycle is trivial i.e.
-$\mathfrak{u}_{s,t}=\mathfrak{1}$ for all $s,t\in\Gamma$, then $(A,\alpha)$ is said to be a $\Gamma$-$\mathrm{C}^{*}$-algebra.
-Definition 1.21 (cf. [40, Definition 1.10]).
-Let $(\alpha,\mathfrak{u}):\Gamma\curvearrowright A$ and $(\beta,\mathfrak{v}):\Gamma\curvearrowright B$ be two twisted actions on $\mathrm{C}^{*}$-algebras $A$ and $B$ respectively.
+$\mathfrak{u}_{g,h}=\mathfrak{1}$ for all $g,h\in\Gamma$, then $(A,\alpha)$ is said to be a $\Gamma$-$\mathrm{C}^{*}$-algebra.
+Definition 1.19 (cf. [41, Definition 1.10]).
+Let $(\alpha,\mathfrak{u}):\Gamma\curvearrowright A$ and $(\beta,\mathfrak{v}):\Gamma\curvearrowright B$ be two twisted actions on $\mathrm{C}^{*}$-algebras AA and $B$ respectively.
   1. (i)
-    A cocycle representation from $(A,\alpha,\mathfrak{u})$ to $(B,\beta,\mathfrak{v})$ is a pair $(\psi,\mathbbm{v})$, where $\psi:A\to\mathcal{M}(B)$ is an extendible $*{}^{*}$-homomorphism and $\mathbbm{v}:\Gamma\to\mathcal{U}(\mathcal{M}(B))$ is a strictly continuous map such that
+    A cocycle representation from $(A,\alpha,\mathfrak{u})$ to $(B,\beta,\mathfrak{v})$ is a pair $(\psi,\mathbbm{v})$, where $\psi:A\to\mathcal{M}(B)$ is an extendible -homomorphism and $\mathbbm{v}:\Gamma\to\mathcal{U}(\mathcal{M}(B))$ is a strictly continuous map such that
     (1.7) $\beta_{g}\circ\psi=\mathrm{Ad}(\mathbbm{v}_{g})\circ\psi\circ\alpha_{g}$
@@
 Due to our change of conventions when defining twisted actions we need to change the definition of a cocycle representation.
-The formula in (1.8) differs from [40, Definition 1.10] precisely by taking adjoints.
-As shown in [40], there exists a category with objects being twisted $\Gamma$-$\mathrm{C}^{*}$-algebras and morphisms being cocycle morphisms with composition defined in [40, Proposition 1.15].
-This category is denoted by $\mathrm{C}^{*}_{\Gamma,t}$ ([40, Definition 1.16]).
+The formula in (1.8) differs from [41, Definition 1.10] precisely by taking adjoints.
+As shown in [41], there exists a category with objects being twisted $\Gamma$-$\mathrm{C}^{*}$-algebras and morphisms being cocycle morphisms with composition defined in [41, Proposition 1.15].
+This category is denoted by $\mathrm{C}^{*}_{\Gamma,t}$ ([41, Definition 1.16]).
 Our attention will now focus on generalising this construction to actions of semisimple $\mathrm{C}^{*}$-tensor categories.
 2. Actions of $\mathrm{C}^{*}$-tensor categories
-For $\mathrm{C}^{*}$-tensor categories $(\mathcal{C},\otimes)$ and $(\mathcal{D},\boxtimes)$, $F:\mathcal{C}\to\mathcal{D}$ is said to be a $\mathrm{C}^{*}$-tensor functor if it is a $\mathrm{C}^{*}$-functor such that $F(1_{\mathcal{C}})=1_{\mathcal{D}}$ there exists a unitary natural isomorphism $J_{X,Y}:F(X)\boxtimes F(Y)\to F(X\otimes Y)$ such that
-[…]
+For $\mathrm{C}^{*}$-tensor categories $(\mathcal{C},\otimes)$ and $(\mathcal{D},\boxtimes)$, $F:\mathcal{C}\to\mathcal{D}$ is said to be a $\mathrm{C}^{*}$-tensor functor if it is a $\mathrm{C}^{*}$-functor such that $F(1_{\mathcal{C}})=1_{\mathcal{D}}$ and there exists a unitary natural isomorphism $J_{X,Y}:F(X)\boxtimes F(Y)\to F(X\otimes Y)$ such that
+(2.1) $(F(X)\boxtimes F(Y))\boxtimes F(Z)$$F(X)\boxtimes(F(Y)\boxtimes F(Z))$$F(X\otimes Y)\boxtimes F(Z)$$F(X)\boxtimes F(Y\otimes Z)$$F((X\otimes Y)\otimes Z)$$F(X\otimes(Y\otimes Z))$$\scriptstyle{\alpha_{F(X),F(Y),F(Z)}}$$\scriptstyle{J_{X,Y}\boxtimes\mathrm{id}_{F(Z)}}$$\scriptstyle{\mathrm{id}_{F(X)}\boxtimes J_{Y,Z}}$$\scriptstyle{J_{X\otimes Y,Z}}$$\scriptstyle{J_{X,Y\otimes Z}}$$\scriptstyle{F(\alpha_{X,Y,Z})}$
 commutes for all $X,Y,Z\in\mathcal{C}$.
 In the following definition we denote by $\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$ the full subcategory of $\mathrm{Corr}_{0}(A)$ consisting of bimodules with countable dense subsets.
+Moreover, for a $\mathrm{C}^{*}$-tensor category $\mathcal{C}$, we denote by $\mathcal{C}^{\mathrm{rev}}$ the $\mathrm{C}^{*}$-tensor category whose underlying category is $\mathcal{C}$, but the tensor product is reversed i.e.
+$X\otimes^{\mathrm{rev}}Y=Y\otimes X$.
 Definition 2.1.
-[…]
+A $\mathrm{C}^{*}$-tensor category $\mathcal{C}$ is said to act on a $\mathrm{C}^{*}$-algebra AA if there exists a $\mathrm{C}^{*}$-tensor functor $F:\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}(A)$.
+If AA is separable we further impose that $F$ is valued in $\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$.
+We will often denote this by $\mathcal{C}\overset{F}{\curvearrowright}A$ or by the triple $(A,F,J)$, where
+$J=\{J_{X,Y}:F(Y)\boxtimes F(X)\to F(X\otimes Y)\}_{X,Y\in\mathcal{C}}$
+is the natural isomorphism associated with the functor $F$. In this case, we say that the triple $(A,F,J)$ is a $\mathcal{C}$-$\mathrm{C}^{*}$-algebra.
+Remark $2.$
+Often in the literature an action of $\mathcal{C}$ on AA is given instead by a C-tensor functor $F:\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$.
+We choose to define it as a functor from $\mathcal{C}^{\mathrm{rev}}$ so that an action of a countable discrete group $\Gamma$ induces an action of the category $\mathrm{Hilb}(\Gamma)$ (see Example 2.4).
 Remark 2.3.
-In the literature, the main interest is actions of unitary tensor categories on $*{}^{*}$-algebras (see e.g.
+In the literature, the main interest is actions of unitary tensor categories on C-algebras (see e.g.
 [1, 7, 24, 26, 43, 35]).
-This is because unitary tensor categories axiomatise the standard invariant in subfactor theory and can be thought of as the mathematical objects encoding the symmetry in finite index inclusions of $*{}^{*}$-algebras.
-If $\mathcal{C}$ is a unitary tensor category, $A$ is separable, and $(F,J):\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$ is a $*{}^{*}$-tensor functor, then the bimodule associated to the correspondence $F(X)$ for any $X\in\mathcal{C}$ has a countable dense subset by [25, Corollary 2.24] and hence $(F,J)$ automatically falls into $\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$.
+This is because unitary tensor categories axiomatise the standard invariant in subfactor theory and can be thought of as the mathematical objects encoding the symmetry in finite index inclusions of C-algebras.
+If $\mathcal{C}$ is a unitary tensor category, AA is separable, and $(F,J):\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$ is a C-tensor functor, then the bimodule associated to the correspondence $F(X)$ for any $X\in\mathcal{C}$ is of finite index and hence it has a countable dense subset by [26, Corollary 2.24].
+Therefore, $(F,J)$ automatically falls into $\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$.
 Example 2.4.
 Let $\Gamma$ be a a countable discrete group and let $(\alpha,\mathfrak{u}):\Gamma\rightarrow\mathrm{Aut}(A)$ be a twisted action.
-The pair $(\alpha,\mathfrak{u})$ will induce a $*{}^{*}$-tensor functor $(\alpha,\mathfrak{u}):\mathrm{Hilb}(\Gamma)\rightarrow\mathrm{Corr}_{0}(A)$ by setting
+The pair $(\alpha,\mathfrak{u})$ will induce a C-tensor functor $(\alpha,\mathfrak{u}):\mathrm{Hilb}(\Gamma)^{\mathrm{rev}}\rightarrow\mathrm{Corr}_{0}(A)$ by setting
 (2.2) $\alpha(\mathbb{C}_{g})={}_{\alpha_{g}}A,$
-(2.3) $\mathfrak{u}_{\mathbb{C}_{g},\mathbb{C}_{h}}(a\boxtimes b)=\mathfrak{u}_{g,h}^{*}\alpha_{g}(a)b.$
-The functor may then be extended by linearity to all of $\mathrm{Hilb}(\Gamma)$ in a similar manner to [14, Proposition 5.6].
-In general, actions of $\mathrm{Hilb}(\Gamma)$ on $A$ correspond to twisted actions of $\Gamma$ on $A\otimes\mathbb{K}$.
-If $A$ is a $\mathrm{C}^{*}$-algebra, then its sequence algebra $A_{\infty}$ is defined by
+(2.3) $\mathfrak{u}_{\mathbb{C}_{g},\mathbb{C}_{h}}(a\boxtimes b)=\mathfrak{u}_{g,h}\alpha_{g}(a)b.$
+The functor may then be extended by linearity to all of $\mathrm{Hilb}(\Gamma)^{\mathrm{rev}}$ in a similar manner to [16, Proposition 5.6].
+In general, actions of $\mathrm{Hilb}(\Gamma)$ on AA correspond to twisted actions of $\Gamma$ on $A\otimes\mathbb{K}$.
+If AA is a $\mathrm{C}^{*}$-algebra, then its sequence algebra $A_{\infty}$ is defined by
 $A_{\infty}=\ell^{\infty}(\mathbb{N},A)/\Big\{(a_{n})_{n\geq 1}:\lim\limits_{n\to\infty}\|a_{n}\|=0\Big\}.$
-[…]
-For any $X\in\mathcal{C}$, we can view $F(X)$ as a non-degenerate Hilbert $A$-bimodule. Define
+We will end this section by showing that if $\mathcal{C}\overset{F}{\curvearrowright}A$ is an action on a separable $\mathrm{C}^{*}$-algebra AA, then we can induce an action of $\mathcal{C}$ on its sequence algebra $A_{\infty}$ whenever the image bimodules are finite rank in the following sense.
+Definition 2.5.
+We call a right Hilbert module $E$ over a C-algebra AA finite rank if there exist $n\in\mathbb{N}$ and a projection $p\in M_{n}(\mathcal{M}(A))$ such that $E\cong pA^{n}$.
+In fact, $E$ is finite rank if and only if there exists a collection of elements $\xi_{i}\in\mathcal{L}(A,E)$ for $1\leq i\leq n$ such that
+(2.4) $\sum_{i=1}^{n}\xi_{i}\xi_{i}^{*}=\mathrm{id}_{E}.$
+We call $\xi_{i}$ a basis for $E$. We say that $(F,J)$ is finite rank if each $F(X)$ for $X\in\mathrm{Irr}(\mathcal{C})$ is finite rank.
+Suppose that $F:\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$ is a $\mathrm{C}^{*}$-tensor functor.
+Our goal is to build a $\mathrm{C}^{*}$-tensor functor $F_{\infty}:\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}(A_{\infty})$.
+For any $X\in\mathcal{C}$, we can view $F(X)$ as a non-degenerate Hilbert AA-bimodule. Define
 (2.5) $F_{\infty}(X)=\ell^{\infty}(\mathbb{N},F(X))/\Big\{(\xi_{n})_{n\geq 1}:\lim_{n\to\infty}\|\xi_{n}\|=0\Big\},$
@@
 Precisely, for any $a\in A_{\infty}$ and any $\xi\in F_{\infty}(X)$ represented by $(a_{n})_{n\geq 1}$ and $(\xi_{n})_{n\geq 1}$, we let
-(2.5) $a\rhd\xi=(a_{n}\rhd\xi_{n})_{n\geq 1}\quad\mathrm{and}\quad\xi\lhd a=(\xi_{n}\lhd a_{n})_{n\geq 1}.$
+[…]
 Similarly, for any $\xi,\eta\in F_{\infty}(X)$ we define
@@
 $\|\langle\xi_{n}\lhd(a_{n}-a_{n}^{\prime}),\xi_{n}\lhd(a_{n}-a_{n}^{\prime})\rangle_{A}\|\leq\|\langle\xi_{n},\xi_{n}\rangle_{A}\|\|a_{n}-a_{n}^{\prime}\|^{2},$
-which converges to 00 as $n\to\infty$.
+which converges to 0 as $n\to\infty$.
 Exactly the same calculation shows that if $(\xi_{n})_{n\geq 1}$ and $(\xi_{n}^{\prime})_{n\geq 1}$ induce the same element in $F_{\infty}(X)$ and $(a_{n})_{n\geq 1}\in A_{\infty}$, then $(\xi_{n}\lhd a_{n})_{n\geq 1}=(\xi_{n}^{\prime}\lhd a_{n})_{n\geq 1}$ as elements in $F_{\infty}(X)$.
@@
 $\langle\xi_{n},\eta_{n}\rangle_{A}-\langle\xi_{n}^{\prime},\eta_{n}^{\prime}\rangle_{A}=\langle\xi_{n}-\xi_{n}^{\prime},\eta_{n}\rangle_{A}+\langle\xi_{n}^{\prime},\eta_{n}-\eta_{n}^{\prime}\rangle_{A},$
-which converges to 00 by Cauchy-Schwarz.
-Moreover, the sequences $\big{(}\langle\xi_{n},\eta_{n}\rangle_{A}\big{)}_{n\geq 1}$ and $\big{(}\langle\xi_{n}^{\prime},\eta_{n}^{\prime}\rangle_{A}\big{)}_{n\geq 1}$ are bounded, so induce elements in $A_{\infty}$.
+which converges to 0 by Cauchy-Schwarz.
+Moreover, the sequences $\big(\langle\xi_{n},\eta_{n}\rangle_{A}\big)_{n\geq 1}$ and $\big(\langle\xi_{n}^{\prime},\eta_{n}^{\prime}\rangle_{A}\big)_{n\geq 1}$ are bounded, so induce the same element in $A_{\infty}$.
 Therefore, the function $\langle\cdot,\cdot\rangle_{A_{\infty}}:F_{\infty}(X)\times F_{\infty}(X)\to A_{\infty}$ is well-defined.
-It is now straightforward to check that this function is right linear, left conjugate linear, and antisymmetric since all this properties are satisfied pointwise for each $n\in\mathbb{N}$.
-Finally, it is clear that $\langle\xi,\xi\rangle_{A_{\infty}}\geq 0$ and that $\langle\xi,\xi\rangle_{A_{\infty}}=0$ if and only if $\langle\xi_{n},\xi_{n}\rangle_{A}$ converges to 00 i.e.
+It is now straightforward to check that this function is right linear, left conjugate linear, and antisymmetric since all these properties are satisfied pointwise for each $n\in\mathbb{N}$.
+Finally, it is clear that $\langle\xi,\xi\rangle_{A_{\infty}}\geq 0$ and that $\langle\xi,\xi\rangle_{A_{\infty}}=0$ if and only if $\langle\xi_{n},\xi_{n}\rangle_{A}$ converges to 0 i.e.
 $\xi=0$ in $F_{\infty}(X)$.
-Since, $F_{\infty}(X)$ is complete with respect to the inner product defined in (2.6), it is a right-Hilbert $A_{\infty}$-bimodule.
+Since, $F_{\infty}(X)$ is complete with respect to the inner product defined in (2.7) (as a quotient of a complete space by a closed subspace), it is a right-Hilbert $A_{\infty}$-bimodule.
 ∎
@@
 As in the proof of Lemma 2.6 we have that (2.6) gives a well-defined left action by $A_{\infty}$.
-Since the left action of $A$ on $F(X)$ is adjointable, the left action of $(a_{n})_{n\geq 1}\in A_{\infty}$ is an adjointable operator with adjoint given by the action of $(a_{n}^{*})_{n\geq 1}$.
+Since the left action of AA on $F(X)$ is adjointable, the left action of $(a_{n})_{n\geq 1}\in A_{\infty}$ is an adjointable operator with adjoint given by the action of $(a_{n}^{*})_{n\geq 1}$.
 Moreover, $F_{\infty}(X)$ is non-degenerate for any $X\in\mathcal{C}$ as $F(X)$ is.
 ∎
-Consider the functor $F_{\infty}:\mathcal{C}\to\mathrm{Corr}_{0}(A_{\infty})$ defined by sending any $X\in\mathcal{C}$ to the correspondence induced by the Hilbert $A_{\infty}$-bimodule $F_{\infty}(X)$, and any $f\in\mathrm{Hom}(X,Y)$ for $X,Y\in\mathcal{C}$ to the intertwiner defined on $F_{\infty}(X)$ by $F_{\infty}(f)((\xi_{n})_{n\geq 1})=(F(f)(\xi_{n}))_{n\geq 1}$.
+Consider the functor $F_{\infty}:\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}(A_{\infty})$ defined by sending any $X\in\mathcal{C}$ to the correspondence induced by the Hilbert $A_{\infty}$-bimodule $F_{\infty}(X)$, and any $f\in\mathrm{Hom}(X,Y)$ for $X,Y\in\mathcal{C}$ to the intertwiner defined on $F_{\infty}(X)$ by $F_{\infty}(f)((\xi_{n})_{n\geq 1})=(F(f)(\xi_{n}))_{n\geq 1}$.
 Also, one can define $J_{X,Y}^{\infty}:F_{\infty}(Y)\boxtimes F_{\infty}(X)\to F_{\infty}(X\otimes Y)$ as the unique continuous extension of the map
-$J_{X,Y}^{\infty}(\xi\boxtimes\eta)=J_{X,Y}(\xi_{n}\boxtimes\eta_{n}),$
+$J_{X,Y}^{\infty}(\xi\boxtimes\eta)=(J_{X,Y}(\xi_{n}\boxtimes\eta_{n}))_{n},$
 for any $\xi\in F_{\infty}(Y)$, any $\eta\in F_{\infty}(X)$ (that this is well defined follows in a similar fashion as the arguments in the proof of Lemma 2.6).
@@
 Lemma 2.8.
-Let $F:\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$ be an action of a $*{}^{*}$-tensor category $\mathcal{C}$ on a unital $*{}^{*}$-algebra $A$ such that $F(X)$ is algebraically finitely generated as a right $A$-module for all $X\in\mathcal{C}$.
-Then the map $F_{\infty}:\mathcal{C}\to\mathrm{Corr}_{0}(A_{\infty})$ defined above is a $C^{*}$-tensor functor and so $\mathcal{C}$ acts on $A_{\infty}$ via the triple $(A_{\infty},F_{\infty},J^{\infty})$.
+Let $(F,J,A)$ be a finite rank action of a C-tensor category $\mathcal{C}$. Then $\mathcal{C}$ acts on $A_{\infty}$ via the triple $(A_{\infty},F_{\infty},J^{\infty})$.
 Proof.
-It follows from construction that $F_{\infty}$ is a $*{}^{*}$-functor.
+It follows from construction that $F_{\infty}$ is a C-functor.
 Moreover, the naturality of $J^{\infty}$ and that $J^{\infty}$ satisfies commuting diagrams as in (2.1) follows from direct computations.
@@
 It remains to show the surjectivity of $\iota_{X,Y}$.
-As each $F(X)$ is algebraically finitely generated, then $F(X)$ is also projective ([44, Corollary 15.4.8]).
-Equivalently for each $X\in\mathcal{C}$ the bimodule $F(X)$ has a right $A$-basis, i.e.
-there exist a finite set $I_{X}$ and $u_{i}^{X}\in F(X)$ for $i\in I_{X}$ such that $\sum_{i\in I_{X}}u_{i}^{X}\lhd\langle u_{i}^{X},\xi\rangle_{A}=\xi$ for all $\xi\in F(X)$ (see for example [26, Lemma 1.11]).
-Similarly $(F(Y)\boxtimes F(X))_{\infty}$ is also finitely generated projective as a right $A_{\infty}$-module with right $A_{\infty}$-basis $u_{i}^{Y}\boxtimes u_{j}^{X}\in(F(Y)\boxtimes F(X))_{\infty}$ for $i\in I_{Y}$ and $j\in I_{X}$.
-As $\iota_{X,Y}(u_{i}^{Y}\boxtimes u_{j}^{X})=u_{i}^{Y}\boxtimes u_{j}^{X}$ and $u_{i}^{Y}\boxtimes u_{j}^{X}$ algebraically generate $(F(Y)\boxtimes F(X))_{\infty}$ as a right $A_{\infty}$-module it follows that $\iota_{X,Y}$ is surjective.
-∎
+Let $X,Y\in\mathcal{C}$ and let $\xi_{i}^{Y}$ with $i\in I_{Y}$ finite such that
+$\sum_{i\in I_{Y}}\xi_{i}^{Y}{\xi_{i}^{Y}}^{*}=\mathrm{id}_{F(Y)}.$
+Any $\zeta\in(F(Y)\boxtimes F(X))_{\infty}$ is represented by a sequence $(\sum_{l=1}^{k_{n}}y_{l}^{(n)}\boxtimes x_{l}^{(n)})_{n}$.
+By a standard reindexation one can choose an approximate unit $e_{n}$ for AA such that
+${e_{n}\xi_{i}^{Y}}^{*}(y_{l}^{(n)})-{\xi_{i}^{Y}}^{*}(y_{l}^{(n)})\rightarrow 0$
+as $n$ tends to infinity for all $i\in I_{Y}$, $1\leq l\leq k_{n}$. Then one has that
+$\zeta$ $=\left(\sum_{l=1}^{k_{n}}y_{l}^{(n)}\boxtimes x_{l}^{(n)}\right)_{n}$
+$=\left(\sum_{l=1}^{k_{n}}\sum_{i\in I_{Y}}\xi_{i}^{Y}(e_{n}{\xi_{i}^{Y}}^{*}(y_{l}^{(n)}))\boxtimes x_{l}^{(n)}\right)_{n}$
+$=\left(\sum_{i\in I_{Y}}\xi_{i}^{Y}(e_{n})\boxtimes\left(\sum_{l=1}^{k_{n}}{\xi_{i}^{Y}}^{*}(y_{l}^{(n)})x_{l}^{(n)}\right)\right)_{n}.$
+Hence, it suffices to show that
+$\sum_{l=1}^{k_{n}}{\xi_{i}^{Y}}^{*}(y_{l}^{(n)})x_{l}^{(n)}$
+is a bounded sequence. This follows precisely as $\zeta$ is a bounded sequence. Indeed,
+$\lVert\sum_{l=1}^{k_{n}}{\xi_{i}^{Y}}^{*}(y_{l}^{(n)})x_{l}^{(n)}\rVert^{2}$ $=\|\sum_{l,l^{\prime}}\langle x_{l}^{(n)},({\xi_{i}^{Y}}^{*}(y_{l}^{(n)}))^{*}{\xi_{i}^{Y}}^{*}(y_{l^{\prime}}^{(n)})x_{l^{\prime}}^{(n)}\rangle\|$
+$=\|\sum_{l,l^{\prime}}\langle x_{l}^{(n)},({y_{l}^{(n)}}^{*}\xi_{i}^{Y}{\xi_{i}^{Y}}^{*}\boxtimes\mathrm{id}_{F(X)})(y_{l^{\prime}}^{(n)}\boxtimes{x_{l^{\prime}}}^{(n)})\rangle\|$
+$\leq\|\sum_{l,l^{\prime}}\langle x_{l}^{(n)},({y_{l}^{(n)}}^{*}\boxtimes\mathrm{id}_{F(X)})(y_{l^{\prime}}^{(n)}\boxtimes{x_{l^{\prime}}}^{(n)})\rangle\|$
+$=\|\sum_{l,l^{\prime}}\langle x_{l}^{(n)},\langle{y_{l}^{(n)}},y_{l^{\prime}}^{(n)}\rangle x_{l^{\prime}}^{(n)}\rangle\|$
+$=\|\zeta_{n}\|^{2}$
+and $\zeta_{n}$ is a bounded sequence. ∎
 Remark 2.9.
-If $F:\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$ is an action of a unitary tensor category (in the sense of [6] for example) on a unital $*{}^{*}$-algebra $A$, then it follows from [25, 26] that $F(X)$ is algebraically finitely generated as a right $A$-module for each $X\in\mathcal{C}$.
+If $\mathcal{C}\overset{F}{\curvearrowright}A$ is an action of a unitary tensor category (in the sense of [7] for example) on a unital C-algebra AA, then it follows from [26, 27] that $F$ is finite rank.
 3. The generalised cocycle category
@@
 Throughout the rest of this paper $\mathcal{C}$ is always assumed to be a $\mathrm{C}^{*}$-tensor category.
-Recall the definition of an action of $\mathcal{C}$ on a $\mathrm{C}^{*}$-algebra $A$ from Definition 2.1.
 Definition 3.1.
-[…]
+Let $\mathcal{C}\overset{F}{\curvearrowright}A$ and $\mathcal{C}\overset{G}{\curvearrowright}B$ be actions of $\mathcal{C}$ on $\mathrm{C}^{*}$-algebras AA and $B$.
   1. (i)
-    A correspondence morphism from $(A,F,J)$ to $(B,G,I)$ is a pair $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}})$, where $\phi:A\to\mathcal{L}(E)$ is an $A$-$B$ correspondence and $\{\mathbbm{v}_{X}:\phi\circ F(X)\to G(X)\circ\phi\}_{X\in\mathcal{C}}$ is a natural family of $A$-$B$-bimodule maps such that $\mathbbm{v}_{X}$ is an isometry (not necessarily adjointable) for any $X\in\mathcal{C}$ [^6] By an isometry, we mean a map which preserves the norm.
-By the polarisation identity, it is equivalent to assume that it preserves the inner product.
+    A correspondence morphism from $(A,F,J)$ to $(B,G,I)$ is a pair $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}})$, where $\phi:A\to\mathcal{L}(E)$ is an AA-$B$-correspondence and $\{\mathbbm{v}_{X}:\phi\circ F(X)\to G(X)\circ\phi\}_{X\in\mathcal{C}}$ is a natural family of AA-$B$-bimodule maps such that $\mathbbm{v}_{X}$ is an isometry (not necessarily adjointable) for any $X\in\mathcal{C}$.
+[^6] By an isometry, we mean a map which preserves the norm.
+By the proof of the theorem in [30], it is equivalent to assume that it preserves the inner product.
 Moreover, for all $X,Y\in\mathcal{C}$ the following pentagon diagram commutes
-    (3.1) {tikzcd}
+    (3.1) $\phi\circ F(X)\circ F(Y)$$G(X)\circ\phi\circ F(Y)$$\phi\circ F(X\otimes Y)$$G(X)\circ G(Y)\circ\phi$$G(X\otimes Y)\circ\phi,$$\scriptstyle{J_{X,Y}\boxtimes\mathrm{id}_{\phi}}$$\scriptstyle{\mathrm{id}_{F(Y)}\boxtimes\mathbbm{v}_{X}}$$\scriptstyle{\mathbbm{v}_{Y}\boxtimes\mathrm{id}_{G(X)}}$$\scriptstyle{\mathbbm{v}_{X\otimes Y}}$$\scriptstyle{\mathrm{id}_{\phi}\boxtimes I_{X,Y}}$
     and $\mathbbm{v}_{1_{\mathcal{C}}}:A\boxtimes{}_{\phi}E\to{}_{\phi}E\boxtimes B\cong{}_{\phi}E$ is given by $\mathbbm{v}_{1_{\mathcal{C}}}(a\boxtimes x)=\phi(a)x$ for any $a\in A$ and $x\in{}_{\phi}E$.
@@
   2. (ii)
-    A cocycle representation $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ is a correspondence morphism for which we further require that $\phi:A\to\mathcal{M}(B)$ is a $*{}^{*}$-homomorphism.
+    A cocycle representation $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ is a correspondence morphism for which we further require that $\phi:A\to\mathcal{M}(B)$ is a -homomorphism.
   3. (iii)
-    A cocycle morphism $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ is a cocycle representation for which we further require that $\phi:A\to B$ is a $*{}^{*}$-homomorphism.
+    A cocycle morphism $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ is a cocycle representation for which we further require that $\phi:A\to B$ is a -homomorphism.
 Remark 3.2.
@@
 In the case of group actions, Definition 3.1 recovers Szabó’s notion of a cocycle morphism (see Definition 1.19).
-Example 3.3.
-Suppose $(A,\alpha)$ and $(B,\beta)$ are actions of a group $\Gamma$ on $\mathrm{C}^{*}$-algebras $A$ and $B$.
+Example $3.$
+Suppose $(A,\alpha)$ and $(B,\beta)$ are actions of a countable discrete group $\Gamma$ on $\mathrm{C}^{*}$-algebras AA and $B$.
 Consider them as actions of $\mathrm{Hilb}(\Gamma)$ as in Example 2.4.
-Let $(\phi,\mathbbm{v}):(A,\alpha)\to(B,\beta)$ be an extendible cocycle morphism as in Definition 3.1 with $\mathbbm{v}_{g}$ being adjointable for all $g\in\Gamma$.
-Fix $g\in\Gamma$ and let $f_{g}:=T_{g}^{\prime}\circ\mathbbm{v}_{\mathbb{C}_{g}}\circ S_{g}:{}_{\phi\circ\alpha_{g}}B\to{}_{\beta_{g}\circ\phi}B$, where $S_{g}:{}_{\phi\circ\alpha_{g}}B\to{}_{\alpha_{g}}A\boxtimes{}_{\phi}B$ is given by $S_{g}(b)=\lim\limits_{\lambda}\xi_{\lambda}\boxtimes b$ for any $b\in B$, where $\xi_{\lambda}$ is an approximate unit for $A$ and $T_{g}^{\prime}:{}_{\phi}B\boxtimes{}_{\beta_{g}}B\rightarrow{}_{\beta_{g}\circ\phi}B$ is given by $T_{g}^{\prime}(b\boxtimes c)=\beta_{g}(b)c$ for $b,c\in B$ [^8] Note that $S_{g}$ is well-defined by Proposition 1.7.
+Let $(\phi,\mathbbm{v}):(A,\alpha)\to(B,\beta)$ be an extendible cocycle morphism as in Definition 3.1 with $\mathbbm{v}_{\mathbb{C}g}$ being adjointable for all $g\in\Gamma$.
+Fix $g\in\Gamma$ and let $f_{g}:=T_{g}^{\prime}\circ\mathbbm{v}_{\mathbb{C}_{g}}\circ S_{g}:{}_{\phi\circ\alpha_{g}}B\to{}_{\beta_{g}\circ\phi}B$, where $S_{g}:{}_{\phi\circ\alpha_{g}}B\to{}_{\alpha_{g}}A\boxtimes{}_{\phi}B$ is given by $S_{g}(b)=\lim\limits_{\lambda}\xi_{\lambda}\boxtimes b$ for any $b\in B$, where $\xi_{\lambda}$ is an approximate unit for AA and $T_{g}^{\prime}:{}_{\phi}B\boxtimes{}_{\beta_{g}}B\rightarrow{}_{\beta_{g}\circ\phi}B$ is given by $T_{g}^{\prime}(b\boxtimes c)=\beta_{g}(b)c$ for $b,c\in B$.
+[^8] Note that $S_{g}$ is well-defined by Proposition 1.6.
 Moreover, let $T_{g}:{}_{\alpha_{g}}A\boxtimes{}_{\phi}B\rightarrow{}_{\phi\circ\alpha_{g}}B$ be given by $T_{g}(a\boxtimes b)=\phi(a)b$ for any $a\in A$ and any $b\in B$.
@@
 As $S_{g}\circ T_{g}$ is the identity map, it follows that the diagram
-[…]
+(3.2) $_{\alpha_{g}}A\boxtimes{}_{\phi}B$$_{\phi}B\boxtimes{}_{\beta_{g}}B$$_{\phi\circ\alpha_{g}}B$$_{\beta_{g}\circ\phi}B$$\scriptstyle{T_{g}}$$\scriptstyle{\mathbbm{v}_{\mathbb{C}_{g}}}$$\scriptstyle{f_{g}}$$\scriptstyle{(T_{g}^{\prime})^{-1}}$
 commutes. Hence,
@@
 Remark 3.4.
-Note that, unlike in [40, Definition 1.10], we do not require $\phi\in\mathrm{Corr}(A,B)$ to be extendible.
+Note that, unlike in [41, Definition 1.10], we do not require $\phi\in\mathrm{Corr}(A,B)$ to be extendible.
 In fact, this is precisely the reason why we consider the maps $\mathbbm{v}_{X}$ to be isometries (possibly non-adjointable) instead of unitaries.
@@
 Furthermore, the morphisms of Definition 3.1 fit into the $\mathcal{C}$-equivariant $\mathrm{KK}$-theory developed in [1] (see [1, Example 3.3]).
-In [1] a correspondence morphism is instead called a $\mathcal{C}$-Hilbert $A$-$B$-bimodule and a cocycle morphism is called a cocycle-$\mathcal{C}$-***-homomorphism.
-[…]
+In [1] a correspondence morphism is instead called a $\mathcal{C}$-Hilbert AA-$B$-bimodule and a cocycle morphism is called a cocycle-$\mathcal{C}$-*-homomorphism.
+We now define composition formulae for the various notions of morphisms in Definition 3.1.
+Using the standard composition between correspondences, we can define composition of correspondence morphisms in the obvious way.
+Let $\mathcal{C}\overset{F}{\curvearrowright}A$, $\mathcal{C}\overset{G}{\curvearrowright}B$, and $\mathcal{C}\overset{H}{\curvearrowright}C$ be actions of $\mathcal{C}$ on $\mathrm{C}^{*}$-algebras AA, $B$, and $C$ respectively.
+If $(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ and $(\psi,\mathbbm{w}):(B,G,I)\to(C,H,K)$ are correspondence morphisms, their composition is denoted by $(\psi\circ\phi,\mathbbm{w}\circ\mathbbm{v})$, where
 (3.3) $(\mathbbm{w}\circ\mathbbm{v})_{X}=(\mathrm{id}_{\phi}\boxtimes\mathbbm{w}_{X})\circ(\mathbbm{v}_{X}\boxtimes\mathrm{id}_{\psi}).$
-By combining the pentagon diagrams for $\mathbbm{v}$ and $\mathbbm{w}$, one obtains the following lemma.
-Lemma 3.5.
-With the notation above, $(\psi\circ\phi,\mathbbm{w}\circ\mathbbm{v}):(A,F,J)\to(C,H,K)$ is a correspondence morphism which denotes the composition of $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$.
+By combining the pentagon diagrams for $\mathbbm{v}$ and $\mathbbm{w}$, one obtains that $(\psi\circ\phi,\mathbbm{w}\circ\mathbbm{v}):(A,F,J)\to(C,H,K)$ is indeed a correspondence morphism.
 However, if $\phi$ and $\psi$ are possibly degenerate cocycle morphisms, the composition formula above will not give a cocycle morphism.
@@
 Let $\mathbbm{w}*\mathbbm{v}$ be the collection of isometries $\{(\mathbbm{w}*\mathbbm{v})_{X}\}_{X\in\mathcal{C}}$ given by
-[…]
-Here $S_{X}(x\boxtimes c)=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}\boxtimes c$ and $T(b\boxtimes c)=\psi(b)c$ for any $X\in\mathcal{C}$, $x\in F(X)$, $b\in B$, and $c\in C$, with $\eta_{\lambda}$ being an approximate unit for $B$ [^10] We have not shown that $S_{X}$ is well-defined at this point.
+(3.4) $F(X)\boxtimes{}_{\psi\circ\phi}C$$_{\psi\circ\phi}C\boxtimes H(X)$$F(X)\boxtimes{}_{\phi}B\boxtimes{}_{\psi}C$$_{\phi}B\boxtimes{}_{\psi}C\boxtimes H(X).$$\scriptstyle{S_{X}}$$\scriptstyle{(\mathbbm{w}*\mathbbm{v})_{X}}$$\scriptstyle{(\mathbbm{w}\circ\mathbbm{v})_{X}}$$\scriptstyle{T\boxtimes\mathrm{id}_{H(X)}}$
+Here $S_{X}(x\boxtimes c)=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}\boxtimes c$ and $T(b\boxtimes c)=\psi(b)c$ for any $X\in\mathcal{C}$, $x\in F(X)$, $b\in B$, and $c\in C$, with $\eta_{\lambda}$ being an approximate unit for BB.
+[^10] We have not shown that $S_{X}$ is well-defined at this point.
 Lemma 3.6.
@@
 $\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c,x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c\rangle_{C}=\langle c,\psi((\eta_{\lambda}-\eta_{\mu})\phi(\langle x,x\rangle_{A})(\eta_{\lambda}-\eta_{\mu}))c\rangle_{C}.$
-By the Cauchy-Schwarz inequality and since $(\eta_{\lambda}-\eta_{\mu})\phi(\langle x,x\rangle_{A})^{1/2}$ converges to 00, it is readily seen that $\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c,x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c\rangle_{C}$ converges to 00, so $S_{X}$ is well-defined.
+By the Cauchy-Schwarz inequality and since $(\eta_{\lambda}-\eta_{\mu})\phi(\langle x,x\rangle_{A})^{1/2}$ converges to 0, it is readily seen that $\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c,x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c\rangle_{C}$ converges to 0, so $S_{X}$ is well-defined.
 Moreover, commutation with the left and right actions are immediate.
@@
 Let $x\in F(X)$ and $c\in C$.
-Since $F(X)$ is a non-degenerate $A$-bimodule, recall from Lemma 1.11 that the map $R_{X}^{-1}:F(X)\boxtimes A\to F(X)$ given by $R_{X}^{-1}(x\boxtimes a)=x\lhd a$ is a bimodule isomorphism.
+Since $F(X)$ is a non-degenerate AA-bimodule, the map $R_{X}^{-1}:F(X)\boxtimes A\to F(X)$ given by $R_{X}^{-1}(x\boxtimes a)=x\lhd a$ is a bimodule isomorphism (see the discussion above Lemma 1.10).
 Then, let $y\in F(X)$ and $a\in A$ such that $y\lhd a=x$.
@@
 ∎
-Note that the proof of Lemma 3.7 also shows that the map $S_{X}$ does not depend on the choice of approximate unit.
+Note that the proof of Lemma 3.6 also shows that the map $S_{X}$ does not depend on the choice of approximate unit. The proof of the following Lemma is routine.
 Lemma 3.7.
 Let $(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ and $(\psi,\mathbbm{w}):(B,G,I)\to(C,H,K)$ be cocycle morphisms. Then $(\psi\circ\phi,\mathbbm{w}*\mathbbm{v})$ is a cocycle morphism.
-Proof.
-Note that $\psi\circ\phi:A\to C$ is a $*{}^{*}$-homomorphism.
-Then, $(\mathbbm{w}*\mathbbm{v})_{X}$ is an isometry as it is a composition of isometries, and it is also an $A$-$C$-bimodule map as it is a composition of bimodule maps.
-Moreover, naturality of $\mathbbm{w}*\mathbbm{v}$ follows from naturality of $\mathbbm{w}\circ\mathbbm{v}$.
-Let us now check that the family of maps $\{(\mathbbm{w}*\mathbbm{v})_{X}\}$ satisfies the diagram
-(3.5) {tikzcd}
-for all $X,Y\in\mathcal{C}$.
-Fix $X,Y\in\mathcal{C}$. By Definition 3.6, composing the two rightmost arrows in (3.5) yields that
-(3.6) $(\mathbbm{w}*\mathbbm{v})_{X\otimes Y}\circ(J_{X,Y}\boxtimes\mathrm{id}_{C})$
-$=(T\boxtimes\mathrm{id}_{H(X\otimes Y)})\circ(\mathbbm{w}\circ\mathbbm{v})_{X\otimes Y}\circ S_{X\otimes Y}\circ(J_{X,Y}\boxtimes\mathrm{id}_{C})$
-(3.7) $=(T\boxtimes\mathrm{id}_{H(X\otimes Y)})\circ(\mathbbm{w}\circ\mathbbm{v})_{X\otimes Y}\circ(J_{X,Y}\boxtimes\mathrm{id}_{B}\boxtimes\mathrm{id}_{C})\circ(\mathrm{id}_{F(Y)}\boxtimes S_{X}).$
-Then, by Lemma 3.5,
-(3.8) $(\mathbbm{w}\circ\mathbbm{v})_{X\otimes Y}\circ(J_{X,Y}\boxtimes\mathrm{id}_{B}\boxtimes\mathrm{id}_{C})$
-(3.9) $=(\mathrm{id}_{B}\boxtimes\mathrm{id}_{C}\boxtimes K_{X,Y})\circ\big{(}(\mathbbm{w}\circ\mathbbm{v})_{Y}\boxtimes\mathrm{id}_{H(X)}\big{)}\circ\big{(}\mathrm{id}_{F(Y)}\boxtimes(\mathbbm{w}\circ\mathbbm{v})_{X}\big{)}.$
-Substituting (3.8) into (3.6), and using Definition 3.6, it follows that the diagram in (3.5) commutes.
-It remains to check that $(\mathbbm{w}*\mathbbm{v})_{1_{\mathcal{C}}}:A\boxtimes{}_{\psi\circ\phi}C\to{}_{\psi\circ\phi}C\boxtimes C$ is given by $(\mathbbm{w}*\mathbbm{v})_{1_{\mathcal{C}}}(a\boxtimes c)=\psi(\phi(a))\boxtimes c.$ By definition,
-$(\mathbbm{w}*\mathbbm{v})_{1_{\mathcal{C}}}=(T\boxtimes\mathrm{id}_{C})\circ(\mathrm{id}_{B}\boxtimes\mathbbm{w}_{1_{\mathcal{C}}})\circ(\mathbbm{v}_{1_{\mathcal{C}}}\boxtimes\mathrm{id}_{C})\circ S_{1_{\mathcal{C}}}.$
-Let $\eta_{\lambda}$ be an approximate unit for $B$. Then, a direct calculation shows that
-$(\mathbbm{w}*\mathbbm{v})_{1_{\mathcal{C}}}(a\boxtimes c)$ $=\lim\limits_{\lambda}(T\boxtimes\mathrm{id}_{C})\circ(\mathrm{id}_{B}\boxtimes\mathbbm{w}_{1_{\mathcal{C}}})\circ(\mathbbm{v}_{1_{\mathcal{C}}}\boxtimes\mathrm{id}_{C})(a\boxtimes\eta_{\lambda}\boxtimes c)$
-$=\lim\limits_{\lambda}(T\boxtimes\mathrm{id}_{C})\circ(\mathrm{id}_{B}\boxtimes\mathbbm{w}_{1_{\mathcal{C}}})(\phi(a)\boxtimes\eta_{\lambda}\boxtimes c)$
-$=\lim\limits_{\lambda}(T\boxtimes\mathrm{id}_{C})(\phi(a)\boxtimes\psi(\eta_{\lambda})\boxtimes c)$
-$=\lim\limits_{\lambda}\psi(\phi(a))\psi(\eta_{\lambda})\boxtimes c$
-$=\psi(\phi(a))\boxtimes c,$
-for any $a\in A$ and any $c\in C$. Hence, $(\psi\circ\phi,\mathbbm{w}*\mathbbm{v})$ is a cocycle morphism. ∎
 Remark 3.8.
-With the same notation as in Definition 3.6, if $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are extendible cocycle representations, then it also follows that $(\psi\circ\phi,\mathbbm{w}*\mathbbm{v})$ is a well-defined cocycle representation.
+We invite the reader to recall Remark 1.5.
+Then, with the same notation as in Definition 3.5, if $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are cocycle representations and $\psi$ is extendible, then it also follows that $(\psi\circ\phi,\mathbbm{w}*\mathbbm{v})$ is a well-defined cocycle representation.
 Remark 3.9.
 If the cocycle morphisms in Definition 3.5 are assumed to be non-degenerate, then the maps $S_{X}$ and $T$ are bimodule isomorphisms for any $X\in\mathcal{C}$.
-Therefore, the composition in Definition 3.6 (denoted ***) corresponds canonically to the composition of correspondence morphisms (denoted $\circ$).
+Therefore, the composition in Definition 3.5 (denoted *) corresponds canonically to the composition of correspondence morphisms (denoted $\circ$).
 The composition considered in Definition 3.5 defines a category.
@@
 This viewpoint will facilitate our constructions and proofs in later sections.
-Our approach is motivated by [6, Lemma 3.8] which introduces this alternative viewpoint in the unital setting.
-First, we recall a way of extending an action $(F,J)$ of a $*{}^{*}$-tensor category $\mathcal{C}$ on $A$ to its matrix amplification $M_{n}(A)$.
-Consider the functor $F^{\mathrm{amp}(n)}:\mathcal{C}\rightarrow\mathrm{Corr}(M_{n}(A))$ that maps objects $X\in\mathcal{C}$ to the correspondence with underlying bimodule $F(X)\otimes M_{n}(\mathbb{C})$ with the right inner product defined by $\langle(x_{ij}),(y_{ij})\rangle_{M_{n}(A)}=(\sum_{l}\langle x_{li},y_{lj}\rangle_{A})$, right $M_{n}(A)$ action given by $(x_{ij})\lhd(a_{ij})=(\sum_{l}x_{il}\lhd a_{lj})$ and with the left action given by $(a_{ij})\rhd(x_{ij})=(\sum_{l}a_{il}\rhd x_{lj})$ for $a_{ij}\in A$ and $x_{ij},y_{ij}\in F(X)$ for $1\leq i,j\leq n$ [^11] This bimodule can be identified with the external tensor product of $F(X)$ and $M_{n}(\mathbb{C})$.
+Our approach is motivated by [7, Lemma 3.8] which introduces this alternative viewpoint in the unital setting.
+First, we recall a way of extending an action $(F,J)$ of a C-tensor category $\mathcal{C}$ on AA to its matrix amplification $M_{n}(A)$.
+Consider the functor $F^{\mathrm{amp}(n)}:\mathcal{C}^{\mathrm{rev}}\rightarrow\mathrm{Corr}(M_{n}(A))$ that maps objects $X\in\mathcal{C}$ to the correspondence with underlying bimodule $F(X)\otimes M_{n}(\mathbb{C})$ with the right inner product defined by $\langle(x_{ij}),(y_{ij})\rangle_{M_{n}(A)}=(\sum_{l}\langle x_{li},y_{lj}\rangle_{A})$, right $M_{n}(A)$ action given by $(x_{ij})\lhd(a_{ij})=(\sum_{l}x_{il}\lhd a_{lj})$ and with the left action given by $(a_{ij})\rhd(x_{ij})=(\sum_{l}a_{il}\rhd x_{lj})$ for $a_{ij}\in A$ and $x_{ij},y_{ij}\in F(X)$ for $\leq i,j\leq n.11$1111This bimodule can be identified with the external tensor product of $F(X)$ and $M_{n}(\mathbb{C})$.
 Hence the inner product defines a Hilbert $M_{n}(A)$-bimodule (see [29]).
 For a morphism $T\in\mathrm{Hom}(X,Y)$ we let $F^{\mathrm{amp}(n)}(T)=T\otimes\mathrm{id}_{M_{n}(\mathbb{C})}$.
-Moreover, letting $J_{X,Y}^{\mathrm{amp}(n)}((y_{ij})\boxtimes(x_{ij})))=(\sum_{l}J_{X,Y}(y_{il}\boxtimes x_{lj}))$ for $x_{ij}\in F(X)$ and $y_{ij}\in F(Y)$ for $1\leq i,j\leq n$, it is a straightforward calculation that $(F^{\mathrm{amp}(n)},J^{\mathrm{amp}(n)})$ is an action of $\mathcal{C}$ on $M_{n}(A)$.
+Moreover, letting $J_{X,Y}^{\mathrm{amp}(n)}((y_{ij})\boxtimes(x_{ij}))=(\sum_{l}J_{X,Y}(y_{il}\boxtimes x_{lj}))$ for $x_{ij}\in F(X)$ and $y_{ij}\in F(Y)$ for $1\leq i,j\leq n$, it is a straightforward calculation that $(F^{\mathrm{amp}(n)},J^{\mathrm{amp}(n)})$ is an action of $\mathcal{C}$ on $M_{n}(A)$.
 Lemma 3.10.
-Let $\mathcal{C}$ be a $\mathrm{C}^{*}$-tensor category acting on $\mathrm{C}^{*}$-algebras $A$ and $B$ via $(A,F,J)$ and $(B,G,I)$ respectively, and let $\phi:A\rightarrow B$ be a $*{}^{*}$-homomorphism.
+Let $\mathcal{C}$ be a $\mathrm{C}^{*}$-tensor category acting on $\mathrm{C}^{*}$-algebras AA and $B$ via $(A,F,J)$ and $(B,G,I)$ respectively, and let $\phi:A\rightarrow B$ be a -homomorphism.
 Then there is a bijection between the families $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ corresponding to a cocycle morphism $(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ and families of linear maps:
@@
     the diagram:
-[…]
+    $F(Y)\boxtimes F(X)$$F(X\otimes Y)$$G(Y)\boxtimes G(X)$$G(X\otimes Y)$$\scriptstyle{h^{Y}\boxtimes h^{X}}$$\scriptstyle{J_{X,Y}}$$\scriptstyle{h^{X\otimes Y}}$$\scriptstyle{I_{X,Y}}$
     commutes;
@@
 Suppose we are given a collection of linear maps $\{h^{X}\}$ satisfying the conditions listed above.
-Fix $\zeta_{\lambda}$ an approximate unit for $A$.
+Fix $\zeta_{\lambda}$ an approximate unit for AA.
 For all $X\in\mathcal{C}$ let $\mathbbm{v}_{X}:F(X)\boxtimes{}_{\phi}B\to{}_{\phi}B\boxtimes G(X)$ be given by
@@
 Then, for any $x\in F(X)$, $\zeta_{\lambda}\rhd x=\sigma(\zeta_{\lambda})x$ converges to $x$.
-Moreover, condition (iii) implies that $h^{X}$ is continuous, so the right hand side of (3.11) converges to zero and hence $\|\phi(\zeta_{\lambda}-\zeta_{\mu})\boxtimes h^{X}(x)\lhd b\|^{2}$ converges to 00.
+Moreover, condition (iii) implies that $h^{X}$ is continuous, so the right hand side of (3.6) converges to zero and hence $\|\phi(\zeta_{\lambda}-\zeta_{\mu})\boxtimes h^{X}(x)\lhd b\|^{2}$ converges to 0.
 Therefore the formula in (3.5) gives a well-defined map $\mathbbm{v}_{X}$ for all $X\in\mathcal{C}$.
@@
 $=a\rhd\mathbbm{v}_{X}(x\boxtimes b).$
-Therefore, $\mathbbm{v}_{X}$ is an $A$-$B$-bimodule map. Moreover, using (v),
+Therefore, $\mathbbm{v}_{X}$ is an AA-$B$-bimodule map. Moreover, using (v),
 $\mathbbm{v}_{1_{\mathcal{C}}}(a\boxtimes b)$ $=\lim\limits_{\lambda}\phi(\zeta_{\lambda})\boxtimes\phi(a)b$
@@
 It remains to check that the diagram
-(3.12) {tikzcd}
+(3.7) $F(Y)\boxtimes F(X)\boxtimes{}_{\phi}B$$F(Y)\boxtimes{}_{\phi}B\boxtimes G(X)$$F(X\otimes Y)\boxtimes{}_{\phi}B$$_{\phi}B\boxtimes G(Y)\boxtimes G(X)$$_{\phi}B\boxtimes G(X\otimes Y).$$\scriptstyle{J_{X,Y}\boxtimes\mathrm{id}_{B}}$$\scriptstyle{\mathrm{id}_{F(Y)}\boxtimes\mathbbm{v}_{X}}$$\scriptstyle{\mathbbm{v}_{Y}\boxtimes\mathrm{id}_{G(X)}}$$\scriptstyle{\mathbbm{v}_{X\otimes Y}}$$\scriptstyle{\mathrm{id}_{B}\boxtimes I_{X,Y}}$
 commutes for all $X,Y\in\mathcal{C}$.
@@
 We claim that $\Psi$ is independent of the choice of approximate unit.
-Indeed let $\zeta_{\lambda}$ and $\xi_{\lambda}$ be two approximate units for $A$.
-Similarly as in (3.11) we have that, $\|\phi(\zeta_{\lambda}-\xi_{\lambda})\boxtimes h^{X}(x)\lhd b\|=\|h^{X}(\zeta_{\lambda}\rhd x-\xi_{\lambda}\rhd x)\lhd b\|$, which converges to 00.
+Indeed let $\zeta_{\lambda}$ and $\xi_{\lambda}$ be two approximate units for AA.
+Similarly as in (3.6) we have that, $\|\phi(\zeta_{\lambda}-\xi_{\lambda})\boxtimes h^{X}(x)\lhd b\|=\|h^{X}(\zeta_{\lambda}\rhd x-\xi_{\lambda}\rhd x)\lhd b\|$, which converges to 0.
 Hence, the map $\mathbbm{v}_{X}$ is independent of the choice of approximate unit, and so $\Psi$ is well-defined.
@@
 where $\iota(x)=\lim_{\lambda}x\boxtimes\eta_{\lambda}$ for some approximate unit $\eta_{\lambda}$ of $B$ and all $x\in F(X)$, and $f$ is the map given by $f(b\boxtimes y)=b\rhd y$ for all $b\in B$ and $y\in G(X)$.
-Note that $f$ is an $A$-$B$-bimodule isomorphism if we see $G(X)$ as a left $A$-module through $\phi$ (i.e.
+Note that $f$ is an AA-$B$-bimodule isomorphism if we see $G(X)$ as a left AA-module through $\phi$ (i.e.
 $\phi(a)\rhd_{B}f(y)=a\rhd_{A}f(y)$ for all $y\in{}_{\phi}B\boxtimes G(X)$ and $a\in A$).
@@
 Precisely, one has that $\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu}),x\boxtimes(\eta_{\lambda}-\eta_{\mu})\rangle_{B}=\langle\eta_{\lambda}-\eta_{\mu},\langle x,x\rangle_{A}\rhd(\eta_{\lambda}-\eta_{\mu})\rangle_{B}=|\phi(\langle x,x\rangle_{A})^{1/2}(\eta_{\lambda}-\eta_{\mu})|^{2}$.
-This converges to 00 since the image of $\phi$ is contained in $B$ and $\eta_{\lambda}$ is an approximate unit for $B$.
+This converges to 0 since the image of $\phi$ is contained in $B$ and $\eta_{\lambda}$ is an approximate unit for $B$.
 We now check that the family $\{h^{X}\}$ defined above satisfies the required compatibility conditions.
@@
 To see (i), note that $f$, $\mathbbm{v}_{X}$ and $\iota$ are left module maps so $\phi(a)\rhd h^{X}(x)=h^{X}(a\rhd x)$ for all $a\in A$ and $x\in F(X)$.
-Moreover, as $f$ and $\mathbbm{v}_{X}$ are right $B$-module maps and $\iota$ is a right $A$-module map
+Moreover, as $f$ and $\mathbbm{v}_{X}$ are right $B$-module maps and $\iota$ is a right AA-module map
 $h^{X}(x)\lhd\phi(a)$ $=\lim\limits_{\lambda}f(\mathbbm{v}_{X}(x\boxtimes\eta_{\lambda}\phi(a)))=h^{X}(x\lhd a).$
@@
 Indeed, let $\eta_{\lambda}$ and $\xi_{\lambda}$ be two approximate units for $B$.
-We show that the net $x\boxtimes(\eta_{\lambda}-\xi_{\lambda})$ converges to 00 for any $x\in F(X)$.
+We show that the net $x\boxtimes(\eta_{\lambda}-\xi_{\lambda})$ converges to 0 for any $x\in F(X)$.
 Note that
@@
 $=|\phi(\langle x,x\rangle_{A})^{1/2}(\eta_{\lambda}-\xi_{\lambda})|^{2},$
-which converges to 00.
+which converges to 0.
 We claim that $\Phi$ and $\Psi$ are inverses to each other. First we show that $\Phi\circ\Psi$ is the identity map. For any $X\in\mathcal{C}$ and $x\in F(X)$, it follows that
@@
 That is because if $(\phi,\mathbbm{v})$ is a cocycle representation, $\phi$ can land in $M(B)\setminus B$.
-As $\eta_{\lambda}$ is an approximate unit for $B$, $\|\phi(\langle x,x\rangle_{A})^{1/2}(\eta_{\lambda}-\eta_{\mu})\|^{2}$ need not converge to 00.
+As $\eta_{\lambda}$ is an approximate unit for $B$, $\|\phi(\langle x,x\rangle_{A})^{1/2}(\eta_{\lambda}-\eta_{\mu})\|^{2}$ need not converge to 0.
+Remark 3.11.
+Note that condition (i) follows from (iv) as $J_{1,X}$ and $J_{X,1}$ correspond to the left and right actions of AA on $F(X)$.
+Similarly from (iv) it is clear that $h^{1_{\mathcal{C}}}$ is a -homomorphism so we may simply define a cocycle morphism satisfying conditions (ii)-(iv) by setting $h^{1_{\mathcal{C}}}=\phi$.
 Remark 3.12.
@@
 In particular, it suffices to check that a family of linear maps $\{h^{X}:F(X)\rightarrow G(X)\}_{X\in\mathrm{Irr}(\mathcal{C})}$ satisfy the conditions of Lemma 3.10 to yield a cocycle morphism, with understanding condition (iv) as $I_{X,Y}\circ(h^{Y}\boxtimes h^{X})=\bigoplus_{i}h^{X_{i}}\circ J_{X,Y}$ for all $X,Y\in\mathrm{Irr}(\mathcal{C})$ and $X\otimes Y\cong\bigoplus_{i}X_{i}$ is the irreducible decomposition.
+Lemma 3.10 shows that any cocycle morphism can be equivalently represented by a pair $(\phi,h)$, where for convenience, we denote a cocycle morphism by $(\phi,h)$, where $h$ denotes the collection of linear maps $\{h^{X}\}_{X\in\mathcal{C}}$.
+Moreover, Lemma 3.15 below shows that in the latter picture, the composition of cocycle morphisms translates to the composition of the underlying -homomorphisms and linear maps.
+From now on, we will freely identify these two pictures.
 Remark 3.13.
@@
 Similarly, for any $u,u^{\prime}\in\mathcal{U}(\mathcal{M}(A))$ it follows that $h^{X}(u\rhd x\lhd u^{\prime})=\phi^{\dagger}(u)\rhd h^{X}(x)\lhd\phi^{\dagger}(u^{\prime})$.
-Indeed, for an approximate unit $e_{\lambda}$ of $A$, $\phi(e_{\lambda})$ converges to $p$ and
+Indeed, for an approximate unit $e_{\lambda}$ of AA, $\phi(e_{\lambda})$ converges to $p$ and
 $p\rhd h^{X}(x)=\lim\limits_{\lambda}\phi(e_{\lambda})\rhd h^{X}(x)=\lim\limits_{\lambda}h^{X}(e_{\lambda}\rhd x)=h^{X}(x).$
-So $\phi^{\dagger}(u)\rhd h^{X}(x)\lhd\phi^{\dagger}(u^{\prime})=\phi_{p}(u)\rhd h^{X}(x)\lhd\phi_{p}(u^{\prime})=h^{X}(u\rhd x\lhd u^{\prime})$.
+Likewise, we have $h^{X}(x)\lhd p=h^{X}(x)$. Therefore,
+$\phi^{\dagger}(u)\rhd h^{X}(x)\lhd\phi^{\dagger}(u^{\prime})=\phi_{p}(u)\rhd h^{X}(x)\lhd\phi_{p}(u^{\prime})=h^{X}(u\rhd x\lhd u^{\prime}).$
 Example 3.14.
-Suppose $\mathcal{C}=\mathrm{Hilb}(\Gamma)$ and $(\phi,\{h^{g}\}_{g\in\Gamma}):(A,\alpha,J)\to(B,\beta,I)$ is an extendible cocycle morphism.
-Recall from Example 3.3 that $\mathbbm{v}_{g}(a\boxtimes b)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes\mathbbm{u}_{g}\phi(a)b,$ where $\eta_{\lambda}$ is an approximate unit for $B$ and $(\mathbbm{u}_{g})_{g\in\Gamma}\subseteq\mathcal{M}(B)$.
+Suppose $(A,\alpha)$ and $(B,\beta)$ are actions of a countable discrete group $\Gamma$ on $\mathrm{C}^{*}$-algebras AA and $B$.
+Consider them as actions of $\mathrm{Hilb}(\Gamma)$ as in Example 2.4 and let $(\phi,\{h^{g}\}_{g\in\Gamma}):(A,\alpha)\to(B,\beta)$ be an extendible cocycle morphism.
+Recall from Example 3.3 that $\mathbbm{v}_{\mathbb{C}g}(a\boxtimes b)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes\mathbbm{u}_{g}\phi(a)b,$ where $\eta_{\lambda}$ is an approximate unit for $B$ and $(\mathbbm{u}_{g})_{g\in\Gamma}\subseteq\mathcal{M}(B)$.
 Then for any $g\in\Gamma$ and $a\in A$,
-$h^{g}(a)=f\left(\mathbbm{v}_{g}\left(\lim\limits_{\lambda}a\boxtimes\eta_{\lambda}\right)\right)=f\left(\lim\limits_{\lambda}\eta_{\lambda}\boxtimes\mathbbm{u}_{g}\phi(a)\right)=\mathbbm{u}_{g}\phi(a).$
-For convenience, we will denote a cocycle morphism by $(\phi,h)$, where $h$ denotes the collection of linear maps $\{h^{X}\}_{X\in\mathcal{C}}$. We now discuss their composition.
+$h^{g}(a)=f\left(\mathbbm{v}_{\mathbb{C}g}\left(\lim\limits_{\lambda}a\boxtimes\eta_{\lambda}\right)\right)=f\left(\lim\limits_{\lambda}\eta_{\lambda}\boxtimes\mathbbm{u}_{g}\phi(a)\right)=\mathbbm{u}_{g}\phi(a).$
+We now discuss the composition of cocycle morphisms.
 Lemma 3.15.
@@
 Proof.
-Clearly $\psi\circ\phi:A\to C$ is a $*{}^{*}$-homomorphism and $\{l^{X}\circ h^{X}:F(X)\to H(X)\}_{X\in\mathcal{C}}$ is a family of linear maps.
+Clearly $\psi\circ\phi:A\to C$ is a -homomorphism and $\{l^{X}\circ h^{X}:F(X)\to H(X)\}_{X\in\mathcal{C}}$ is a family of linear maps.
 Conditions (i), (ii), (iii), and (v) are immediate.
@@
 where $(\mathbbm{w}\circ\mathbbm{v})_{X}=(\mathrm{id}_{\phi}\boxtimes\mathbbm{w}_{X})\circ(\mathbbm{v}_{X}\boxtimes\mathrm{id}_{\psi})$ (see (3.3)).
-Start with an elementary tensor $x\boxtimes c$ and let $\zeta_{\mu}$ be an approximate unit of $A$ and $\eta_{\lambda}$ be an approximate unit of $B$.
+Start with an elementary tensor $x\boxtimes c$ and let $\zeta_{\mu}$ be an approximate unit of AA and $\eta_{\lambda}$ be an approximate unit of $B$.
 Following the composition of maps defining $(\mathbbm{w}*\mathbbm{v})_{X}$, we get that
@@
 ∎
-Lemma 3.11 shows that any cocycle morphism can be equivalently represented by a pair $(\phi,h)$.
-Moreover Lemma 3.15 shows that in the latter picture the composition of cocycle morphisms translates to the composition of the underlying $*{}^{*}$-homomorphisms and linear maps.
-From now on we will freely identify these two pictures.
 Lemma 3.16.
@@
 ∎
-In the spirit of [40, Definition 1.16], we have constructed a category of $\mathcal{C}$-$\mathrm{C}^{*}$-algebras.
+In the spirit of [41, Definition 1.16], we have constructed a category of $\mathcal{C}$-$\mathrm{C}^{*}$-algebras.
 Definition 3.17.
@@
 In this section, we construct inductive limits in $\mathrm{C}^{*}_{\mathcal{C}}$ for semisimple $\mathcal{C}$.
-This is done in [6, Proposition 4.4], when restricted to unital $\mathrm{C}^{*}$-algebras and unital, injective connecting maps.
+This is done in [7, Proposition 4.4], when restricted to unital $\mathrm{C}^{*}$-algebras and unital, injective connecting maps.
 Our approach is slightly different and does not need these assumptions.
@@
 be a sequence of cocycle morphisms. Recall that the $\mathrm{C}^{*}$-inductive limit $A=\lim\limits_{\longrightarrow}\{A_{n},\phi_{n}\}$ is defined as the completion of
-$A^{(0)}=\left\{(a_{n})_{n\geq 1}\in\frac{\bigoplus_{\ell^{\infty}}A_{n}}{\bigoplus_{c_{0}}A_{n}}:\exists N\in\mathbb{N}\ \text{s.t.}\ \forall n\geq N,\phi_{n}(a_{n})=a_{n+1}\right\}$
+$A^{(0)}=\frac{\left\{(a_{n})_{n\geq 1}\in\bigoplus_{\ell^{\infty}}A_{n}:\lim_{n\to\infty}\|\phi_{n}(a_{n})-a_{n+1}\|=0\right\}}{\bigoplus_{c_{0}}A_{n}}$
 with respect to the topology induced by the norm $\|(a_{n})_{n\geq 1}\|=\lim\limits_{n\to\infty}\|a_{n}\|_{A_{n}}$. Recall that the connecting maps $\phi_{n,\infty}:A_{n}\to A$ are given by
 (4.2) $(\phi_{n,\infty}(a_{n}))_{k}=\begin{cases}\phi_{n,k}(a_{n}),\ k\geq n\\ 0,\quad\quad\quad k<n\end{cases}$
-for all $n\geq 1$, where we adopt the standard notation $\phi_{n,m}:=\phi_{m-1}\circ\ldots\circ\phi_{n}$ for any $m>n\geq 1$.
+for all $n\geq 1$, where we adopt the standard notation $\phi_{n,m}:=\phi_{m-1}\circ\ldots\circ\phi_{n}$ and $\phi_{n,n}=\mathrm{id}_{A_{n}}$ for any $m>n\geq 1$.
 Similarly, for any $X\in\mathcal{C}$ and any $m>n\geq 1$, we consider the natural family $h_{n,m}^{X}:F_{n}(X)\to F_{m}(X)$ obtained by composition, with the convention that $h_{n,n+1}^{X}=h_{n}^{X}$.
-To build an action on $A$, we start by constructing bimodules that will form the image of the functor.
-Essentially, for any $X\in\mathcal{C}$, we can build a Hilbert $A$-$A$-bimodule as an inductive limit of $F_{n}(X)$.
+To build an action on AA, we start by constructing bimodules that will form the image of the functor.
+Essentially, for any $X\in\mathcal{C}$, we can build a Hilbert AA-AA-bimodule as an inductive limit of $F_{n}(X)$.
 The construction is very similar to the one in (2.5).
 Define
-$F^{(0)}(X)=\Big{\{}(x_{n})_{n\geq 1}\in\frac{\bigoplus_{\ell^{\infty}}F_{n}(X)}{\bigoplus_{c_{0}}F_{n}(X)}:\exists N\in\mathbb{N}\ \text{s.t.}\ \forall n\geq N,h_{n}^{X}(x_{n})=x_{n+1}\Big{\}},$
+$F^{(0)}(X)=\frac{\{(x_{n})_{n\geq 1}\in\bigoplus_{\ell^{\infty}}F_{n}(X):\lim_{n\to\infty}\|h_{n}^{X}(x_{n})-x_{n+1}\|=0\}}{\bigoplus_{c_{0}}F_{n}(X)},$
 where the norm on $F_{n}(X)$ is induced by the right inner product.
@@
 It is now straightforward to check that this function is right linear, left conjugate linear, and antisymmetric.
-Finally, it is clear that $\langle x,x\rangle_{A^{(0)}}\geq 0$ and $\langle x,x\rangle_{A^{(0)}}=0$ if and only if $\langle x_{n},x_{n}\rangle_{A_{n}}$ converges to 00 i.e.
+Finally, it is clear that $\langle x,x\rangle_{A^{(0)}}\geq 0$ and $\langle x,x\rangle_{A^{(0)}}=0$ if and only if $\langle x_{n},x_{n}\rangle_{A_{n}}$ converges to 0 i.e.
 $x=0$ in $F^{(0)}(X)$.
@@
 ∎
-For any $X\in\mathcal{C}$, since $A^{(0)}$ is a dense $*{}^{*}$-subalgebra of $A$, combining Lemma 4.1 and [37, Lemma 2.16], we form the completion of $F^{(0)}(X)$, denoted by $F(X)$.
-This is a right-Hilbert $A$-module.
+For any $X\in\mathcal{C}$, since $A^{(0)}$ is a dense -subalgebra of AA, combining Lemma 4.1 and [39, Lemma 2.16], we form the completion of $F^{(0)}(X)$, denoted by $F(X)$.
+This is a right-Hilbert AA-module.
 Lemma 4.2.
-For any $X\in\mathcal{C}$, with the notation above, $F(X)$ is a non-degenerate right Hilbert $A$-$A$-bimodule.
+For any $X\in\mathcal{C}$, with the notation above, $F(X)$ is a non-degenerate right Hilbert AA-AA-bimodule.
 Proof.
-We start by defining a left $A$-action on $F(X)$. For any $x=(x_{n})_{n\geq 1}\in F^{(0)}(X)$ and any $a=(a_{n})_{n\geq 1}\in A^{(0)}$, we can define
+We start by defining a left AA-action on $F(X)$. For any $x=(x_{n})_{n\geq 1}\in F^{(0)}(X)$ and any $a=(a_{n})_{n\geq 1}\in A^{(0)}$, we can define
 (4.5) $a\rhd x=(a_{n}\rhd x_{n})_{n\geq 1}.$
@@
 Moreover, the left $A^{(0)}$-action is adjointable, since the left action of $A_{n}$ on $F_{n}(X)$ is adjointable for any $n\geq 1$.
-Thus, using [37, Lemma 2.16], it extends to a left action of $A$ on $F(X)$ by density.
+Thus, using [39, Lemma 2.16], it extends to a left action of AA on $F(X)$ by density.
 ∎
-We now show that the assignment $X\to F(X)$ extends to an action of $\mathcal{C}$ on $A$.
+We now show that the assignment $X\to F(X)$ extends to an action of $\mathcal{C}$ on AA.
 Lemma 4.3.
-With the notation above, $F:\mathcal{C}\to\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$ is a $\mathrm{C}^{*}$-functor. Moreover, there exists a unitary natural isomorphism
+With the notation above, $F:\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$ is a $\mathrm{C}^{*}$-functor. Moreover, there exists a unitary natural isomorphism
 $J:=\{J_{X,Y}:F(Y)\boxtimes F(X)\to F(X\otimes Y):X,Y\in\mathcal{C}\}$
-such that the pair $(F,J)$ is an action of $\mathcal{C}$ on $A$.
+such that the pair $(F,J)$ is an action of $\mathcal{C}$ on AA.
 Proof.
@@
 We define $F(f)(x)=(F_{n}(f)(x_{n}))_{n\geq 1}$ for any $x\in F^{(0)}(X)$.
-As $\mathcal{C}$ is semisimple, by decomposing $X$ and $Y$ into irreducibles one has that $f=\bigoplus\lambda_{i}\mathrm{id}_{X_{i}}$ for finitely many non-zero scalars $\lambda_{i}$ and $X_{i}\in\mathrm{Irr}(\mathcal{C})$.
-Therefore, as each $F_{n}$ is a $*{}^{*}$-functor, $F_{n}(f)$ is uniformly bounded by the supremum of $\lambda_{i}$ and so $F(f)$ is bounded on $F^{(0)}(X)$.
-Therefore, one may extend $F(f)$ to a bounded $A$-$A$-bimodule map on $F(X)$.
+As each $F_{n}$ is a C-functor, the assignment $f\mapsto F_{n}(f)$ is contractive (see for example [20, Definition 1.5]) and so $F(f)$ is bounded on $F^{(0)}(X)$.
+Therefore, one may extend $F(f)$ to a bounded AA-AA-bimodule map on $F(X)$.
 This operator is further adjointable as each $F_{n}(f)$ is adjointable.
@@
 $J_{X,Y}(y\boxtimes x)=(J_{X,Y}^{(n)}(y_{n}\boxtimes x_{n}))_{n\geq 1}.$
-We claim that $J_{X,Y}$ is well-defined. Firstly, note that $J_{X,Y}$ is independent of the choice of representative sequences by continuity of $J_{X,Y}^{(n)}$ for any $n\in\mathbb{N}$.
+We claim that $J_{X,Y}$ is well-defined.
+Firstly, note that $J_{X,Y}$ is independent of the choice of representative sequences as $J_{X,Y}^{(n)}$ is an isometry for any $n\in\mathbb{N}$ and $\lim_{n\to\infty}(F_{n}(Y)\boxtimes F_{n}(X))\cong F(Y)\boxtimes F(X)$.
 Secondly, there exists $N\in\mathbb{N}$ such that for any $n\geq N$,
@@
 Finally, since each family $\{J_{X,Y}^{(n)}:X,Y\in\mathcal{C}\}$ satisfies a commuting diagram as in (2.1), so does the collection of maps $\{J_{X,Y}:X,Y\in\mathcal{C}\}$.
-Hence, the pair $(F,J)$ gives an action of $\mathcal{C}$ on $A$.
+Hence, the pair $(F,J)$ gives an action of $\mathcal{C}$ on AA.
 ∎
 We now show the existence of sequential inductive limits in $\mathrm{C}^{*}_{\mathcal{C}}$.
-Proposition 4.4.
+Proposition $4.$
 The triple $(A,F,J)$ constructed above defines the inductive limit of the inductive system in (4.1).
@@
 commutes.
-For any $n\geq 1$, we have a $*{}^{*}$-homomorphism $\phi_{n,\infty}:A_{n}\to A$.
+Recalling (4.2), for any $n\geq 1$, we have a -homomorphism $\phi_{n,\infty}:A_{n}\to A$.
 To define a cocycle morphism from $(A_{n},F_{n},J^{(n)})$ to $(A,F,J)$, it suffices to find a collection of linear maps $\{h_{n,\infty}^{X}:F_{n}(X)\to F(X)\}_{X\in\mathcal{C}}$ satisfying the conditions in Lemma 3.10.
@@
 We now check that the triple $(A,F,J)$ is the inductive limit of the sequence of triples $(A_{n},F_{n},J^{(n)})$, together with connecting morphisms $(\phi_{n},h_{n})$.
-Firstly, we claim that (4.6) commutes.
-Since $A$ is an inductive limit for the system $(A_{n},\phi_{n})_{n\geq 1}$, it follows that for any $n\geq 1$, $\phi_{n,\infty}=\phi_{n+1,\infty}\circ\phi_{n}$.
-Therefore, it suffices to check that for any $X\in\mathcal{C}$ and any $n\geq 1$, we have that $h_{n,\infty}^{X}=h_{n+1,\infty}^{X}\circ h_{n}^{X}$.
-This follows by (4.7) since for any $k>n+1$, $h_{n,k}^{X}=h_{n+1,k}^{X}\circ h_{n}^{X}$.
+Firstly, it is clear that (4.6) commutes.
 It remains to check that $(A,F,J)$ satisfies the universal property.
@@
 $(\Phi,r)\circ(\phi,h)_{n,\infty}=(\psi,l)_{n,\infty}.$
-By the universal property of the inductive limit $A$, there exists a unique $*{}^{*}$-homomorphism $\Phi:A\to B$ such that $\Phi\circ\phi_{n,\infty}=\psi_{n,\infty}$ for all $n\geq 1$.
-Then, for any $X\in\mathcal{C}$, the union $\bigcup_{k\geq 1}h_{k,\infty}^{X}(F_{k}(X))$ is dense in $F(X)$.
-Since $l_{k,\infty}^{X}$ is uniformly bounded in $k$, (4.8) yields that $\mathrm{Ker}\ h_{k,\infty}^{X}\subseteq\mathrm{Ker}\ l_{k,\infty}^{X}$ for any $X\in\mathcal{C}$ and any $k\in\mathbb{N}$.
+For any $X\in\mathcal{C}$, the union $\bigcup_{k\geq 1}h_{k,\infty}^{X}(F_{k}(X))$ is dense in $F(X)$.
+Since $l_{k,\infty}^{X}$ is contractive, (4.8) yields that $\mathrm{Ker}\ h_{k,\infty}^{X}\subseteq\mathrm{Ker}\ l_{k,\infty}^{X}$ for any $X\in\mathcal{C}$ and any $k\in\mathbb{N}$.
 Therefore, for any $x_{k}\in F_{k}(X)$, define $r^{X}(h_{k,\infty}^{X}(x_{k}))=l_{k,\infty}^{X}(x_{k})$.
-As $l_{k,\infty}^{X}$ is uniformly bounded in $k$, we may extend $r^{X}$ to a well-defined linear map $r^{X}:F(X)\to G(X)$.
+As $l_{k,\infty}^{X}$ is contractive, we may extend $r^{X}$ to a well-defined linear map $r^{X}:F(X)\to G(X)$.
 The fact that $r^{X}$ satisfies the conditions appearing in Lemma 3.10 is routine and essentially follows from the fact that for each $k\geq 1$, $(\psi_{k,\infty},l_{k,\infty})$ and $(\phi_{k,\infty},h_{k,\infty})$ are cocycle morphisms.
-For the convenience of the reader, we check condition (iii) of Lemma 3.11.
-Let $X\in\mathcal{C}$, $x_{k},y_{k}\in F_{k}(X)$.
-Then,
-$\langle r^{X}(h_{k,\infty}^{X}(x_{k})),r^{X}(h_{k,\infty}^{X}(y_{k}))\rangle$ $=\langle l_{k,\infty}^{X}(x_{k}),l_{k,\infty}^{X}(y_{k})\rangle$
-$=\psi_{k,\infty}(\langle x_{k},y_{k}\rangle)$
-$=\Phi\circ\phi_{k,\infty}(\langle x_{k},y_{k}\rangle)$
-$=\Phi(\langle h_{k,\infty}^{X}(x_{k}),h_{k,\infty}^{X}(y_{k})\rangle).$
-The remaining conditions follow similarly.
 Therefore, $(\Phi,r)$ is the unique cocycle morphism such that $(\Phi,r)\circ(\phi,h)_{n,\infty}=(\psi,l)_{n,\infty}$.
@@
 Note that if the connecting maps $\phi_{n}:A_{n}\to A_{n+1}$ of the inductive system are extendible, then the inductive limit connecting map $\phi_{k,\infty}:A_{k}\to A$ is extendible for any $k\geq 1$.
-4.1. Topology induced on cocycle representations
-[…]
+5. Approximate unitary equivalence
+Let $\mathcal{C}$ be a semisimple $\mathrm{C}^{*}$-tensor category with countably many isomorphism classes of simple objects.
+To define approximate unitary equivalence, we first introduce a topology on the space of cocycle representations between two actions $\mathcal{C}\overset{F}{\curvearrowright}A$ and $\mathcal{C}\overset{G}{\curvearrowright}B$.
+For this we employ the same approach as in [41, Section 2.1]; introducing a family of pseudometrics which measure the distance between cocycle representations.
 Remark 5.1.
@@
 Note that $F(1_{\mathcal{C}})=A$ and for any $a\in A$ and $b\in B$ $\mathbbm{v}_{1_{\mathcal{C}}}(a\boxtimes_{\phi}b)=\phi(a)b$ and $\mathbbm{w}_{1_{\mathcal{C}}}(a\boxtimes_{\psi}b)=\psi(a)b$.
-Therefore, the family of pseudometrics $d_{\mathcal{F}}$ bounds the pointwise distance between $\phi$ and $\psi$ in the strict topology.
 We will only use the topology generated by the family of pseudometrics $d_{\mathcal{F}}$ in the setting of cocycle morphisms, so let us derive the relevant subset topology.
@@
 Suppose that for any $X\in\mathrm{Irr}(\mathcal{C})$, $h_{\lambda}^{X}$ converges pointwise to $h^{X}$ in the norm induced by the right inner product.
-Let $\zeta_{\mu}$ be an approximate unit of $A$ and recall from (3.10) that for any $X\in\mathcal{C}$, $x\in F(X)$, and $b\in B$
+Let $\zeta_{\mu}$ be an approximate unit of AA and recall from (3.5) that for any $X\in\mathcal{C}$, $x\in F(X)$, and $b\in B$
 $\mathbbm{v}_{X}(x\boxtimes_{\phi}b)=\lim\limits_{\mu}\phi(\zeta_{\mu})\boxtimes h^{X}(x)\lhd b$
@@
 ∎
-Example 4.10.
+Example $5.$
 Let $\Gamma$ be a countable discrete group.
@@
 (5.5) $d_{\mathcal{F}^{A}\times K}\big((\phi,\mathbbm{v}),(\psi,\mathbbm{w})\big)=\max\limits_{a_{g}\in\mathcal{F}}\|\mathbbm{v}_{g}^{\prime}\phi(a_{g})-\mathbbm{w}_{g}^{\prime}\psi(a_{g})\|.$
-In [40, Definition 2.5], Szabó defines a topology on cocycle morphisms, which using the conventions of this paper, is generated by the family of pseudometrics
+In [41, Definition 2.5], Szabó defines a topology on the space of cocycle morphisms which is generated by the family of pseudometrics
 (5.6) $d_{\mathcal{F}}\big((\phi,\mathbbm{v}^{\prime}),(\psi,\mathbbm{w}^{\prime})\big)=\max\limits_{a\in\mathcal{F}^{A}}\|\phi(a)-\psi(a)\|+\max\limits_{g\in K}\max\limits_{b\in\mathcal{F}^{B}}\|b((\mathbbm{v}_{g}^{\prime})^{*}-(\mathbbm{w}_{g}^{\prime})^{*})\|.$
@@
 However, these topologies are different outside the non-degenerate setting, with the topology induced by the family of pseudometrics in (5.4) being coarser than the topology induced by the family of pseudometrics in (5.6).
-Finally, we would like to point out that the topology induced by (4.13) coincides with that used in [31].
+Finally, we would like to point out that the topology induced by (5.5) coincides with that used in [33, Section 5].
 We finish our discussion by noticing that the composition of cocycle morphisms is jointly continuous with respect to the topology defined above.
@@
 Let $(\phi,h):(A,F,J)\to(B,G,I)$ and $(\psi,l):(B,G,I)\to(C,H,K)$ be two cocycle morphisms. Then the composition map given by
-$[(\phi,h),(\psi,l)]\mapsto(\psi,l)\circ(\phi,h)$
+$(\phi,h),(\psi,l)]\mapsto(\psi,l)\circ(\phi,h)$
 is jointly continuous.
-Proof.
-Let $(\phi_{\lambda},h_{\lambda}):(A,F,J)\to(B,G,I)$ and $(\psi_{\lambda},l_{\lambda}):(B,G,I)\to(C,H,K)$ be two convergent nets with limits $(\phi,h)$ and $(\psi,l)$ respectively.
-First, note that for any $X\in\mathrm{Irr}(\mathcal{C})$, $l^{X}\circ h^{X}=\lim\limits_{\lambda\to\infty}l_{\lambda}^{X}\circ h_{\lambda}^{X}$ in the point-norm topology.
-Then, the case $X=1_{\mathcal{C}}$ yields that $\psi\circ\phi=\lim\limits_{\lambda\to\infty}\psi_{\lambda}\circ\phi_{\lambda}$ holds in the point-norm topology.
-Using the composition formula in Lemma 3.15 and Lemma 4.9, the conclusion follows.
-∎
-4.2. Approximate unitary equivalence
 We can now introduce a notion of approximate unitary equivalence that will be crucial to perform equivariant Elliott intertwining arguments.
@@
 We denote this action by the triple $(B,G,I)$.
-For any $u$ unitary in $\mathcal{M}(B)$, we consider $\mathrm{Ad}(u):B\to B$ to be the $*{}^{*}$-homomorphism given by $b\mapsto ubu^{*}$.
+For any $u$ unitary in $\mathcal{M}(B)$, we consider $\mathrm{Ad}(u):B\to B$ to be the -homomorphism given by $b\mapsto ubu^{*}$.
 Then $\mathrm{Ad}(u)$ induces a Hilbert $B$-bimodule $_{\mathrm{Ad}(u)}B$.
@@
 But for any $X\in\mathcal{C}$, $G(X)\boxtimes B\cong B\boxtimes G(X)$, so there exists a unitary isomorphism $(\mathbbm{v}_{u})_{X}:G(X)\boxtimes{}_{\mathrm{Ad}(u)}B\to{}_{\mathrm{Ad}(u)}B\boxtimes G(X)$.
-It follows that $(\mathrm{Ad}(u),\mathbbm{v}_{u}):(B,G,I)\to(B,G,I)$ is a cocycle morphism.
+It follows that $(\mathrm{Ad}(u),\mathbbm{v}_{u}):(B,G,I)\to(B,G,I)$ is a cocycle morphism..
 We denote by $h_{u}=\{h_{u}^{X}:G(X)\to G(X)\}_{X\in\mathcal{C}}$ the collection of linear maps corresponding to the cocycle morphism induced by $\mathrm{Ad}(u)$.
@@
 Proof.
-For any $X\in\mathcal{C}$, consider the bimodule maps $L_{X}:G(X)\to B\boxtimes G(X)$ and $R_{X}:G(X)\to G(X)\boxtimes B$ given by $L_{X}(x)=\lim\limits_{\mu}\eta_{\mu}\boxtimes x$ and $R_{X}(x)=\lim\limits_{\mu}x\boxtimes\eta_{\mu}$ for some $\eta_{\mu}$ quasicentral approximate unit of $B$ with respect to $\mathcal{M}(B)$.
+For any $X\in\mathcal{C}$, consider the bimodule maps $L_{X}:G(X)\to B\boxtimes G(X)$ and $R_{X}:G(X)\to G(X)\boxtimes B$ given by $L_{X}(x)=\lim\limits_{\mu}\eta_{\mu}\boxtimes x$ and $R_{X}(x)=\lim\limits_{\mu}x\boxtimes\eta_{\mu}$ for some $\eta_{\mu}$ quasicentral approximate unit of $B$ with respect to $\mathcal{M}(B)$ (i.e.
+$\eta_{\mu}x-x\eta_{\mu}\to 0$ for all $x\in\mathcal{M}(B)$).
 As the bimodule $G(X)$ is non-degenerate for any $X\in\mathcal{C}$, it follows from Lemma 1.10 that
-$u\rhd x=L_{X}^{-1}(u\rhd L_{X}(x))\quad\mathrm{and}\quad x\lhd u^{*}=R_{X}^{-1}(R_{X}(x)\lhd u^{*}).$
+[…]
 Then, consider the map
-$\iota:G(X)\to G(X)\boxtimes{}_{\mathrm{Ad}(u)}B\to G(X)\boxtimes B$
-given by $x\mapsto\lim\limits_{\mu}x\boxtimes\eta_{\mu}\mapsto\lim\limits_{\mu}x\boxtimes u^{*}\eta_{\mu}$. Since $\eta_{\mu}$ is quasicentral,
+$\iota:G(X)\to G(X)\boxtimes B$
+given by $x\mapsto\lim\limits_{\mu}x\boxtimes u^{*}\eta_{\mu}$. Since $\eta_{\mu}$ is quasicentral,
 $\iota(x)=\lim\limits_{\mu}x\boxtimes\eta_{\mu}u^{*}=R_{X}(x)\lhd u^{*}.$
@@
 Finally, we consider the map
-$f:B\boxtimes G(X)\to{}_{\mathrm{Ad}(u)}B\boxtimes G(X)\to G(X)$
+$f:B\boxtimes G(X)\to G(X)$
 given by $f(b\boxtimes x)=L_{X}^{-1}(u\rhd(b\boxtimes x))$.
@@
 Definition 5.8.
+Let $\mathcal{C}\overset{F}{\curvearrowright}A$ and $\mathcal{C}\overset{G}{\curvearrowright}B$ be actions of $\mathcal{C}$ on $\mathrm{C}^{*}$-algebras AA and $B$ and let $(\phi,\mathbbm{v}),(\psi,\mathbbm{w}):(A,F,J)\to(B,G,I)$ be cocycle representations.
   1. (i)
-[…]
+    We say that the pairs $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are unitarily equivalent if there exists a unitary $u\in\mathcal{U}(\mathcal{M}(B))$ such that
     $(\mathrm{Ad}(u),\mathbbm{v}_{u})\circ(\phi,\mathbbm{v})=(\psi,\mathbbm{w}).$
   2. (ii)
-    Let $(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ be a cocycle representation and $u\in\mathcal{M}(B)$ a unitary.
     We say that $(\psi,\mathbbm{w})$ is an approximate unitary conjugate of $(\phi,\mathbbm{v})$ if there exists a net of unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
@@
     and
-[…]
+    $\|\mathbbm{w}_{X}(x\boxtimes_{\psi}b)-(\mathbbm{v}_{u_{\lambda}}*\mathbbm{v})_{X}(x\boxtimes_{\mathrm{Ad}(u_{\lambda})\phi}b)\|\overset{\lambda}{\longrightarrow}0$
     for all $a\in A$, $b\in B$, $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$.
@@
 [^14] Note that this means precisely that $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})\circ(\phi,\mathbbm{v})$ converges to $(\psi,\mathbbm{w})$ with respect to the topology in Definition 5.2.
-  3. (iii)
-    We say that $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are approximately unitarily equivalent if $(\psi,\mathbbm{w})\precapprox_{u}(\phi,\mathbbm{v})$ and $(\phi,\mathbbm{v})\precapprox_{u}(\psi,\mathbbm{w})$.
-This will be denoted by $(\phi,\mathbbm{v})\approx_{u}(\psi,\mathbbm{w})$.
-In general, we will be interested in these notions when $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are cocycle morphisms, so let us record the following lemma.
+The analogous notion of approximate unitary conjugacy in [41, Definition 2.8] is only a subequivalence relation in general.
+Indeed, symmetry fails when the cocycle representations considered are not extendible ([41, Remark 2.9]).
+However, the topology that we consider does indeed make it an equivalence relation.
 Lemma 5.9.
-If $(\phi,\mathbbm{v}),(\psi,\mathbbm{w}):(A,F,J)\to(B,G,I)$ are cocycle morphisms with associated families of linear maps $\{h^{X}\}$ and $\{l^{X}\}$ respectively, then $(\psi,\mathbbm{w})$ is an approximate unitary conjugate of $(\phi,\mathbbm{v})$ if and only if there exists a net of unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
-[…]
+Let $(\phi,\mathbbm{v}),(\psi,\mathbbm{w}):(A,F,J)\rightarrow(B,G,I)$ be cocycle representations between $\mathcal{C}$-C-algebras.
+Then if $(\psi,\mathbbm{w})\precapprox_{u}(\phi,\mathbbm{v})$, one also has that $(\phi,\mathbbm{v})\precapprox_{u}(\psi,\mathbbm{w})$.
+Proof.
+It can be computed from the definitions that for any unitary $u\in\mathcal{U}(\mathcal{M}(B))$, $X\in\mathcal{C}$ and $x\in F(X)$
+$(\mathbbm{v}_{u}*\mathbbm{v})_{X}(x\boxtimes_{\mathrm{Ad}(u)\phi}b)=u\rhd\mathbbm{v}_{X}(x\boxtimes_{\phi}u^{*}b).$
+In particular
+$\|\mathbbm{w}_{X}(x\boxtimes_{\psi}b)-(\mathbbm{v}_{u_{\lambda}}*\mathbbm{v})_{X}(x\boxtimes_{\mathrm{Ad}(u_{\lambda})\phi}b)\|$
+$=\lVert\mathbbm{w}_{X}(x\boxtimes_{\psi}b)-u_{\lambda}\rhd\mathbbm{v}_{X}(x\boxtimes_{\phi}u_{\lambda}^{*}b)\rVert$
+$=\lVert u_{\lambda}^{*}\rhd\mathbbm{w}_{X}(x\boxtimes_{\psi}b)-\mathbbm{v}_{X}(x\boxtimes_{\phi}u_{\lambda}^{*}b)\rVert$
+$=\lVert(\mathbbm{v}_{u_{\lambda}^{*}}*\mathbbm{w})_{X}(x\boxtimes_{\mathrm{Ad}(u_{\lambda}^{*})\psi}u_{\lambda}^{*}b)-\mathbbm{v}_{X}(x\boxtimes_{\phi}u_{\lambda}^{*}b)\rVert.$
+Thus the result follows. ∎
+In light of Lemma 5.9, we simply call two cocycle representations
+$(\psi,\mathbbm{w}),(\phi,\mathbbm{v}):(A,F,J)\rightarrow(B,G,I)$
+such that $(\psi,\mathbbm{w})\precapprox_{u}(\phi,\mathbbm{v})$ approximately unitarily equivalent and we denote it rather by $(\phi,\mathbbm{v})\approx_{u}(\psi,\mathbbm{w})$.
+In general, we will be interested in these notions when $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are cocycle morphisms, so let us record the following immediate consequence of Lemma 5.7.
+Lemma 5.10.
+If $(\phi,\mathbbm{v}),(\psi,\mathbbm{w}):(A,F,J)\to(B,G,I)$ are cocycle morphisms with associated families of linear maps $\{h^{X}\}$ and $\{l^{X}\}$ respectively, then $(\psi,\mathbbm{w})$ is approximately unitarily equivalent to $(\phi,\mathbbm{v})$ if and only if there exists a net of unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
+$\|l^{X}(x)-u_{\lambda}\rhd h^{X}(x)\lhd u_{\lambda}^{*}\|\overset{\lambda}{\longrightarrow}0$
 for any $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$.
-Proof.
-By definition, $(\psi,\mathbbm{w})$ is an approximate unitary conjugate of $(\phi,\mathbbm{v})$ if and only if there exists a net of unitaries $u_{\lambda}$ in $\mathcal{M}(B)$ inducing a net of cocycle morphisms $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})$ such that $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})\circ(\phi,\mathbbm{v})$ converges to $(\psi,\mathbbm{w})$.
-Suppose the cocycle morphisms $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})$ have associated families of linear maps $\{h_{u_{\lambda}}^{X}\}$.
-By Lemma 4.9, $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})\circ(\phi,\mathbbm{v})$ converges to $(\psi,\mathbbm{w})$ if and only if $h_{u_{\lambda}}^{X}\circ h^{X}$ converges pointwise to $l^{X}$ for any $X\in\mathrm{Irr}(\mathcal{C})$.
-Moreover, by Lemma 4.12,
-$h_{u_{\lambda}}^{X}(h^{X}(x))=u_{\lambda}\rhd h^{X}(x)\lhd u_{\lambda}^{*}$
-for all $X\in\mathcal{C}$ and all $x\in F(X)$, so the conclusion follows. ∎
-Remark 4.15.
-There is a key difference to Szabó’s notion of approximate unitary conjugacy in [40, Definition 2.8].
-The notion of approximate unitary conjugacy in Definition 4.13 is symmetric, even when restricted to possibly degenerate morphisms.
-In Lemma 4.16, we show that approximate unitary conjugacy implies approximate unitary equivalence.
-This is a consequence of the difference in topologies (see Example 4.10).
-Lemma 4.16.
-Let $(\phi,h),(\psi,l):(A,F,J)\to(B,G,I)$ be cocycle morphisms. If $(\psi,l)$ is an approximate unitary conjugate of $(\phi,h)$, then $(\phi,h)$ and $(\psi,l)$ are approximately unitarily equivalent.
-Proof.
-Suppose there exist unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
-[…]
-for any $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$. We show that
-[…]
-for any $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$.
-Note that for any $X\in\mathcal{C}$ and any $x\in F(X)$, we have that
-$h_{u_{\lambda}^{*}}^{X}(h_{u_{\lambda}}^{X}(h^{X}(x)))=u_{\lambda}^{*}\rhd h_{u_{\lambda}}^{X}(h^{X}(x))\lhd u_{\lambda}=1_{\mathcal{M}(B)}\rhd h^{X}(x)\lhd 1_{\mathcal{M}(B)}.$
-But $G(X)$ is non-degenerate, so $1_{\mathcal{M}(B)}\rhd h^{X}(x)\lhd 1_{\mathcal{M}(B)}=h^{X}(x)$ by Lemma 1.11. Therefore, for any $X\in\mathcal{C}$ and any $x\in F(X)$,
-$\|h_{u_{\lambda}^{*}}^{X}\circ l^{X}(x)-h^{X}(x)\|=\|h_{u_{\lambda}^{*}}^{X}(l^{X}(x)-h_{u_{\lambda}}^{X}\circ h^{X}(x))\|.$
-Take $y_{\lambda}=l^{X}(x)-h_{u_{\lambda}}^{X}\circ h^{X}(x)$ and note that
-$\langle h_{u_{\lambda}^{*}}^{X}(y_{\lambda}),h_{u_{\lambda}^{*}}^{X}(y_{\lambda})\rangle_{B}=\mathrm{Ad}(u_{\lambda}^{*})(\langle y_{\lambda},y_{\lambda}\rangle_{B}).$
-[…]
 We finish this section by defining asymptotic unitary equivalence for cocycle representations.
 Definition 5.11.
-[…]
+Let $\mathcal{C}\overset{F}{\curvearrowright}A$ and $\mathcal{C}\overset{G}{\curvearrowright}B$ be actions of $\mathcal{C}$ on $\mathrm{C}^{*}$-algebras AA and $B$ respectively and let $(\phi,\mathbbm{v}),(\psi,\mathbbm{w}):(A,F,J)\to(B,G,I)$ be cocycle representations.
+We say that $(\psi,\mathbbm{w})$ is asymptotically unitarily equivalent to $(\phi,\mathbbm{v})$ if there exists a strictly continuous map $u:[0,\infty)\to\mathcal{U}(\mathcal{M}(B))$ such that
 $\psi(a)=\lim\limits_{t\to\infty}u_{t}\phi(a)u_{t}^{*},$
 and
-$\|\mathbbm{w}_{X}(x\boxtimes_{\psi}b)-(\mathbbm{v}_{u_{t}}*\mathbbm{v})_{X}(x\boxtimes_{\phi}b)\|\overset{t\to\infty}{\longrightarrow}0$
-for all $a\in A$, $b\in B$, $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$.
-This will be denoted by $(\phi,\mathbbm{v})\cong_{u}(\psi,\mathbbm{w})$.
-[^15] Following Definition 4.13, we could have defined asymptotic unitary conjugacy.
-However, this implies asymptotic unitary equivalence as in Lemma 4.16.
-The same argument as in Lemma 4.14 gives the following equivalent characterisation for asymptotic unitary equivalence of cocycle morphisms.
+$\|\mathbbm{w}_{X}(x\boxtimes_{\psi}b)-(\mathbbm{v}_{u_{t}}*\mathbbm{v})_{X}(x\boxtimes_{\mathrm{Ad}(u_{t})\phi}b)\|\overset{t\to\infty}{\longrightarrow}0$
+for all $a\in A$, $b\in B$, $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$. This will be denoted by $(\phi,\mathbbm{v})\cong_{u}(\psi,\mathbbm{w})$.
+The same argument as in Lemma 5.9 shows that asymptotic unitary equivalence is an equivalence relation.
+Moreover, the same argument as in Lemma 5.10 gives the following equivalent characterisation for asymptotic unitary equivalence of cocycle morphisms.
 Lemma 5.12.
@@
 Then, $v\psi_{1}^{\dagger}(u)\rhd l_{1}^{X}(h_{1}^{X}(x))\lhd(v\psi_{1}^{\dagger}(u))^{*}=l_{2}^{X}(h_{2}^{X}(x))$.
-In general, the vertical composition in the $2$-category of correspondences is not well defined outside of the extendible setting.
+In general, the vertical composition in the 22-category of correspondences is not well defined outside of the extendible setting.
 We will finish by recording the following lemma which will be used in Section 7.
@@
 Suppose that $(\Phi,H)$ and $(\Psi,L)$ are cocycle conjugacies such that
-$(\Phi,H)\cong_{u}(\phi,h)\quad\mathrm{and}\quad(\Psi,L)\cong_{u}(\psi,l).$
+[…]
 By Lemma 5.6, composition is jointly continuous. Moreover, since $\Psi$ is extendible, $(\Psi\circ\Phi,L\circ H)\cong_{u}(\psi\circ\phi,l\circ h)$ by Remark 5.13. ∎
@@
 We check the conditions of the second theorem in [13].
-Our underlying category is the subcategory of $\mathrm{C}^{*}_{\mathcal{C}}$ of separable $*{}^{*}$-algebras, together with extendible cocycle morphisms (see Definition 3.17).
+Our underlying category is the subcategory of $\mathrm{C}^{*}_{\mathcal{C}}$ of separable C-algebras, together with extendible cocycle morphisms (see Definition 3.17).
 In this category, Lemma 5.7 gives us a notion of inner automorphisms in the sense of [13].
@@
 However, we also decide to give the technical argument in full, picking up some more general results in the process.
-We start by introducing the setup for performing approximate intertwining arguments in the spirit of [40, Definition 3.3].
+We start by introducing the setup for performing approximate intertwining arguments in the spirit of [41, Definition 3.3].
 Definition 6.1.
@@
 fitting into the family of not necessarily commutative diagrams
-[…]
+(6.1) $\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${F_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s_{n}^{X}}$$\scriptstyle{h_{n}^{X}}$${F_{n+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\dots}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{n}^{X}}$$\scriptstyle{l_{n}^{X}}$${G_{n+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{n+1}^{X}}$${\dots\quad.}$
 We will call the collection of diagrams (6.1) an approximate cocycle intertwining, if the following hold: There exist an increasing sequence of finite sets $K_{n}\subset\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$, finite sets $\mathcal{F}_{n}^{X}\subset F_{n}(X)$ and $\mathcal{G}_{n}^{X}\subset G_{n}(X)$ for any $X\in K_{n}$, and numbers $\delta_{n}>0$ satisfying
@@
 Remark 6.2.
-The conditions listed above are in the spirit of Elliott approximate intertwining arguments, as seen for example in [38, Definition $2.3.1$] or [40, Definition $3.3$].
+The conditions listed above are in the spirit of Elliott approximate intertwining arguments, as seen for example in [40, Definition $2.3.1$] or [41, Definition $3.3$].
 Note that applying $X=1_{\mathcal{C}}$ in conditions (i)-(iv) above, together with the (vi) recover the usual assumptions in the two-sided Elliott intertwining argument.
@@
 This is why conditions (i)-(iv) above are phrased for irreducible elements in the category.
-These conditions resemble the assumptions needed in [40, Definition 3.3] in the case of a twisted group action, although recall that the topology on cocycle morphisms is different.
-Moreover, (i)-(ii) encode that the diagrams in 5.1 approximately commute in the topology described in Lemma 4.9.
+These conditions resemble the assumptions needed in [41, Definition 3.3] in the case of a twisted group action, although recall that the topology on cocycle morphisms is different.
+Moreover, (i)-(ii) encode that the diagrams in 6.1 approximately commute in the topology described in Lemma 5.4.
 Furthermore, in (iv) we are making crucial use of our definition of the functors $F_{n}$ and $G_{n}$ to assume that $F_{n}(X)$ and $G_{n}(X)$ have a countable dense subset for any $X\in\mathcal{C}$ and any $n\in\mathbb{N}$.
@@
 The limits in the formulae (6.3) and (6.5) are taken in the topologies induced by the respective right inner products.
-Remark 5.4.
-To deduce the formulae above, one can work backwards. If the cocycle conjugacy $\theta$ was to exist, the fact that $A$ and $B$ are inductive limits would force the diagrams
-[…]
-to commute for all $k\in\mathbb{N}$.
-This immediately gives (5.2).
-Considering the associated families of linear maps for the cocycle morphisms in (5.2), we would need a similar diagram involving the linear maps to commute, which gives (5.3).
-The formulae in (5.4) and (5.5) can be obtained in exactly the same fashion.
 Proof.
 We will show that the limits in (6.2) and (6.3) exist and that the pair $(\theta,\{s^{X}\}_{X\in\mathcal{C}})$ is a cocycle morphism.
-Firstly, note that the limit in (5.2) exists and $\theta:A\to B$ is a $*{}^{*}$-homomorphism by the general Elliott intertwining, as can be found for example in [38, Proposition 2.3.2].
-We first show that the limit in (5.3) exists for all $X\in\mathrm{Irr}(\mathcal{C}).$ We can employ a similar argument to the one in [38, Proposition 2.3.2].
+Firstly, note that the limit in (6.2) exists and $\theta:A\to B$ is a -homomorphism by the general Elliott intertwining, as can be found for example in [40, Proposition 2.3.2].
+We first show that the limit in (6.3) exists for all $X\in\mathrm{Irr}(\mathcal{C}).$ We can employ a similar argument to the one in [40, Proposition 2.3.2].
 Let $X\in K_{n}$ and given $x_{n}\in F_{n}(X)$, by condition (iv) in Definition 6.1, we may assume $x_{n}$ is contained in $\bigcup_{m>n}(h_{n,m}^{X})^{-1}(\mathcal{F}_{m}^{X})$.
@@
 Consider the possibly non-commutative diagram
-[…]
+(6.6) ${F_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h_{n,m}^{X}}$${F_{m}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s_{m}^{X}}$$\scriptstyle{h_{m}^{X}}$${F_{m+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s_{m+1}^{X}}$${G_{m+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{m+1}^{X}}$$\scriptstyle{l_{m+1}^{X}}$${G_{m+2}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\quad l_{m+2,\infty}^{X}}$${G(X).}$
 Using triangle inequality, it follows that
@@
 Also, by (iii) $s_{m}^{X}(\mathcal{F}_{m}^{X})\subset\mathcal{G}_{m+1}^{X}$, so we may combine (i), (ii), and that $s_{m+1}^{X}$ is contractive to get that
-$\|(s_{m+1}^{X}\circ h_{m}^{X})(x_{m})-(l_{m+1}^{X}\circ s_{m}^{X})(x_{m})\|\leq\delta_{m}+\delta_{m+1},\quad\forall\quad x_{m}\in\mathcal{F}_{m}^{X}\subseteq F_{m}(X).$
+$\|(s_{m+1}^{X}\circ h_{m}^{X})(x_{m})-(l_{m+1}^{X}\circ s_{m}^{X})(x_{m})\|\leq\delta_{m}+\delta_{m+1},\quad\forall\ x_{m}\in\mathcal{F}_{m}^{X}\subseteq F_{m}(X).$
 Moreover, since $l_{m+2,\infty}^{X}$ is contractive, it follows that
@@
 For this, we check the conditions in Lemma 3.10.
-For any $n\in\mathbb{N}$, $X\in\mathrm{Irr}(\mathcal{C})$, $x_{n}\in F_{n}(X)$, and any $a_{n},a_{n}^{\prime}\in A_{n}$, as $(\phi_{n,\infty},h_{n,\infty})$ is a cocycle morphism, it follows that
-$s^{X}(\phi_{n,\infty}(a_{n})\rhd h_{n,\infty}^{X}(x_{n})\lhd\phi_{n,\infty}(a_{n}^{\prime}))=s^{X}(h_{n,\infty}^{X}(a_{n}\rhd x_{n}\lhd a_{n}^{\prime})).$
-Since the limit in (5.3) exists and $l_{k+1,\infty}^{X},s_{k}^{X},h_{n,k}^{X}$ are cocycle morphisms, a direct calculation shows that
-$s^{X}(h_{n,\infty}^{X}(a_{n}\rhd x_{n}\lhd a_{n}^{\prime}))=\theta(\phi_{n,\infty}(a_{n}))\rhd s^{X}(h_{n,\infty}^{X}(x_{n}))\lhd\theta(\phi_{n,\infty}(a_{n}^{\prime})).$
 Given $X,Y\in\mathrm{Irr}(\mathcal{C})$, consider a morphism $f\in\mathrm{Hom}(X,Y)$.
-Since $\mathrm{Hom}(X,Y)$ is 00 if $X\neq Y$ and $\mathrm{Hom}(X,X)\cong\mathbb{C}$, it follows immediately that $G(f)\circ s^{X}=s^{Y}\circ F(f)$.
+Since $\mathrm{Hom}(X,Y)$ is 0 if $X\neq Y$ and $\mathrm{Hom}(X,X)\cong\mathbb{C}$, it follows immediately that $G(f)\circ s^{X}=s^{Y}\circ F(f)$.
 Next, we want to show that $\theta(\langle x,y\rangle_{A})=\langle s^{X}(x),s^{X}(y)\rangle_{B}$ for any $x,y\in F(X)$.
@@
 Then, we check (iv) of Lemma 3.10. Let $X,Y\in\mathcal{C}$. Since $(\psi_{k+1,\infty}\circ\theta_{k}\circ\phi_{n,k},l_{k+1,\infty}\circ s_{k}\circ h_{n,k})$ is a cocycle morphism, the diagram
-[…]
+$F_{n}(Y)\boxtimes F_{n}(X)$$F_{n}(X\otimes Y)$$G(Y)\boxtimes G(X)$$G(X\otimes Y)$$\scriptstyle{(l_{k+1,\infty}^{Y}\circ s_{k}^{Y}\circ h_{n,k}^{Y})\boxtimes(l_{k+1,\infty}^{X}\circ s_{k}^{X}\circ h_{n,k}^{X})}$$\scriptstyle{J_{X,Y}^{(n)}}$$\scriptstyle{l_{k+1,\infty}^{X\otimes Y}\circ s_{k}^{X\otimes Y}\circ h_{n,k}^{X\otimes Y}}$$\scriptstyle{I_{X,Y}}$
 commutes.
@@
 By density it now follows that $I_{X,Y}\circ(s^{Y}\boxtimes s^{X})=s^{X\otimes Y}\circ J_{X,Y}$.
-Finally, it follows from (5.2) that for any $a\in A$, $s^{1_{\mathcal{C}}}(a)=\theta(a)$.
 Hence, as $\mathcal{C}$ is semisimple, $(\theta,\{s^{X}\}_{X\in\mathrm{Irr}(\mathcal{C})})$ induces a cocycle morphism (see Remark 3.12).
@@
 It follows in the same way that $(\kappa,\{r^{X}\}_{X\in\mathcal{C}})$ given by the formulae in (6.4) and (6.5) yields a well-defined cocycle morphism.
-Moreover, the fact that $\theta$ and $\kappa$ are mutually inverse isomorphisms follows from [38, Proposition 2.3.2].
+Moreover, the fact that $\theta$ and $\kappa$ are mutually inverse isomorphisms follows from [40, Proposition 2.3.2].
 Finally, it remains to check that for any $X\in\mathcal{C}$, $r^{X}\circ s^{X}=\mathrm{id}_{F(X)}$ and $s^{X}\circ r^{X}=\mathrm{id}_{G(X)}$.
@@
 (6.7) $r^{X}\circ l_{k+1,\infty}^{X}\circ s_{k}^{X}\circ h_{n,k}^{X}=\lim\limits_{m\to\infty}(h_{m,\infty}^{X}\circ r_{m}^{X}\circ l_{k+1,m}^{X}\circ s_{k}^{X}\circ h_{n,k}^{X}).$
-Using conditions (i) and (ii) of Definition 5.1, it follows that the right hand side evaluated at $x\in(h_{n,k}^{X})^{-1}(\mathcal{F}_{k}^{X})$ converges to $h_{n,\infty}(x)$ uniformly as $k$ tends to infinity.
+Using conditions (i),(ii) and (vi) of Definition 6.1, it follows that the right hand side evaluated at $x\in(h_{n,k}^{X})^{-1}(\mathcal{F}_{k}^{X})$ converges to $h_{n,\infty}(x)$ uniformly on $k$ as $k$ tends to infinity.
 Now, taking the limit as $k$ goes to infinity in (6.7) and using condition (iv) of Definition 6.1, we get that $r^{X}\circ s^{X}\circ h_{n,\infty}^{X}=h_{n,\infty}^{X}$.
@@
 We now use Theorem 6.3 to show that if we assume that the diagrams in (6.1) commute up to approximate unitary equivalence, then there exist mutually inverse cocycle conjugacies as in Theorem 6.3.
-The proof follows in a similar fashion to [40, Theorem 3.6] and [38, Corollary 2.3.3].
+The proof follows in a similar fashion to [41, Theorem 3.6] and [40, Corollary 2.3.3].
 Theorem 6.4.
-Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects.
+Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects.
 Let $(F_{n},J^{(n)}):\mathcal{C}\curvearrowright A_{n}$ and $(G_{n},I^{(n)}):\mathcal{C}\curvearrowright B_{n}$ be sequences of actions on separable $\mathrm{C}^{*}$-algebras.
@@
 fitting into the not necessarily commutative collection of diagrams
-[…]
+(6.8) $\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${F_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s_{n}^{X}}$$\scriptstyle{h_{n}^{X}}$${F_{n+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\dots}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{n}^{X}}$$\scriptstyle{l_{n}^{X}}$${G_{n+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{n+1}^{X}}$${\dots\quad.}$
 Suppose that
@@
 For any $n\in\mathbb{N}$, and any $X\in K_{n}$, we can choose $t_{n}^{X}\in\mathbb{N}$ and finite sets $\{f_{m,n}^{X}\}_{1\leq m\leq t_{n}^{X}}\subset F_{n}(X)$ and $\{g_{m,n}^{X}\}_{1\leq m\leq t_{n}^{X}}\subset G_{n}(X)$ such that the inclusions
-(5.9) $\bigcup\limits_{k>n}(h_{n,k}^{X})^{-1}(\{f_{m,k}^{X}\}_{1\leq m\leq t_{k}^{X}})\subset F_{n}(X)\quad\mathrm{and}\quad\bigcup\limits_{k>n}(l_{n,k}^{X})^{-1}(\{g_{m,k}^{X}\}_{1\leq m\leq t_{k}^{X}})\subset G_{n}(X)$
+[…]
 are dense for all $n\in\mathbb{N}$. To simplify notation, we will write $\{f_{m,n}^{X}\}$ to denote the set $\{f_{m,n}^{X}\}_{1\leq m\leq t_{n}^{X}}$.
@@
 We claim that the diagram
-[…]
+(6.16) $\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${F_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{S_{n}^{X}}$$\scriptstyle{h_{n}^{X}}$${F_{n+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\dots}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G_{n}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R_{n}^{X}}$$\scriptstyle{l_{n}^{X}}$${G_{n+1}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R_{n+1}(X)}$${\dots\quad.}$
 is an approximate cocycle intertwining in the sense of Definition 6.1.
@@
 and $(\zeta_{m},S_{m})=\mathrm{Ad}(u_{m})\circ(\theta_{m},s_{m})$ for any $m\in\mathbb{N}$. Then, for all $n\geq 1$ and all $k>n$ we have that
-$\begin{array}[]{cl}(\psi_{n+1,k+1},l_{n+1,k+1})\circ(\theta_{n},s_{n})\hfil\\ \approx_{u}&(\theta_{k},s_{k})\circ(\kappa_{k},r_{k})\circ(\theta_{k-1},s_{k-1})\circ\ldots\circ(\kappa_{n+1},r_{n+1})\circ(\theta_{n},s_{n})\\ \approx_{u}&(\theta_{k},s_{k})\circ(\phi_{n,k},h_{n,k})\\ \approx_{u}&(\zeta_{k},S_{k})\circ(\phi_{n,k},h_{n,k}).\end{array}$
-Composing with $(\psi_{k+1,\infty},l_{k+1,\infty})$ yields that
+[…]
+Composing with $(\psi_{k+1,\infty},l_{k+1,\infty})$ then by 4.5 we have that
 $(\psi_{n+1,\infty},l_{n+1,\infty})\circ(\theta_{n},s_{n})\approx_{u}(\psi_{k+1,\infty},l_{k+1,\infty})\circ(\zeta_{k},S_{k})\circ(\phi_{n,k},h_{n,k}).$
@@
 Corollary 6.5.
-Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects.
+Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects.
 Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras.
@@
 This is a direct application of Theorem 6.4 with $A_{n}=A$, $B_{n}=B$, $\phi_{n}=\mathrm{id}_{A}$, $\psi_{n}=\mathrm{id}_{B}$, $k_{n}=\psi$, and $\theta_{n}=\phi$ for all $n\in\mathbb{N}$. ∎
-Remark 5.7.
+Remark $6.$
 Note that in Corollary 6.5, if the approximate unitary equivalences are realised by unitaries in the minimal unitisations, then we may drop the assumption of extendibility on the connecting morphisms.
@@
 This type of result appears for example in [18, 41] and it is used in successful classification results in [18, 17, 6].
-If $A$ is a separable $C^{*}$-algebra and $\eta:\mathbb{N}\to\mathbb{N}$ is any map with $\lim\limits_{n\to\infty}\eta(n)=\infty$, then it induces an endomorphism $\eta^{*}$ on $A_{\infty}$ via $\eta^{*}((a_{n})_{n})=(a_{\eta(n)})_{n}$.
-Moreover, if a $*{}^{*}$-tensor category $\mathcal{C}$ acts on a unital $*{}^{*}$-algebra $A$ by the triple $(A,F,J)$ with $F(X)$ algebraically finitely generated for all $X\in\mathcal{C}$, it follows from Lemma 2.6 that there exists an induced action on $A_{\infty}$ by the triple $(A_{\infty},F_{\infty},J^{\infty})$ (see also Remark 2.7).
+If AA is a separable $C^{*}$-algebra and $\eta:\mathbb{N}\to\mathbb{N}$ is any map with $\lim\limits_{n\to\infty}\eta(n)=\infty$, then it induces an endomorphism $\eta^{*}$ on $A_{\infty}$ via $\eta^{*}((a_{n})_{n})=(a_{\eta(n)})_{n}$.
+Moreover, if a C-tensor category $\mathcal{C}$ acts on a C-algebra AA by a finite rank triple $(A,F,J)$, it follows from Lemma 2.8 that there is an induced action on $A_{\infty}$ by the triple $(A_{\infty},F_{\infty},J^{\infty})$ (see also Remark 2.9).
 Then, a straightforward checking of the conditions in Lemma 3.10 shows that $\eta$ induces a cocycle morphism $(\eta^{*},r):(A_{\infty},F_{\infty},J^{\infty})\to(A_{\infty},F_{\infty},J^{\infty})$, where $r^{X}:F_{\infty}(X)\to F_{\infty}(X)$ is given by $r^{X}((\xi_{n})_{n})=((\xi_{\eta(n)})_{n})$ for any $X\in\mathcal{C}$ and any $(\xi_{n})_{n}\in F_{\infty}(X)$.
@@
 Lemma 6.7.
-Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions of a $*{}^{*}$-tensor category $\mathcal{C}$ on separable $\mathrm{C}^{*}$-algebras with $B$ unital and $G(X)$ be algebraically finitely generated as a right $B$-module for all $X\in\mathcal{C}$.
+Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions of a C-tensor category $\mathcal{C}$ on separable $\mathrm{C}^{*}$-algebras with $B$ unital and $(G,I)$ finite rank.
 Suppose that $(\phi,h):(A,F,J)\to(B_{\infty},G_{\infty},I^{\infty})$ is a cocycle morphism such that for any map $\eta:\mathbb{N}\to\mathbb{N}$ with $\lim\limits_{n\to\infty}\eta(n)=\infty$, the cocycle morphisms $(\phi,h)$ and $(\eta^{*},r)\circ(\phi,h)$ are approximately unitarily equivalent.
-For each $X\in\mathcal{C}$, let $(h_{n}^{X})_{n}$ be any lift of $h^{X}$.
+For each $X\in\mathcal{C}$, let $(h_{n}^{X})_{n}\colon F(X)\rightarrow l^{\infty}(\mathbb{N},G(X))$ be a linear lift of $h^{X}$.
 Then, for every finite set $K\subseteq\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$, finite sets $\mathcal{F}^{X}\subseteq F(X)$ for $X\in K$, $\epsilon>0$, and $m\in\mathbb{N}$, there is an integer $k\geq m$ such that for every integer $n\geq k$ there is a unitary $u\in B$ for which
@@
 Suppose that there exists a finite set $K\subseteq\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$, finite sets $\mathcal{F}^{X}\subseteq F(X)$ for $X\in K$, $\epsilon>0$, and $m\in\mathbb{N}$ such that for every $k\geq m$, there exists $n_{k}\geq k$ for which
-(5.19) $\max\limits_{x\in\mathcal{F}^{X}}\|u_{k}\rhd h_{n_{k}}^{X}(x)\lhd u_{k}-h_{k}^{X}(x)\|\geq\epsilon,$
+(6.19) $\max\limits_{x\in\mathcal{F}^{X}}\|u_{k}\rhd h_{n_{k}}^{X}(x)\lhd u_{k}^{*}-h_{k}^{X}(x)\|\geq\epsilon,$
 for any $X\in K$, $x\in\mathcal{F}^{X}$, and every unitary $u_{k}\in B$.
 Let $\eta:\mathbb{N}\to\mathbb{N}$ be the map $\eta(k)=n_{k}$ whenever $k\geq m$ and $\eta(k)=1$ for $k<m$.
-As $\eta_{k}\geq k$ for $k\geq m$, it follows that $\lim\limits_{k\to\infty}\eta(k)=\infty$.
+As $\eta(k)\geq k$ for $k\geq m$, it follows that $\lim\limits_{k\to\infty}\eta(k)=\infty$.
 Moreover, $(\phi,h)$ and $(\eta^{*},r)\circ(\phi,h)$ are approximately unitarily equivalent, so there exists a unitary $u\in B_{\infty}$ for which
@@
 Theorem 6.8.
-Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects.
-Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras with $B$ unital and $G(X)$ be algebraically finitely generated as a right $B$-module for all $X\in\mathcal{C}$.
+Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects.
+Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras with $B$ unital and $(G,I)$ finite rank.
 Let $(\phi,h):(A,F,J)\to(B_{\infty},G_{\infty},I^{\infty})$ be a cocycle morphism.
@@
 Let $K_{n}\subseteq\mathrm{Irr}(\mathcal{C})$ be an increasing sequence of finite sets containing $1_{\mathcal{C}}$ such that $\bigcup_{n=1}^{\infty}K_{n}=\mathrm{Irr}(\mathcal{C})$, and $\mathcal{F}_{n}^{X}\subseteq F(X)$ be an increasing sequence of finite sets such that $\bigcup_{n=1}^{\infty}\mathcal{F}_{n}^{X}$ is dense in $F(X)$ for any $X\in\mathrm{Irr}(\mathcal{C})$.
+Let $(h_{n}^{X})_{n}:F(X)\rightarrow l^{\infty}(\mathbb{N},G(X))$ be a linear lifting.
 Recursively applying Lemma 6.7 to $K=K_{n}$, $\mathcal{F}^{X}=\mathcal{F}_{n}^{X}$, $\epsilon=\frac{1}{2^{n}}$, one may pick $k_{0}=1<k_{1}<k_{2}<\ldots$ and unitaries $u_{n}\in B$ for $n\in\mathbb{N}$ such that
@@
 Let $v_{n}=u_{1}u_{2}\ldots u_{n}$, we claim that $l^{X}(x):=\lim\limits_{n\to\infty}v_{n}\rhd h_{k_{n}}^{X}(x)\lhd v_{n}^{*}$ is well-defined for any $X\in\mathrm{Irr}(\mathcal{C})$ and any $x\in F(X)$.
-First note that $h^{X}$ is contractive by (iii) of Lemma 3.11, so $\limsup_{n\to\infty}\|h^{X}_{n}(x)\|\leq\|x\|$ for any $x\in F(X)$.
-In particular, the sequence $h^{X}_{k_{n}}$ is norm bounded.
-We will now show that the sequence $(v_{n}\rhd h_{k_{n}}^{X}(x)\lhd v_{n}^{*})_{n\geq 1}$ is Cauchy for any $X\in\mathrm{Irr}(\mathcal{C})$ and any $x\in F(X)$.
-Given $\epsilon>0$, $n\in\mathbb{N}$, $X\in K_{n}$, and $x\in F(X)$, there exists $m\in\mathbb{N}$ and $y\in\mathcal{F}_{m}^{X}$ such that $\|x-y\|<\epsilon/3$.
-Then, for $j>l\geq m$, we may use (5.20) repeatedly, together with triangle inequality, to achieve that
-$\|v_{j}\rhd h_{k_{j}}^{X}(y)\lhd v_{j}^{*}-v_{l}\rhd h_{k_{l}}^{X}(y)\lhd v_{l}^{*}\|$ $\leq\sum\limits_{i=l+1}^{j}2^{-i}<\epsilon/3$
-for any $l$ greater than some large enough $N$.
-Moreover, since $h_{k_{l}}^{X}$ is a contractive linear map for any $l\in\mathbb{N}$ and any $X\in\mathrm{Irr}(\mathcal{C})$, it follows that $\|h_{k_{l}}^{X}(x)-h_{k_{l}}^{X}(y)\|<\epsilon/3$ Therefore,
-$\|v_{j}\rhd h_{k_{j}}^{X}(x)\lhd v_{j}^{*}-v_{l}\rhd h_{k_{l}}^{X}(x)\lhd v_{l}^{*}\|<\epsilon,$
-for all $l>N$, so the sequence is Cauchy. Thus, $l^{X}:F(X)\to G(X)$ is a well-defined map for every $X\in\mathrm{Irr}(\mathcal{C})$ and it is linear.
-We let $\psi:A\to B$ be the linear map given by $\psi=l^{1_{\mathcal{C}}}$.
-If $v\in B_{\infty}$ is the unitary given by $(v_{n})_{n}$ and $\eta:\mathbb{N}\to\mathbb{N}$ is the map $\eta(n)=k_{n}$, we have that
-$(\psi,l^{X})=(\mathrm{Ad}(v),h_{v}^{X})\circ(\eta^{*},r^{X})\circ(\phi,h^{X})$
+First let $y\in\bigcup_{n\in\mathbb{N}}\mathcal{F}_{n}^{X}$ for $X\in\mathrm{Irr}(\mathcal{C})$ and $\varepsilon>0$.
+Then there exists $n\in\mathbb{N}$ such that $X\in K_{n}$, $y\in\mathcal{F}_{n}^{X}$ and $2^{-n}<\varepsilon$.
+In particular, for any $j>l\geq n$ we may use (6.20) repeatedly, together with triangle inequality, to achieve that
+$\|v_{j}\rhd h_{k_{j}}^{X}(y)\lhd v_{j}^{*}-v_{l}\rhd h_{k_{l}}^{X}(y)\lhd v_{l}^{*}\|$ $\leq\sum\limits_{i=l+1}^{j}2^{-i}<\epsilon.$
+As $\varepsilon$ is arbitrary $v_{n}\rhd h_{k_{n}}^{X}(y)\lhd v_{n}^{*}$ is Cauchy for any $X\in\mathrm{Irr}(\mathcal{C})$ and any $y\in\bigcup_{n\in\mathbb{N}}\mathcal{F}_{n}^{X}$.
+This implies that, letting $(v_{n})_{n}=V\in\mathcal{U}(B_{\infty})$, $\eta:\mathbb{N}\rightarrow\mathbb{N}$ be given by $\eta(n)=k_{n}$ and their induced cocycle morphisms by $(\mathrm{Ad}(V),h_{V})$ and $(\eta^{*},r)$,
+$(\mathrm{Ad}(V),h_{V})\circ(\eta^{*},r)\circ(\phi,h):(A,F,J)\rightarrow(B_{\infty},G_{\infty},I^{\infty})$
+is a cocycle morphism such that it sends a dense subspace of $F(X)$ into the closed subspace $G(X)\subset G(X)_{\infty}$ for each $X\in\mathrm{Irr}(\mathcal{C})$. Thus
+$(h_{V})^{X}\circ r^{X}\circ h^{X}:F(X)\to G(X)\subset G(X)_{\infty}$
+by continuity for any $X\in\mathrm{Irr}(\mathcal{C})$. This precisely shows that
+$\lim_{n\rightarrow\infty}v_{n}\rhd h_{k_{n}}^{X}(x)\lhd v_{n}^{*}$
+exists for all $x\in F(X)$. Let $\psi:A\to B$ be the linear map given by $\psi=l^{1_{\mathcal{C}}}$. By construction we have that
+$(\psi,l)=(\mathrm{Ad}(V),h_{V})\circ(\eta^{*},r)\circ(\phi,h)$
 for any $X\in\mathrm{Irr}(\mathcal{C})$.
-Hence, $\psi$ is a $*{}^{*}$-homomorphism and the pair $(\psi,l)$ induces a cocycle morphism.
+Hence, $\psi$ is a -homomorphism and the pair $(\psi,l)$ induces a cocycle morphism.
 Extending by linearity, we can define $l^{X}$ in the same way for any $X\in\mathcal{C}$ and it is straightforward to see that these maps are linear.
-Moreover, $(\psi,l)$ is approximately unitarily equivalent to $(\phi,h)$, which finishes the proof.
+Moreover, $(\psi,l)$ is approximately unitarily equivalent to $(\phi,h)$.
+Finally, to show that the approximate unitary equivalence of $(\psi,l)$ and $(\phi,h)$ implies unitary equivalence follows as in [18, Lemma 4.1].
 ∎
 Remark 6.9.
-As noted in Remark 2.7 the assumption that $G(X)$ are algebraically finitely generated in Theorem 5.9 is immediate whenever the acting category $\mathcal{C}$ is a unitary tensor category.
+Note that in Theorem 6.8, if the approximate unitary equivalences are realised by unitaries in the minimal unitisations, then we may drop the assumption of unitality on $B$.
+Moreover, as noted in Remark 2.9, the assumption that $G(X)$ are finite rank in Theorem 6.8 is immediate whenever the acting category $\mathcal{C}$ is a unitary tensor category and $B$ is unital.
 7. One sided intertwining arguments
-We start this section by showing a tensor category equivariant adaptation of the classical one-sided intertwining argument (see [38, Proposition 2.3.5]).
-A group equivariant version of this result can be found in [40, Proposition 4.3].
+We start this section by showing a tensor category equivariant adaptation of the classical one-sided intertwining argument (see [40, Proposition 2.3.5]).
+A group equivariant version of this result can be found in [41, Proposition 4.3].
 We end this section by proving an asymptotic Elliott two-sided intertwining (see Theorem 7.4).
+First note that as $\mathrm{Ad}(v)$ is an isometric -homomorphism for any unitary $v\in\mathcal{U}(\mathcal{M}(A))$, injectivity of a cocycle morphism, which coincides with the injectivity of it’s underlying -homomorphism, is preserved under both approximate and asymptotic unitary equivalence.
 Theorem 7.1.
-Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects.
-Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $*{}^{*}$-algebras and $(\varphi,h):(A,F,J)\rightarrow(B,G,I)$ an injective cocycle morphism.
+Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects.
+Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable C-algebras and $(\varphi,h):(A,F,J)\rightarrow(B,G,I)$ an injective cocycle morphism.
 Then $(\varphi,h)$ is asymptotically unitarily equivalent to a cocycle conjugacy if and only if:
@@
     $\mathrm{dist}(z_{1}^{*}\rhd y\lhd z_{1},h^{X}(F(X)))\leq\varepsilon$ for all $X\in K$ and $y\in\mathcal{G}^{X}$.
+[^18] Note that the injectivity assumption is redundant for the only if statement as it follows from the hypothesis.
+However, this assumption it is still required for the if direction.
 Proof.
-The proof of this Theorem will follow the proof of [40, Proposition 4.3] closely.
+The proof of this Theorem will follow the proof of [41, Proposition 4.3] closely.
 We start by showing the “only if” statement.
@@
 (7.3) $\sup\limits_{t\geq n_{2}}\|H^{X}(x)-w_{t}\rhd h^{X}(x)\lhd w_{t}^{*}\|\leq\varepsilon,\ X\in K,\ x\in(H^{X})^{-1}(w_{n_{1}}\rhd\mathcal{G}^{X}\lhd w_{n_{1}}^{*}).$
-We claim that the unitary path $z_{t}=w_{n_{1}}^{*}w_{(1-t)n_{1}+tn_{2}}$ satisfies A.1 and A.2.
+We claim that the unitary path $z_{t}=w_{n_{1}}^{*}w_{(1-t)n_{1}+tn_{2}}$ satisfies A.1 and 18.
 Indeed, it is a strictly continuous path $z:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ with $z_{0}=1$.
 Moreover, for $X\in K$, $t\in[0,1]$ and $x\in\mathcal{F}^{X}$, using (7.2) we have that
-$\|h^{X}(x)-z_{t}$ $*\displaystyle\rhd h^{X}(x)\lhd z_{t}^{*}\|$
+$\|h^{X}(x)-z_{t}$ $\rhd h^{X}(x)\lhd z_{t}^{*}\|$
 $=\|h^{X}(x)-w_{n_{1}}^{*}w_{(1-t)n_{1}+tn_{2}}\rhd h^{X}(x)\lhd w_{(1-t)n_{1}+tn_{2}}^{*}w_{n_{1}}\|$
@@
 $\leq\varepsilon.$
-This shows condition A.1. We now turn to condition A.2. Let $X\in K$, $y\in\mathcal{G}^{X}$, and $x=(H^{X})^{-1}(w_{n_{1}}\rhd y\lhd w_{n_{1}}^{*})$. Then, we get that
+This shows condition A.1. We now turn to condition 18. Let $X\in K$, $y\in\mathcal{G}^{X}$, and $x=(H^{X})^{-1}(w_{n_{1}}\rhd y\lhd w_{n_{1}}^{*})$. Then, we get that
 $\|z_{1}^{*}\rhd y\lhd z_{1}-h^{X}(x)\|$
@@
 $=\|H^{X}(x)-w_{n_{2}}\rhd h^{X}(x)\lhd w_{n_{2}}^{*}\|$
-(6.3)ϵ.superscriptitalic-(6.3italic-)absentitalic-${{\leq}}\epsilon.$
-As $x\in F(X)$, condition A.2 holds.
+[…]
+As $x\in F(X)$, condition 18 holds.
 We now turn to the if direction.
-To prove this statement we will use A.1-A.2 to construct a strictly continuous path of unitaries $v_{t}\in\mathcal{U}(\mathcal{M}(B))$ for $t\in[0,\infty)$ such that $(\mathrm{Ad}(v_{t}),h_{v_{t}})\circ(\varphi,h)$ converges to a cocycle conjugacy.
+To prove this statement we will use A.1-18 to construct a strictly continuous path of unitaries $v_{t}\in\mathcal{U}(\mathcal{M}(B))$ for $t\in[0,\infty)$ such that $(\mathrm{Ad}(v_{t}),h_{v_{t}})\circ(\varphi,h)$ converges to a cocycle conjugacy.
 Let $\{x_{n}^{X}\}_{n\in\mathbb{N}}\subset F(X)$, $\{y_{n}^{X}\}_{n\in\mathbb{N}}\subset G(X)$ be countable dense subsets for any $X\in\mathrm{Irr}(\mathcal{C})$ and $K_{n}$ increasing finite subsets of $\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$ such that $\mathrm{Irr}(\mathcal{C})=\bigcup_{n\in\mathbb{N}}K_{n}$.
-Firstly, use A.1-A.2 to find $x_{1,1}^{X}\in F(X)$ for $X\in K_{1}$ and $z^{(1)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that $z_{0}^{(1)}=1$ and for $0\leq t\leq 1$
+Firstly, use A.1-18 to find $x_{1,1}^{X}\in F(X)$ for $X\in K_{1}$ and $z^{(1)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that $z_{0}^{(1)}=1$ and for $0\leq t\leq 1$
   1. (i)
@@
     $\|(z_{1}^{(1)})^{*}\rhd y_{1}^{X}\lhd z_{1}^{(1)}-h^{X}(x_{1,1}^{X})\|\leq 1/2$ for $X\in K_{1}$.
-Again use A.1-A.2 to find $x_{2,1}^{X}$, $x_{2,2}^{X}$ in $F(X)$ for $X\in K_{2}$, $z^{(2)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that $z^{(2)}_{0}=1$ and for every $0\leq t\leq 1$
+Again use A.1-18 to find $x_{2,1}^{X}$, $x_{2,2}^{X}$ in $F(X)$ for $X\in K_{2}$, $z^{(2)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that $z^{(2)}_{0}=1$ and for every $0\leq t\leq 1$
   1. (i)
@@
   3. S.3
-    $\|(z_{1}^{(n)})^{*}\ldots(z_{1}^{(1)})^{*}\rhd y_{j}^{X}\lhd z_{1}^{(1)}\ldots z_{1}^{(n)}-h^{X}(x_{n,j}^{X})\|\leq 2^{-n}$ for $X\in K_{n}$ and $1\leq j\leq n$.
-Then use A.1-A.2 to get $\{x_{n+1,j}^{X}\}_{j\leq n+1}\in F(X)$ for $X\in K_{n+1}$ and $z^{(n+1)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that for all $t\in[0,1]$
+    $\lVert\left((z_{1}^{(n)})^{*}\ldots(z_{1}^{(1)})^{*}\rhd y_{j}^{X}\lhd z_{1}^{(1)}\ldots z_{1}^{(n)}\right)-h^{X}(x_{n,j}^{X})\rVert\leq 2^{-n}$ for $X\in K_{n}$ and $1\leq j\leq n$.
+Then use A.1-18 to get $\{x_{n+1,j}^{X}\}_{j\leq n+1}\in F(X)$ for $X\in K_{n+1}$ and $z^{(n+1)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that for all $t\in[0,1]$
   1. (i)
-    $\|z_{t}^{(n+1)}\rhd h^{X}(x_{j}^{X})\lhd(z_{t}^{(n+1)})^{*}-h^{X}(x_{j}^{X})\|\leq 2^{-(n+1)}$ for $X\in K_{n+1}$ and $11\leq j\leq n+$ 1,
+    $\|z_{t}^{(n+1)}\rhd h^{X}(x_{j}^{X})\lhd(z_{t}^{(n+1)})^{*}-h^{X}(x_{j}^{X})\|\leq 2^{-(n+1)}$ for $X\in K_{n+1}$ and $1\leq j\leq n+1$,
   2. (ii)
-    $\|z_{t}^{(n+1)}\rhd h^{X}(x_{m,j}^{X})\lhd(z_{t}^{(n+1)})^{*}-h^{X}(x_{m,j}^{X})\|\leq 2^{-(n+1)}$ for $X\in K_{n+1}$ and $11\leq j\leq m<n+$ 1,
+    $\|z_{t}^{(n+1)}\rhd h^{X}(x_{m,j}^{X})\lhd(z_{t}^{(n+1)})^{*}-h^{X}(x_{m,j}^{X})\|\leq 2^{-(n+1)}$ for $X\in K_{n+1}$ and $1\leq j\leq m<n+1$,
   3. (iii)
-    $\|(z_{1}^{(n+1)})^{*}\ldots(z_{1}^{(1)})^{*}\rhd y_{j}^{X}\lhd z_{1}^{(1)}\ldots z_{1}^{(n+1)}-h^{X}(x_{n+1,j}^{X})\|\leq 2^{-(n+1)}$ for $X\in K_{n+1}$ and $11\leq j\leq n+$ 1.
-We carry on inductively to construct $z^{(n)}$ and $x_{m,j}$ for every $n,m\in\mathbb{N}$ and $j\leq m$ satisfying S.1-S.3.
+    $\|\left((z_{1}^{(n+1)})^{*}\ldots(z_{1}^{(1)})^{*}\rhd y_{j}^{X}\lhd z_{1}^{(1)}\ldots z_{1}^{(n+1)}\right)-h^{X}(x_{n+1,j}^{X})\|\leq 2^{-(n+1)}$ for $X\in K_{n+1}$ and $1\leq j\leq n+1$.
+We carry on inductively to construct $z^{(n)}$ and $x_{m,j}^{X}$ for every $n,m\in\mathbb{N}$ and $j\leq m$ satisfying S.1-S.3.
 We may now define the path $v_{t}:[0,\infty)\rightarrow\mathcal{U}(\mathcal{M}(B))$ by $v_{t}=z_{1}^{(1)}\ldots z_{1}^{(n)}z_{t-n}^{(n+1)}$ for every $t\in[n,n+1]$.
@@
 The maps $H^{X}:F(X)\rightarrow G(X)$ are linear as inherited by the linearity of $h_{t}^{X}$.
-Letting $X=1_{\mathcal{C}}$ the map $h^{1_{\mathcal{C}}}$ coincides with $\varphi$ and we get that $\Psi(a):=\lim_{t\rightarrow\infty}\mathrm{Ad}(v_{t})\circ\varphi(a)=\lim_{t\rightarrow\infty}\psi_{t}(a)$ exists for all $a\in A$.
-As each $\psi_{t}$ is a $*{}^{*}$-homomorphism, then so is $\Psi$.
-In light of Remark 3.12, it suffices to check that the family of maps $\{H^{X}\}_{X\in\mathrm{Irr}(\mathcal{C})}$ satisfies the conditions (i)-(v) of Lemma 3.11 to conclude that $(\Psi,H)$ is a cocycle morphism.
-Condition (v) follows by construction, whereas conditions (i)-(iv) are easily verified as $(\psi_{t},h_{t})$ is a cocycle morphism for every $t\in[0,\infty)$ and one may approximate $\Psi$ pointwise by $\psi_{t}$ and $H$ pointwise by $h_{t}$ for some large enough $t$.
-It remains to check that $\Phi$ is bijective and $H^{X}$ is bijective for every $X\in\mathcal{C}$.
-By Remark 3.12 and by taking $X=1_{\mathcal{C}}$, it suffices to check that $H^{X}$ is bijective for every $X\in\mathrm{Irr}(\mathcal{C})$.
+In light of Remark 3.12, it suffices to check that the family of maps $\{H^{X}\}_{X\in\mathrm{Irr}(\mathcal{C})}$ satisfies the conditions (ii)-(iv) of Lemma 3.10 to conclude that $(\Psi,H)$ is a cocycle morphism.
+They are easily verified as $(\psi_{t},h_{t})$ is a cocycle morphism for every $t\in[0,\infty)$ and one may approximate $H$ pointwise by $h_{t}$ for some large enough $t$.
+It remains to check that $H^{X}$ is bijective for every $X\in\mathcal{C}$.
+By Remark 3.12, it suffices to check that $H^{X}$ is bijective for every $X\in\mathrm{Irr}(\mathcal{C})$.
 Firstly, each $H^{X}$ is isometric.
-To see this, note that $\Psi$ is an injective $*{}^{*}$-homomorphism and hence isometric, so as $(\Psi,H)$ is a cocycle morphism
+To see this, note that $\Psi$ is an injective -homomorphism and hence isometric, so as $(\Psi,H)$ is a cocycle morphism
 $\|H^{X}(x)\|^{2}$ $=\|\langle H^{X}(x),H^{X}(x)\rangle_{B}\|$
@@
 $\leq\sum_{k=n}^{\infty}2^{-(k+1)}$
-$\leq 2^{1-n},$
+$=2^{-n},$
 for any $n\geq n_{0}$.
-So $\|y_{j}^{X}-H^{X}(x_{n,j}^{X})\|\leq 2^{1-n}+2^{-n}$.
+So $\|y_{j}^{X}-H^{X}(x_{n,j}^{X})\|\leq 2^{1-n}$.
 Therefore, as $n$ may be chosen arbitrarily and $\{y_{j}^{X}\}_{j\in\mathbb{N}}$ is dense in $G(X)$, it follows that $H^{X}$ is surjective.
@@
 Theorem 7.2.
-Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects.
-Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $*{}^{*}$-algebras and $(\varphi,h):(A,F,J)\rightarrow(B,G,I)$ an injective cocycle morphism.
+Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects.
+Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable C-algebras and $(\varphi,h):(A,F,J)\rightarrow(B,G,I)$ an injective cocycle morphism.
 Then $(\varphi,h)$ is approximately unitarily equivalent to a cocycle conjugacy if and only if:
@@
 Lemma 7.3.
-Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects.
+Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects.
 Let $(F_{n},J^{(n)}):\mathcal{C}\curvearrowright A_{n}$ be a sequence of actions on separable $\mathrm{C}^{*}$-algebras.
@@
   2. (ii)
-    $\mathrm{dist}(y_{1}^{*}\rhd\eta\lhd y_{1},h_{1,n}^{X}(F_{1}(X)))\leq\varepsilon$ for any $X\in K$, $\eta\in\mathcal{F}_{n}^{X}$
+    $\mathrm{dist}(y_{1}^{*}\rhd\eta\lhd y_{1},h_{1,n}^{X}(F_{1}(X)))\leq\varepsilon$ for any $X\in K$, $\eta\in\mathcal{F}_{n}^{X}$ .
 Let $z_{t}=\phi_{n,\infty}^{\dagger}(y_{t})$ be a unitary in $\mathcal{M}(A)$ for any $t\in[0,1]$, where $\phi_{n,\infty}^{\dagger}:\mathcal{U}(\mathcal{M}(A_{n}))\to\mathcal{U}(\mathcal{M}(A))$ is as in Definition 1.4.
@@
 Theorem 7.4.
-Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects.
+Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects.
 Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras.
@@
 fitting into the family of commuting diagrams
-[…]
+(7.4) $\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${F(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h^{X}}$$\scriptstyle{r^{X}}$${F(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h^{X}}\textstyle{\dots}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{l^{X}}$$\scriptstyle{s^{X}}$${G(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{l^{X}}$$\scriptstyle{s^{X}}$${\dots\quad.}$
 We can then form the inductive limits
@@
 such that $(\phi_{\infty},h_{\infty})\circ(\kappa_{\infty},r_{\infty})=(\theta_{\infty},s_{\infty})\circ(\phi,h)$ and $(\psi_{\infty},l_{\infty})\circ(\theta_{\infty},s_{\infty})=(\kappa_{\infty},r_{\infty})\circ(\psi,l)$.
-Furthermore, the cocycle morphism $(\kappa,r)$ is asymptotically unitarily equivalent to the cocycle morphism induced by the identity map on $A$, so Lemma 6.3 gives that $(\kappa_{\infty},r_{\infty})$ is asymptotically unitarily equivalent to a cocycle conjugacy $(K,R):(A,F,J)\to(A^{(\infty)},F^{(\infty)},J^{(\infty)})$.
+Furthermore, the cocycle morphism $(\kappa,r)$ is asymptotically unitarily equivalent to the cocycle morphism induced by the identity map on AA, which in particular implies that $(\kappa,r)$ is injective and thus Lemma 7.3 gives that $(\kappa_{\infty},r_{\infty})$ is asymptotically unitarily equivalent to a cocycle conjugacy $(K,R):(A,F,J)\to(A^{(\infty)},F^{(\infty)},J^{(\infty)})$.
 Likewise, $(\theta_{\infty},s_{\infty})$ is asymptotically unitarily equivalent to a cocycle conjugacy $(\Theta,S):(B,G,I)\to(B^{(\infty)},G^{(\infty)},I^{(\infty)})$.
@@
 Similarly, if we take $(\Psi,L)=(\Phi,H)^{-1}=(K,R)^{-1}\circ(\psi_{\infty},l_{\infty})\circ(\Theta,S)$, we get that $(\Psi,L)\cong_{u}(\psi,l)$, which finishes the proof. ∎
+Remark 7.5.
+Note that in Theorem 7.4, if the asymptotic unitary equivalences are realised by unitaries in the minimal unitisations, then we may drop the assumption of extendibility on the connecting morphisms.
+This follows as extendibility is only required to evaluate a morphism on a given unitary in the multiplier algebra.
+However, there is no issue in doing so when the unitary is in the minimal unitisation.
 Acknowledgements
 We would like to thank Stuart White and Samuel Evington for their supervision on this project.
-We would also like to thank George Elliott and Gabór Szabó for useful discussions on the topic of this paper.
+We would also like to thank Corey Jones, George Elliott, and Gabór Szabó for useful discussions on the topic of this paper.
 Part of this work was completed during the authors’ stay at the Fields Institute for Research in Mathematical Sciences for the ’Thematic Program on Operator Algebras and Applications’ in Autumn 2023.
 We thank the Fields Institute and the organisers for the hospitality.
-The first named author was supported by the Ioan and Rosemary James Scholarship awarded by St John’s College and the Mathematical Institute, University of Oxford, as well as by project G085020N funded by the Research Foundation Flanders (FWO).
-The second named author was supported by the EPSRC grant EP/R513295/1.
+The first named author was supported by the Ioan and Rosemary James Scholarship awarded by St John’s College and the Mathematical Institute, University of Oxford, as well as by projects G085020N and 1249225N funded by the Research Foundation Flanders (FWO).
+The second named author was supported by the EPSRC grant EP/R513295/1, by the European Research Council under the European Union’s Horizon Europe research and innovation programme (ERC grant AMEN–101124789), and by the postdoctoral fellowship 1204626N of the Research Foundation Flanders (FWO).
 The authors’ stay at the Fields Institute was partially funded by a Special Grant awarded by St John’s College, Oxford.
 The first named author’s stay was also partially supported by the Fields Institute while the travel costs of the second named author were supported by the EPSRC grant EP/X026647/1.
+For the purpose of open access, the authors have applied a $C$ BY public copyright license to any author accepted manuscript version arising from this submission.